A Complete Bibliography of Publications in The Computer Journal: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 2.04

Title word cross-reference

2 [?, ?]. 7 [?]. ^{link} [?]. A_{k} [?]. β [?]. F [?]. h [?]. k [?, ?, ?, ?, ?]. L [?]. LU [?]. m [?]. μ [?, ?]. n [?]. O(1) [?]. p [?]. π [?]. τ [?]. X [?]. Z [?].

-adic [?]. -ary [?, ?, ?]. -bases [?]. -bit [?]. -calculus [?]. -code [?]. -cubes [?]. -D [?]. -Dimensional [?]. -restricted [?]. -Set [?]. -system [?]. -Tree [?]. -Way [?].

1 [?]. 1-2-3 [?]. 1990 [?]. 1991 [?].

2000 [?].

3rd [?].

77 [?].

95 [?, ?].

absolute [?]. Abstract [?, ?]. Abstraction [?]. Abstractions [?]. Abstracts [?]. Access [?, ?, ?, ?, ?, ?]. Accesses [?, ?, ?]. accommodate [?]. Achieving [?]. ACLE [?]. Action [?, ?]. Actions [?, ?]. Active [?]. actual [?]. acyclic [?]. Ada [?, ?, ?, ?]. Adapting [?]. Adaptive [?, ?, ?, ?, ?, ?, ?]. Adding [?]. Addition [?]. Additional [?, ?, ?, ?, ?, ?, ?, ?]. Additions [?]. Addressing [?, ?]. Adele [?]. adic [?]. adjusting [?]. Admissible [?]. Admission [?]. Advanced [?, ?]. Affine [?, ?]. against [?]. aggregation [?]. agreement [?]. AI [?, ?]. Aid [?]. Air [?]. Air-Storage [?]. Algebra [?, ?, ?, ?, ?, ?, ?, ?, ?]. Algebraic [?, ?, ?, ?, ?, ?]. Algebras [?, ?, ?, ?, ?]. Algol [?]. Algorithm [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Algorithmic [?, ?]. Algorithms [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Alignment [?]. Alive [?]. All-sharing [?]. Allocation [?, ?]. Alphabets [?]. Alternative [?]. Amoeba [?]. Amoeba-Prolog [?]. analog [?]. Analogy [?]. Analysing [?]. Analysis [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Analytic [?]. Analyzing [?]. AND-Parallel [?]. Animating [?]. Animation [?]. Announcement [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Announcements [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Answering [?]. Answers [?]. Application [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Applications [?, ?, ?, ?, ?, ?, ?, ?, ?]. Applying [?, ?, ?, ?]. Appraisal [?]. Approach [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Approaches [?, ?, ?]. Approximate [?, ?, ?, ?, ?]. Approximation [?]. Arbitrary [?]. ARC [?]. Architects [?]. Architectural [?]. Architecture [?, ?, ?, ?, ?, ?, ?, ?]. Architectures [?, ?, ?, ?, ?]. Area [?, ?, ?]. Arithmetic [?, ?]. ARP [?]. Arrangement [?]. arrangements [?]. Array [?, ?, ?, ?, ?, ?, ?]. Arrays [?, ?, ?, ?]. Art [?, ?, ?]. Artificial [?]. Artificially [?]. ary [?, ?, ?]. Aspects [?, ?]. assertion [?]. assertion-guided [?]. Assignment [?]. Assistance [?]. Assisted [?]. Associated [?]. Association [?]. Associative [?, ?, ?]. Associative-Commutative [?]. Assurance [?]. Asymmetrical [?]. Asynchronous [?, ?]. Atomic [?]. Attribute [?, ?, ?, ?]. Attributes [?]. Australian [?]. Authenticating [?]. authentication [?]. Author [?]. Authoring [?]. Autodescriptivity [?]. Automata [?, ?, ?, ?]. Automate [?]. Automated [?, ?, ?, ?, ?, ?]. Automatic [?, ?, ?]. Automating [?]. automaton [?]. automaton-based [?]. Availability [?]. Available [?, ?, ?, ?]. Average [?, ?]. Award [?, ?]. axioms [?].

B [?, ?, ?, ?]. B-tree [?]. B-Trees [?, ?]. B. [?]. Backbone [?]. Backward [?]. Balanced [?, ?]. Balancing [?, ?, ?, ?]. Bandwidth [?]. Base [?, ?]. Based [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Baseline [?]. Bases [?, ?]. Basic [?, ?, ?, ?]. Basis [?]. Bayesian [?, ?, ?]. BD [?]. be [?]. become [?, ?]. Beginners [?]. beginning [?]. Behaviour [?, ?]. Behavioural [?, ?]. Behaviours [?]. Belief [?, ?]. Believed [?]. Belong [?]. Benchmarks [?]. Bendix [?]. Beta [?]. Beta-splines [?]. Between [?, ?, ?]. beware [?]. Bidirectional [?]. Binary [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Binary-Residue [?]. biology [?]. Bipartite [?]. bit [?, ?, ?]. Block [?, ?, ?, ?, ?, ?, ?]. Block-Structured [?, ?]. Blue [?]. Board [?]. Book [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Book [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Book [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Boolean [?, ?]. boundary [?]. boundary-restricted [?]. Bounded [?]. Bounding [?]. Bounds [?, ?]. Boyer [?]. Branch [?, ?, ?, ?]. branching [?, ?]. bridge [?]. bridge- [?]. British [?, ?, ?]. BRMQ [?]. Broadcast [?, ?]. Broadcasting [?, ?]. Bruijn [?]. buffered [?]. Building [?, ?, ?, ?, ?]. Burrows [?]. Bus [?, ?]. Bus-Based [?]. bus-connected [?]. Business [?, ?, ?, ?, ?]. bypass [?]. Byrne [?]. Byrne-Jones [?].

C [?, ?, ?, ?, ?, ?, ?, ?, ?]. C-Based [?]. C-ODA [?]. Cache [?, ?, ?]. Caching [?, ?]. calculating [?]. Calculational [?]. Calculations [AB90]. Calculator [?, ?, ?]. Calculi [?]. Calculus [?, ?]. Call [?, ?, ?]. calls [?]. candidate [?]. capacity [?]. Cards [?]. Carlo [?]. carousel [?]. cartridges [?]. Case [?, ?, ?, ?, ?, ?, ?, ?, ?]. categorical [?, ?]. Causal [?, ?]. Causality [?]. Causality-Based [?]. CCS [?]. CD [?, ?]. CD-I [?]. CD-ROM [?]. CD-ROMs [?]. CDTV [?]. cell [?]. Cellular [?]. Central [?]. Chain [?]. Chaining [?]. Changes [?, ?]. Channel [?]. channels [?]. Character [?]. characterisation [?]. Characters [?]. Charles [?]. Charts [?, ?]. checked [?]. checking [?]. Chemical [?]. chip [?]. chordal [?]. Church [?]. Cilk [?]. Ciphers [?]. Circle [?]. Circuit [?, ?, ?]. circuits [?, ?]. City [?]. Clashes [?, ?]. Class [?, ?, ?, ?, ?]. Classical [?, ?]. Classification [?, ?, ?, ?, ?]. Classifications [?]. Clause [?]. Client [?, ?]. Client-Server [?, ?]. Clocks [?]. Clubs [?]. Cluster [?, ?, ?]. cluster-based [?]. Clustering [?, ?, ?, ?, ?, ?]. Clusters [?, ?, ?]. CMCM [?]. co [?]. co-simulation [?]. coated [?]. Code [?, ?, ?, ?, ?, ?, ?]. codec [?]. Codewords [?]. CODIL [?]. Coding [?, ?, ?, ?, ?, ?]. Codings [?]. Cognitive [?]. coherence [?]. Colantonio [?]. Collaboration [?]. collaborative [?]. Collection [?]. Collections [?]. Collins [?]. COMA [?]. Comandos [?, ?]. combinational [?]. Combinator [?]. Combinatorial [?, ?]. combinators [?]. Combining [?, ?]. command [?]. Comment [?, ?, ?]. Comments [?]. Committing [?]. Commonsense [?]. communicating [?]. Communication [?, ?, ?, ?, ?, ?]. Communication-Minimal [?]. Communications [?, ?, ?]. Commutative [?]. Comparative [?, ?, ?, ?, ?]. Comparing [?]. Comparison [?, ?, ?, ?, ?, ?, ?]. Compatible [?]. Competition [?]. Competitive [?, ?]. Compilation [?]. Compile [?, ?]. Compile-Time [?, ?]. Compiler [?, ?, ?]. compiling [?]. complete [?, ?]. Completeness [?, ?, ?, ?]. Completion [?]. Complex [?, ?, ?, ?, ?, ?]. Complexity [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. components [?]. Composition [?, ?]. Compositional [?]. Comprehension [?]. compressed [?]. Compressing [?, ?]. Compression [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. computability [?]. Computation [?, ?, ?, ?, ?, ?, ?]. Computational [?, ?, ?, ?]. Computations [?, ?]. compute [?]. Computer [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Computers [?, ?, ?, ?]. Computing [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Concept [?, ?]. Concepts [?, ?]. Conceptual [?, ?, ?, ?]. Concurrency [?, ?, ?, ?, ?]. Concurrent [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. condition [?]. conditioned [?]. Conflicts [?, ?, ?]. connected [?, ?]. Connectivity [?]. Conquer [?]. Consequences [?]. Considerations [?]. Consistency [?, ?]. Consistent [?]. Constant [?, ?, ?]. Constrained [?, ?, ?, ?]. Constraint [?, ?, ?, ?, ?, ?, ?]. Constraint-Based [?]. constraint-driven [?]. Constraint-oriented [?]. Constraints [?, ?, ?, ?, ?]. Constructing [?, ?, ?]. Construction [?, ?, ?, ?, ?, ?, ?, ?]. Constructive [?]. Containing [?]. Contemporary [?]. Content [AA99]. contents [?]. Context [?, ?, ?, ?, ?, ?]. Context-Based [?, ?]. Context-Free [?]. Contexts [?]. contextual [?]. Continuations [?]. continuity [?]. Continuous [?]. Contouring [?]. Control [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Controlled [?]. controlling [?]. Convergence [?]. conversion [?]. Converters [?, ?]. Cooperative [?, ?, ?, ?]. Coordination [?, ?]. Copper [?]. Copyright [?]. Core [?]. Corporate [?]. Correct [?]. Corrective [?]. Correctness [?, ?, ?, ?]. Correlated [?]. Correspondence [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Corrigendum [?]. COSMOS [AA99]. COSMOS-Content [AA99]. Cost [?, ?]. Cost-Effectiveness [?]. Counterfactual [?]. Counting [?, ?, ?]. Course [?, ?]. Crash [?]. Creation [?]. Creations [?, ?]. Criterion [?]. Critical [?, ?, ?, ?, ?]. CRL [?, ?]. ‘Cross’ [?]. Crossword [?]. crunching [?]. Cryptanalysis [?]. Cryptosystem [?]. CSG [?]. CSMA [?]. CSMA/CD [?]. Cubes [?, ?, ?]. cubic [?]. culture [?]. Curricula [?, ?]. Custom [?]. cutting [?]. CyberCrime [?]. Cycle [?]. Cycle-Sort [?]. Cyclic [?, ?]. cylindrical [?].

D [?, ?, ?]. D. [?]. DAP [?]. DAPLEX [?]. Data [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Database [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Databases [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. datalog [?]. dBASE [?]. deadlock [?, ?]. deadlock-free [?]. Debugging [?]. decay [?]. Decentralising [?]. Decentralized [?]. Decidable [?]. Decision [?, ?]. Decisions [?]. declarative [?]. Decomposition [?, ?, ?, ?, ?]. Deductive [?, ?, ?]. Default [?]. defer [?]. definite [?]. definition [?]. definitions [?]. Degeneracy [?]. Degradable [?]. degrading [?]. degree [?]. degrees [?]. Demonstrating [?]. Dependable [?, ?]. dependences [?]. Dependencies [?, ?]. Dependency [?]. Depth [?, ?]. depth-first [?]. Derivation [?, ?]. Derived [?]. Describing [?]. Description [?, ?, ?, ?]. Design [?, AB90, ?, ACDP91, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Designing [?, ?, ?, ?, ?]. Designs [?]. Detecting [?, ?, ?, ?]. Detection [?, ?]. determination [?]. Determining [?]. Determinism [?]. deterministic [?, ?]. Develop [?]. developing [?]. Developmen [?]. Development [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Developments [?]. device [?]. Diagnosis [?, ?]. Diagrams [?]. Diakoptics [?]. Diameter [?]. DiCE [?]. dictation [?]. Dictionary [?, ?, ?]. difference [?]. digital [?, ?, ?]. digitised [?]. digraphs [?]. Dimensional [?, ?, ?]. Dimensions [?]. direct [?]. Directed [?]. directory [?]. directory-based [?]. Disc [?, ?]. Disclosure [?]. Discourse [?, ?, ?]. Discovering [?, ?]. Discretizer [?]. discs [?]. Discussion [?, ?, ?, ?]. disequations [?]. Disk [?, ?]. Dispatching [?]. Display [?, ?]. Dissemination [?, ?]. Dissimilarity [?]. Distance [?]. Distributed [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Distributed-Application [?]. Distributed-Memory [?]. Distributing [?]. distribution [?]. Distributions [?]. Divide [?]. Divide-and-Conquer [?]. Division [?, ?]. Division-free [?]. divisionless [?]. Document [?, ?]. Documents [?, ?]. Does [?, ?]. Doesn’t [?]. domain [?, ?]. dotted [?]. Double [?, ?, ?]. Double-ended [?, ?]. doubly [?]. doubly-linked [?]. Dowe [?, ?, ?]. Down [?, ?]. Downdating [?]. Drawability [?]. Drawings [?]. driven [?]. Drivers [?]. DROL [?]. drown [?]. duration [?]. Dynamic [?, ?, ?, ?, ?, ?, ?].

Early [?]. ECLAIR [?]. edge [?]. edge-solvable [?]. edition [?]. editor [?]. Editorial [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. edn [?]. eds [?]. Educating [?]. education [?]. effect [?]. Effective [?]. Effectiveness [?]. Effects [?]. Efficiency [?, ?, ?]. Efficient [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Eiffel [?]. Electrical [?]. Electronic [?, ?, ?, ?]. Eliminating [?].

Elimination [?, ?, ?, ?, ?, ?, ?, ?, ?]. ElipSys [?]. ELLA [?]. embedded [?]. embedding [?, ?]. EMPA [?]. Empirical [?, ?]. Encoding [?]. Encounter [?]. ended [?, ?]. Ends [?]. enforce [?]. Engineering [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. English [?]. Enhance [?]. Enhanced [?]. Enhancing [?]. Enough [?]. Ensembles [?]. entities [?]. entrant [?]. Enumeration [?]. Environment [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Environmental [?]. Environments [?, ?, ?, ?, ?, ?, ?, ?]. Envisioning [?]. equational [?]. equations [?, ?, ?]. Equipment [?]. Equivalence [?, ?]. Equivalent [?]. Era [?]. Ernest [?]. Errata [?]. Erratum [?, ?, ?]. Error [?, ?, ?]. ES-IS [?]. Essentials [?]. Establishing [?]. Estimation [?, ?]. Ethernet [?, ?]. European [?]. evaluate [?]. Evaluating [?]. Evaluation [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Events [?, ?]. Evidential [?]. Evolution [?, ?, ?, ?]. Evolution-Based [?]. exact [?]. examination [?, ?]. Examples [?, ?]. exception [?]. Exceptions [?, ?]. Exchange [?]. exchanges [?]. Exclusion [?, ?, ?]. Executable [?, ?]. Executing [?, ?]. Executing-Committing [?]. Execution [?, ?, ?, ?]. Executions [?]. exercise [?]. existential [?, ?]. Expansion [?]. Expected [?, ?]. Experiences [?]. Experimental [?]. Expert [?, ?]. Explanation [?]. Exploiting [?, ?, ?, ?, ?]. exploration [?]. Exploratory [?]. Exploring [?]. Exposing [?]. Expression [?]. Expressions [?]. Expressive [?]. Expressiveness [?]. Extended [?, ?, ?, ?, ?, ?, ?]. Extending [?, ?, ?]. Extensible [?]. External [?]. extra [?]. extraction [?]. extremal [?]. Extreme [?]. Extreme-value [?].

facilitate [?]. facility [?]. factorisation [?]. Factors [?]. Facts [?]. Failure [?, ?, ?]. failures [?]. Fair [?]. Family [?, ?]. Fast [?, ?, ?, ?, ?, ?]. faster [?]. Fat [?]. Fault [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Fault-based [?]. Fault-Secure [?]. Fault-Tolerant [?, ?, ?, ?, ?]. Faults [?]. faulty [?]. FC [?]. Fddi [?, ?, ?]. FDDI-M [?]. FDL [?]. Feature [?, ?, ?, ?]. Features [?]. Federating [?]. Feedback [?]. Fibonacci [?]. Fields [?]. FIFO [?]. File [?, ?, ?, ?]. Files [?, ?, ?, ?]. Filter [?]. Filtering [?]. Finding [?, ?, ?, ?, ?, ?, ?]. Finite [?, ?, ?, ?, ?]. Finite-State [?, ?]. First [?, ?]. First-Order [?]. Fixed [?]. Flat [?]. Flexibility [?, ?]. Flexible [?, ?, ?, ?, ?, ?]. Floating [?]. Flow [?, ?, ?]. Flowers [?]. Foresight [?]. Foreword [?]. form [?]. Formal [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. formulas [?]. formulation [?]. FORTRAN [?, ?, ?]. Forward [?, ?]. Forwarding [?]. Foundation [?, ?, ?]. Foundations [?, ?]. four [?, ?]. Fourier [?]. Fractal [?, ?, ?, ?]. Framework [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Frameworks [?]. Free [?, ?, ?, ?, ?]. Full [?]. Full-text [?]. fully [?]. Function [?, ?]. Functional [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Functions [?, ?, ?, ?, ?, ?, ?, ?]. Fundamentals [?]. FunZ [?]. Fusion [?]. Fuzzy [?].

G [?]. Garbage [?]. Gassilloud [?]. gate [?]. Gaussian [?, ?]. Gem [?].

General [?, ?, ?, ?, ?]. generalized [?, ?]. Generate [?, ?]. Generating [?, ?]. Generation [?, ?, ?, ?, ?]. generative [?]. generator [?]. Generators [?, ?]. Genetic [?, ?]. Geographic [?]. Geometric [?, ?, ?, ?]. geometry [?]. GIS [?]. Glass [?]. Global [?]. globe [?]. Goldfarb [?]. goto [?]. gracefully [?]. Grammar [?, ?]. Grammar-Oriented [?]. Grammars [?, ?, ?]. Graph [?, ?, ?, ?, ?, ?]. Graph-Based [?]. Graph-Reducer [?]. Graphical [?, ?, ?, ?]. Graphics [?, ?, ?]. Graphs [?, ?, ?, ?, ?]. Gray [?]. Greedy [?]. Greeting [?]. Grid [?, ?, ?]. Grid-based [?, ?]. GRIDS [?]. Grossetie [?]. groups [?]. growth [?]. growth-decay [?]. Guaranteeing [?]. Guard [?]. Guide [?, ?, ?, ?]. guided [?]. Guidelines [?].

Hairy [?]. Handbook [?]. Handling [?, ?, ?, ?]. Hanoi [?]. hard [?]. Hardware [?, ?, ?, ?]. hardware/software [?]. HASE [?]. Hash [?, ?, ?]. hash-based [?, ?]. Hashing [?, ?, ?, ?, ?, ?]. HDG [?]. HDG-Machine [?]. headed [?]. Heap [?]. Heaps [?]. Heapsort [?, ?, ?]. Height [?]. Helene [?]. Hermitian [?]. Heterogeneity [?, ?]. Heterogeneous [?, ?, ?, ?, ?, ?]. Heuristic [?, ?]. heuristics [?]. HIDEL [ACDP91]. Hierarchic [?]. Hierarchical [ACDP91, ?, ?, ?, ?, ?, ?, ?]. High [?, ?, ?, ?]. high-level [?]. Higher [?, ?, ?, ?, ?, ?, ?]. Higher-Order [?, ?, ?, ?, ?]. Highly [?, ?]. Historical [?]. History [?, ?]. HOL [?, ?, ?, ?, ?]. holds [?]. Hope [?]. Hopfield [?]. Horn [?]. HPP [?]. Hubs [?]. Huffman [?, ?]. Human [?, ?, ?, ?, ?]. Human-Computer [?]. Humans [?]. Hybrid [?, ?, ?, ?, ?, ?]. Hypercube [?, ?]. Hypercube-Labeling [?]. Hypercubes [?, ?, ?, ?]. Hypermedia [?, ?]. hypermesh [?]. hypermeshes [?]. Hypernet [?, ?]. Hypertext [?, ?]. Hypothesis [?].

I-Pigs [?]. IC [?]. ICL [?]. ICN [?]. Icon [?]. Idea [?]. Ideas [?, ?]. Identification [?, ?]. Identifiers [?]. ill [?]. ill-conditioned [?]. Illingworth [?]. ILP [?]. Image [?, ?, ?, ?, ?, ?, ?]. images [?, ?, ?, ?]. immediate [?]. Impact [?, ?]. Imperfect [?]. Implement [?]. Implementation [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Implementations [?, ?]. Implementing [?]. implications [?]. implicit [?, ?]. Imply [?]. Imprecise [?, ?]. Improve [?]. Improved [?, ?, ?]. Improvement [?, ?]. Improvements [?, ?]. Improving [?, ?, ?]. In-Place [?]. Incomplete [?, ?]. Incompressibility [?]. Incremental [?, ?]. Independent [?, ?]. index [?, ?]. Indexing [?, ?, ?]. Indicators [?]. Induction [?, ?]. Inductive [?]. industrial [?]. Industry [?, ?]. inequalities [?]. Infection [?, ?, ?, ?]. inference [?]. Informatics [?]. Information [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Informatization [?]. inherent [?]. injection [?]. Innovation [?]. Innovations [?]. instances [?]. institution [?]. Integers [?]. Integrated [?, ?, ?, ?, ?, ?]. Integrating [?, ?, ?]. Integration [?, ?]. Integrity [?]. Intelligence [?]. Intelligent [?]. inter [?]. inter-media [?]. Interaction [?, ?, ?, ?, ?, ?]. Interactive [?, ?, ?, ?]. Interconnect [?]. Interconnection [?, ?, ?, ?, ?, ?]. Interface [?, ?, ?]. Interfaces [?]. Interfacing [?]. Interference [?]. Interior [?]. Intermediate [?]. Internal [?]. International [?]. Internet [?]. Interpolation [?]. Interpretation [?]. Interpreting [?]. Intersection [?, ?, ?, ?, ?]. Interval [?, ?, ?, ?, ?, ?]. intervals [?]. intra [?]. intra-media [?]. Intrinsic [?]. Introducing [?]. Introduction [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Inversion [?]. Invertibility [?, ?]. Investment [?]. Investments [?]. invocation [?]. irregular [?]. IRS [?]. ISO [?]. Issue [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Issues [?, ?]. Iteration [?, ?, ?]. Iterative [?, ?, ?, ?, ?]. IV [?].

J [?]. Jackson [?, ?, ?]. Jape [?]. Jigsaw [?]. Job [?]. Join [?, ?]. Joins [?]. Jones [?]. Just [?].

Karp [?]. kernel [?]. Kershner [?]. key [?, ?]. keys [?]. Kinds [?]. km [?]. knitting [?]. Knowledge [?, ?, ?, ?]. Knuth [?, ?]. Kolmogorov [?, ?, ?, ?, ?]. Kron [?].

L [?]. Label [?]. Labeling [?]. Labelling [?]. Laboratory [?]. Lambda [?]. Language [ACDP91, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Languages [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. LANs [?]. Large [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Large-Scale [?, ?, ?]. Latent [?]. Lattices [?]. Law [?, ?]. Layered [?]. layout [?]. Lazy [?]. LCF [?]. Learning [?, ?, ?]. Least [?]. Least-Squares [?]. Lectures [?]. left [?]. left-to-right [?]. Legibility [?]. Length [?, ?, ?, ?, ?]. Length-limited [?]. lengths [?]. Lenient [?]. LEO [?]. Letters [?, ?, ?]. Level [?, ?]. Lexical [?]. lexicographical [?, ?]. Librarian [?]. Library [?]. limited [?]. line [?]. Linear [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. lines [?]. Link [?]. linkage [?]. Linked [?, ?]. Linked-List [?]. Links [?]. Lisp [?]. List [?, ?, ?]. literacy [?, ?]. Literal [?]. live [?]. Load [?, ?, ?]. Load-balancing [?]. Locality [?]. locally [?]. Location [?, ?, ?, ?]. Locking [?, ?]. Lof [?]. logarithmic [?]. Logbooks [?]. Logic [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Logic-based [?]. Logical [?, ?, ?, ?, ?, ?]. long [?]. long-duration [?]. Longest [?]. Loop [?, ?, ?, ?]. Loops [?, ?, ?, ?]. loss [?]. Lossless [?, ?, ?, ?, ?]. lossy [?]. lost [?]. Lotus [?]. Luby [?].

M [?, ?]. M-way [?]. Machine [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. machine-checked [?]. Machine-Independent [?]. Machines [?, ?, ?]. macro [?]. magic [?]. Mainframe [?]. Mainstream [?]. Maintaining [?]. Maintenance [?, ?, ?, ?]. Majority [?, ?, ?]. Mäkinen [?, ?]. Management [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Managerial [?, ?]. Managing [?, ?]. Manipulating [?]. Manipulation [?, ?]. manufacturing [?]. Many [?, ?]. Many-Sorted [?]. Mapping [?]. Mark [?]. Markov [?]. Markovian [?, ?]. Martin [?]. Martin-Lof [?]. mass [?, ?]. Masses [?]. Match [?, ?]. Matching [?, ?, ?, ?, ?, ?, ?, ?]. Mathematical [?, ?]. Mathematics [?]. matrix [?, ?]. Maturing [?]. McPOPLOG [?]. MDL [?, ?]. Meaning [?]. means [?]. Measure [?]. Measurement [?]. Measurements [?, ?]. Measures [?]. Measuring [?]. Mechanical [?, ?]. mechanically [?]. Mechanisms [?, ?]. Mechanizing [?]. Media [?, ?, ?]. Medium [?]. Memories [?, ?]. Memory [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Mergeable [?]. Merging [?, ?, ?]. merits [?]. Meshes [?, ?, ?, ?]. Message [?, ?, ?, ?, ?]. Messages [?, ?, ?]. Messaging [?]. Meta [?]. Meta-Message [?]. metal [?]. Method [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Methodologies [?, ?, ?]. Methods [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Metri [?]. Microsoft [?]. Millar [?, ?]. Minimal [?, ?, ?]. Minimizing [?, ?]. Minimum [?, ?]. miss [?]. Mixed [?, ?]. Ml [?, ?, ?]. MML [?]. Mobile [?, ?, ?, ?, ?, ?, ?, ?, ?]. Mode [?, ?]. Model [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Model-Based [?]. Modeling [?]. Modelling [AA99, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Models [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. modification [?, ?]. Modular [?, ?]. Module [?]. modulo [?]. molecular [?]. Monitoring [?]. Monte [?]. Moore [?]. Motif [?]. Motion [?]. Motivated [?]. mouse [?]. MP1 [?]. Mu [?]. Mu-balancing [?]. Multi [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. multi-attribute [?]. Multi-Combinator [?]. Multi-Label [?]. multi-media [?]. Multi-Partition [?]. multi-stage [?, ?]. Multi-Transputer [?]. multi-user [?]. Multi-valued [?, ?]. Multi-View [?]. multiattribute [?]. Multicast [?, ?]. Multicomputers [?, ?, ?]. Multikey [?]. Multilanguage [?]. multilevel [?]. multilist [?]. Multimedia [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Multipath [?, ?]. Multiple [?, ?, ?, ?, ?, ?, ?]. MultiPolynomial [?]. Multiprocessor [?, ?, ?, ?, ?, ?, ?, ?]. Multiprocessors [?, ?]. multiscale [?]. Multithreading [?]. Muse [?]. Mutual [?, ?, ?, ?]. Myths [?].

Narrowing [?]. Natural [?, ?, ?, ?, ?]. natural-language [?]. Navigation [?]. Near [?]. Necessary [?]. Neighbour [?]. Neighbour-Greeting [?]. Nested [?, ?, ?, ?, ?]. Nesting [?]. Net [?, ?, ?]. nets [?, ?]. Network [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Networks [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Neural [?, ?, ?, ?, ?, ?]. neurons [?]. No [?, ?]. node [?, ?, ?]. node-to-node [?, ?]. Nodes [?, ?]. noisy [?, ?]. Non [?, ?, ?, ?, ?, ?, ?]. Non-Constant [?]. Non-Determinism [?]. Non-Executable [?]. non-linear [?]. Non-repudiation [?]. non-strict [?]. non-uniform [?]. nondeterministic [?, ?]. Nonmonotonic [?]. Notation [?]. Note [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Notes [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Notice [?]. Novel [?]. NRDC [?]. Number [?, ?, ?, ?]. numerical [?].

Object [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Object-Oriented [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Objects [?, ?, ?, ?, ?]. Occam [?, ?]. Occam- [?]. Occupancy [?]. octrees [?, ?]. Octtree [?]. ODA [?, ?]. Odds [?]. Office [?, ?, ?]. Offices [?, ?]. offs [?]. Oggetto [?]. One [?, ?]. One-bit [?, ?]. OO [?]. Open [?, ?, ?]. Operating [?, ?]. Operations [?, ?, ?]. operators [?]. OPLA [?]. Optimal [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Optimality [?, ?]. Optimally [?]. Optimisation [?, ?, ?]. Optimising [?, ?]. Optimization [?, ?, ?]. Optimized [?]. Optimizing [?, ?]. options [?]. Ord [?]. Ord-Smith [?]. Order [?, ?, ?, ?, ?, ?, ?, ?, ?]. Order-sorted [?]. ordered [?]. Organising [?]. organizational [?]. organizing [?]. Oriented [AA99, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Origins [?]. Orthogonal [?, ?, ?]. OSSAD [?]. Overcoming [?]. overflow [?]. Overlap [?]. Overlapping [?, ?]. Overlay [AA99]. Overview [?, ?].

Package [?]. packet [?]. Packetized [?]. Page [?]. Paige [?]. paper [?]. Papers [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Paradigm [?, ?]. Parallel [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Parallelism [?, ?, ?, ?]. Parallelization [?, ?, ?]. Parameter [?]. parameterised [?]. Parameters [?, ?]. Parametric [?]. Parametrized [?]. Parsing [?]. Part [?, ?]. Partial [?, ?]. Partial-Match [?]. Partiality [?]. partially [?]. Partition [?, ?]. Partitioning [?, ?, ?, ?, ?]. Partitions [?]. Pascal [?, ?]. Pascal-FC [?]. password [?]. Path [?, ?, ?, ?]. Paths [?]. Pattern [?, ?, ?]. Pattern-Matching [?]. PCTE [?]. Pebble [?]. Pebble-crunching [?]. Penalty [?]. PEPA [?]. Perfect [?, ?, ?, ?]. Perfectly [?]. Performability [?]. Performance [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Performance-Based [?]. Periodic [?, ?]. Permutation [?, ?, ?]. Permutations [?]. persistence [?]. Persistent [?, ?]. personal [?]. Perspective [?, ?, ?]. perturbation [?]. Peter [?]. Peterson [?]. Petri [?, ?, ?]. Pexlib [?]. Philosophy [?]. Physical [?]. pictures [?]. Pigs [?]. pipe [?]. pipeline [?, ?]. Pipelined [?, ?]. Pipelines [?]. Pipelining [?, ?, ?]. Place [?]. Plagiarism [?]. Planning [?, ?, ?]. Platform [?, ?]. plus [?]. Poet [?]. Point [?, ?, ?]. points [?]. Policies [?, ?]. Polygon [?]. Polyhedron [?, ?]. Polymorphism [?]. Polynomial [?, ?, ?]. polynomial-time [?]. Polytechnics [?]. Populations [?]. Portable [?]. Position [?]. Postmodern [?]. Power [?, ?]. Powerful [?]. PPM [?, ?]. PPMB [?]. Practical [?, ?, ?, ?, ?, ?, ?]. Practice [?, ?, ?]. Practitioner [?]. pragmatic [?]. PRAM [?]. PRAMs [?]. PRAXIS [AB90]. Precedence [?]. Precedence-Constrained [?]. Precise [?]. Precondition [?]. Predicates [?]. Predictable [?]. Prediction [?, ?, ?]. Predictive [?]. preferences [?]. Prefix [?]. preprocessing [?]. prescribed [?]. Presentations [?, ?]. Presumed [?]. Pricing [?]. primality [?]. prime [?]. Principle [?, ?]. principles [?]. Print [?]. Prioritizing [?]. Priority [?, ?, ?, ?, ?]. Probabilistic [?, ?, ?, ?, ?, ?, ?]. probability [?]. Probe [?]. Probing [?, ?]. Problem [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Problems [?, ?, ?, ?, ?, ?]. Procedural [?]. Procedure [?, ?, ?]. procedures [?]. Proceedings [?]. Process [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Processes [?, ?]. Processin [?]. Processing [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Processor [?, ?, ?]. Processors [?, ?, ?]. Produce [?]. Produced [?, ?]. Product [?, ?, ?]. production [?, ?]. Proforma [AB90]. Program [AB90, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Programme [?]. Programmers [?]. Programming [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Programs [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Progress [?]. Project [?]. projections [?]. Prolog [?, ?, ?, ?, ?, ?, ?]. PROM [?]. Proof [?, ?, ?, ?]. Proofs [?]. Properties [?, ?, ?, ?, ?]. Property [?, ?]. Proportional [?, ?]. Proposed [?, ?]. Prosthesis [?]. Protecting [?]. protection [?]. Protocol [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Protocols [?, ?, ?, ?]. Prototyping [?, ?]. Prototyping-oriented [?]. Provably [?]. Proving [?]. provision [?]. Proxies [?]. Proximity [?, ?]. PS [?]. PS-Algol [?]. pseudo [?]. pseudo-lexicographical [?]. Public [?]. purely [?]. Puzzle [?, ?, ?].

quadratic [?]. Quadtree [?, ?, ?, ?]. Quadtree-Based [?]. Quadtrees [?, ?]. Quality [?, ?, ?]. Quantifier [?, ?, ?, ?, ?, ?, ?]. Quantitative [?]. Quasi [?, ?]. Quasi-perfect [?]. quasi-reversible [?]. Quaternions [?]. Queries [?, ?, ?, ?, ?, ?]. Query [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Queue [?, ?, ?, ?]. queues [?]. Quorum [?]. Quorum-Based [?].

Racing [?]. Radix [?, ?, ?]. RAID [?]. Ramacher [?]. Random [?, ?, ?, ?, ?, ?]. Randomized [?]. Range [?, ?]. Range-Join [?]. raster [?]. raster-scanned [?]. rat [?]. Ray [?]. Re [?, ?, ?, ?, ?, ?]. Re-Authoring [?]. Re-Engineering [?, ?]. Re-entrant [?]. re-evaluation [?]. re-organizing [?]. Read [?]. Read-miss [?]. Real [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Real-Time [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Real-Valued [?, ?]. reals [?]. Reasoning [?, ?, ?, ?, ?, ?, ?]. Receipt [?]. Recipes [?]. Recognition [?, ?, ?, ?]. recommendations [?]. recomposition [?]. Recompression [?]. Reconciling [?]. Reconfigurable [?, ?]. Reconstruction [?]. Reconstructions [?]. Record [?, ?]. Record-Oriented [?]. Recovery [?, ?]. Rectangle [?]. Recursion [?, ?]. recursive [?]. Red [?]. Redistribution [?]. Reducer [?]. Reducing [?, ?]. Reduction [?, ?, ?]. Redundant [?]. Reference [?, ?]. Refined [?]. Refinement [?, ?, ?]. Refinements [?]. Reflections [?]. Regular [?]. Reification [?]. Reified [?]. Rejoinder [?, ?]. Related [?]. relation [?]. Relational [?, ?, ?, ?].

Relations [?, ?, ?]. Relationships [?, ?, ?]. relative [?, ?]. relevance [?]. Relevant [?]. Reliability [?, ?]. Reliable [?, ?]. Remote [?, ?]. Removal [?]. Removal/Introduction [?]. Removing [?]. Rendering [?]. Reorganization [?]. Reorganize [?]. Repair [?]. Replicated [?, ?, ?, ?]. Replication [?]. Reply [?, ?, ?]. Report [?]. Reporting [?]. represent [?]. Representation [?, ?, ?, ?]. representations [?]. Representing [?, ?, ?]. Reproducing [AB90]. repudiation [?]. Requirements [?, ?, ?, ?]. Research [?, ?]. Residue [?, ?]. resolution [?]. Resolving [?]. Resource [?]. Resource-Directed [?]. Response [?, ?]. restricted [?, ?]. Resultant [?]. resultants [?]. Results [?]. Retrieval [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Retrieving [?, ?]. Retrospective [?]. Reuse [?, ?]. Reusing [?]. Reve [?, ?, ?, ?]. Reverse [?, ?, ?]. reversible [?]. Reversing [?]. Review [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Review [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Review [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Reviews [?, ?, ?, ?, ?, ?, ?, ?]. revised [?]. Revision [?]. Revisited [?, ?]. rewrite [?]. Rewriting [?, ?, ?, ?]. Right [?, ?]. rigorous [?]. Ring [?, ?, ?]. RISC [?, ?]. RISC-Style [?]. Rissanen [?, ?]. RMIT [?]. RNS [?]. Road [?, ?]. Roadmaps [?]. Robots [?]. Robust [?]. Role [?]. ROM [?]. ROMs [?]. rooted [?]. Rotation [?]. Rotational [?]. Rotational-Admissible [?]. Route [?]. Routines [?]. Routing [?, ?, ?, ?, ?, ?, ?]. Ruckert [?]. Rudiments [?]. Rule [?, ?]. Rules [?, ?]. Run [?]. Run-Time [?]. running [?].

S [?, ?]. Safe [?]. Safety [?, ?, ?, ?]. Safety-Critical [?, ?]. Sale [?]. Sampling [?, ?]. SAMPO [?, ?]. Samways [?]. satisfaction [?]. satisfiability [?]. Scalable [?, ?, ?, ?, ?]. Scale [?, ?, ?, ?]. scaling [?]. scanned [?]. Scanning [?]. Scenes [?]. Scheduling [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Scheduling-Oriented [?]. schema [?]. Scheme [AA99, ?, ?, ?, ?, ?, ?, ?, ?]. Schemes [?, ?, ?, ?, ?, ?, ?, ?]. Science [?]. Scientific [?]. Scientists [?]. Scoping [?, ?]. Screen [?]. Search [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. searcher [?]. Searches [?, ?]. searching [?]. Secret [?, ?]. Sector [?]. Secure [?, ?, ?, ?]. Securing [?]. Security [?, ?, ?, ?]. Selected [?]. Selection [?, ?, ?, ?, ?]. Self [?, ?, ?, ?]. Self-adjusting [?]. self-affine [?]. Self-Organising [?]. Semantic [AA99, ?, ?, ?, ?, ?, ?, ?, ?]. Semantically [?]. Semantico [?]. Semantico-pragmatic [?]. Semantics [?, ?, ?, ?, ?, ?]. Semaphore [?].

Semi [?, ?, ?]. Semi- [?]. Semi-Algebraic [?]. Semi-Custom [?]. Semiotics [?]. sentences [?]. Separating [?]. Sequence [?, ?, ?, ?, ?]. Sequences [?, ?, ?, ?]. Sequential [?]. Server [?, ?, ?, ?]. Servers [?]. Service [?, ?, ?, ?]. Services [?]. Sessions [?]. Set [?, ?, ?, ?]. Sets [?, ?]. Seven [?]. several [?]. SGML [?]. Shared [?, ?, ?, ?, ?, ?]. Shared-Memory [?]. Shares [?, ?]. Sharing [?, ?, ?, ?]. Shellsort [?]. Short [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Shortest [?]. Shortest-Path [?]. sign [?]. Signal [?]. Signature [?, ?]. significantly [?]. SIMD [?]. Similar [?]. Similarity [?, ?, ?]. simple [?, ?, ?]. Simplicity [?]. Simstrict [?]. Simulating [?, ?, ?, ?]. Simulation [?, ?, ?, ?, ?, ?, ?, ?]. Simulator [?, ?]. Simulators [?]. simultaneous [?]. Single [?, ?]. Single-Probe [?]. Site [?]. Size [?]. Sliding [?, ?]. slope [?, ?]. Slotted [?]. Slotted-FIFO [?]. Smart [?]. Smith [?]. Soft [?]. Software [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Solution [?, ?, ?, ?, ?, ?, ?]. solutions [?]. solvable [?]. Solving [?, ?, ?, ?, ?, ?]. Some [?, ?, ?, ?, ?]. Sort [?, ?, ?, ?, ?, ?]. Sort-Based [?]. Sorted [?, ?]. Sorter [?]. Sorting [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Source [?, ?, ?]. Sources [?]. Sp [?, ?]. SPA [?]. Space [?, ?, ?, ?, ?, ?]. Space-time [?]. Spanning [?, ?]. sparse [?, ?]. Spatial [?, ?, ?, ?, ?, ?, ?, ?, ?]. Spatially [?]. Special [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Specification [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Specifications [?, ?, ?, ?, ?, ?, ?, ?]. specified [?]. Speech [?]. Speed [?]. Spiral [?]. splines [?]. splitting [?]. Spotting [?]. Spurious [?]. SQL [?]. Squares [?]. SQUIRREL [?]. Stabbing [?]. stability [?, ?]. Stable [?, ?]. Stack [?]. stage [?, ?, ?]. Standard [?, ?, ?, ?]. Standards [?, ?]. Starvation [?]. Starvation-Free [?]. State [?, ?, ?, ?, ?]. Statements [?]. Static [?, ?]. Statistical [?]. still [?]. Stochastic [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Storage [?, ?, ?, ?, ?, ?]. Store [?]. Strategies [?, ?, ?, ?, ?, ?, ?]. Strategy [?, ?, ?, ?]. Stratified [?]. Streams [?]. Strict [?, ?, ?]. String [?, ?, ?]. strings [?]. structural [?]. Structure [?, ?, ?, ?, ?, ?, ?, ?, ?]. Structured [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Structures [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Structuring [?, ?]. student [?]. Studies [?]. Studio [?]. Studio-Structured [?]. Study [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Studying [?]. Style [?, ?]. Styles [?]. Sub [?]. Sub-knowledge [?]. Subject [?]. subroutine [?]. Subroutines [?, ?]. Subset [?]. Subsets [?]. Substitutions [?]. Subsystems [?]. successful [?]. successive [?]. Suite [?]. SuperPascal [?]. superscalar [?]. Supplement [?]. Support [?, ?, ?]. Supporting [?, ?, ?, ?]. supports [?]. Surface [?]. Survey [?, ?, ?, ?]. Suspected [?]. Switching [?]. SWN [?]. Symbolic [?, ?]. Symmetric [?, ?]. Symmetries [?]. Synchronisation [?]. synchronization [?]. Synthesis [?, ?, ?]. Synthesized [?]. System [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Systematic [?, ?, ?]. Systems [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Systolic [?, ?, ?, ?, ?].

T [?, ?]. Tables [?]. Tagged [?]. tailoring [?]. Talk [?]. TANE [?]. Target [?]. Target-System [?]. Task [?]. Tasks [?, ?]. Teaching [?, ?, ?]. tearing [?]. Technical [?, ?]. Technique [?, ?, ?, ?, ?, ?, ?]. Techniques [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Technology [?, ?, ?]. template [?]. Temporal [?, ?, ?, ?, ?, ?, ?, ?]. Ten [?]. Term [?, ?, ?]. Terrain [?]. Tesseral [?]. Test [?, ?, ?]. Testing [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Tests [?]. Text [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Their [?]. Themselves [?]. Theorem [?]. theoretic [?]. Theoretical [?, ?]. Theories [?]. Theory [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Thesis [?]. thinning [?]. Third [?, ?]. Thread [?]. Three [?, ?]. Three-Dimensional [?]. three-track [?]. Threshold [?, ?, ?]. Threshold-based [?]. Tight [?]. Time [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Time-constrained [?]. time-oriented [?]. timestamps [?]. Timetabling [?, ?]. Timing [?]. Timothy [?]. Token [?, ?]. Tolerant [?, ?, ?, ?, ?, ?]. Tompkins [?]. Tompkins-Paige [?]. tool [?, ?]. Toolbox [?]. toolkit [?]. Tools [?, ?, ?]. Toolset [?, ?]. Top [?]. Top-down [?]. topographic [?]. Topological [?]. Topology [?, ?, ?]. Total [?]. Towers [?]. Tracing [?]. track [?]. Tracking [?, ?]. tracts [?]. trade [?]. trade-offs [?]. traffic [?]. Train [?]. Transaction [?, ?, ?, ?, ?, ?, ?, ?, ?]. Transactions [?, ?, ?, ?]. Transfer [?, ?]. Transfers [?]. Transform [?, ?, ?]. Transformation [?, ?, ?, ?, ?, ?, ?]. Transformational [?]. Transformations [?, ?]. Transforming [?]. translation [?]. Transmission [?]. Transposition [?]. Transpositions [?]. transputer [?, ?, ?, ?, ?, ?]. transputer-based [?]. travel [?]. Traversal [?]. Traversals [?, ?]. Treating [?]. Tree [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Tree-Based [?]. Trees [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Triangulated [?, ?]. Tricky [?]. Triple [?]. Trojans [?, ?, ?, ?]. truncation [?]. tuples [?]. Turing [?, ?]. Tutorial [?]. Two [?, ?, ?, ?, ?, ?, ?, ?]. Two-Dimensional [?]. Two-Disc [?]. two-headed [?]. two-stage [?]. Type [?, ?, ?, ?]. Typed [?]. Types [?, ?, ?].

U.K. [?]. UET [?]. UK [?]. Ulrich [?]. UML [?]. Unbounded [?]. Undefined [?]. undergraduates [?]. underlying [?]. Understanding [?]. unfoldable [?]. Unification [?]. Unified [?]. Uniform [?, ?]. Unifying [?]. Unique [?]. Uniqueness [?]. Uniquest [?]. Unity [?]. universal [?]. Universities [?, ?, ?, ?]. University [?]. Unix [?]. unpredictability [?]. unranking [?]. unsuccessful [?]. Up/Down [?]. Update [?, ?]. updating [?]. URPC [?]. Usability [?, ?]. Usable [?]. Use [?, ?, ?, ?, ?, ?]. User [?, ?, ?, ?, ?, ?, ?]. User-interface [?]. Users [?]. Using [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Utilisation [?].

Valerie [?]. Validation [?, ?, ?]. value [?]. Valued [?, ?, ?, ?]. Values [?]. Variable [?, ?]. Variables [?]. variant [?]. Variants [?]. varying [?]. VDM [?, ?]. Verification [?, ?, ?, ?, ?, ?, ?, ?, ?]. verified [?]. versus [?, ?]. vertical [?]. Via [?, ?, ?, ?, ?]. Video [?, ?]. View [?, ?]. Viewpoint [?]. views [?]. Virtual [?, ?, ?, ?]. Virus [?, ?, ?, ?]. Vision [?, ?, ?]. Visputer [?]. Visual [?, ?]. visualization [?]. Visualizations [?, ?]. Vlsi [?, ?, ACDP91, ?, ?, ?, ?, ?]. Volume [?, ?]. Voting [?].

Waiting [?]. walk [?]. Wallace [?, ?, ?]. Warp [?]. was [?]. Way [?, ?]. Ways [?]. Weakest [?]. Web [?]. Wegner [?, ?]. Weighted [?, ?]. Weighting [?]. Welcome [?]. Wheeler [?]. Which [?]. Wide [?, ?]. wide-area [?]. Wilkes [?, ?]. Window [?, ?, ?]. Wireless [?, ?]. Without [?]. Wordperfect [?]. Words [?]. Work [?, ?, ?]. work-optimal [?]. Working [?]. Workloads [?]. Workshop [?]. Workstations [?, ?, ?]. World [?]. Writer [?]. Writing [?]. WWW [?].

X [?, ?]. X.400 [?]. X.400-1988 [?]. Xunrang [?].

Year [?]. Yearbook [?]. Years [?]. Yuzhang [?].

Z [?]. Zero [?]. Zero-Knowledge [?].

References

Agius:1999:CCO

[AA99] Harry W. Agius and Marios C. Angelides. COSMOS-content oriented semantic modelling overlay scheme. The Computer Journal, 42(3):153–176, ???? 1999. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

            /comjnl.oxfordjournals.org/content/42/3/153.full.pdf+
            html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
            42/Issue_03/420153.sgm.abs.html;   http://www3.oup.co.
            uk/computer_journal/hdb/Volume_42/Issue_03/pdf/420153.
            pdf.

Alcock:1990:PPR

[AB90] D. Alcock and D. Brown. PRAXIS: a program for reproducing proforma design calculations. The Computer Journal, 33(4):337– 343, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/33/4/337.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/337.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/338.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/339.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/340.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/341.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/342.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/343.tif.

Ancona:1991:HLH

[ACDP91] M. Ancona, A. Clematis, L. De Floriani, and E. Puppo. HIDEL: a language for hierarchical VLSI design. The Computer Journal, 34(3):195–206, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

\bibitem[AD95]{Ahmad:1995:WGP} Imtiaz Ahmad and Muhammad~
K. Dhodhi. \newblock On the $m$-way graph partitioning
problem. \newblock {\em The Computer Journal}, 38\penalty0
(3):\penalty0 237--244, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/237.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_03/Vol38_03.body.html#AbstractAhmad|}
\fi.
\bibitem[ADD90]{Ancona:1990:SST} M.~Ancona, L.~{De Floriani}
, and J.~S. Deogun. \newblock Structured spanning trees.
\newblock {\em The Computer Journal}, 33\penalty0 (4)
:\penalty0 344--355,   August 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/344.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/344.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/345.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/346.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/347.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/348.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_33/Issue_04/tiff/349.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/350.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/351.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/353.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/354.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/355.
t.
\bibitem[ADK{\etalchar{+}}93]{Abolhassan:1993:PDP} F.
~Abolhassan, R.~Drefenstedt, J.~Keller, W.~J. Paul, and
D.~Scheerer. \newblock On the physical design of {PRAMs}
. \newblock {\em The Computer Journal}, 36\penalty0 (
8):\penalty0 756--762, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/8/756.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractAbolhassan|}
\fi.
\bibitem[AE90a]{Abramson:1990:RDF} D.~Abramson and G.
~Egan. \newblock The {RMIT} data flow computer: a hybrid
architecture. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 230--240, June   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/3/
230.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/230.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/231.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/232.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/234.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/235.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/236.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/237.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/239.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_33/Issue_03/tiff/240.tif|}\fi.
\bibitem[AE90b]{Agrawal:1990:ISF} D.~Agrawal and A.~{
El Abbadi}. \newblock Integrating security with fault-
tolerant distributed databases. \newblock {\em The Computer
Journal}, 33\penalty0 (1):\penalty0 71--78,   February
1990. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/33/1/71.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/71.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/73.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/76.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/78.tif|}\fi.
\bibitem[AG98]{Abdel-Ghaffar:1998:DST} Khaled A.~A. Abdel-
Ghaffar. \newblock Detecting substitutions and transpositions
of characters. \newblock {\em The Computer Journal}, 41\
penalty0 (4):\penalty0 270--277, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/4/
270.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_04/Vol41_04.body.html#AbstractAbdel-Ghaffar|}
\fi.
\bibitem[Aga96]{Agar:1996:HCF} Jon Agar. \newblock History
of computing feature: The provision of digital computers
to   {British} universities up to the {Flowers Report}
(1966). \newblock {\em The Computer Journal}, 39\penalty0
(7):\penalty0 630--642, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/7/630.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_07/390630.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_07/Vol39_
07.body.html#AbstractAgar|}\fi.

\bibitem[Age95]{Agerholm:1995:LEH} Sten Agerholm. \newblock
{LCF} examples in {HOL}. \newblock {\em The Computer Journal}
, 38\penalty0 (2):\penalty0 121--130, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/2/
121.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_02/Vol38_02.body.html#AbstractAgerholm|}
\fi.
\bibitem[AGNS91]{Alagar:1991:IMS} V.~S. Alagar, P.~Goyal,
P.~S. Nair, and F.~Sadri. \newblock Integrated magic set
method: a rule rewrite scheme for optimising   linear
datalog programs. \newblock {\em The Computer Journal}
, 34\penalty0 (6):\penalty0 A121--A129,   December 1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}.
\bibitem[AHL92a]{Auramaki:1992:MOTb} E.~Auram{\"a}ki,
R.~Hirschheim, and K.~Lyytinen. \newblock Modelling offices
through discourse analysis: a comparison and   evaluation
of {SAMPO} with {OSSAD} and {ICN}. \newblock {\em The
Computer Journal}, 35\penalty0 (5):\penalty0 492--500,
 October 1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/5/492.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractAuramaki|}\fi.
\bibitem[AHL92b]{Auramaki:1992:MOTa} E.~Auram{\"a}ki,
R.~Hirschheim, and K.~Lyytinen. \newblock Modelling offices
through discourse analysis: the {SAMPO} approach. \newblock
{\em The Computer Journal}, 35\penalty0 (4):\penalty0
342--352,   August 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/4/342.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractAuramakil|}\fi.
\bibitem[Ait96]{Aitken:1996:BR} S.~Aitken. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(9):\penalty0 816, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.

oxfordjournals.org/content/39/9/816.1.full.pdf+html|}
\fi.
\bibitem[AK94a]{Arun-Kumar:1994:BRa} S.~Arun-Kumar. \
newblock Book review. \newblock {\em The Computer Journal}
, 37\penalty0 (1):\penalty0 77, ???? 1994. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/1/
77.1.full.pdf+html|}\fi.
\bibitem[AK94b]{Arun-Kumar:1994:BRb} S.~Arun-Kumar. \
newblock Book review. \newblock {\em The Computer Journal}
, 37\penalty0 (7):\penalty0 647, ???? 1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
647.1.full.pdf+html|}\fi.
\bibitem[AK95]{Arun-Kumar:1995:BR} S.~Arun-Kumar. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (9):\penalty0 749--750, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/9/
749.full.pdf+html|}\fi.
\bibitem[Akm91]{Akman:1991:SNI} V.~Akman. \newblock Short
notes: Implementation of {Karp--Luby Monte Carlo} method:
an   exercise in approximate counting. \newblock {\em
The Computer Journal}, 34\penalty0 (3):\penalty0 279-
-282, June   1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/3/279.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/279.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/280.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/281.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/282.tif|}\fi.
\bibitem[Aky92]{Akyildiz:1992:PAM} I.~F. Akyildiz. \newblock
Performance analysis of a multiprocessor system model
with process   communication. \newblock {\em The Computer
Journal}, 35\penalty0 (1):\penalty0 52--61,   February
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{

0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/1/52.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_01/Vol35_01.body.
html#AbstractAkyildiz|}\fi.
\bibitem[All90]{Allison:1990:CIG} L.~Allison. \newblock
Continuations implement generators and streams. \newblock
{\em The Computer Journal}, 33\penalty0 (5):\penalty0
460--465,   October 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/5/460.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/460.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/461.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/462.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/463.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/464.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/465.tif|}\fi.
\bibitem[Alm93]{Alm:1993:BR} N.~Alm. \newblock Book review.
\newblock {\em The Computer Journal}, 36\penalty0 (3)
:\penalty0 295--296, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/3/295.
full.pdf+html|}\fi.
\bibitem[AM90]{Alia:1990:SNV} G.~Alia and E.~Martinelli.
\newblock Short note: {VLSI} binary-residue converters
for pipelined   processing. \newblock {\em The Computer
Journal}, 33\penalty0 (5):\penalty0 473--475,   October
1990. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/33/5/473.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/473.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/474.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/475.tif|}\fi.
\bibitem[AM91]{Agrawal:1991:EDI} D.~Agrawal and A.~Malpani.
\newblock Efficient dissemination of information in computer
networks. \newblock {\em The Computer Journal}, 34\penalty0

(6):\penalty0 534--541,   December 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/6/534.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/534.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/535.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/536.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/537.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/538.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/539.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/540.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/541.tif|}\fi.
\bibitem[AM92]{Agosti:1992:UNI} Maristella Agosti and
Pier~Giorgio Marchetti. \newblock User navigation in the
{IRS} conceptual structure through a semantic   association
function. \newblock {\em The Computer Journal}, 35\penalty0
(3):\penalty0 194--199, June   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/194.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_03/Vol35_03.body.html#AbstractAgosti|}
\fi.
\bibitem[AM97]{Arnavut:1997:LPS} Z.~Arnavut and S.~S.
Magliveras. \newblock Lexical permutation sorting algorithm.
\newblock {\em The Computer Journal}, 40\penalty0 (5)
:\penalty0 292--295, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/5/292.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_05/Vol40_05.body.html#AbstractArnavut;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
05/Vol40_05.body.html#AbstractArnavut|}\fi.
\bibitem[AM98]{Alia:1998:LDF} Giuseppe Alia and Enrico
Martinelli. \newblock Logic design of a fast circuit for
iterative additions in redundant   hybrid number systems.
\newblock {\em The Computer Journal}, 41\penalty0 (1)
:\penalty0 45--51, ????   1998. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/1/45.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_01/Vol41_01.body.html#AbstractAlia|}\
fi.
\bibitem[AM99]{Al-Mouhamed:1999:EBS} Mayez Al-Mouhamed.
\newblock Evolution-based scheduling of computations and
communications on   distributed-memory multicomputers.
\newblock {\em The Computer Journal}, 42\penalty0 (5)
:\penalty0 373--390, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/5/373.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_05/420373.sgm.abs.html|}\fi.
\bibitem[And90]{Andersson:1990:SNN} A.~Andersson. \newblock
Short note: a note on the expected behaviour of binary
tree   traversals. \newblock {\em The Computer Journal}
, 33\penalty0 (5):\penalty0 471--472,   October 1990.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/5/471.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/471.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/472.tif|}\fi.
\bibitem[And96]{Anderson:1996:BR} R.~Anderson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (5):\penalty0 452--453, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/5/
452.full.pdf+html|}\fi.
\bibitem[Ano90a]{Anonymous:1990:Aa} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (1):\penalty0 10, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/10.
full.pdf+html|}\fi.

\bibitem[Ano90b]{Anonymous:1990:Ab} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (1):\penalty0 30, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/30.
full.pdf+html|}\fi.
\bibitem[Ano90c]{Anonymous:1990:Ad} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (1):\penalty0 70, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/70.
full.pdf+html|}\fi.
\bibitem[Ano90d]{Anonymous:1990:Af} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (1):\penalty0 93, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/93.
2.full.pdf+html|}\fi.
\bibitem[Ano90e]{Anonymous:1990:Ak} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 163, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/163.
2.full.pdf+html|}\fi.
\bibitem[Ano90f]{Anonymous:1990:Aw} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (4):\penalty0 329, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/329.
full.pdf+html|}\fi.
\bibitem[Ano90g]{Anonymous:1990:Ax} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (4):\penalty0 355, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/355.
full.pdf+html|}\fi.

\bibitem[Ano90h]{Anonymous:1990:Ay} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (4):\penalty0 364, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/364.
full.pdf+html|}\fi.
\bibitem[Ano90i]{Anonymous:1990:Az} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (4):\penalty0 374, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/374.
full.pdf+html|}\fi.
\bibitem[Ano90j]{Anonymous:1990:Aab} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 33\
penalty0 (5):\penalty0 465, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/465.
full.pdf+html|}\fi.
\bibitem[Ano90k]{Anonymous:1990:Ac} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (1):\penalty0 48, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/48.
full.pdf+html|}\fi.
\bibitem[Ano90l]{Anonymous:1990:Ae} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (1):\penalty0 78, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/78.
full.pdf+html|}\fi.
\bibitem[Ano90m]{Anonymous:1990:Ag} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 125, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/125.
full.pdf+html|}\fi.

\bibitem[Ano90n]{Anonymous:1990:Ah} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 139, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/139.
full.pdf+html|}\fi.
\bibitem[Ano90o]{Anonymous:1990:Ai} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 146, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/146.
full.pdf+html|}\fi.
\bibitem[Ano90p]{Anonymous:1990:Aj} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 154, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/154.
full.pdf+html|}\fi.
\bibitem[Ano90q]{Anonymous:1990:Al} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 184, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/184.
full.pdf+html|}\fi.
\bibitem[Ano90r]{Anonymous:1990:Am} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 187, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/187.
full.pdf+html|}\fi.
\bibitem[Ano90s]{Anonymous:1990:An} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 214, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/214.
full.pdf+html|}\fi.

\bibitem[Ano90t]{Anonymous:1990:Ao} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 229, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/229.
full.pdf+html|}\fi.
\bibitem[Ano90u]{Anonymous:1990:Ap} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 246, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/246.
full.pdf+html|}\fi.
\bibitem[Ano90v]{Anonymous:1990:Aq} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 265, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/265.
full.pdf+html|}\fi.
\bibitem[Ano90w]{Anonymous:1990:Ar} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 273, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/273.
full.pdf+html|}\fi.
\bibitem[Ano90x]{Anonymous:1990:As} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 278, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/278.
full.pdf+html|}\fi.
\bibitem[Ano90y]{Anonymous:1990:At} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 280, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/280.
full.pdf+html|}\fi.

\bibitem[Ano90z]{Anonymous:1990:Au} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 282, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/282.
full.pdf+html|}\fi.
\bibitem[Ano90-27]{Anonymous:1990:Av} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (4):\penalty0 313, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/313.
full.pdf+html|}\fi.
\bibitem[Ano90-28]{Anonymous:1990:Aaa} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (5):\penalty0 416, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/416.
full.pdf+html|}\fi.
\bibitem[Ano90-29]{Anonymous:1990:Aac} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (5):\penalty0 476--477, ????   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/5/
476.full.pdf+html|}\fi.
\bibitem[Ano90-30]{Anonymous:1990:Aad} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (6):\penalty0 500, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/500.
full.pdf+html|}\fi.
\bibitem[Ano90-31]{Anonymous:1990:Aae} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (6):\penalty0 507, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/507.
full.pdf+html|}\fi.

\bibitem[Ano90-32]{Anonymous:1990:Aaf} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (6):\penalty0 517, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/517.
full.pdf+html|}\fi.
\bibitem[Ano90-33]{Anonymous:1990:Aag} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (6):\penalty0 546, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/546.
full.pdf+html|}\fi.
\bibitem[Ano90-34]{Anonymous:1990:Aah} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (6):\penalty0 561, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/561.
full.pdf+html|}\fi.
\bibitem[Ano90-35]{Anonymous:1990:Aai} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 33\
penalty0 (6):\penalty0 572, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/572.
full.pdf+html|}\fi.
\bibitem[Ano90-36]{Anonymous:1990:ED} Anonymous. \newblock
Editorial --- databases. \newblock {\em The Computer Journal}
, 33\penalty0 (1):\penalty0 1, ???? 1990. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/1/
1.full.pdf+html|}\fi.
\bibitem[Ano90-37]{Anonymous:1990:EPP} Anonymous. \newblock
Editorial --- procedural programming. \newblock {\em The
Computer Journal}, 33\penalty0 (2):\penalty0 97, ????
1990. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/33/2/97.full.pdf+html|}\fi.

\bibitem[Ano90-38]{Anonymous:1990:WA} Anonymous. \newblock
Wilkes award. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 163, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/163.
1.full.pdf+html|}\fi.
\bibitem[Ano91a]{Anonymous:1991:APa} Anonymous. \newblock
Additional papers. \newblock {\em The Computer Journal}
, 34\penalty0 (6):\penalty0 571--572,   December 1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/6/571.full.pdf+html|}\fi.
\bibitem[Ano91b]{Anonymous:1991:AS} Anonymous. \newblock
Algorithms supplement. \newblock {\em The Computer Journal}
, 34\penalty0 (4):\penalty0 358--379, ????   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/4/
358.full.pdf+html|}\fi.
\bibitem[Ano91c]{Anonymous:1991:Ac} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 121, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/121.
full.pdf+html|}\fi.
\bibitem[Ano91d]{Anonymous:1991:Ad} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 142, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/142.
full.pdf+html|}\fi.
\bibitem[Ano91e]{Anonymous:1991:Ae} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 152, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/152.
full.pdf+html|}\fi.

\bibitem[Ano91f]{Anonymous:1991:Af} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 163, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/163.
full.pdf+html|}\fi.
\bibitem[Ano91g]{Anonymous:1991:Ag} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 181, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/181.
full.pdf+html|}\fi.
\bibitem[Ano91h]{Anonymous:1991:Ah} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 185, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/185.
full.pdf+html|}\fi.
\bibitem[Ano91i]{Anonymous:1991:Al} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (4):\penalty0 344, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/344.
full.pdf+html|}\fi.
\bibitem[Ano91j]{Anonymous:1991:Am} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (4):\penalty0 381, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/381.
full.pdf+html|}\fi.
\bibitem[Ano91k]{Anonymous:1991:Ao} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 437, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/437.
full.pdf+html|}\fi.

\bibitem[Ano91l]{Anonymous:1991:Ap} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 454, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/454.
full.pdf+html|}\fi.
\bibitem[Ano91m]{Anonymous:1991:Aq} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 468, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/468.
2.full.pdf+html|}\fi.
\bibitem[Ano91n]{Anonymous:1991:Ar} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 476, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/476.
full.pdf+html|}\fi.
\bibitem[Ano91o]{Anonymous:1991:Ax} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 34\
penalty0 (6):\penalty0 565, December   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/6/
565.full.pdf+html|}\fi.
\bibitem[Ano91p]{Anonymous:1991:Aa} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (1):\penalty0 87, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/87.
full.pdf+html|}\fi.
\bibitem[Ano91q]{Anonymous:1991:Ab} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (1):\penalty0 91, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/91.
full.pdf+html|}\fi.

\bibitem[Ano91r]{Anonymous:1991:Ai} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (3):\penalty0 238, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/238.
full.pdf+html|}\fi.
\bibitem[Ano91s]{Anonymous:1991:Aj} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (3):\penalty0 244, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/244.
full.pdf+html|}\fi.
\bibitem[Ano91t]{Anonymous:1991:Ak} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (3):\penalty0 271, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/271.
full.pdf+html|}\fi.
\bibitem[Ano91u]{Anonymous:1991:An} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 414, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/414.
full.pdf+html|}\fi.
\bibitem[Ano91v]{Anonymous:1991:At} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (6):\penalty0 492, December   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/6/
492.full.pdf+html|}\fi.
\bibitem[Ano91w]{Anonymous:1991:Au} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (6):\penalty0 541, December   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/6/
541.full.pdf+html|}\fi.

\bibitem[Ano91x]{Anonymous:1991:Av} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (6):\penalty0 550, December   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/6/
550.full.pdf+html|}\fi.
\bibitem[Ano91y]{Anonymous:1991:Aw} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 34\
penalty0 (6):\penalty0 558, December   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/6/
558.full.pdf+html|}\fi.
\bibitem[Ano91z]{Anonymous:1991:WA} Anonymous. \newblock
Wilkes award. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 112, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/112.
full.pdf+html|}\fi.
\bibitem[Ano92a]{Anonymous:1992:APa} Anonymous. \newblock
Additional papers. \newblock {\em The Computer Journal}
, 35\penalty0 (1):\penalty0 93, February   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/1/
93.full.pdf+html|}\fi.
\bibitem[Ano92b]{Anonymous:1992:APb} Anonymous. \newblock
Additional papers. \newblock {\em The Computer Journal}
, 35\penalty0 (2):\penalty0 188, April   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/2/
188.full.pdf+html|}\fi.
\bibitem[Ano92c]{Anonymous:1992:APc} Anonymous. \newblock
Additional papers. \newblock {\em The Computer Journal}
, 35\penalty0 (3):\penalty0 311--312, June   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/3/
311.full.pdf+html|}\fi.

\bibitem[Ano92d]{Anonymous:1992:APd} Anonymous. \newblock
Additional papers. \newblock {\em The Computer Journal}
, 35\penalty0 (4):\penalty0 412--413,   August 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/4/412.full.pdf+html|}\fi.
\bibitem[Ano92e]{Anonymous:1992:APe} Anonymous. \newblock
Additional papers. \newblock {\em The Computer Journal}
, 35\penalty0 (5):\penalty0 538, October   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
538.full.pdf+html|}\fi.
\bibitem[Ano92f]{Anonymous:1992:APf} Anonymous. \newblock
Additional papers. \newblock {\em The Computer Journal}
, 35\penalty0 (6):\penalty0 669--670,   December 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/669.full.pdf+html|}\fi.
\bibitem[Ano92g]{Anonymous:1992:Aa} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (1):\penalty0 51, February   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/1/
51.full.pdf+html|}\fi.
\bibitem[Ano92h]{Anonymous:1992:Ab} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (1):\penalty0 61, February   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/1/
61.full.pdf+html|}\fi.
\bibitem[Ano92i]{Anonymous:1992:Ae} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (3):\penalty0 220, June 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/220.
full.pdf+html|}\fi.

\bibitem[Ano92j]{Anonymous:1992:Af} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (3):\penalty0 242, June 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/242.
full.pdf+html|}\fi.
\bibitem[Ano92k]{Anonymous:1992:Ag} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (3):\penalty0 278a--278, June   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/3/
278-a.full.pdf+html|}\fi.
\bibitem[Ano92l]{Anonymous:1992:Ah} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (3):\penalty0 305, June 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/305.
full.pdf+html|}\fi.
\bibitem[Ano92m]{Anonymous:1992:Aj} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (4):\penalty0 352, August   1992. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/4/
352.full.pdf+html|}\fi.
\bibitem[Ano92n]{Anonymous:1992:Ak} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (4):\penalty0 394, August   1992. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/4/
394.full.pdf+html|}\fi.
\bibitem[Ano92o]{Anonymous:1992:Ao} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 477a--477,   October 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
477-a.full.pdf+html|}\fi.

\bibitem[Ano92p]{Anonymous:1992:Ap} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 500, October   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
500.full.pdf+html|}\fi.
\bibitem[Ano92q]{Anonymous:1992:Aq} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 537, October   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
537.full.pdf+html|}\fi.
\bibitem[Ano92r]{Anonymous:1992:Ar} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 539a--539,   October 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
539-a.full.pdf+html|}\fi.
\bibitem[Ano92s]{Anonymous:1992:As} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 599, December   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
599.full.pdf+html|}\fi.
\bibitem[Ano92t]{Anonymous:1992:At} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 610a--610,   December 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
610-a.full.pdf+html|}\fi.
\bibitem[Ano92u]{Anonymous:1992:Au} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 635, December   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
635.full.pdf+html|}\fi.

\bibitem[Ano92v]{Anonymous:1992:Av} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 642, December   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
642.full.pdf+html|}\fi.
\bibitem[Ano92w]{Anonymous:1992:Aw} Anonymous. \newblock
Announcement. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 668a--668,   December 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
668-a.full.pdf+html|}\fi.
\bibitem[Ano92x]{Anonymous:1992:Ac} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 35\
penalty0 (2):\penalty0 107, April   1992. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/2/
107.full.pdf+html|}\fi.
\bibitem[Ano92y]{Anonymous:1992:Ad} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 35\
penalty0 (2):\penalty0 137, April   1992. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/2/
137.full.pdf+html|}\fi.
\bibitem[Ano92z]{Anonymous:1992:Ai} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 35\
penalty0 (3):\penalty0 310, June 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/310.
full.pdf+html|}\fi.
\bibitem[Ano92-27]{Anonymous:1992:Al} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 35\
penalty0 (4):\penalty0 411, August   1992. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/4/
411.full.pdf+html|}\fi.

\bibitem[Ano92-28]{Anonymous:1992:Am} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 423, October   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
423.full.pdf+html|}\fi.
\bibitem[Ano92-29]{Anonymous:1992:An} Anonymous. \newblock
Announcements. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 513, October   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
513.full.pdf+html|}\fi.
\bibitem[Ano92-30]{Anonymous:1992:C} Anonymous. \newblock
Correspondence. \newblock {\em The Computer Journal},
35\penalty0 (3):\penalty0 298, June 1992. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/3/
298.full.pdf+html|}\fi.
\bibitem[Ano92-31]{Anonymous:1992:Ea} Anonymous. \newblock
Editorial. \newblock {\em The Computer Journal}, 35\penalty0
(2):\penalty0 97, April 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/2/97.full.pdf+html|}\fi.
\bibitem[Ano92-32]{Anonymous:1992:EMA} Anonymous. \newblock
Editorial --- models and architectures. \newblock {\em
The Computer Journal}, 35\penalty0 (4):\penalty0 321,
August   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/4/321.full.pdf+html|}\fi.
\bibitem[Ano92-33]{Anonymous:1992:Ed} Anonymous. \newblock
Errata. \newblock {\em The Computer Journal}, 35\penalty0
(6):\penalty0 670, December   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/35/6/670.
full.pdf+html|}\fi.
\bibitem[Ano92-34]{Anonymous:1992:Eb} Anonymous. \newblock
Erratum. \newblock {\em The Computer Journal}, 35\penalty0
(2):\penalty0 107a--107,   April 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/2/107-
a.full.pdf+html|}\fi.
\bibitem[Ano92-35]{Anonymous:1992:Ec} Anonymous. \newblock
Erratum. \newblock {\em The Computer Journal}, 35\penalty0
(3):\penalty0 290a--290, June   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/290-
a.full.pdf+html|}\fi.
\bibitem[Ano92-36]{Anonymous:1992:N} Anonymous. \newblock
Notice. \newblock {\em The Computer Journal}, 35\penalty0
(6):\penalty0 578, December   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/578.
full.pdf+html|}\fi.
\bibitem[Ano92-37]{Anonymous:1992:SISa} Anonymous. \newblock
Special issue on spatial data --- call for papers. \newblock
{\em The Computer Journal}, 35\penalty0 (5):\penalty0
467, October   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/5/467.full.pdf+html|}\fi.
\bibitem[Ano92-38]{Anonymous:1992:SISb} Anonymous. \newblock
Special issue on spatial data --- call for papers. \newblock
{\em The Computer Journal}, 35\penalty0 (6):\penalty0
554a--554,   December 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/6/554-a.full.pdf+html|}
\fi.
\bibitem[Ano93]{Anonymous:1993:BR} Anonymous. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 292--293, ????   1993. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/3/
292.2.full.pdf+html|}\fi.
\bibitem[Ano94]{Anonymous:1994:EN} Anonymous. \newblock
Editorial note. \newblock {\em The Computer Journal},
37\penalty0 (9):\penalty0 744, ???? 1994. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/9/
744.full.pdf+html|}\fi.
\bibitem[Ano95]{Anonymous:1995:L} Anonymous. \newblock
Letters. \newblock {\em The Computer Journal}, 38\penalty0
(3):\penalty0 257--258, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/257.
full.pdf+html|}\fi.
\bibitem[Ano96a]{Anonymous:1996:BRa} Anonymous. \newblock
Book reviews. \newblock {\em The Computer Journal}, 39\
penalty0 (7):\penalty0 ??, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://www3.oup.co.uk/computer_journal/Volume_39/
Issue_07/Vol39_07.body.html#AbstractReviews397|}\fi.
\bibitem[Ano96b]{Anonymous:1996:BRb} Anonymous. \newblock
Book reviews. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 723--730, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www.oup.co.uk/jnls/list/comjnl/hdb/Volume_
39/Issue_08/390723.sgm.abs.html|}\fi.
\bibitem[Ano97a]{Anonymous:1997:AI} Anonymous. \newblock
Author index for 1997. \newblock {\em The Computer Journal}
, 40\penalty0 (10):\penalty0 ??, ???? 1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www3.oup.co.uk/computer_journal/Volume_
40/Issue_10/Vol40_10.body.html#AbstractAuthors|}\fi.
\bibitem[Ano97b]{Anonymous:1997:SI} Anonymous. \newblock
Subject index for 1997. \newblock {\em The Computer Journal}
, 40\penalty0 (10):\penalty0 ??, ???? 1997. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www3.oup.co.uk/computer_journal/Volume_
40/Issue_10/Vol40_10.body.html#AbstractSubjects|}\fi.
\bibitem[Ant94a]{Anthony:1994:BRa} D.~Anthony. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 229, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/229.
1.full.pdf+html|}\fi.
\bibitem[Ant94b]{Anthony:1994:BRb} D.~Anthony. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 352, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/352.
1.full.pdf+html|}\fi.
\bibitem[APD99]{Allison:1999:CAM} L.~Allison, D.~Powell,
and T.~I. Dix. \newblock Compression and approximate matching.
\newblock {\em The Computer Journal}, 42\penalty0 (1)
:\penalty0 1--10, ????   1999. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/1/1.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
01/420001.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_01/pdf/420001.pdf|}\fi.
\bibitem[Ara94]{Arazi:1994:PTU} Benjamin Arazi. \newblock
On primality testing using purely divisionless operations.
\newblock {\em The Computer Journal}, 37\penalty0 (3)
:\penalty0 219--222, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/219.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_03/Vol37_03.body.html#AbstractArazi|}
\fi.
\bibitem[Arg91]{Argo:1991:WWW} G.~Argo. \newblock Weighting
without waiting: the weighted path length tree. \newblock
{\em The Computer Journal}, 34\penalty0 (5):\penalty0
444--449,   October 1991. \newblock \showCODEN{CMPJA6}

. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/5/444.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/444.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/445.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/446.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/447.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/448.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/449.tif|}\fi.
\bibitem[Ash95a]{Ashton:1995:AS} Michael Ashton. \newblock
Analysing systems. \newblock {\em The Computer Journal}
, 38\penalty0 (1):\penalty0 82, ???? 1995. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
82.full.pdf+html|}\fi.
\bibitem[Ash95b]{Ashton:1995:SQA} Michael Ashton. \newblock
Software quality assurance --- a practical approach. \
newblock {\em The Computer Journal}, 38\penalty0 (1):
\penalty0 82a--83, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/1/82-
a.full.pdf+html|}\fi.
\bibitem[Ash96]{Ashworth:1996:BR} T.~Ashworth. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (9):\penalty0 816--817, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
816.2.full.pdf+html|}\fi.
\bibitem[Atk93]{Atkinson:1993:UGR} M.~D. Atkinson. \newblock
Uniform generation of rooted ordered trees with prescribed
degrees. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 593--594,   December 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/593.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

Volume_36/Issue_06/Vol36_06.body.html#AbstractAtkinson|}
\fi.
\bibitem[ATL93]{Appelt:1993:FSI} W.~Appelt and N.~Tetteh-
Lartey. \newblock The formal specification of the {ISO}
open document architecture   ({ODA}) standard. \newblock
{\em The Computer Journal}, 36\penalty0 (3):\penalty0
269--279, June   1993. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/3/269.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractAppelt|}\fi.
\bibitem[AWH91]{Avison:1991:ISD} D.~E. Avison and A.~
T. Wood-Harper. \newblock Information systems development
research: an exploration of ideas in   practice. \newblock
{\em The Computer Journal}, 34\penalty0 (2):\penalty0
98--112, April   1991. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/2/98.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/100.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/101.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/102.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/103.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/105.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/106.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/107.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/108.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/109.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/110.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/111.ti.
\bibitem[Axf90]{Axford:1990:RCC} T.~H. Axford. \newblock
Reference counting of cyclic graphs for functional programs.
\newblock {\em The Computer Journal}, 33\penalty0 (5)
:\penalty0 466--470,   October 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/33/5/466.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/466.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/467.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/468.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/469.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/470.tif|}\fi.
\bibitem[Azm90]{Azmoodeh:1990:BDI} M.~Azmoodeh. \newblock
{BRMQ}: a database interface facility based on graph traversals
and   extended relationships on groups of entities. \
newblock {\em The Computer Journal}, 33\penalty0 (1):
\penalty0 31--39,   February 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/31.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/31.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/32.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/34.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/35.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/36.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/37.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/38.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/39.tif|}\fi.
\bibitem[B{\Dbar}96]{Borger:1996:CCO} Egon. B{\"o}rger
and Igor {\Dbar}ur{\dbar}anovic. \newblock Correctness
of compiling {Occam} to {transputer} code. \newblock {
\em The Computer Journal}, 39\penalty0 (1):\penalty0 52-
-92, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/1/52.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
01/390052.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_01/Vol39_01.body.html#AbstractEgon|}
\fi.
\bibitem[Bae94]{Baeten:1994:SIE} J.~C.~M. Baeten. \newblock

Special issue editorial process algebra. \newblock {\
em The Computer Journal}, 37\penalty0 (5):\penalty0 474-
-475, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/5/474.full.pdf+html|}\fi.
\bibitem[Bak94]{Bakker:1994:SAE} J.~A. Bakker. \newblock
A semantic approach to enforce correctness of data distribution
 schemes. \newblock {\em The Computer Journal}, 37\penalty0
(7):\penalty0 561--575, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/7/561.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_07/Vol37_07.body.html#AbstractBakker|}
\fi.
\bibitem[BB94]{Berrisford:1994:ROT} G.~Berrisford and
M.~Burrows. \newblock Reconciling {OO} with {Turing} machines.
\newblock {\em The Computer Journal}, 37\penalty0 (10)
:\penalty0 888--906, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/10/888.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_10/Vol37_10.body.html#AbstractBerrisford|}
\fi.
\bibitem[BBF{\etalchar{+}}94]{Benford:1994:SCW} Steve
Benford, John Bowers, Lennart~E. Fahl{\'e}n, John Mariani,
and Tom   Rodden. \newblock Supporting cooperative work
in virtual environments. \newblock {\em The Computer Journal}
, 37\penalty0 (8):\penalty0 653--668, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
653.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_08/Vol37_08.body.html#AbstractBenford|}
\fi.
\bibitem[BBG95]{Bernardo:1995:DSE} Marco Bernardo, Nadia
Busi, and Roberto Gorrieri. \newblock A distributed semantics
for {EMPA} based on stochastic contextual   nets. \newblock
{\em The Computer Journal}, 38\penalty0 (7):\penalty0

492--509, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/7/492.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractBernardo|}\fi.
\bibitem[BBP94]{Bergstra:1994:PAI} Jan~A. Bergstra, Inge
Bethke, and Alban Ponse. \newblock Process algebra with
iteration and nesting. \newblock {\em The Computer Journal}
, 37\penalty0 (4):\penalty0 243--258, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
243.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_04/Vol37_04.body.html#AbstractBergstra|}
\fi.
\bibitem[BBP98]{Baldoni:1998:SFC} Roberto Baldoni, Roberto
Beraldi, and Ravi Prakash. \newblock Slotted-{FIFO} communication
for asynchronous distributed systems. \newblock {\em The
Computer Journal}, 41\penalty0 (5):\penalty0 337--348,
????   1998. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/41/5/337.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_41/Issue_05/Vol41_05.body.
html#AbstractBaldoni|}\fi.
\bibitem[BCH91]{Bastani:1991:MSA} F.~B. Bastani, I.-R.
Chen, and W.~Hilal. \newblock A model for the stability
analysis of maintenance strategies for   linear list.
\newblock {\em The Computer Journal}, 34\penalty0 (1)
:\penalty0 80--87,   February 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/80.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/80.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/81.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/84.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/85.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/86.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/87.tif|}\fi.
\bibitem[BCV93]{Bergsten:1993:OPA} B.~Bergsten, M.~Couprie,
and P.~Valduriez. \newblock Overview of parallel architectures
for databases. \newblock {\em The Computer Journal}, 36\
penalty0 (8):\penalty0 734--740, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/8/
734.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractBergsten|}
\fi.
\bibitem[BD98]{Barendregt:1998:BR} H.~Barendregt and W.
~Dekkers. \newblock Book review. \newblock {\em The Computer
Journal}, 41\penalty0 (2):\penalty0 135, ???? 1998. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/2/135.full.pdf+html|}\fi.
\bibitem[BDB92]{Burmester:1992:EZK} M.~V.~D. Burmester,
Y.~G. Desmedt, and T.~Beth. \newblock Efficient zero-
knowledge identification schemes for smart cards. \newblock
{\em The Computer Journal}, 35\penalty0 (1):\penalty0
21--29,   February 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/1/21.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractBurmester|}\fi.
\bibitem[BDG99]{Blundo:1999:RKA} Carlo Blundo, P.~D'Arco,
and A.~G. Gaggia. \newblock A $\tau$-restricted key agreement
scheme. \newblock {\em The Computer Journal}, 42\penalty0
(1):\penalty0 51--61, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/1/51.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_01/420051.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_01/
pdf/420051.pdf|}\fi.

\bibitem[BDP93]{Berrington:1993:GUP} Neil Berrington,
David {De Roure}, and Julian Padget. \newblock Guaranteeing
unpredictability (programming model). \newblock {\em The
Computer Journal}, 36\penalty0 (8):\penalty0 723--733,
????   1993. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/8/723.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_08/Vol36_08.body.
html#AbstractBerrington|}\fi.
\bibitem[Ben95]{Bennett:1995:IG} Simon Bennett. \newblock
Innovations in {GIS 1}. \newblock {\em The Computer Journal}
, 38\penalty0 (1):\penalty0 87a--88, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
87-a.full.pdf+html|}\fi.
\bibitem[BENM98]{Bai:1998:FTB} L.~Q. Bai, H.~Ebara, H.
~Nakano, and H.~Maeda. \newblock Fault-tolerant broadcasting
on the arrangement graph. \newblock {\em The Computer
Journal}, 41\penalty0 (3):\penalty0 171--184, ????   1998.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/3/171.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_41/Issue_03/Vol41_03.body.html#AbstractBai|}
\fi.
\bibitem[Ber92]{Berry:1992:MWP} Dave Berry. \newblock
Ml for the working programme. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 450a--450,   October
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/450-a.full.pdf+html|}\fi.
\bibitem[BG94a]{Bezem:1994:CCP} M.~Bezem and J.~F. Groote.
\newblock Corrigendum to a correctness proof of a one-
bit sliding window   protocol in {$\mu$CRL}. \newblock
{\em The Computer Journal}, 37\penalty0 (7):\penalty0
651, ???? 1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/7/651.full.pdf+html|}\fi.
\newblock See \cite{Bezem:1994:CPO}.

\bibitem[BG94b]{Bezem:1994:CPO} M.~A. Bezem and J.~F.
Groote. \newblock A correctness proof of a one-bit sliding
window protocol in   {$\mu$CRL}. \newblock {\em The Computer
Journal}, 37\penalty0 (4):\penalty0 289--307, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/4/289.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_04/Vol37_04.body.html#AbstractBezem|}
\fi. \newblock See corrigendum \cite{Bezem:1994:CCP}.
\bibitem[BGMW96]{Barringer:1996:PAF} Howard Barringer,
Graham Gough, Brian Monahan, and Alan Williams. \newblock
A process algebra foundation for reasoning about {Core
ELLA}. \newblock {\em The Computer Journal}, 39\penalty0
(4):\penalty0 303--324, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/4/303.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_04/390303.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_04/Vol39_
04.body.html#AbstractBarringer|}\fi.
\bibitem[BGO94]{Bhagavathi:1994:GHL} D.~Bhagavathi, C.
~E. Grosch, and S.~Olariu. \newblock A greedy hypercube-
labeling algorithm. \newblock {\em The Computer Journal}
, 37\penalty0 (2):\penalty0 124--128, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
124.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_02/Vol37_02.body.html#AbstractBhagavathi|}
\fi.
\bibitem[BGR98]{Benjamin:1998:PIS} R.~Benjamin, B.~Gladman,
and B.~Randell. \newblock Protecting {IT} systems from
{CyberCrime}. \newblock {\em The Computer Journal}, 41\
penalty0 (7):\penalty0 429--443, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/7/
429.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_07/Vol41_07.body.html#AbstractBenjamin|}
\fi.

\bibitem[BGS99]{Bickmore:1999:WPF} Timothy Bickmore, Andreas
Girgensohn, and Joseph~W. Sullivan. \newblock {Web} page
filtering and re-authoring for mobile users. \newblock
{\em The Computer Journal}, 42\penalty0 (6):\penalty0
534--546, ????   1999. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/6/534.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
06/420534.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_06/pdf/420534.pdf|}\fi.
\bibitem[Bha91]{Bhaskar:1991:BR} K.~Bhaskar. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 396, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/396.
full.pdf+html|}\fi.
\bibitem[Bhu91]{Bhunia:1991:C} C.~T. Bhunia. \newblock
Correspondence. \newblock {\em The Computer Journal},
34\penalty0 (5):\penalty0 427, ???? 1991. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/5/
427.full.pdf+html|}\fi.
\bibitem[Bis92a]{Bishop:1992:CWT} J.~Bishop. \newblock
A compiler {Writer}'s toolbox. \newblock {\em The Computer
Journal}, 35\penalty0 (6):\penalty0 573, December   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/573.full.pdf+html|}\fi.
\bibitem[Bis92b]{Bishop:1992:IPS} Judy Bishop. \newblock
Introduction to {Pascal} and structured design; {Pascal}
plus data   structures, algorithms and advanced programming.
\newblock {\em The Computer Journal}, 35\penalty0 (3)
:\penalty0 255, June 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/3/255.full.pdf+html|}\fi.

\bibitem[BJ98]{Bodorik:1998:MVM} P.~Bodorik and D.~Jutla.
\newblock Multi-view memory support to operating systems
in locking for   transaction and database systems. \newblock
{\em The Computer Journal}, 41\penalty0 (2):\penalty0
84--97, ????   1998. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/2/84.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_02/
Vol41_02.body.html#AbstractBodorik|}\fi.
\bibitem[BJKW97]{Burke:1997:AUT} E.~Burke, K.~Jackson,
J.~H. Kingston, and R.~Weare. \newblock Automated university
timetabling: the state of the art. \newblock {\em The
Computer Journal}, 40\penalty0 (9):\penalty0 565--571,
????   1997. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/9/565.full.pdf+html|}\fi.
\bibitem[BK92]{Born:1992:TPB} R.~G. B{\"o}rn and J.~R.
Kenevan. \newblock Theoretical performance-based cost-
effectiveness of multicomputers. \newblock {\em The Computer
Journal}, 35\penalty0 (1):\penalty0 62--70,   February
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/1/62.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_01/Vol35_01.body.
html#AbstractBorn|}\fi.
\bibitem[BKLL95]{Brinksma:1995:SCB} Ed~Brinksma, Joost-
Pieter Katoen, Rom Langerak, and Diego Latella. \newblock
A stochastic causality-based process algebra. \newblock
{\em The Computer Journal}, 38\penalty0 (7):\penalty0
552--565, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/7/552.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractBrinksma|}\fi.
\bibitem[BKMK90]{Burton:1990:SNI} F.~W. Burton, J.~G.
Kollias, D.~G. Matsakis, and V.~G. Kollias. \newblock
Short note: Implementation of overlapping {B}-trees for
time and   space efficient representation of collections

of similar files. \newblock {\em The Computer Journal}
, 33\penalty0 (3):\penalty0 279--280, June   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/3/
279.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/279.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/280.
tif|}\fi.
\bibitem[BKMT96]{Bozanis:1996:RBI} Panayiotis Bozanis,
Nectarios Kitsios, Christos Makris, and Athanasios   Tsakalidis.
\newblock Red--blue intersection reporting for objects
of non-constant size. \newblock {\em The Computer Journal}
, 39\penalty0 (6):\penalty0 541--546, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/6/
541.full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_06/390541.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_06/Vol39_
06.body.html#AbstractBozanis|}\fi.
\bibitem[BKMT97]{Bozanis:1997:NRI} P.~Bozanis, N.~Kitsios,
C.~Makris, and A.~Tsakalidis. \newblock New results on
intersection query problems. \newblock {\em The Computer
Journal}, 40\penalty0 (1):\penalty0 22--29, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/1/22.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_01/Vol40_01.body.html#AbstractBozanis;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
01/Vol40_01.body.html#AbstractBozanis|}\fi.
\bibitem[BL90]{Bilodeau:1990:DSM} M.~Bilodeau and D.~
Laguitton. \newblock Development of screen management
subroutines compatible with   {Microsoft FORTRAN} compiler.
\newblock {\em The Computer Journal}, 33\penalty0 (5)
:\penalty0 444--452,   October 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/444.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/444.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/445.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/446.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/447.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/448.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/449.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/450.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/451.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/452.tif|}\fi.
\bibitem[BL91]{Burns:1991:FBD} A.~Burns and A.~M. Lister.
\newblock A framework for building dependable systems.
\newblock {\em The Computer Journal}, 34\penalty0 (2)
:\penalty0 173--181, April   1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/173.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/173.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/174.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/176.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/177.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/178.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/179.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/180.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/181.tif|}\fi.
\bibitem[Bla92]{Blair:1992:IRP} D.~C. Blair. \newblock
Information retrieval and the philosophy of language.
\newblock {\em The Computer Journal}, 35\penalty0 (3)
:\penalty0 200--207, June   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/200.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_03/Vol35_03.body.html#AbstractBlair|}
\fi.
\bibitem[BLPL92]{Bertossi:1992:SNP} A.~A. Bertossi, F.
~Luccio, L.~Pagli, and E.~Lodi. \newblock Short notes:
a parallel solution to the approximate string matching

 problem. \newblock {\em The Computer Journal}, 35\penalty0
(5):\penalty0 524--526,   October 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/5/524.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/350524.sgm.abs.html; http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/tiff/524.tif;
 http://www3.oup.co.uk/computer_journal/Volume_35/Issue_
05/tiff/525.tif;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/tiff/526.tif|}\fi.
\bibitem[BLR94]{Balter:1994:GL} Roland Balter, Serge Lacourte,
and Michel Riveill. \newblock The {Guide} language. \
newblock {\em The Computer Journal}, 37\penalty0 (6):
\penalty0 519--530, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/6/519.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_06/Vol37_06.body.html#AbstractBalter|}
\fi.
\bibitem[BM90a]{Baik:1990:TAT} K.~H. Baik and L.~L. Miller.
\newblock Topological approach for testing equivalence
in heterogeneous   relational databases. \newblock {\
em The Computer Journal}, 33\penalty0 (1):\penalty0 2-
-10, February   1990. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/1/2.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
01/tiff/10.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/3.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/4.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/
6.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/7.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/
9.tif|}\fi.
\bibitem[BM90b]{Bailey:1990:PDC} R.~L. Bailey and R.~

Mukkamala. \newblock Pipelining data compression algorithms.
\newblock {\em The Computer Journal}, 33\penalty0 (4)
:\penalty0 308--313,   August 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/308.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/308.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/309.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/310.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/311.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/312.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/313.tif|}\fi.
\bibitem[BM90c]{Bell:1990:TQO} S.~B.~M. Bell and D.~C.
Mason. \newblock Tesseral quaternions for the octtree.
\newblock {\em The Computer Journal}, 33\penalty0 (5)
:\penalty0 386--397,   October 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/386.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/386.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/387.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/388.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/390.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/391.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/392.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/393.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/395.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/396.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/397.
t.
\bibitem[BM92]{Barroca:1992:FMU} L.~M. Barroca and J.
~A. McDermid. \newblock Formal methods: use and relevance
for the development of   safety-critical systems. \newblock

{\em The Computer Journal}, 35\penalty0 (6):\penalty0
579--599,   December 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/6/579.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractBarroca|}\fi. \newblock See
editorial \cite{Cooke:1992:EFMa}.
\bibitem[BM93]{Burdorf:1993:LBS} C.~Burdorf and J.~Marti.
\newblock Load balancing strategies for {Time Warp} on
multi-user workstations. \newblock {\em The Computer Journal}
, 36\penalty0 (2):\penalty0 168--176, April   1993. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/2/168.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/Vol36_02.body.html#AbstractBurdorf|}
\fi.
\bibitem[BM94a]{Belkhatir:1994:SSD} Noureddine Belkhatir
and Walcelio~L. Melo. \newblock Supporting software development
processes in {Adele} 2. \newblock {\em The Computer Journal}
, 37\penalty0 (7):\penalty0 621--628, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
621.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_07/Vol37_07.body.html#AbstractBelkhatir|}
\fi.
\bibitem[BM94b]{Bundy:1994:NSC} Alan Bundy and Hector
MacQueen. \newblock The new software copyright law. \
newblock {\em The Computer Journal}, 37\penalty0 (2):
\penalty0 79--82, ????   1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/2/79.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_02/
Vol37_02.body.html#AbstractBundy|}\fi.
\bibitem[BMD92]{Burns:1992:MSS} A.~Burns, J.~McDermid,
and J.~Dobson. \newblock On the meaning of safety and
security. \newblock {\em The Computer Journal}, 35\penalty0
(1):\penalty0 3--15, February   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/35/1/3.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_01/Vol35_01.body.html#AbstractBurns|}
\fi.
\bibitem[BML94]{Bassiouni:1994:DIE} Mostafa~A. Bassiouni,
Amar Mukherjee, and Mark~J. Llewellyn. \newblock Design
and implementation of extended {Boolean} and comparison
 operators for time-oriented query languages. \newblock
{\em The Computer Journal}, 37\penalty0 (7):\penalty0
576--587, ????   1994. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/7/576.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_07/
Vol37_07.body.html#AbstractBassiouni|}\fi.
\bibitem[BMRS90]{Burton:1990:AUS} F.~W. Burton, G.~P.
McKeown, and V.~J. Rayward-Smith. \newblock Applications
of {UET} scheduling theory to the implementation of
declarative languages. \newblock {\em The Computer Journal}
, 33\penalty0 (4):\penalty0 330--336,   August 1990. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/4/330.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/330.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/331.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/332.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/333.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/334.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/335.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/336.tif|}\fi.
\bibitem[BMT95]{Bozanis:1995:PWF} Panayiotis Bozanis,
Christos Makris, and Athanasios Tsakalidis. \newblock
Parametric weighted filter: an efficient dynamic manipulation
of   signature files. \newblock {\em The Computer Journal}
, 38\penalty0 (6):\penalty0 479--488, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/6/
479.full.pdf+html;   http://www3.oup.co.uk/computer_journal/

Volume_38/Issue_06/Vol38_06.body.html#AbstractBozanis|}
\fi.
\bibitem[BMW93]{Bowyer:1993:C} A.~Bowyer, R.~Martin, and
J.~Woodwark. \newblock Correspondence. \newblock {\em
The Computer Journal}, 36\penalty0 (7):\penalty0 681,
???? 1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/7/681.1.full.pdf+html|}\fi.
\bibitem[Bol97]{Bolognesi:1997:EFC} T.~Bolognesi. \newblock
Expressive flexibility of a constraint-oriented notation.
\newblock {\em The Computer Journal}, 40\penalty0 (5)
:\penalty0 259--277, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/5/259.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_05/Vol40_05.body.html#AbstractBolognesi;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
05/Vol40_05.body.html#AbstractBolognesi|}\fi.
\bibitem[Bot94]{Bott:1994:SSC} Frank Bott. \newblock Software
systems construction with examples in ada. \newblock {
\em The Computer Journal}, 37\penalty0 (6):\penalty0 545-
-546, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/545.full.pdf+html|}\fi.
\bibitem[Bot95]{Bott:1995:PPS} M.~F. Bott. \newblock Programming
and problem solving with ada. \newblock {\em The Computer
Journal}, 38\penalty0 (3):\penalty0 263, ???? 1995. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/3/263.full.pdf+html|}\fi.
\bibitem[Bow90]{Bowden:1990:KMT} K.~Bowden. \newblock
{Kron}'s method of tearing on a {Transputer} array. \
newblock {\em The Computer Journal}, 33\penalty0 (5):
\penalty0 453--459,   October 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/453.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/453.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/454.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/455.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/456.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/457.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/458.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/459.
tif|}\fi.
\bibitem[Bow97]{Bowen:1997:BR} J.~Bowen. \newblock Book
review. \newblock {\em The Computer Journal}, 40\penalty0
(5):\penalty0 296--297, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/5/296.
2.full.pdf+html|}\fi.
\bibitem[BP92]{Barsi:1992:AFH} F.~Barsi and M.~C. Pinotti.
\newblock Adding flexibility to hybrid number systems.
\newblock {\em The Computer Journal}, 35\penalty0 (6)
:\penalty0 630--635,   December 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/630.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_06/Vol35_06.body.html#AbstractBarsi|}
\fi.
\bibitem[BP94]{Barsi:1994:TOM} Ferruccio Barsi and M.
~Cristina Pinotti. \newblock Time optimal mixed radix
conversion for residue number applications. \newblock
{\em The Computer Journal}, 37\penalty0 (10):\penalty0
907--916, ????   1994. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/10/907.full.pdf+html;
http://www3.oup.co.uk/computer_journal/Volume_37/Issue_
10/Vol37_10.body.html#AbstractBarsi|}\fi.
\bibitem[BPTV92]{Boiten:1992:HPC} E.~A. Boiten, H.~A.
Partsch, D.~Tuijnman, and N.~V{\"o}lker. \newblock How
to produce correct software --- an introduction to formal
 specification and program development by transformations.
\newblock {\em The Computer Journal}, 35\penalty0 (6)

:\penalty0 547--554,   December 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/547.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_06/Vol35_06.body.html#AbstractBoiten|}
\fi.
\bibitem[BR90]{Berghel:1990:PSM} H.~Berghel and R.~Rankin.
\newblock A proposed standard for measuring crossword
compilation efficiency. \newblock {\em The Computer Journal}
, 33\penalty0 (2):\penalty0 181--184, April   1990. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/2/181.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/182.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/183.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/184.
tif|}\fi.
\bibitem[BR91]{Bodorik:1991:TVP} P.~Bodorik and J.~S.
Riordon. \newblock Threshold values for processing distributed
queries. \newblock {\em The Computer Journal}, 34\penalty0
(6):\penalty0 551--558,   December 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/6/551.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/551.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/552.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/553.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/554.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/555.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/556.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/557.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/558.tif|}\fi.
\bibitem[Bra94]{Bradley:1994:CSS} James Bradley. \newblock
A complete ${L}$-system specification for generating an
exact   self-affine growth-decay function with a random

walk scaling property. \newblock {\em The Computer Journal}
, 37\penalty0 (1):\penalty0 58--66, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/1/
58.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractBradley|}
\fi.
\bibitem[Bre93]{Breene:1993:QHG} L.~A. Breene. \newblock
Quadtrees and hypercubes: grid embedding strategies based
on spatial   data structure addressing. \newblock {\em
The Computer Journal}, 36\penalty0 (6):\penalty0 562-
-569,   December 1993. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/562.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_06/
Vol36_06.body.html#AbstractBreene|}\fi.
\bibitem[Bri96]{Briscoe:1996:BR} E.~J. Briscoe. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (4):\penalty0 359--360, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/4/
359.full.pdf+html|}\fi.
\bibitem[Bro90]{Brokate:1990:C} K.~Brokate. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (2):\penalty0 188, ???? 1990. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/2/
188.1.full.pdf+html|}\fi.
\bibitem[Bro94a]{Brown:1994:BRb} C.~Brown. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(8):\penalty0 721--722, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/721.
full.pdf+html|}\fi.
\bibitem[Bro94b]{Brown:1994:BRa} H.~Brown. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(5):\penalty0 469--470, ????   1994. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/5/469.
2.full.pdf+html|}\fi.
\bibitem[Bro95]{Browne:1995:EMT} R.~F. Browne. \newblock
The embedding of meshes and trees into degree four chordal
ring   networks. \newblock {\em The Computer Journal}
, 38\penalty0 (1):\penalty0 71--77, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
71.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_01/Vol38_01.body.html#AbstractBrowne|}
\fi.
\bibitem[BRS91]{Brink:1991:SNA} C.~Brink, I.~M. Rewitzky,
and R.~A. Schmidt. \newblock Short notes: Autodescriptivity:
beware! \newblock {\em The Computer Journal}, 34\penalty0
(4):\penalty0 380--381,   August 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/380.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/380.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/381.
tif|}\fi.
\bibitem[Bru92]{Bruha:1992:AMS} I.~Bruha. \newblock {
AI} multilanguage system {McPOPLOG}: The power of communication
 between its subsystems. \newblock {\em The Computer Journal}
, 35\penalty0 (6):\penalty0 625--629,   December 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/625.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/Vol35_06.body.html#AbstractBruha|}
\fi.
\bibitem[BS99]{Bornat:1999:AFP} Richard Bornat and Bernard
Sufrin. \newblock Animating formal proof at the surface:
The {Jape} proof calculator. \newblock {\em The Computer
Journal}, 42\penalty0 (3):\penalty0 177--192, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/3/177.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_42/Issue_03/420177.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_03/pdf/420177.pdf|}\fi.
\bibitem[BSH90]{Bell:1990:IDD} D.~A. Bell, J.~Shao, and
M.~Elizabeth~C. Hull. \newblock Integrated deductive database
system implementation. {A} systematic   study. \newblock
{\em The Computer Journal}, 33\penalty0 (1):\penalty0
40--48,   February 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/1/40.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
01/tiff/40.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/41.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/42.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/43.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/44.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/45.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/47.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/48.tif|}\fi.
\bibitem[BTD99]{Bojovic:1999:IDT} Miroslav Bojovic, Milo
Tomasevic, and Jovan Djordjevic. \newblock The interactive
development and testing system for a {RISC}-style   processor.
\newblock {\em The Computer Journal}, 42\penalty0 (5)
:\penalty0 422--429, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/5/422.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_05/420422.sgm.abs.html|}\fi.
\bibitem[Bun97]{Bunton:1997:SMI} S.~Bunton. \newblock
Semantically motivated improvements for {PPM} variants.
\newblock {\em The Computer Journal}, 40\penalty0 (2-
-3):\penalty0 76--93, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/2_and_
3/76.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_02/Vol40_02.body.html#AbstractBunton;

 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_02.body.html#AbstractBunton;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_02/Vol40_03.body.
html#AbstractBunton|}\fi.
\bibitem[Bur90]{Burgess:1990:C} C.~G. Burgess. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (2):\penalty0 188, ???? 1990. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/2/
188.3.full.pdf+html|}\fi.
\bibitem[Bur92a]{Burnett:1992:YLC} Rachel Burnett. \newblock
Yearbook of law computers and technology, volume 5. \
newblock {\em The Computer Journal}, 35\penalty0 (1):
\penalty0 15, February   1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/1/15.full.pdf+html|}\fi.
\bibitem[Bur92b]{Burns:1992:CPR} Alan Burns. \newblock
Constructing predictable real-time systems. \newblock
{\em The Computer Journal}, 35\penalty0 (4):\penalty0
368a--368,   August 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/4/368-a.full.pdf+html|}
\fi.
\bibitem[But97]{Butler:1997:BR} M.~Butler. \newblock Book
review. \newblock {\em The Computer Journal}, 40\penalty0
(1):\penalty0 59--61, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/1/59.
full.pdf+html|}\fi.
\bibitem[BV92]{Bruza:1992:SHS} P.~D. Bruza and T.~P. {
Van Der Weide}. \newblock Stratified hypermedia structures
for information disclosure. \newblock {\em The Computer
Journal}, 35\penalty0 (3):\penalty0 208--220, June   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/3/208.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/Volume_35/Issue_03/Vol35_03.body.html#AbstractBruza|}
\fi.
\bibitem[BZN98]{Biasizzo:1998:SDA} Anton Biasizzo, Alenka
Zuzek, and Franc Novak. \newblock Sequential diagnosis
with asymmetrical tests. \newblock {\em The Computer Journal}
, 41\penalty0 (3):\penalty0 163--170, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/3/
163.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_03/Vol41_03.body.html#AbstractBiasizzo|}
\fi.
\bibitem[Cam94]{Caminer:1994:BR} D.~Caminer. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (5):\penalty0 467, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/5/467.
1.full.pdf+html|}\fi.
\bibitem[Cam95]{Campbell:1995:PEP} D.~K.~G. Campbell.
\newblock Parallel evolution of parallel processors. \
newblock {\em The Computer Journal}, 38\penalty0 (1):
\penalty0 81, ???? 1995. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/1/81.full.pdf+html|}\fi.
\bibitem[Cam96a]{Campbell:1996:BRa} D.~K.~G. Campbell.
\newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (6):\penalty0 556--557, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/6/
556.2.full.pdf+html|}\fi.
\bibitem[Cam96b]{Campbell:1996:BRb} D.~K.~G. Campbell.
\newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (7):\penalty0 644--645, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/7/
644.2.full.pdf+html|}\fi.

\bibitem[Cam96c]{Campbell:1996:BRc} D.~K.~G. Campbell.
\newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (8):\penalty0 723--724, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/8/
723.2.full.pdf+html|}\fi.
\bibitem[Cam96d]{Campbell:1996:BRd} D.~K.~G. Campbell.
\newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (8):\penalty0 726, ???? 1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/8/
726.1.full.pdf+html|}\fi.
\bibitem[Cam96e]{Campbell:1996:BRe} D.~K.~G. Campbell.
\newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (10):\penalty0 892--893, ????   1996. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/10/892.2.full.pdf+html|}\fi.
\bibitem[{Cam}97]{Caminer:1997:LAB} David~Tresman {Caminer,
OBE}. \newblock {LEO} and its applications: the beginning
of business computing. \newblock {\em The Computer Journal}
, 40\penalty0 (10):\penalty0 585--597, ????   1997. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/10/585.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_10/Vol40_10.body.html#AbstractCaminer|}
\fi.
\bibitem[Can93a]{Canny:1993:CRG} John Canny. \newblock
Computing roadmaps of general semi-algebraic sets. \newblock
{\em The Computer Journal}, 36\penalty0 (5):\penalty0
504--514,   October 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/5/504.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractCanny2|}\fi.
\bibitem[Can93b]{Canny:1993:IAS} John Canny. \newblock
Improved algorithms for sign determination and existential

quantifier   elimination. \newblock {\em The Computer
Journal}, 36\penalty0 (5):\penalty0 409--418,   October
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/5/409.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_05/Vol36_05.body.
html#AbstractCanny|}\fi.
\bibitem[Car92]{Carlsson:1992:NH} S.~Carlsson. \newblock
A note on heapsort. \newblock {\em The Computer Journal}
, 35\penalty0 (4):\penalty0 410--411,   August 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/4/410.full.pdf+html|}\fi.
\bibitem[Caw96]{Cawsey:1996:BR} A.~Cawsey. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(3):\penalty0 251, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/3/251.1.full.pdf+html|}
\fi.
\bibitem[CBH{\etalchar{+}}94]{Cahill:1994:CDA} Vinny Cahill,
Roland Balter, David Harper, Neville Harris, Xavier Rousset~
De   Pina, and Pedro Sousa. \newblock The {Comandos} distributed
application platform. \newblock {\em The Computer Journal}
, 37\penalty0 (6):\penalty0 477--486, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
477.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_06/Vol37_06.body.html#AbstractCahill|}
\fi.
\bibitem[CC92]{Chiaramella:1992:ARM} Y.~Chiaramella and
J.~P. Chevallet. \newblock About retrieval models and
logic. \newblock {\em The Computer Journal}, 35\penalty0
(3):\penalty0 233--242, June   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/233.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_03/Vol35_03.body.html#AbstractChiaramella|}
\fi.

\bibitem[CC93]{Chang:1993:NLA} Henry Ker-Chang Chang and
Shing~Hong Chen. \newblock A new locally adaptive data
compression scheme using multilist   structure. \newblock
{\em The Computer Journal}, 36\penalty0 (6):\penalty0
570--578,   December 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/570.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_06/
Vol36_06.body.html#AbstractChang|}\fi.
\bibitem[CC96]{Chen:1996:TBA} Ing-Ray Chen and Chi-Ming
Chen. \newblock Threshold-based admission control policies
for multimedia servers. \newblock {\em The Computer Journal}
, 39\penalty0 (9):\penalty0 757--766, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
757.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_39/Issue_09/Vol39_09.body.html#AbstractChen1;
http://www3.oup.co.uk/computer_journal/Volume_39/Issue_
09/Vol39_09.body.html#AbstractChen1|}\fi.
\bibitem[CCB95]{Clarke:1995:AST} Charles L.~A. Clarke,
G.~V. Cormack, and F.~J. Burkowski. \newblock An algebra
for structured text search and a framework for its   implementation.
\newblock {\em The Computer Journal}, 38\penalty0 (1)
:\penalty0 43--56, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/1/43.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_01/Vol38_01.body.html#AbstractClarke|}
\fi.
\bibitem[CCC92]{Chang:1992:SGC} C.~C. Chang, H.~Y. Chen,
and C.~Y. Chen. \newblock Symbolic {Gray} code as a data
allocation scheme for two-disc   systems. \newblock {
\em The Computer Journal}, 35\penalty0 (3):\penalty0 299-
-305, June   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/3/299.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractChang|}\fi.

\bibitem[CCJ91]{Chang:1991:DMI} C.~C. Chang, C.~Y. Chen,
and J.~K. Jan. \newblock On the design of a machine-independent
perfect hashing scheme. \newblock {\em The Computer Journal}
, 34\penalty0 (5):\penalty0 469--474,   October 1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/5/469.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/469.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/470.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/471.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/472.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/473.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/474.tif|}\fi.
\bibitem[CCL98]{Chen:1998:PEF} Ing-Ray Chen, Tsong-Min
Chen, and Chiang Lee. \newblock Performance evaluation
of forwarding strategies for location   management in
mobile networks. \newblock {\em The Computer Journal}
, 41\penalty0 (4):\penalty0 243--253, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/4/
243.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_04/Vol41_04.body.html#AbstractChen|}\
fi.
\bibitem[CCMS97]{Chan:1997:SPO} Fun~Ting Chan, Tsong~
Yueh Chen, Ieng~Kei Mark, and Shir~Ming Shen. \newblock
On some properties of the optimally refined proportional
sampling   strategy. \newblock {\em The Computer Journal}
, 40\penalty0 (4):\penalty0 194--199, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/4/
194.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_03/Vol40_04.body.html#AbstractChan;
http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
04/Vol40_04.body.html#AbstractChan|}\fi.
\bibitem[CDCL97]{Chan:1997:TPF} E.~Chan, C.~Daoxu, J.
~Cao, and C.-H. Lee. \newblock Timing properties of the
{FDDI-M} medium access protocol. \newblock {\em The Computer
Journal}, 40\penalty0 (1):\penalty0 43--49, ????   1997.

\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/1/43.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_01/Vol40_01.body.html#AbstractChan;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
01/Vol40_01.body.html#AbstractChan|}\fi.
\bibitem[CDHL93]{Charlton:1993:AGR} C.~C. Charlton, P.
~E. Dunne, K.~Halewood, and P.~H. Leng. \newblock An algorithm
to generate random large combinational circuits. \newblock
{\em The Computer Journal}, 36\penalty0 (2):\penalty0
200--206, April   1993. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/2/200.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractCharlton|}\fi.
\bibitem[CF96]{Cifuentes:1996:ARC} Cristina Cifuentes
and Anne Fitzgerald. \newblock {Australian} recommendations
on computer software protection. \newblock {\em The Computer
Journal}, 39\penalty0 (7):\penalty0 566--576, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/7/566.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_07/390566.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_07/
Vol39_07.body.html#AbstractCifuentes|}\fi.
\bibitem[CG92]{Conlon:1992:DTC} T.~Conlon and S.~Gregory.
\newblock Debugging tools for concurrent logic programming.
\newblock {\em The Computer Journal}, 35\penalty0 (2)
:\penalty0 157--169, April   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/2/157.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_02/Vol35_02.body.html#AbstractConlon|}
\fi.
\bibitem[CG99]{Chinchwadkar:1999:OVP} G.~S. Chinchwadkar
and A.~Goh. \newblock An overview of vertical partitioning
in object oriented databases. \newblock {\em The Computer
Journal}, 42\penalty0 (1):\penalty0 39--50, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-

4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/1/39.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_01/420039.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_01/pdf/420039.pdf|}\fi.
\bibitem[CGRJ92]{Cawsey:1992:ALB} Alison Cawsey, Julia
Galliers, Steven Reece, and Karen~Sparck Jones. \newblock
Automating the librarian: Belief revision as a base for
system action   and communication with the user. \newblock
{\em The Computer Journal}, 35\penalty0 (3):\penalty0
221--232, June   1992. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/3/221.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractCawsey|}\fi.
\bibitem[CH99]{Chuang:1999:EFD} Po-Jen Chuang and Huey-
Ming Hsu. \newblock Effective fault diagnosis approaches
for multipath networks. \newblock {\em The Computer Journal}
, 42\penalty0 (5):\penalty0 409--421, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/5/
409.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_05/420409.sgm.abs.html|}\fi.
\bibitem[Cha90a]{Chaitin:1990:C} G.~J. Chaitin. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (1):\penalty0 92, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/92.
1.full.pdf+html|}\fi.
\bibitem[Cha90b]{Chaudhuri:1990:FUD} P.~Chaudhuri. \newblock
Finding and updating depth-first spanning trees of acyclic
digraphs   in parallel. \newblock {\em The Computer Journal}
, 33\penalty0 (3):\penalty0 247--251, June   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/3/
247.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/247.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/248.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/249.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/250.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/251.tif|}\fi.
\bibitem[Cha91]{Challab:1991:IFA} D.~J. Challab. \newblock
Implementation of flexible arrays using balanced trees.
\newblock {\em The Computer Journal}, 34\penalty0 (5)
:\penalty0 386--396,   October 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/386.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/386.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/387.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/388.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/389.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/390.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/391.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/392.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/393.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/394.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/395.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/396.tif|}\fi.
\bibitem[Cha92]{Challab:1992:SNF} D.~J. Challab. \newblock
Short notes: Flexible arrays: evaluation of the {B}-tree
based   system. \newblock {\em The Computer Journal},
35\penalty0 (5):\penalty0 527--529,   October 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/5/527.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/350527.sgm.abs.html; http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_05/
tiff/527.tif;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/tiff/528.tif;   http://www3.oup.co.
uk/computer_journal/Volume_35/Issue_05/tiff/529.tif|}
\fi.
\bibitem[Cha93a]{Challis:1993:TNT} M.~F. Challis. \newblock

Two new techniques for computing extremal $h$-bases $
{A}_k$. \newblock {\em The Computer Journal}, 36\penalty0
(2):\penalty0 117--126, April   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/2/117.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_02/Vol36_02.body.html#AbstractChallis|}
\fi.
\bibitem[Cha93b]{Chandru:1993:VEL} Vijay Chandru. \newblock
Variable elimination in linear constraints. \newblock
{\em The Computer Journal}, 36\penalty0 (5):\penalty0
463--472,   October 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/5/463.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractChandru|}\fi.
\bibitem[Cha97]{Chaudhuri:1997:ODA} Pranay Chaudhuri.
\newblock An optimal distributed algorithm for computing
bridge- connected   components. \newblock {\em The Computer
Journal}, 40\penalty0 (4):\penalty0 200--207, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/4/200.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_03/Vol40_04.body.html#AbstractChaudhuri;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
04/Vol40_04.body.html#AbstractChaudhuri|}\fi.
\bibitem[Che90]{Chen:1990:CMV} T.~Y. Chen. \newblock On
the consistency of multi-valued functions. \newblock {
\em The Computer Journal}, 33\penalty0 (6):\penalty0 570-
-572,   December 1990. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/6/570.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/570.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/571.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/572.
tif|}\fi.
\bibitem[Che95a]{Chen:1995:AIC} Ing-Ray Chen. \newblock
On applying imprecise computation to real-time {AI} systems.

\newblock {\em The Computer Journal}, 38\penalty0 (6)
:\penalty0 434--442, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/6/434.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_06/Vol38_06.body.html#AbstractChen|}\
fi.
\bibitem[Che95b]{Chen:1995:SPN} Ing-Ray Chen. \newblock
Stochastic {Petri} net analysis of deadlock detection
algorithms in   transaction database systems with dynamic
locking. \newblock {\em The Computer Journal}, 38\penalty0
(9):\penalty0 717--733, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/9/717.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_09/Vol38_09.body.html#AbstractChen|}\
fi.
\bibitem[Che95c]{Chen:1995:DBT} Ing-Ray~Ray Chen. \newblock
A degradable {B}$^{\rm link}$-tree with periodic data
reorganization. \newblock {\em The Computer Journal},
38\penalty0 (3):\penalty0 245--252, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/3/
245.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_03/Vol38_03.body.html#AbstractChen|}\
fi.
\bibitem[Che95d]{Chen:1995:ECA} Jingsen Chen. \newblock
An efficient construction algorithm for a class of implicit
 double-ended priority queues. \newblock {\em The Computer
Journal}, 38\penalty0 (10):\penalty0 818--821, ????
1995. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/10/818.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_38/Issue_10/Vol38_10.body.
html#AbstractChen|}\fi.
\bibitem[CHM94]{Christensen:1994:DSC} S{\o}ren Christensen,
Yoram Hirshfeld, and Faron Moller. \newblock Decidable
subsets of {CCS}. \newblock {\em The Computer Journal}
, 37\penalty0 (4):\penalty0 233--242, ????   1994. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
233.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_04/Vol37_04.body.html#AbstractChristensen|}
\fi.
\bibitem[Cho95]{Chou:1995:MVD} Ching-Tsun Chou. \newblock
Mechanical verification of distributed algorithms in higher-
order   logic. \newblock {\em The Computer Journal}, 38\
penalty0 (2):\penalty0 152--161, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/2/
152.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_02/Vol38_02.body.html#AbstractChou|}\
fi.
\bibitem[Cho99]{Choy:1999:CIR} M.~Choy. \newblock Consistent
implementations of replicated objects. \newblock {\em
The Computer Journal}, 42\penalty0 (1):\penalty0 24--
38, ????   1999. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/1/24.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
01/420024.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_01/pdf/420024.pdf|}\fi.
\bibitem[CK93]{Cooper:1993:SRI} Tim~B. Cooper and Jeffrey~
H. Kingston. \newblock The solution of real instances
of the timetabling problem. \newblock {\em The Computer
Journal}, 36\penalty0 (7):\penalty0 645--653, ????   1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/7/645.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_07/Vol36_07.body.html#AbstractCooper|}
\fi.
\bibitem[CK94]{Chu:1994:ASH} Jiang-Hsing Chu and Gary~
D. Knott. \newblock An analysis of spiral hashing. \newblock
{\em The Computer Journal}, 37\penalty0 (8):\penalty0
715--719, ????   1994. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/8/715.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/Volume_37/Issue_08/
Vol37_08.body.html#AbstractChu|}\fi.
\bibitem[CK95]{Campbell-Kelly:1995:IEB} M.~Campbell-Kelly.
\newblock {ICL} and the evolution of the {British} mainframe.
\newblock {\em The Computer Journal}, 38\penalty0 (5)
:\penalty0 400--412, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/5/400.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_05/Vol38_05.body.html#AbstractKelly|}
\fi.
\bibitem[CL99a]{Chen:1999:AMP} Guan-Chi Chen and Suh-
Yin Lee. \newblock An analytic model for performance analysis
of concurrency control   strategies in mobile environments.
\newblock {\em The Computer Journal}, 42\penalty0 (6)
:\penalty0 511--521, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/6/511.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_06/420511.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_06/
pdf/420511.pdf|}\fi.
\bibitem[CL99b]{Chen:1999:CTS} Tsong~Yueh Chen and Man~
Fai Lau. \newblock On the completeness of a test suite
reduction strategy. \newblock {\em The Computer Journal}
, 42\penalty0 (5):\penalty0 430--440, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/5/
430.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_05/420430.sgm.abs.html|}\fi.
\bibitem[Cla90]{Clark:1990:PLP} K.~L. Clark. \newblock
Parallel logic programming. \newblock {\em The Computer
Journal}, 33\penalty0 (6):\penalty0 482--493,   December
1990. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/33/6/482.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/482.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/483.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_06/tiff/484.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/485.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/486.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/487.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/488.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/489.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/490.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/491.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/492.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/493.t.
\bibitem[Cla92a]{Clarkson:1992:BRB} Trevor Clarkson. \
newblock Book review: {{\booktitle{VLSI Design of Neural
Networks}}, by Ulrich   Ramacher and Ulrich Ruckert}.
\newblock {\em The Computer Journal}, 35\penalty0 (6)
:\penalty0 624, December   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://www3.oup.co.uk/computer_journal/hdb/Volume_
35/Issue_06/tiff/624.tif|}\fi.
\bibitem[Cla92b]{Clarkson:1992:VDN} Trevor Clarkson. \
newblock Vlsi design of neural networks. \newblock {\
em The Computer Journal}, 35\penalty0 (6):\penalty0 624,
December   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/6/624.full.pdf+html|}\fi.
\bibitem[Cla94]{Clark:1994:PRN} A.~N. Clark. \newblock
Pattern recognition of noisy sequences of behavioural
events using   functional combinators. \newblock {\em
The Computer Journal}, 37\penalty0 (5):\penalty0 385-
-398, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/5/385.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractClark|}\fi.
\bibitem[Cla99]{Clarke:1999:DPR} B.~Clarke. \newblock
Discussion of the papers by {Rissanen}, and by {Wallace}

and {Dowe}. \newblock {\em The Computer Journal}, 42\
penalty0 (4):\penalty0 338--339, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/4/
338.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420338.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420338.pdf|}\fi.
\bibitem[CLC92]{Chang:1992:OPD} C.~C. Chang, H.~F. Lin,
and C.~Y. Chen. \newblock Optimality properties of the
disc modulo allocation method for   $k$-ary multiple key
hashing files. \newblock {\em The Computer Journal}, 35\
penalty0 (4):\penalty0 390--394,   August 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/4/
390.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_04/Vol35_04.body.html#AbstractChang|}
\fi.
\bibitem[Cle90]{Clement:1990:C} B.~E.~P. Clement. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (5):\penalty0 475, ???? 1990. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/5/
475.full.pdf+html|}\fi.
\bibitem[Cle92]{Clement:1992:RDR} T.~Clement. \newblock
The role of data reification in program refinement: Origins,
 synthesis and appraisal. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 451--459,   October
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/451.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/Vol35_05.body.
html#AbstractClement|}\fi.
\bibitem[Cli94]{Clippingdale:1994:PCV} Simon Clippingdale.
\newblock Practical computer vision using {C}. \newblock
{\em The Computer Journal}, 37\penalty0 (6):\penalty0
543--544, ????   1994. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.

oxfordjournals.org/content/37/6/543.full.pdf+html|}\fi.
\bibitem[CLW98]{Chu:1998:EHB} K.~W. Chu, S.~K. Lam, and
M.~H. Wong. \newblock An efficient hash-based algorithm
for sequence data searching. \newblock {\em The Computer
Journal}, 41\penalty0 (6):\penalty0 402--415, ????   1998.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/6/402.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_41/Issue_06/Vol41_06.body.html#AbstractChu|}
\fi.
\bibitem[CM91a]{Campbell:1991:FMW} D.~Campbell and T.
~McNeill. \newblock Finding a majority when sorting is
not available. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 186, April   1991. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/186.tif|}\fi. \newblock See correspondence
\cite{Gries:1992:CFM}.
\bibitem[CM91b]{Campbell:1991:SNF} D.~Campbell and T.
~McNeill. \newblock Short note: Finding a majority when
sorting is not available. \newblock {\em The Computer
Journal}, 34\penalty0 (2):\penalty0 186, ???? 1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/2/186.full.pdf+html|}\fi.
\bibitem[CM93]{Czech:1993:LTA} Zbigniew~J. Czech and Bohdan~
S. Majewski. \newblock A linear time algorithm for finding
minimal perfect hash functions. \newblock {\em The Computer
Journal}, 36\penalty0 (6):\penalty0 579--587,   December
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/6/579.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_06/Vol36_06.body.
html#AbstractCzech|}\fi.
\bibitem[CMW98]{Cockshott:1998:HPO} W.~Paul Cockshott,
Douglas McGregor, and John Wilson. \newblock High performance
operations using a compressed database architecture. \

newblock {\em The Computer Journal}, 41\penalty0 (5):
\penalty0 283--296, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/5/283.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_05/Vol41_05.body.html#AbstractCockshott|}
\fi.
\bibitem[CNTR97]{Cahill:1997:OMD} Vinny Cahill, Paddy
Nixon, Brendan Tangney, and Fethi Rabhi. \newblock Object
models for distributed or persistent programming. \newblock
{\em The Computer Journal}, 40\penalty0 (8):\penalty0
513--527, ????   1997. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/8/513.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_40/Issue_08/
Vol40_08.body.html#AbstractCahill|}\fi.
\bibitem[CO96]{Cardell-Oliver:1996:BR} R.~Cardell-Oliver.
\newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (10):\penalty0 892, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/10/
892.1.full.pdf+html|}\fi.
\bibitem[Coc97]{Cochrane:1997:CGR} P.~Cochrane. \newblock
From copper to glass: the right idea, decisions and investments
at   the right time\ldots{}. \newblock {\em The Computer
Journal}, 40\penalty0 (1):\penalty0 1--11, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/1/1.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_01/Vol40_01.body.html#AbstractCochrane;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
01/Vol40_01.body.html#AbstractCochrane|}\fi.
\bibitem[Coh01]{Cohen:2001:RCF} Fred Cohen. \newblock
Reply to {``Comment on `A Framework for Modelling Trojans
and   Computer Virus Infection' by E. M{\"a}kinen''}.
\newblock {\em The Computer Journal}, 44\penalty0 (4)
:\penalty0 326--327, ????   2001. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://www3.oup.co.uk/computer_journal/hdb/Volume_
44/Issue_04/440326.sgm.abs.html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_44/Issue_04/pdf/440326.
pdf|}\fi. \newblock See \cite{Thimbleby:1998:FMT,Makinen:
2001:CFM,Thimbleby:2001:RCF}.
\bibitem[Col97]{Colomb:1997:ISH} R.~M. Colomb. \newblock
Impact of semantic heterogeneity on federating databases.
\newblock {\em The Computer Journal}, 40\penalty0 (5)
:\penalty0 235--244, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/5/235.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_05/Vol40_05.body.html#AbstractColomb;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
05/Vol40_05.body.html#AbstractColomb|}\fi.
\bibitem[Coo92a]{Cooke:1992:FMM} J.~Cooke. \newblock Formal
methods --- mathematics, theory, recipes or what? \newblock
{\em The Computer Journal}, 35\penalty0 (5):\penalty0
419--423,   October 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/5/419.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractCooke|}\fi. \newblock See editorial
\cite{Cooke:1992:EFMa,Cooke:1992:EFMb}.
\bibitem[Coo92b]{Cooke:1992:EFMa} John Cooke. \newblock
Editorial --- formal methods: What? why? and when? \newblock
{\em The Computer Journal}, 35\penalty0 (5):\penalty0
417--418,   October 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/5/417.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_35/Issue_
05/tiff/417.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_35/Issue_05/tiff/418.tif|}\fi.
\bibitem[Coo92c]{Cooke:1992:EFMb} John Cooke. \newblock
Editorial --- formal methods: What? why? and when? \newblock
{\em The Computer Journal}, 35\penalty0 (6):\penalty0
545--546,   December 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.

oxfordjournals.org/content/35/6/545.full.pdf+html|}\fi.
\bibitem[Coo92d]{Cooley:1992:IES} R.~E. Cooley. \newblock
The integration of expert systems into mainstream software.
\newblock {\em The Computer Journal}, 35\penalty0 (6)
:\penalty0 610, December   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/610.
full.pdf+html|}\fi.
\bibitem[Coo94]{Cooke:1994:BR} D.~J. Cooke. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (7):\penalty0 644, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/7/644.
1.full.pdf+html|}\fi.
\bibitem[Coo98]{Cooper:1998:BR} R.~Cooper. \newblock Book
review. \newblock {\em The Computer Journal}, 41\penalty0
(1):\penalty0 64--65, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/1/64.
full.pdf+html|}\fi.
\bibitem[Cor93]{Corne:1993:BR} D.~Corne. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 604--605, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/604.
full.pdf+html|}\fi.
\bibitem[CPV93]{Catania:1993:MPE} V.~Catania, A.~Puliafito,
and L.~Vita. \newblock A model for performance evaluation
of gracefully degrading systems. \newblock {\em The Computer
Journal}, 36\penalty0 (2):\penalty0 177--185, April
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/2/177.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_02/Vol36_02.body.
html#AbstractCatania|}\fi.

\bibitem[CR97]{Chandy:1997:SCD} K.~Mandy Chandy and Adam
Rikfin. \newblock Systematic composition of distributed
objects: Processes and   sessions. \newblock {\em The
Computer Journal}, 40\penalty0 (8):\penalty0 465--478,
????   1997. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/8/465.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_08/Vol40_08.body.
html#AbstractChandy|}\fi.
\bibitem[Cra93]{Craig:1993:BR} I.~Craig. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 601, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/601.2.full.pdf+html|}
\fi.
\bibitem[Cra94a]{Craig:1994:BRa} I.~Craig. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(2):\penalty0 145, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/2/145.1.full.pdf+html|}
\fi.
\bibitem[Cra94b]{Craig:1994:BRb} I.~Craig. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(3):\penalty0 224--225, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/224.
full.pdf+html|}\fi.
\bibitem[Cra94c]{Craig:1994:BR} I.~Craig. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(4):\penalty0 347--348, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/347.
full.pdf+html|}\fi.
\bibitem[Cra94d]{Craig:1994:CLP} Iain Craig. \newblock
Constraint logic programming: Selected research {\&} constraint-
based   reasoning. \newblock {\em The Computer Journal}
, 37\penalty0 (6):\penalty0 557a--559, ????   1994. \

newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/6/557-a.full.pdf+html|}\fi.
\bibitem[Cre96a]{Crestani:1996:BRa} F.~Crestani. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (4):\penalty0 358, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/4/358.
1.full.pdf+html|}\fi.
\bibitem[Cre96b]{Crestani:1996:BRb} F.~Crestani. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 728--729, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/8/
728.2.full.pdf+html|}\fi.
\bibitem[Cre98]{Crestani:1998:BR} F.~Crestani. \newblock
Book review. \newblock {\em The Computer Journal}, 41\
penalty0 (6):\penalty0 428, ???? 1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/6/428.
full.pdf+html|}\fi.
\bibitem[Cro94]{Crole:1994:CR} Roy~L. Crole. \newblock
Computation and reasoning. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 552a--553, ????
1994. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/6/552-a.full.pdf+html|}\fi.
\bibitem[CRS{\etalchar{+}}93]{Clark:1993:SLC} D.~A. Clark,
C.~J. Rawlings, J.~Shirazi, L.-L. Li, K.~Schuerman, M.
~Reeve, and   A.~V{\'e}ron. \newblock Solving large combinatorial
problems in molecular biology using the   {ElipSys} parallel
constraint logic programming system. \newblock {\em The
Computer Journal}, 36\penalty0 (8):\penalty0 690--701,
????   1993. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/8/690.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/Volume_36/Issue_08/Vol36_08.body.
html#AbstractClark|}\fi.
\bibitem[CS95]{Chatterjee:1995:DPS} Saurav Chatterjee
and Jay Strosnider. \newblock Distributed pipeline scheduling:
a framework for distributed,   heterogeneous real-time
system design. \newblock {\em The Computer Journal}, 38\
penalty0 (4):\penalty0 271--285, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/4/
271.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_04/Vol38_04.body.html#AbstractChatterjee|}
\fi.
\bibitem[CT92]{Czejdo:1992:IIS} B.~Czejdo and M.~C. Taylor.
\newblock Integration of information systems using an
object-oriented approach. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 501--513,   October
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/501.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/Vol35_05.body.
html#AbstractCzejdo|}\fi.
\bibitem[CT94]{Cai:1994:ART} Wentong Cai and Stephen~
J. Turner. \newblock An approach to the run-time monitoring
of parallel programs. \newblock {\em The Computer Journal}
, 37\penalty0 (4):\penalty0 333--345, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
333.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_04/Vol37_04.body.html#AbstractCai|}\fi.
\bibitem[CT97]{Cleary:1997:ULC} J.~G. Cleary and W.~J.
Teahan. \newblock Unbounded length contexts for {PPM}
. \newblock {\em The Computer Journal}, 40\penalty0 (
2--3):\penalty0 67--75, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/2_and_
3/67.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_02/Vol40_02.body.html#AbstractCleary;

 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_02.body.html#AbstractCleary|}\fi.
\bibitem[CTH99]{Chi:1999:FMM} Kuang-Hwei Chi, Chien-Chao
Tseng, and Ting-Lu Huang. \newblock A framework for mobile
multicast using dynamic route reconstructions. \newblock
{\em The Computer Journal}, 42\penalty0 (6):\penalty0
522--533, ????   1999. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/6/522.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
06/420522.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_06/pdf/420522.pdf|}\fi.
\bibitem[CTW94]{Chan:1994:EOO} D.~K.~C. Chan, P.~W. Trinder,
and R.~C. Welland. \newblock Evaluating object-oriented
query languages. \newblock {\em The Computer Journal}
, 37\penalty0 (10):\penalty0 858--872, ????   1994. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/10/858.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_10/tiff/858.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_37/Issue_
10/tiff/859.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_37/Issue_10/tiff/860.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_37/Issue_10/tiff/861.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
37/Issue_10/tiff/862.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_10/tiff/863.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_37/Issue_
10/tiff/864.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_37/Issue_10/tiff/865.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_37/Issue_10/tiff/866.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
37/Issue_10/tiff/867.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_10/tiff/868.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_37/Issue_
10/tiff/869.
\bibitem[Cuc93]{Cucker:1993:CQE} Felipe Cucker. \newblock
On the complexity of quantifier elimination: the structural
approach. \newblock {\em The Computer Journal}, 36\penalty0
(5):\penalty0 400--408,   October 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/5/400.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_05/Vol36_05.body.html#AbstractCucker|}
\fi.
\bibitem[Cul96]{Culwin:1996:BR} F.~Culwin. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(3):\penalty0 251--252, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/3/251.
2.full.pdf+html|}\fi.
\bibitem[Cur95]{Curzon:1995:TDC} Paul Curzon. \newblock
Tracking design changes with formal machine-checked proof.
\newblock {\em The Computer Journal}, 38\penalty0 (2)
:\penalty0 91--100, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/2/91.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_02/Vol38_02.body.html#AbstractCurzon|}
\fi.
\bibitem[Cur96]{Curzon:1996:BR} P.~Curzon. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(5):\penalty0 449, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/5/449.1.full.pdf+html|}
\fi.
\bibitem[CW96]{Chen:1996:ARD} I.-R. Chen and D.-C. Wang.
\newblock Analysis of replicated data with repair dependency.
\newblock {\em The Computer Journal}, 39\penalty0 (9)
:\penalty0 767--779, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/9/767.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_39/Issue_09/Vol39_09.body.html#AbstractChen2;
http://www3.oup.co.uk/computer_journal/Volume_39/Issue_
09/Vol39_09.body.html#AbstractChen2|}\fi.
\bibitem[CW98]{Cavalcanti:1998:WPS} Ana Cavalcanti and
Jim Woodcock. \newblock A weakest precondition semantics

for {$Z$}. \newblock {\em The Computer Journal}, 41\penalty0
(1):\penalty0 1--15, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/1/1.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_01/Vol41_01.body.html#AbstractCavalcanti|}
\fi.
\bibitem[CWH93]{Chao:1993:CTM} Daniel~Y. Chao, David~
T. Wang, and Douglas D.~C. Hung. \newblock Convergence
time and memory capacity of higher-order {Hopfield}
associative memory with multi-valued neurons. \newblock
{\em The Computer Journal}, 36\penalty0 (6):\penalty0
554--561,   December 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/554.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_06/
Vol36_06.body.html#AbstractChao|}\fi.
\bibitem[CY95]{Chen:1995:APE} Ing-Ray Chen and I.-Ling
Yen. \newblock Analysis of probabilistic error checking
procedures on storage   systems. \newblock {\em The Computer
Journal}, 38\penalty0 (5):\penalty0 348--354, ????   1995.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/5/348.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_38/Issue_05/Vol38_05.body.html#AbstractChen|}
\fi.
\bibitem[CY96]{Chen:1996:CSP} Tsong~Yueh Chen and Yuen~
Tak Yu. \newblock Constraints for safe partition testing
strategies. \newblock {\em The Computer Journal}, 39\
penalty0 (7):\penalty0 619--625, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/7/
619.full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_07/390619.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_07/Vol39_
07.body.html#AbstractChen|}\fi.
\bibitem[Cza94a]{Czarnecki:1994:BRa} C.~Czarnecki. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 725, ???? 1994. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/725.
1.full.pdf+html|}\fi.
\bibitem[Cza94b]{Czarnecki:1994:BRb} C.~Czarnecki. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 730, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/730.
1.full.pdf+html|}\fi.
\bibitem[Cza96]{Czarnecki:1996:BR} C.~A. Czarnecki. \
newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (1):\penalty0 94, ???? 1996. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/1/
94.full.pdf+html|}\fi.
\bibitem[Cze98]{Czech:1998:QPH} Zbigniew~J. Czech. \newblock
Quasi-perfect hashing. \newblock {\em The Computer Journal}
, 41\penalty0 (6):\penalty0 416--421, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/6/
416.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_06/Vol41_06.body.html#AbstractCzech|}
\fi.
\bibitem[CZL97]{Cierniak:1997:CTS} Michal Cierniak, Mohammed~
Javeed Zaki, and Wei Li. \newblock Compile-time scheduling
algorithms for a heterogeneous network of   workstations.
\newblock {\em The Computer Journal}, 40\penalty0 (6)
:\penalty0 356--372, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/6/356.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_06/Vol40_06.body.html#AbstractCierniak;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
06/Vol40_06.body.html#AbstractCierniak|}\fi.
\bibitem[CZW94]{Chao:1994:EKT} Daniel~Y. Chao, Mengchu~
C. Zhou, and David~T. Wang. \newblock Extending the knitting
technique to {Petri Net} synthesis of   automated manufacturing

systems. \newblock {\em The Computer Journal}, 37\penalty0
(1):\penalty0 67--76, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/1/67.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractChao|}\
fi.
\bibitem[Dai94]{Dain:1994:BR} J.~A. Dain. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(2):\penalty0 144, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/2/144.full.pdf+html|}\fi.
\bibitem[Dam95]{Damodaran:1995:SSU} Leela Damodaran. \
newblock Studio-structured user-interface design for interaction
optimisation. \newblock {\em The Computer Journal}, 38\
penalty0 (3):\penalty0 267--268, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/3/
267.full.pdf+html|}\fi.
\bibitem[Dan92]{Dandeker:1992:IPV} K.~N. Dandeker. \newblock
Implementation of parameterised views in relational query
languages   with a macro processor. \newblock {\em The
Computer Journal}, 35\penalty0 (5):\penalty0 A425--??
,   October 1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/35000A425.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/Volume_
35/Issue_05/tiff/A425.tif|}\fi.
\bibitem[Dan93]{Dancik:1993:BR} V.~Danc{\'\i}k. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (7):\penalty0 679, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/7/679.
1.full.pdf+html|}\fi.
\bibitem[Dan94a]{Danck:1994:BRa} V.~Danck. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0

(3):\penalty0 223, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/3/223.1.full.pdf+html|}
\fi.
\bibitem[Dan94b]{Danck:1994:BRb} V.~Danck. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(4):\penalty0 346, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/4/346.1.full.pdf+html|}
\fi.
\bibitem[Das91]{Daszczuk:1991:SSD} W.~B. Daszczuk. \newblock
A structured semantic design of distributed operating
systems. \newblock {\em The Computer Journal}, 34\penalty0
(6):\penalty0 482--492,   December 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/6/482.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/482.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/483.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/484.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/485.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/486.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/487.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/488.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/489.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/490.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/491.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/492.tif|}\fi.
\bibitem[Dav92]{Davis:1992:FRS} I.~J. Davis. \newblock
A fast radix sort. \newblock {\em The Computer Journal}
, 35\penalty0 (6):\penalty0 636--642,   December 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/636.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/Volume_36/Issue_06/Vol35_06.body.html#AbstractDavis|}
\fi.
\bibitem[Dav93]{Davies:1993:BR} C.~Davies. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(3):\penalty0 291--292, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/3/291.
full.pdf+html|}\fi.
\bibitem[Dav95]{Davie:1995:BR} A.~Davie. \newblock Book
review. \newblock {\em The Computer Journal}, 38\penalty0
(9):\penalty0 747, ???? 1995. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/9/747.1.full.pdf+html|}
\fi.
\bibitem[Daw99]{Dawid:1999:DPR} A.~P. Dawid. \newblock
Discussion of the papers by {Rissanen} and by {Wallace}
and {Dowe}. \newblock {\em The Computer Journal}, 42\
penalty0 (4):\penalty0 323--326, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/4/
323.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420323.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420323.pdf|}\fi.
\bibitem[Day94]{Day:1994:BR} A.~M. Day. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(7):\penalty0 649, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/7/649.1.full.pdf+html|}
\fi.
\bibitem[DB90]{Davies:1990:TLP} G.~L. Davies and A.~Burns.
\newblock The teaching language {Pascal-FC}. \newblock
{\em The Computer Journal}, 33\penalty0 (2):\penalty0
147--154, April   1990. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/2/147.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_

02/tiff/147.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/148.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/149.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/150.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/152.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/153.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/154.
tif|}\fi.
\bibitem[DBCRW92]{Dietrich:1992:PID} S.~W. Dietrich, M.
~Brown, E.~Cortes-Rello, and S.~Wunderlin. \newblock A
practitioner's introduction to database performance benchmarks
and   measurements. \newblock {\em The Computer Journal}
, 35\penalty0 (4):\penalty0 322--331,   August 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/4/322.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_04/Vol35_04.body.html#AbstractDietrich|}
\fi.
\bibitem[DC90a]{Dromey:1990:SCA} R.~G. Dromey and T.~
A. Chorvat. \newblock Structure clashes --- an alternative
to program inversion. \newblock {\em The Computer Journal}
, 33\penalty0 (2):\penalty0 126--132, April   1990. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/2/126.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/126.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/127.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/128.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/129.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/131.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/132.tif|}\fi.
\bibitem[DC90b]{Dueck:1990:MAG} Gerald D.~P. Dueck and
Gordon~V. Cormack. \newblock Modular attribute grammars.
\newblock {\em The Computer Journal}, 33\penalty0 (2)

:\penalty0 164--172, April   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/164.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/164.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/165.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/166.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/167.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/168.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/169.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/170.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/171.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/172.tif|}\fi. \newblock
See also: research report CS-88-19, University of Waterloo
(May   1988).
\bibitem[DD91]{Dvorak:1991:DMS} S.~Dvorak and B.~Durian.
\newblock Decomposition merging with symmetric exchanges.
\newblock {\em The Computer Journal}, 34\penalty0 (6)
:\penalty0 A029--A033,   December 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}.
\bibitem[DD98]{DePrisco:1998:DEH} Roberto {De Prisco}
and Alfredo {De Santis}. \newblock On the data expansion
of the {Huffman} compression algorithm. \newblock {\em
The Computer Journal}, 41\penalty0 (3):\penalty0 137-
-144, ????   1998. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/3/137.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_03/
Vol41_03.body.html#AbstractDePrisco|}\fi.
\bibitem[Dea98]{Dearden:1998:BR} A.~Dearden. \newblock
Book review. \newblock {\em The Computer Journal}, 41\
penalty0 (6):\penalty0 427--428, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/6/
427.full.pdf+html|}\fi.

\bibitem[Dei95]{Deighton:1995:ATT} Ken Deighton. \newblock
Average time to travel on two-headed, non-linear discs.
\newblock {\em The Computer Journal}, 38\penalty0 (10)
:\penalty0 811--817, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/10/811.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_10/Vol38_10.body.html#AbstractDeighton|}
\fi.
\bibitem[DF90]{Dehne:1990:ECG} F.~Dehne and L.~Ficocelli.
\newblock An efficient computational geometry method for
detecting dotted lines   in noisy images. \newblock {
\em The Computer Journal}, 33\penalty0 (5):\penalty0 424-
-428,   October 1990. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/5/424.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/424.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/425.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/426.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/427.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/428.tif|}\fi.
\bibitem[dGHM93]{deGuzman:1993:PDC} I.~P. de~Guzm{\'a}
n, P.~G. Harrison, and E.~Medina. \newblock Pipelines
for divide-and-conquer functions. \newblock {\em The Computer
Journal}, 36\penalty0 (3):\penalty0 254--268, June   1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/3/254.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/Vol36_03.body.html#AbstractGuzman|}
\fi.
\bibitem[DH96]{DePaul:1996:BR} J.~L. DePaul and T.~T.
Hewett. \newblock Book review. \newblock {\em The Computer
Journal}, 39\penalty0 (8):\penalty0 729--730, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/8/729.full.pdf+html|}\fi.

\bibitem[DH97]{Dearden:1997:SEM} Andrew~M. Dearden and
Michael~D. Harrison. \newblock A software engineering
model for case memory systems. \newblock {\em The Computer
Journal}, 40\penalty0 (4):\penalty0 167--182, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/4/167.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_03/Vol40_04.body.html#AbstractDearden;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
04/Vol40_04.body.html#AbstractDearden|}\fi.
\bibitem[dHZ97]{deFerreiraRezende:1997:TIN} F.~{de Ferreira
Rezende}, T.~H{\"a}rder, and J.~Zielinski. \newblock Transaction
identifiers in nested transactions: Implementation   schemes
and performance. \newblock {\em The Computer Journal}
, 40\penalty0 (5):\penalty0 245--258, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/5/
245.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_05/Vol40_05.body.html#AbstractRezende;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
05/Vol40_05.body.html#AbstractRezende|}\fi.
\bibitem[Dia92]{Diaper:1992:HCI} Dan Diaper. \newblock
Human-computer interaction: Theory and practice. \newblock
{\em The Computer Journal}, 35\penalty0 (4):\penalty0
331, August   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/4/331.full.pdf+html|}\fi.
\bibitem[Dic91]{Dick:1991:IKB} A.~J.~J. Dick. \newblock
An introduction to {Knuth--Bendix} completion. \newblock
{\em The Computer Journal}, 34\penalty0 (1):\penalty0
2--15, February   1991. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/1/2.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
01/tiff/10.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/12.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

34/Issue_01/tiff/13.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/3.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/4.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/
6.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/7.tif;   http://www3.oup.
\bibitem[Dil91]{Dillon:1991:AHC} Andrew Dillon. \newblock
Abstracts in human computer interaction. \newblock {\
em The Computer Journal}, 34\penalty0 (6):\penalty0 533,
December   1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/6/533.full.pdf+html|}\fi.
\bibitem[Dil94]{Diller:1994:OOS} Antoni Diller. \newblock
Object-oriented specification case studies. \newblock
{\em The Computer Journal}, 37\penalty0 (6):\penalty0
544a--545, ????   1994. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/544-a.full.pdf+html|}
\fi.
\bibitem[Dil95]{Diller:1995:BR} A.~Diller. \newblock Book
review. \newblock {\em The Computer Journal}, 38\penalty0
(9):\penalty0 750--751, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/9/750.
2.full.pdf+html|}\fi.
\bibitem[DJ95]{Dawson:1995:PFI} Scott Dawson and Farnam
Jahanian. \newblock Probing and fault injection of dependable
distributed protocols. \newblock {\em The Computer Journal}
, 38\penalty0 (4):\penalty0 286--300, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/4/
286.full.pdf+html;   http://www3.oup.co.uk/computer_journal/

Volume_38/Issue_04/Vol38_04.body.html#AbstractDawson|}
\fi.
\bibitem[DL92]{Dufrene:1992:EES} W.~R. {Dufrene, Jr.}
and F.~C. Lin. \newblock An efficient external sort algorithm
with no additional space. \newblock {\em The Computer
Journal}, 35\penalty0 (3):\penalty0 308--310, June   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/3/308.full.pdf+html|}\fi.
\bibitem[DL97]{Delis:1997:CAS} Alex Delis and Quang LeViet.
\newblock Contemporary access structures under mixed workloads.
\newblock {\em The Computer Journal}, 40\penalty0 (4)
:\penalty0 183--193, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/4/183.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_03/Vol40_04.body.html#AbstractDelis;
http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
04/Vol40_04.body.html#AbstractDelis|}\fi.
\bibitem[dLSA92]{deLemos:1992:TSC} R.~de~Lemos, A.~Saeed,
and T.~Anderson. \newblock A train set as a case study
for the requirements analysis of   safety-critical systems.
\newblock {\em The Computer Journal}, 35\penalty0 (1)
:\penalty0 30--40,   February 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/1/30.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_01/Vol35_01.body.html#AbstractLemos|}
\fi.
\bibitem[DMX91]{Dollimore:1991:DSD} J.~Dollimore, E.~
Miranda, and Wang Xu. \newblock The design of a system
for distributing shared objects. \newblock {\em The Computer
Journal}, 34\penalty0 (6):\penalty0 514--521,   December
1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/34/6/514.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/514.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/515.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_06/tiff/516.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/517.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/518.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/519.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/520.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/521.tif|}\fi.
\bibitem[Dob91]{Dobosiewicz:1991:ALP} W.~Dobosiewicz.
\newblock Algorithm 126: Linear probing sort. \newblock
{\em The Computer Journal}, 34\penalty0 (4):\penalty0
370--373,   August 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://www3.
oup.co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/
370.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/371.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_34/Issue_04/tiff/372.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_34/
Issue_04/tiff/373.tif|}\fi.
\bibitem[Dou91a]{Douglas:1991:BRa} A.~S. Douglas. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (4):\penalty0 319, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/319.
1.full.pdf+html|}\fi.
\bibitem[Dou91b]{Douglas:1991:BRb} A.~S. Douglas. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (4):\penalty0 319, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/319.
2.full.pdf+html|}\fi.
\bibitem[DPT94]{Diaz:1994:DDR} M.~D{\'\i}az, E.~Pimentel,
and J.~M. Troya. \newblock {DROL}: a distributed and real-
time object-oriented logic   environment. \newblock {
\em The Computer Journal}, 37\penalty0 (5):\penalty0 407-
-421, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/5/407.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/Volume_37/Issue_05/
Vol37_05.body.html#AbstractDiaz|}\fi.
\bibitem[Dra91]{Drapkin:1991:BR} T.~Drapkin. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (1):\penalty0 41, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/41.
full.pdf+html|}\fi.
\bibitem[DRT96]{Diaz:1996:DPL} M.~D{\'\i}az, B.~Rubio,
and J.~M. Troya. \newblock Distributed programming with
a logic channel based coordination   model. \newblock
{\em The Computer Journal}, 39\penalty0 (10):\penalty0
876--889, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/10/876.full.pdf+html;
http://www.oup.co.uk/computer_journal/Volume_39/Issue_
10/Vol39_10.body.html#AbstractDiaz;   http://www3.oup.
co.uk/computer_journal/Volume_39/Issue_10/Vol39_10.body.
html#AbstractDiaz|}\fi.
\bibitem[DS90]{Dey:1990:NPS} S.~Dey and P.~K. Srimani.
\newblock A new parallel sorting algorithm and its efficient
{VLSI}   implementation. \newblock {\em The Computer Journal}
, 33\penalty0 (3):\penalty0 241--246, June   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/3/
241.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/241.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/242.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/243.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/244.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/246.tif|}\fi.
\bibitem[DS91]{Dandamudi:1991:IPM} S.~P. Dandamudi and
P.~G. Sorenson. \newblock Improved partial-match search
algorithms for {BD} trees. \newblock {\em The Computer
Journal}, 34\penalty0 (5):\penalty0 415--422,   October
1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock

\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/34/5/415.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/415.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/416.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/417.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/418.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/419.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/420.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/421.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/422.tif|}\fi.
\bibitem[DS99]{Das:1999:ALP} Sajal~K. Das and Sanjoy~
K. Sen. \newblock Adaptive location prediction strategies
based on a hierarchical   network model in a cellular
mobile environment. \newblock {\em The Computer Journal}
, 42\penalty0 (6):\penalty0 473--486, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/6/
473.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_06/420473.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_06/
pdf/420473.pdf|}\fi.
\bibitem[DSB97]{Dahlgren:1997:RRM} Fredrick~H. Dahlgren,
Per Stenstr{\"o}m, and Marten Bj{\"o}rkman. \newblock
Reducing the read-miss penalty for flat {COMA} protocols.
\newblock {\em The Computer Journal}, 40\penalty0 (4)
:\penalty0 208--219, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/4/208.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_03/Vol40_04.body.html#AbstractDahlgren;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
04/Vol40_04.body.html#AbstractDahlgren|}\fi.
\bibitem[DT96]{Djordjevic:1996:CTS} Goran~Lj. Djordjevi{
\'c} and Milorad~B. To{\v{s}}i{\'c}. \newblock A compile-
time scheduling heuristic for multiprocessor architectures.
\newblock {\em The Computer Journal}, 39\penalty0 (8)
:\penalty0 663--674, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/8/663.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_08/390663.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_08/Vol39_
08.body.html#AbstractDjord|}\fi.
\bibitem[Du96]{Du:1996:IMG} C.~Du. \newblock An interpolation
method for grid-based terrain modelling. \newblock {\
em The Computer Journal}, 39\penalty0 (10):\penalty0 837-
-843, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/10/837.full.pdf+html;
http://www.oup.co.uk/computer_journal/Volume_39/Issue_
10/Vol39_10.body.html#AbstractDu;   http://www3.oup.co.
uk/computer_journal/Volume_39/Issue_10/Vol39_10.body.
html#AbstractDu|}\fi.
\bibitem[Dub93]{Dubhashi:1993:QEA} Devdatt~P. Dubhashi.
\newblock Quantifier elimination in $p$-adic fields. \
newblock {\em The Computer Journal}, 36\penalty0 (5):
\penalty0 419--426,   October 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/5/419.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_05/Vol36_05.body.html#AbstractDubhashi|}
\fi.
\bibitem[Dud93]{Dudman:1993:BR} K.~Dudman. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 603--604, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/603.
2.full.pdf+html|}\fi.
\bibitem[Dui96a]{Duin:1996:BRa} R.~P.~W. Duin. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (6):\penalty0 557--558, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/6/
557.1.full.pdf+html|}\fi.

\bibitem[Dui96b]{Duin:1996:BRb} R.~P.~W. Duin. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (6):\penalty0 557--558, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/6/
557.2.full.pdf+html|}\fi.
\bibitem[Dun94]{Dunne:1994:BR} P.~E. Dunne. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 146, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/146.
1.full.pdf+html|}\fi.
\bibitem[Dun95]{Dunne:1995:BR} P.~E. Dunne. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (9):\penalty0 747, ???? 1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/9/747.
2.full.pdf+html|}\fi.
\bibitem[Dun96]{Dunlop:1996:BR} M.~D. Dunlop. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (9):\penalty0 813, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/9/813.
1.full.pdf+html|}\fi.
\bibitem[Dun98]{Dunlop:1998:BR} M.~D. Dunlop. \newblock
Book review. \newblock {\em The Computer Journal}, 41\
penalty0 (2):\penalty0 134--135, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/2/
134.full.pdf+html|}\fi.
\bibitem[Dvo90]{Dvorak:1990:C} S.~Dvor{\'a}k. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (2):\penalty0 188, ???? 1990. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/2/
188.2.full.pdf+html|}\fi.

\bibitem[DW91]{Dick:1991:OST} A.~J.~J. Dick and P.~Watson.
\newblock Order-sorted term rewriting. \newblock {\em
The Computer Journal}, 34\penalty0 (1):\penalty0 16--
19,   February 1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/1/16.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
01/tiff/16.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/17.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/18.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/19.tif|}\fi.
\bibitem[Dwy91]{Dwyer:1991:C} B.~Dwyer. \newblock Correspondence.
\newblock {\em The Computer Journal}, 34\penalty0 (1)
:\penalty0 72, ???? 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/1/72.full.pdf+html|}\fi.
\bibitem[Dyc92]{Dyckhoff:1992:BR} Roy Dyckhoff. \newblock
Book review. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 539, October   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
539.full.pdf+html|}\fi.
\bibitem[Dye94]{Dyer:1994:SDE} David~C. Dyer. \newblock
Simulation in the design of electronic systems. \newblock
{\em The Computer Journal}, 37\penalty0 (6):\penalty0
542b--543, ????   1994. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/542-b.full.pdf+html|}
\fi.
\bibitem[Eke95]{Eker:1995:ACM} S.~M. Eker. \newblock Associative-
commutative matching via bipartite graph matching. \newblock
{\em The Computer Journal}, 38\penalty0 (5):\penalty0
381--399, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/5/381.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/Volume_38/Issue_05/
Vol38_05.body.html#AbstractEker|}\fi.
\bibitem[Ell91]{Elliot:1991:C} W.~S. Elliot. \newblock
Correspondence. \newblock {\em The Computer Journal},
34\penalty0 (3):\penalty0 206, ???? 1991. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/3/
206.full.pdf+html|}\fi.
\bibitem[EMCO92]{Ehrig:1992:IASa} H.~Ehrig, B.~Mahr, I.
~Classen, and F.~Orejas. \newblock Introduction to algebraic
specification. part 1: Formal methods for   software development.
\newblock {\em The Computer Journal}, 35\penalty0 (5)
:\penalty0 460--467,   October 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/5/460.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/Vol35_05.body.html#AbstractEhrig1|}
\fi.
\bibitem[EMO92]{Ehrig:1992:IASb} H.~Ehrig, B.~Mahr, and
F.~Orejas. \newblock Introduction to algebraic specification.
part 2: From classical view   to foundations of system
specifications. \newblock {\em The Computer Journal},
35\penalty0 (5):\penalty0 468--477,   October 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/5/468.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/Vol35_05.body.html#AbstractEhrig2|}
\fi.
\bibitem[Eng96]{England:1996:BR} D.~England. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (7):\penalty0 646, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/7/646.
full.pdf+html|}\fi.
\bibitem[Eng97]{England:1997:BR} D.~England. \newblock
Book review. \newblock {\em The Computer Journal}, 40\
penalty0 (4):\penalty0 231, ???? 1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/4/231.
1.full.pdf+html|}\fi.
\bibitem[Er91]{Er:1991:NOR} M.~C. Er. \newblock A note
on the optimality of a reve algorithm. \newblock {\em
The Computer Journal}, 34\penalty0 (6):\penalty0 513,
December   1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/6/513.full.pdf+html|}\fi.
\bibitem[Er92]{Er:1992:EGK} M.~C. Er. \newblock Efficient
generation of $k$-ary trees in natural order. \newblock
{\em The Computer Journal}, 35\penalty0 (3):\penalty0
306--308, June   1992. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/3/306.full.pdf+html|}\fi.
\bibitem[Er93]{Er:1993:C} M.~Er. \newblock Correspondence.
\newblock {\em The Computer Journal}, 36\penalty0 (2)
:\penalty0 207, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/2/207.full.pdf+html|}\fi.
\bibitem[ESK95]{Evered:1995:FOI} Mark Evered, Axel Schmolitzky,
and Michael K{\"o}lling. \newblock A flexible object invocation
language based on object-oriented   language definition.
\newblock {\em The Computer Journal}, 38\penalty0 (3)
:\penalty0 181--192, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/181.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_03/Vol38_03.body.html#AbstractEvered|}
\fi.
\bibitem[ET92]{Eastman:1992:OME} C.~M. Eastman and R.
~P. Trueblood. \newblock Occupancy models for the estimation
of block accesses. \newblock {\em The Computer Journal}
, 35\penalty0 (6):\penalty0 654--658,   December 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-

4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/654.full.pdf+html|}\fi.
\bibitem[Eva96]{Evans:1996:BR} H.~Evans. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(6):\penalty0 559, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/6/559.full.pdf+html|}\fi.
\bibitem[Fai98a]{Fairtlough:1998:BRa} M.~Fairtlough. \
newblock Book review. \newblock {\em The Computer Journal}
, 41\penalty0 (4):\penalty0 279--280, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/4/
279.1.full.pdf+html|}\fi.
\bibitem[Fai98b]{Fairtlough:1998:BRb} M.~Fairtlough. \
newblock Book review. \newblock {\em The Computer Journal}
, 41\penalty0 (4):\penalty0 279--280, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/4/
279.2.full.pdf+html|}\fi.
\bibitem[FB99]{Froese:1999:CMM} Kevin~W. Froese and Richard~
B. Bunt. \newblock Cache management for mobile file service.
\newblock {\em The Computer Journal}, 42\penalty0 (6)
:\penalty0 442--454, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/6/442.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_06/420442.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_06/
pdf/420442.pdf|}\fi.
\bibitem[FCHK96]{Fitzpatrick:1996:TAF} Stephen Fitzpatrick,
M.~Clint, T.~J. Harmer, and P.~Kilpatrick. \newblock The
tailoring of abstract functional specifications of numerical
 algorithms for sparse data structures through automated
program derivation   and transformation. \newblock {\
em The Computer Journal}, 39\penalty0 (2):\penalty0 145-
-168, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock

\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/2/145.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
02/390145.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_02/Vol39_02.body.html#AbstractFitzpatrick|}
\fi.
\bibitem[Fen90]{Fenwick:1990:PMK} P.~Fenwick. \newblock
Performance measurements on a 39 km extended {Ethernet}
. \newblock {\em The Computer Journal}, 33\penalty0 (
4):\penalty0 314--322,   August 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/314.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/314.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/315.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/316.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/318.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/319.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/320.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/321.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/322.tif|}\fi.
\bibitem[Fen91]{Fenwick:1991:NTS} P.~M. Fenwick. \newblock
A new technique for self-organising linear searches. \
newblock {\em The Computer Journal}, 34\penalty0 (5):
\penalty0 450--454,   October 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/450.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/450.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/451.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/452.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/453.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/454.tif|}\fi.
\bibitem[Fen94]{Fenton:1994:DVS} Norman Fenton. \newblock

Derivation and validation of software metri. \newblock
{\em The Computer Journal}, 37\penalty0 (6):\penalty0
553a--554, ????   1994. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/553-a.full.pdf+html|}
\fi.
\bibitem[Fen96]{Fenwick:1996:BWT} P.~M. Fenwick. \newblock
The {Burrows--Wheeler Transform} for block sorting text
compression:   principles and improvements. \newblock
{\em The Computer Journal}, 39\penalty0 (9):\penalty0
731--740, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/9/731.full.pdf+html;   http:
//www.oup.co.uk/computer_journal/Volume_39/Issue_09/Vol39_
09.body.html#AbstractFenwick;   http://www3.oup.co.uk/
computer_journal/Volume_39/Issue_09/Vol39_09.body.html#
AbstractFenwick|}\fi.
\bibitem[FFP99]{Fontana:1999:CMI} Francesca~Arcelli Fontana,
Ferrante Formato, and Remo Pareschi. \newblock Computational
models for information reuse. \newblock {\em The Computer
Journal}, 42\penalty0 (7):\penalty0 582--591, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/7/582.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_07/420582.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_07/pdf/420582.pdf|}\fi.
\bibitem[FG90]{Feldman:1990:OAS} R.~Feldman and M.~C.
Golumbic. \newblock Optimization algorithms for student
scheduling via constraint   satisfiability. \newblock
{\em The Computer Journal}, 33\penalty0 (4):\penalty0
356--364,   August 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/4/356.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/356.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/357.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/358.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

33/Issue_04/tiff/359.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/361.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/362.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/364.tif|}\fi.
\bibitem[FH97]{Frey:1997:ESS} B.~J. Frey and G.~E. Hinton.
\newblock Efficient stochastic source coding and an application
to a {Bayesian}   network source model. \newblock {\em
The Computer Journal}, 40\penalty0 (2--3):\penalty0 157-
-165,   ???? 1997. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/2_and_3/157.full.pdf+html;
 http://www.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_02.body.html#AbstractFrey;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_02/Vol40_03.body.
html#AbstractFrey|}\fi.
\bibitem[FH98]{Franti:1998:CMA} Pasi Fr{\"a}nti and Timo
Hatakka. \newblock Context model automata for text compression.
\newblock {\em The Computer Journal}, 41\penalty0 (7)
:\penalty0 474--485, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/7/474.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_07/Vol41_07.body.html#AbstractFranti|}
\fi.
\bibitem[Fie93]{Field:1993:BR} M.~Field. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 598--599, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/598.
2.full.pdf+html|}\fi.
\bibitem[Fin93]{Finkelstein:1993:ECC} Anthony Finkelstein.
\newblock {European} computing curricula: a guide and
comparative analysis. \newblock {\em The Computer Journal}
, 36\penalty0 (4):\penalty0 299--319,   August 1993. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL

{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/4/299.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_04/Vol36_04.body.html#AbstractFinklestein1|}
\fi.
\bibitem[FK90]{Fraenkel:1990:BHC} A.~S. Fraenkel and S.
~T. Klein. \newblock Bidirectional {Huffman} coding. \
newblock {\em The Computer Journal}, 33\penalty0 (4):
\penalty0 296--307,   August 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/296.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/296.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/297.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/298.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/299.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/301.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/302.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/304.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/305.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/306.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/307.
t.
\bibitem[FK93]{Fraenkel:1993:BDS} A.~S. Fraenkel and S.
~T. Klein. \newblock Bounding the depth of search trees.
\newblock {\em The Computer Journal}, 36\penalty0 (7)
:\penalty0 668--678, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/7/668.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_07/Vol36_07.body.html#AbstractFraenkel|}
\fi.
\bibitem[FKA{\etalchar{+}}93]{Finkelstein:1993:IES} Anthony
Finkelstein, Jeff Kramer, Samson Abramsky, Krysia Broda,
Sophia   Drossopoulou, and Susan Eisenbach. \newblock
An integrated engineering study scheme in computing. \

newblock {\em The Computer Journal}, 36\penalty0 (4):
\penalty0 320--334,   August 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/4/320.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_04/Vol36_04.body.html#AbstractFinklestein2|}
\fi.
\bibitem[FKKN97]{Franti:1997:GAL} P.~Fr{\"a}nti, J.~Kivij{
\"a}rvi, T.~Kaukoranta, and O.~Nevalainen. \newblock Genetic
algorithms for large-scale clustering problems. \newblock
{\em The Computer Journal}, 40\penalty0 (9):\penalty0
547--554, ????   1997. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/9/547.full.pdf+html|}\fi.
\bibitem[FKT90]{Foster:1990:CSC} I.~Foster, C.~Kesselman,
and S.~Taylor. \newblock Concurrency: {Simple} concepts
and powerful tools. \newblock {\em The Computer Journal}
, 33\penalty0 (6):\penalty0 501--507,   December 1990.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/6/501.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/501.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/502.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/503.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/504.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/505.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/506.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/507.tif|}\fi.
\bibitem[FLW99]{Fong:1999:DCB} Cedric C.~F. Fong, John
C.~S. Lui, and M.~H. Wong. \newblock Distributed caching
and broadcast in a wireless mobile computing   environment.
\newblock {\em The Computer Journal}, 42\penalty0 (6)
:\penalty0 455--472, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/6/455.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_06/420455.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_06/
pdf/420455.pdf|}\fi.
\bibitem[FN93]{Franti:1993:TSM} Pasi Fr{\"a}nti and Olli
Nevalainen. \newblock A two-stage modelling method for
compressing binary images by   arithmetic coding. \newblock
{\em The Computer Journal}, 36\penalty0 (7):\penalty0
615--622, ????   1993. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/7/615.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_07/
Vol36_07.body.html#AbstractFranti|}\fi.
\bibitem[FNK94]{Franti:1994:CDI} Pasi Fr{\"a}nti, Olli
Nevalainen, and Timo Kaukoranta. \newblock Compression
of digital images by block truncation coding: a survey.
\newblock {\em The Computer Journal}, 37\penalty0 (4)
:\penalty0 308--332, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/308.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_04/Vol37_04.body.html#AbstractFranti|}
\fi.
\bibitem[Fok96]{Fokkinga:1996:ETA} Maarten Fokkinga. \
newblock Expressions that talk about themselves. \newblock
{\em The Computer Journal}, 39\penalty0 (5):\penalty0
408--412, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/5/408.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
05/390408.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_05/Vol39_05.body.html#AbstractFokkinga|}
\fi.
\bibitem[For96]{Forrest:1996:BR} A.~R. Forrest. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (1):\penalty0 93, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/1/93.
1.full.pdf+html|}\fi.

\bibitem[For97]{Forster:1997:BR} T.~Forster. \newblock
Book review. \newblock {\em The Computer Journal}, 40\
penalty0 (1):\penalty0 61--63, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/1/
61.full.pdf+html|}\fi.
\bibitem[Fox94]{Fox:1994:BR} M.~Fox. \newblock Book review.
\newblock {\em The Computer Journal}, 37\penalty0 (9)
:\penalty0 818--821, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/9/818.
2.full.pdf+html|}\fi.
\bibitem[FR97]{Feijs:1997:LFP} L.~M.~G. Feijs and M.~
A. Reniers. \newblock Lazy functional programs in a concurrent
environment. \newblock {\em The Computer Journal}, 40\
penalty0 (9):\penalty0 572--584, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/9/
572.full.pdf+html|}\fi.
\bibitem[FR98]{Fraley:1998:HMC} Chris Fraley and Adrian~
E. Raftery. \newblock How many clusters? which clustering
method? answers via model-based   cluster analysis. \
newblock {\em The Computer Journal}, 41\penalty0 (8):
\penalty0 578--588, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/8/578.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_08/Vol41_08.body.html#AbstractFraley|}
\fi.
\bibitem[Fre92]{French:1992:CS} Carl French. \newblock
Computer science. \newblock {\em The Computer Journal}
, 35\penalty0 (4):\penalty0 404, August   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/4/
404.full.pdf+html|}\fi.
\bibitem[Fro92]{Frost:1992:CPE} R.~A. Frost. \newblock
Constructing programs as executable attribute grammars.

\newblock {\em The Computer Journal}, 35\penalty0 (4)
:\penalty0 376--389,   August 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/4/376.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_04/Vol35_04.body.html#AbstractFrost|}
\fi.
\bibitem[FS94]{Firby:1994:ASR} P.~A. Firby and D.~J. Stone.
\newblock Analysis of the structure and resolution of
a graphics device. \newblock {\em The Computer Journal}
, 37\penalty0 (5):\penalty0 367--376, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/5/
367.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_05/Vol37_05.body.html#AbstractFirby|}
\fi.
\bibitem[Fuh92]{Fuhr:1992:PMI} Norbert Fuhr. \newblock
Probabilistic models in information retrieval. \newblock
{\em The Computer Journal}, 35\penalty0 (3):\penalty0
243--255, June   1992. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/3/243.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractFuhr|}\fi.
\bibitem[Fur92]{Furber:1992:SNS} David Furber. \newblock
Short notes: a survey of the teaching of programming to
computing   undergraduates in {U.K. Universities and Polytechnics}
. \newblock {\em The Computer Journal}, 35\penalty0 (
5):\penalty0 530--533,   October 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/5/530.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/350530.sgm.abs.html; http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/tiff/530.tif;
 http://www3.oup.co.uk/computer_journal/Volume_35/Issue_
05/tiff/531.tif;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/tiff/532.tif;   http://www3.oup.co.
uk/computer_journal/Volume_35/Issue_05/tiff/533.tif|}
\fi.

\bibitem[FvLMS98]{Flammini:1998:CIR} Michele Flammini,
Jan van Leeuwen, and Alberto Marchetti-Spaccamela. \newblock
The complexity of interval routing on random graphs. \
newblock {\em The Computer Journal}, 41\penalty0 (1):
\penalty0 16--25, ????   1998. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/1/16.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_01/
Vol41_01.body.html#AbstractFlammini|}\fi.
\bibitem[FZ94]{Fokkink:1994:BPA} Wan Fokkink and Hans
Zantema. \newblock Basic process algebra with iteration:
completeness of its equational   axioms. \newblock {\
em The Computer Journal}, 37\penalty0 (4):\penalty0 259-
-267, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/4/259.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_04/
Vol37_04.body.html#AbstractFokkink|}\fi.
\bibitem[Gal92a]{Galton:1992:CLC} A.~Galton. \newblock
Classical logic: a crash course for beginners. \newblock
{\em The Computer Journal}, 35\penalty0 (5):\penalty0
424--430,   October 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/5/424.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractGalton1|}\fi.
\bibitem[Gal92b]{Galton:1992:LFM} A.~Galton. \newblock
Logic as a formal method. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 431--440,   October
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/431.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/Vol35_05.body.
html#AbstractGalton2|}\fi.
\bibitem[Ged92]{Gedeon:1992:RPI} T.~D. Gedeon. \newblock
The {Reve}'s puzzle: an iterative solution produced by
 transformation. \newblock {\em The Computer Journal}
, 35\penalty0 (2):\penalty0 186, April   1992. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/2/
186.full.pdf+html|}\fi.
\bibitem[Ged96]{Gedeon:1996:CTH} T.~D. Gedeon. \newblock
The {Cyclic Towers of Hanoi}: an iterative solution produced
by   transformation. \newblock {\em The Computer Journal}
, 39\penalty0 (4):\penalty0 353--356, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/4/
353.full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_04/390353.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_04/Vol39_
04.body.html#AbstractGedeon|}\fi.
\bibitem[Gee91]{Geesink:1991:CDA} L.~H. Geesink. \newblock
The coordination of distributed active messages in a dynamic
network   topology. \newblock {\em The Computer Journal}
, 34\penalty0 (6):\penalty0 542--550,   December 1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/6/542.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/542.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/543.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/544.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/545.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/546.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/547.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/548.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/549.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/550.
tif|}\fi.
\bibitem[Geo95]{Georgatos:1995:LIF} Konstantinos Georgatos.
\newblock Logic and information flow. \newblock {\em The
Computer Journal}, 38\penalty0 (3):\penalty0 259--260,
????   1995. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/3/259.full.pdf+html|}\fi.

\bibitem[GF91]{Ghafoor:1991:DCC} A.~Ghafoor and F.~Y.
Farhat. \newblock Dynamic concurrency control algorithms
for large distributed database   systems. \newblock {
\em The Computer Journal}, 34\penalty0 (6):\penalty0 A095-
-A104,   December 1991. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}.
\bibitem[GH94]{Ganley:1994:ORP} Joseph~L. Ganley and Lenwood~
S. Heath. \newblock Optimal and random partitions of random
graphs. \newblock {\em The Computer Journal}, 37\penalty0
(7):\penalty0 641--643, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/7/641.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_07/Vol37_07.body.html#AbstractGanley|}
\fi.
\bibitem[GH95]{Gilmore:1995:PAT} Stephen Gilmore and Jane
Hillston. \newblock Process algebras and their application
to performance modelling:   {Proceedings of the Third
Workshop on Process Algebra and Performance   Modelling}
. \newblock {\em The Computer Journal}, 38\penalty0 (
7):\penalty0 489--491, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/7/489.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_07/Vol38_07.body.html#AbstractGilmore|}
\fi. \newblock Proceedings of the Third Workshop on Process
Algebra and Performance   Modelling, Edinburgh, Scotland.
\bibitem[GHR{\etalchar{+}}93]{Geuder:1993:GPP} U.~Geuder,
M.~H{\"a}rdtner, A.~Reuter, B.~W{\"o}rner, and R.~Zink.
\newblock {GRIDS} --- a parallel programming system for
grid-based algorithms. \newblock {\em The Computer Journal}
, 36\penalty0 (8):\penalty0 702--711, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/8/
702.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractGeuder|}
\fi.

\bibitem[GIB90]{Gulati:1990:PCM} S.~Gulati, S.~S. Iyengar,
and J.~Barhen. \newblock The pebble-crunching model for
fault-tolerant load balancing in   hypercube ensembles.
\newblock {\em The Computer Journal}, 33\penalty0 (3)
:\penalty0 204--214, June   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/204.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/204.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/205.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/206.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/208.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/209.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/210.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/211.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/213.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/214.tif|}\fi.
\bibitem[Gib94a]{Gibbins:1994:BRa} N.~M. Gibbins. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 231--232, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/3/
231.full.pdf+html|}\fi.
\bibitem[Gib94b]{Gibbins:1994:BRb} N.~M. Gibbins. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 354--355, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
354.full.pdf+html|}\fi.
\bibitem[Gib94c]{Gibson:1994:BRa} M.~B.~M. Gibson. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 227, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/37/3/227.
1.full.pdf+html|}\fi.
\bibitem[Gib94d]{Gibson:1994:BRb} M.~B.~M. Gibson. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 350, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/350.
1.full.pdf+html|}\fi.
\bibitem[Gil91]{Giles:1991:EDS} D.~Giles. \newblock Editorial
--- data structures. \newblock {\em The Computer Journal}
, 34\penalty0 (5):\penalty0 385, ???? 1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/5/
385.full.pdf+html|}\fi.
\bibitem[Gil92]{Gilles:1992:ACS} A.~Gilles. \newblock
The art of computer systems performance analysis (techniques
for   experimental design, measurement, simulation, and
modeling). \newblock {\em The Computer Journal}, 35\penalty0
(3):\penalty0 207, June 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/3/207.full.pdf+html|}\fi.
\bibitem[Gil96]{Gilmore:1996:BR} S.~Gilmore. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (7):\penalty0 643, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/7/643.
1.full.pdf+html|}\fi.
\bibitem[GK96]{Geschiere:1996:PDP} J.~P. Geschiere and
W.~H. F.~J. K{\"o}rver. \newblock A practical development
process for parallel large-scale applications   and its
underlying formal framework. \newblock {\em The Computer
Journal}, 39\penalty0 (5):\penalty0 386--407, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/5/386.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_05/390386.sgm.abs.html;   http:

//www3.oup.co.uk/computer_journal/Volume_39/Issue_05/
Vol39_05.body.html#AbstractGeschiere|}\fi.
\bibitem[GL94]{Gunther:1994:OOT} Oliver G{\"u}nther and
Johannes Lamberts. \newblock Object-oriented techniques
for the management of geographic and   environmental data.
\newblock {\em The Computer Journal}, 37\penalty0 (1)
:\penalty0 16--25, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/1/16.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractGunther|}
\fi.
\bibitem[GLR93]{Gonzalez-Lopez:1993:PTM} M.~J. Gonz{\
'a}lez-L{\'o}pez and T.~Recio. \newblock Path tracking
in motion planning. \newblock {\em The Computer Journal}
, 36\penalty0 (5):\penalty0 515--524,   October 1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/5/515.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/Vol36_05.body.html#AbstractLopez|}
\fi.
\bibitem[GM95a]{Ghosh:1995:TDC} P.~K. Ghosh and S.~P.
Mudur. \newblock Three-dimensional computer vision: a
geometric viewpoint. \newblock {\em The Computer Journal}
, 38\penalty0 (1):\penalty0 85--86, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
85.full.pdf+html|}\fi.
\bibitem[GM95b]{Gunter:1995:SMM} Elsa Gunter and Savi
Maharaj. \newblock Studying the {ML} module system in
{HOL}. \newblock {\em The Computer Journal}, 38\penalty0
(2):\penalty0 142--151, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/2/142.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_02/Vol38_02.body.html#AbstractGunter|}
\fi.

\bibitem[GN94]{Giddy:1994:ACT} J.~P. Giddy and R.~Safavi
Naini. \newblock Automated cryptanalysis of transposition
ciphers. \newblock {\em The Computer Journal}, 37\penalty0
(5):\penalty0 429--436, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/5/429.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_05/Vol37_05.body.html#AbstractGiddy|}
\fi.
\bibitem[God96]{Godwin:1996:BSD} Anthony~N. Godwin. \
newblock Building a semantic description for a data modelling
language. \newblock {\em The Computer Journal}, 39\penalty0
(4):\penalty0 275--290, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/4/275.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_04/390275.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_04/Vol39_
04.body.html#AbstractGodwin|}\fi.
\bibitem[Gol93a]{Goldberg:1993:FSA} Robert~R. Goldberg.
\newblock Finite state automata from regular expression
trees. \newblock {\em The Computer Journal}, 36\penalty0
(7):\penalty0 623--630, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/7/623.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_07/Vol36_07.body.html#AbstractGoldberg|}
\fi.
\bibitem[Gol93b]{Gollmann:1993:BR} D.~Gollmann. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 296, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/3/296.
full.pdf+html|}\fi.
\bibitem[Gol94]{Goldson:1994:SCN} Doug Goldson. \newblock
A symbolic calculator for non-strict functional programs.
\newblock {\em The Computer Journal}, 37\penalty0 (3)
:\penalty0 177--187, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/177.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_03/Vol37_03.body.html#AbstractGoldson|}
\fi.
\bibitem[Goo90]{Gooch:1990:C} B.~G. Gooch. \newblock Correspondence.
\newblock {\em The Computer Journal}, 33\penalty0 (1)
:\penalty0 92, ???? 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/1/92.2.full.pdf+html|}\
fi.
\bibitem[GP90]{Govindarajan:1990:SNL} R.~Govindarajan
and L.~M. Patnaik. \newblock Short notes: Lenient execution
and concurrent execution of re-entrant   routines: Efficient
implementation in data flow systems. \newblock {\em The
Computer Journal}, 33\penalty0 (2):\penalty0 185--187,
April   1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/2/185.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/185.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/186.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/187.
tif|}\fi.
\bibitem[GP96a]{Gu:1996:EAN} Qian-Ping Gu and Shietung
Peng. \newblock An efficient algorithm for node-to-node
routing in hypercubes with   faulty clusters. \newblock
{\em The Computer Journal}, 39\penalty0 (1):\penalty0
14--19, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/1/14.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
01/390014.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_01/Vol39_01.body.html#AbstractGu|}
\fi.
\bibitem[GP96b]{Gu:1996:OAN} Qian-Ping Gu and Shietung
Peng. \newblock Optimal algorithms for node-to-node fault
tolerant routing in   hypercubes. \newblock {\em The Computer
Journal}, 39\penalty0 (7):\penalty0 626--629, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-

4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/7/626.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_07/390626.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_07/
Vol39_07.body.html#AbstractGu|}\fi.
\bibitem[GRB93]{Goldson:1993:RSP} Doug Goldson, Steve
Reeves, and Richard Bornat. \newblock A review of several
programs for the teaching of logic. \newblock {\em The
Computer Journal}, 36\penalty0 (4):\penalty0 373--386,
 August 1993. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/4/373.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_04/
Vol36_04.body.html#AbstractGoldson|}\fi.
\bibitem[Gre90]{Gregory:1990:ECL} S.~Gregory. \newblock
Editorial --- concurrent logic programming. \newblock
{\em The Computer Journal}, 33\penalty0 (6):\penalty0
481, ???? 1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/6/481.full.pdf+html|}\fi.
\bibitem[Gre93]{Gregory:1993:BR} S.~Gregory. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 294, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/3/294.
1.full.pdf+html|}\fi.
\bibitem[Gre94a]{Green:1994:BRa} P.~Green. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(3):\penalty0 229--230, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/229.
2.full.pdf+html|}\fi.
\bibitem[Gre94b]{Green:1994:BRb} P.~Green. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(4):\penalty0 352--353, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/352.
2.full.pdf+html|}\fi.
\bibitem[Gri90]{Griswold:1990:SSI} R.~E. Griswold. \newblock
String scanning in the {Icon} programming language. \
newblock {\em The Computer Journal}, 33\penalty0 (2):
\penalty0 98--107, April   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/98.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/100.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/101.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/105.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/106.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/107.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/
tiff/99.tif|}\fi.
\bibitem[Gri92]{Gries:1992:CFM} David Gries. \newblock
Correspondence: On {``Finding a Majority When Sorting
is Not   Available''}. \newblock {\em The Computer Journal}
, 35\penalty0 (3):\penalty0 298, June 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www3.oup.co.uk/computer_journal/hdb/Volume_
35/Issue_03/tiff/298.tif|}\fi. \newblock See \cite{Campbell:
1991:FMW}. Points out that the supposedly new   algorithm
was previously published by Boyer and Moore (1982), Misra
and Gries   (1982), and Gries (1990).
\bibitem[Gru96]{Grundy:1996:THR} Jim Grundy. \newblock
Transformational hierarchical reasoning. \newblock {\
em The Computer Journal}, 39\penalty0 (4):\penalty0 291-
-302, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/4/291.full.pdf+html;   http:

//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
04/390291.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_04/Vol39_04.body.html#AbstractGrundy|}
\fi.
\bibitem[Gup94]{Gupta:1994:BR} A.~Gupta. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(8):\penalty0 731, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/8/731.1.full.pdf+html|}
\fi.
\bibitem[Gup95]{Gupta:1995:NEC} P.~P. Gupta. \newblock
A new era in computation. \newblock {\em The Computer
Journal}, 38\penalty0 (1):\penalty0 87, ???? 1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
87.full.pdf+html|}\fi.
\bibitem[GV91]{Goel:1991:OFC} V.~Goel and Y.~V. Venkatesh.
\newblock On an optimal and faster construction of linear
quadtrees from   raster-scanned images. \newblock {\em
The Computer Journal}, 34\penalty0 (6):\penalty0 A073-
-A083,   December 1991. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}.
\bibitem[GV99a]{Gammerman:1999:KCS} Alexander Gammerman
and Vladimir Vovk. \newblock {Kolmogorov} complexity:
Sources, theory and applications. \newblock {\em The Computer
Journal}, 42\penalty0 (4):\penalty0 252--255, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/4/252.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_04/420252.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_04/pdf/420252.pdf|}\fi.
\bibitem[GV99b]{Gammerman:1999:SIE} Alexander Gammerman
and Vladimir Vovk. \newblock Special issue editorial:
{Kolmogorov} complexity. \newblock {\em The Computer Journal}
, 42\penalty0 (4):\penalty0 251, ???? 1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/42/4/
251.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420251.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420251.pdf|}\fi.
\bibitem[GWWB98]{Gillet:1998:SDM} Valerie~J. Gillet, David~
J. Wild, Peter Willett, and John Bradshaw. \newblock Similarity
and dissimilarity methods for processing chemical   structure
databases. \newblock {\em The Computer Journal}, 41\penalty0
(8):\penalty0 547--558, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/8/547.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_08/Vol41_08.body.html#AbstractGillet|}
\fi.
\bibitem[HAAZ99]{Hiasat:1999:SCV} Ahmad~A. Hiasat and
Hoda Abdel-Aty-Zohdy. \newblock Semi-custom {VLSI} design
and implementation of a new efficient {RNS}   division
algorithm. \newblock {\em The Computer Journal}, 42\penalty0
(3):\penalty0 232--240, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/3/232.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_03/420232.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_03/
pdf/420232.pdf|}\fi.
\bibitem[Had90]{Haddon:1990:CSL} B.~K. Haddon. \newblock
{Cycle-Sort}: a linear sorting method. \newblock {\em
The Computer Journal}, 33\penalty0 (4):\penalty0 365-
-367,   August 1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/4/365.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/365.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/366.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/367.
tif|}\fi.
\bibitem[Hal93a]{Hall:1993:BRa} P.~Hall. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0

(6):\penalty0 595, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/595.1.full.pdf+html|}
\fi.
\bibitem[Hal93b]{Hall:1993:BRb} P.~Hall. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 596--597, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/596.
2.full.pdf+html|}\fi.
\bibitem[Hal93c]{Hall:1993:BR} P.~Hall. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 597, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/597.1.full.pdf+html|}
\fi.
\bibitem[Ham91a]{Hammersley:1991:AET} A.~P. Hammersley.
\newblock Algorithm 125: Efficient two-dimensional fast
{Fourier} transform   subroutines for real-valued or {
Hermitian} data. \newblock {\em The Computer Journal}
, 34\penalty0 (4):\penalty0 362--370,   August 1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/362.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/366.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/367.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/368.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/369.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/370.tif|}\fi.
\bibitem[Ham91b]{Hammersley:1991:ISD} P.~Hammersley. \
newblock Information systems design methodologies ---
book reviews. \newblock {\em The Computer Journal}, 34\
penalty0 (2):\penalty0 182--185, April   1991. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/2/
182.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/182.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/183.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/184.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/185.tif|}\fi.
\bibitem[Ham93a]{Hammersley:1993:E} P.~Hammersley. \newblock
Editorial. \newblock {\em The Computer Journal}, 36\penalty0
(4):\penalty0 297--298, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/4/297.
full.pdf+html|}\fi.
\bibitem[Ham93b]{Hammersley:1993:SIE} Peter Hammersley.
\newblock Special issue editorial: Computing curricula.
\newblock {\em The Computer Journal}, 36\penalty0 (4)
:\penalty0 297--??, August   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://www3.oup.co.uk/computer_journal/Volume_36/
Issue_04/Vol36_04.body.html#AbstractHammersley|}\fi.
\bibitem[Ham94a]{Hamilton:1994:BRa} R.~Hamilton. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 148, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/148.
1.full.pdf+html|}\fi.
\bibitem[Ham94b]{Hamilton:1994:BRb} R.~Hamilton. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (7):\penalty0 649--650, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
649.2.full.pdf+html|}\fi.
\bibitem[Ham96a]{Hamilton:1996:BR} A.~G. Hamilton. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 727--728, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)

, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/8/
727.full.pdf+html|}\fi.
\bibitem[Ham96b]{Hammond:1996:BRaa} K.~Hammond. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (9):\penalty0 813--815, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
813.2.full.pdf+html|}\fi.
\bibitem[Ham96c]{Hammond:1996:BRbb} K.~Hammond. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (9):\penalty0 813--815, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
813.3.full.pdf+html|}\fi.
\bibitem[Ham96d]{Hammond:1996:BR} K.~Hammond. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (9):\penalty0 813--815, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
813.4.full.pdf+html|}\fi.
\bibitem[HAM98]{Hubert:1998:RSP} Lawrence Hubert, Phipps
Arabie, and Jacqueline Meulman. \newblock The representation
of symmetric proximity data: Dimensions and   classifications.
\newblock {\em The Computer Journal}, 41\penalty0 (8)
:\penalty0 566--577, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/8/566.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_08/Vol41_08.body.html#AbstractHubert|}
\fi.
\bibitem[Han91]{Hankins:1991:BR} C.~Hankins. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (3):\penalty0 194, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/194.
2.full.pdf+html|}\fi.

\bibitem[Han94a]{Hanna:1994:BR} F.~K. Hanna. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 720--721, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
720.2.full.pdf+html|}\fi.
\bibitem[Han94b]{Hansen:1994:ICS} Per~Brinch Hansen. \
newblock Interference control in {SuperPascal} --- a block-
structured parallel   language. \newblock {\em The Computer
Journal}, 37\penalty0 (5):\penalty0 399--406, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/5/399.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/Vol37_05.body.html#AbstractHansen|}
\fi.
\bibitem[Han95]{Hankin:1995:BR} C.~L. Hankin. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (10):\penalty0 831--832, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/10/
831.2.full.pdf+html|}\fi.
\bibitem[Har92]{Harrison:1992:HOA} P.~G. Harrison. \newblock
A higher-order approach to parallel algorithms. \newblock
{\em The Computer Journal}, 35\penalty0 (6):\penalty0
555--566,   December 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/6/555.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractHarrison|}\fi.
\bibitem[Har95]{Harrison:1995:BDD} John Harrison. \newblock
Binary decision diagrams as a {HOL} derived rule. \newblock
{\em The Computer Journal}, 38\penalty0 (2):\penalty0
162--170, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/2/162.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractHarrison|}\fi.

\bibitem[Har96]{Harvey:1996:BR} I.~Harvey. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(2):\penalty0 169, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/2/169.1.full.pdf+html|}
\fi.
\bibitem[Hau94]{Haughton:1994:BR} H.~P. Haughton. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (7):\penalty0 646--647, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
646.2.full.pdf+html|}\fi.
\bibitem[Hav95]{Haverkort:1995:SPM} Boudewijn~R. Haverkort.
\newblock In search of probability mass: probabilistic
evaluation of high-level   specified {Markov} models.
\newblock {\em The Computer Journal}, 38\penalty0 (7)
:\penalty0 521--529, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/7/521.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_07/Vol38_07.body.html#AbstractHaverkort|}
\fi.
\bibitem[Hay94]{Hayes:1994:BR} G.~Hayes. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(8):\penalty0 722--723, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/722.
2.full.pdf+html|}\fi.
\bibitem[HCH96a]{Hsu:1996:GNS} Wen-Jing Hsu, Moon~Jung
Chung, and Zejun Hu. \newblock {Gaussian} networks for
scalable distributed systems. \newblock {\em The Computer
Journal}, 39\penalty0 (5):\penalty0 417--426, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/5/417.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_05/390417.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_05/
Vol39_05.body.html#AbstractHsu|}\fi.

\bibitem[HCH96b]{Huang:1996:RPV} Chung-Ming Huang, Ye-
In Chang, and Duen-Tay Huang. \newblock Reverse protocol
verification: concept, algorithm and application. \newblock
{\em The Computer Journal}, 39\penalty0 (6):\penalty0
511--524, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/6/511.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
06/390511.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_06/Vol39_06.body.html#AbstractHuang1|}
\fi.
\bibitem[Hel93]{Hellwagner:1993:DCS} Hermann Hellwagner.
\newblock Design considerations for scalable parallel
file systems. \newblock {\em The Computer Journal}, 36\
penalty0 (8):\penalty0 741--755, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/8/
741.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractHellwagner|}
\fi.
\bibitem[Hen91]{Henderson:1991:C} J.~Henderson. \newblock
Correspondence. \newblock {\em The Computer Journal},
34\penalty0 (4):\penalty0 310, ???? 1991. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/4/
310.full.pdf+html|}\fi.
\bibitem[Her91]{Herbert:1991:EDS} Andrew Herbert. \newblock
Editorial: Distributed systems. \newblock {\em The Computer
Journal}, 34\penalty0 (6):\penalty0 481, December   1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/6/481.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/481.tif|}\fi.
\bibitem[HGL93]{Hui:1993:MIS} S.~C. Hui, A.~Goh, and L.
~H. Lau. \newblock A multimedia information system for
{IC} failure analysis. \newblock {\em The Computer Journal}
, 36\penalty0 (1):\penalty0 97--105,   February 1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL

{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/1/97.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/Vol36_01.body.html#AbstractHui|}
\fi.
\bibitem[HH94a]{Hamdi:1994:ECI} M.~Hamdi and R.~W. Hall.
\newblock An efficient class of interconnection networks
for parallel   computations. \newblock {\em The Computer
Journal}, 37\penalty0 (3):\penalty0 206--218, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/3/206.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_03/Vol37_03.body.html#AbstractHamdi|}
\fi.
\bibitem[HH94b]{Huang:1994:IPV} Chung-Ming Huang and Jenq-
Muh Hsu. \newblock An incremental protocol verification
method. \newblock {\em The Computer Journal}, 37\penalty0
(8):\penalty0 698--710, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/698.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_08/Vol37_08.body.html#AbstractHuang|}
\fi.
\bibitem[HH95]{Harrison:1995:EQR} Peter Harrison and Jane
Hillston. \newblock Exploiting quasi-reversible structures
in {Markovian} process algebra   models. \newblock {\
em The Computer Journal}, 38\penalty0 (7):\penalty0 510-
-520, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/7/510.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractHarrison|}\fi.
\bibitem[Hie96]{Hierons:1996:ETS} R.~M. Hierons. \newblock
Extending test sequence overlap by invertibility. \newblock
{\em The Computer Journal}, 39\penalty0 (4):\penalty0
325--330, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/4/325.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
04/390325.sgm.abs.html;   http://www3.oup.co.uk/computer_

journal/Volume_39/Issue_04/Vol39_04.body.html#AbstractHierons|}
\fi.
\bibitem[Hie97]{Hierons:1997:TFS} Robert~M. Hierons. \
newblock Testing from a finite-state machine: Extending
invertibility to   sequences. \newblock {\em The Computer
Journal}, 40\penalty0 (4):\penalty0 220--230, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/4/220.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_03/Vol40_04.body.html#AbstractHierons;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
04/Vol40_04.body.html#AbstractHierons|}\fi.
\bibitem[Hie98]{Hierons:1998:ATD} R.~M. Hierons. \newblock
Adaptive testing of a deterministic implementation against
a   nondeterministic finite state machine. \newblock {
\em The Computer Journal}, 41\penalty0 (5):\penalty0 349-
-355, ????   1998. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/5/349.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_05/
Vol41_05.body.html#AbstractHierons|}\fi.
\bibitem[Hil91]{Hill:1991:BRV} I.~D. Hill. \newblock Book
review: {Valerie Illingworth (editor), {\booktitle{Dictionary
of   Computing (third edition)}}}. \newblock {\em The
Computer Journal}, 34\penalty0 (6):\penalty0 521, December
 1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/34/6/521.full.pdf+html|}\fi.
\bibitem[Hil92]{Hill:1992:BRB} I.~D. Hill. \newblock Book
review: {B. Samways and T. Byrne-Jones, \booktitle{Collins
Gem   Basic Facts: Computers, 3rd edn}}. \newblock {\
em The Computer Journal}, 35\penalty0 (2):\penalty0 128,
April   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/2/128.full.pdf+html|}\fi.
\bibitem[HJ99]{Huang:1999:ITB} Chung-Ming Huang and Ming-
Yuhe Jang. \newblock Interactive temporal behaviours and

modelling for multimedia   presentations in the {WWW}
environment. \newblock {\em The Computer Journal}, 42\
penalty0 (2):\penalty0 112--128, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/2/
112.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_02/420112.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_02/
pdf/420112.pdf|}\fi.
\bibitem[HK92]{Harrison:1992:MTD} P.~G. Harrison and H.
~Khoshnevisan. \newblock The mechanical transformation
of data types. \newblock {\em The Computer Journal}, 35\
penalty0 (2):\penalty0 138--147, April   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/2/
138.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_02/Vol35_02.body.html#AbstractHarrison|}
\fi.
\bibitem[HKK91]{Hermann:1991:ITR} M.~Hermann, C.~Kirchner,
and H.~Kirchner. \newblock Implementations of term rewriting
systems. \newblock {\em The Computer Journal}, 34\penalty0
(1):\penalty0 20--33,   February 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/20.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/21.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/24.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/28.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/29.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/30.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/31.
tif;   http://ww.
\bibitem[HKPT99]{Huhtala:1999:TEA} Yk{\"a} Huhtala, Juha
K{\"a}rkk{\"a}inen, Pasi Porkka, and Hannu Toivonen. \
newblock {TANE}: An efficient algorithm for discovering
functional and   approximate dependencies. \newblock {
\em The Computer Journal}, 42\penalty0 (2):\penalty0 100-
-111, ????   1999. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/2/100.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
02/420100.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_02/pdf/420100.pdf|}\fi.
\bibitem[HKT95]{Halang:1995:EGP} Wolfgang~A. Halang, Bernd
Kr{\"a}mer, and Leszek Trybus. \newblock Exploiting a
graphical programming paradigm to facilitate rigorous
 verification of embedded software. \newblock {\em The
Computer Journal}, 38\penalty0 (4):\penalty0 301--309,
????   1995. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/4/301.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_38/Issue_04/Vol38_04.body.
html#AbstractHalang|}\fi.
\bibitem[HL92]{Huang:1992:FSM} B.-C. Huang and M.~A. Langston.
\newblock Fast stable merging and sorting in constant
extra space. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 643--650,   December 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
643.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_06/Vol35_06.body.html#AbstractHuang|}
\fi.
\bibitem[HLK94]{Hao:1994:ASL} Ying Hao, Jyh-Charn~S. Liu,
and Junguk~L. Kim. \newblock An all-sharing load-balancing
scheme on the {CSMA\slash CD} network   and its analysis.
\newblock {\em The Computer Journal}, 37\penalty0 (9)
:\penalty0 779--794, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/9/779.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_09/Vol37_09.body.html#AbstractHao|}\fi.
\bibitem[HM93]{Hains:1993:PFP} G.~Hains and L.~M.~R. Mullin.
\newblock Parallel functional programming with arrays.
\newblock {\em The Computer Journal}, 36\penalty0 (3)
:\penalty0 238--245, June   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/3/238.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_03/Vol36_03.body.html#AbstractHains|}
\fi.
\bibitem[HM95a]{Hillston:1995:STS} Jane Hillston and Vassilis
Mertsiotakis. \newblock A simple time scale decomposition
technique for stochastic process   algebras. \newblock
{\em The Computer Journal}, 38\penalty0 (7):\penalty0
566--577, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/7/566.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractHillston|}\fi.
\bibitem[HM95b]{Homeier:1995:MVV} Peter~V. Homeier and
David~F. Martin. \newblock A mechanically verified verification
condition generator. \newblock {\em The Computer Journal}
, 38\penalty0 (2):\penalty0 131--141, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/2/
131.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_02/Vol38_02.body.html#AbstractHomeier|}
\fi.
\bibitem[HM96]{Hui:1996:RCA} Lucas Chi~Kwong Hui and Charles~
U. Martel. \newblock Randomized competitive algorithms
for successful and unsuccessful   search. \newblock {
\em The Computer Journal}, 39\penalty0 (5):\penalty0 427-
-438, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/5/427.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
05/390427.sgm.abs.html;   http://www3.oup.co.uk/computer_

journal/Volume_39/Issue_05/Vol39_05.body.html#AbstractHui|}
\fi.
\bibitem[HM98]{Hoggar:1998:FCJ} Stuart~G. Hoggar and Lindsey
Menzies. \newblock Fractal compression and the jigsaw
property {I}. \newblock {\em The Computer Journal}, 41\
penalty0 (5):\penalty0 319--336, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/5/
319.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_05/Vol41_05.body.html#AbstractHoggar|}
\fi.
\bibitem[HO93]{Hull:1993:FRB} M.~Elizabeth~C. Hull and
P.~G. O'Donoghue. \newblock Family relationships between
requirements and design specification   methods. \newblock
{\em The Computer Journal}, 36\penalty0 (2):\penalty0
153--167, April   1993. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/2/153.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractHull|}\fi.
\bibitem[HOH91]{Hull:1991:DMR} M.~Elizabeth~C. Hull, P.
~G. O'Donoghue, and B.~J. Hagan. \newblock Development
methods for real-time systems. \newblock {\em The Computer
Journal}, 34\penalty0 (2):\penalty0 164--172, April
1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/34/2/164.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/164.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/165.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/166.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/167.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/168.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/169.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/170.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/171.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/172.tif|}\fi.

\bibitem[Hol94]{Holcombe:1994:BR} M.~Holcombe. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (5):\penalty0 469, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/5/469.
1.full.pdf+html|}\fi.
\bibitem[Hol95]{Holton:1995:PSI} D.~R.~W. Holton. \newblock
A {PEPA} specification of an industrial production cell.
\newblock {\em The Computer Journal}, 38\penalty0 (7)
:\penalty0 542--551, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/7/542.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_07/Vol38_07.body.html#AbstractHolton|}
\fi.
\bibitem[Hon93a]{Hong:1993:QEF} Hoon Hong. \newblock Quantifier
elimination for formulas constrained by quadratic   equations
via slope resultants. \newblock {\em The Computer Journal}
, 36\penalty0 (5):\penalty0 439--449,   October 1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/5/439.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/Vol36_05.body.html#AbstractHong|}
\fi.
\bibitem[Hon93b]{Hong:1993:SIE} Hoon Hong. \newblock Special
issue editorial: Computational quantifier elimination.
\newblock {\em The Computer Journal}, 36\penalty0 (5)
:\penalty0 399, October   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/5/399.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_05/Vol36_05.body.html#AbstractHong|}\
fi.
\bibitem[Hor93]{Horng:1993:CTA} S.~J. Horng. \newblock
Constant time algorithm for template matching on a reconfigurable
 array of processors. \newblock {\em The Computer Journal}
, 36\penalty0 (3):\penalty0 246--253, June   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/36/3/
246.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_03/Vol36_03.body.html#AbstractHorng|}
\fi.
\bibitem[How97]{Howard:1997:TIC} P.~G. Howard. \newblock
Text image compression using soft pattern matching. \
newblock {\em The Computer Journal}, 40\penalty0 (2--
3):\penalty0 146--156,   ???? 1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/2_and_
3/146.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_02/Vol40_02.body.html#AbstractHoward;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_03.body.html#AbstractHoward|}\fi.
\bibitem[HP91]{Hussain:1991:C} F.~Hussain and M.~L.~V.
Pitteway. \newblock Correspondence. \newblock {\em The
Computer Journal}, 34\penalty0 (5):\penalty0 405, ???
? 1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/34/5/405.full.pdf+html|}\fi.
\bibitem[HR96]{Huang:1996:UTP} Yen-Min Huang and Chinya~
V. Ravishankar. \newblock {URPC}: a toolkit for prototyping
remote procedure calls. \newblock {\em The Computer Journal}
, 39\penalty0 (6):\penalty0 525--540, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/6/
525.full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_06/390525.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_06/Vol39_
06.body.html#AbstractHuang2|}\fi.
\bibitem[HR97]{Harris:1997:OCR} Evan~P. Harris and Kotagiri
Ramamohanarao. \newblock Optimal clustering of relations
to improve sorting and partitioning   for joins. \newblock
{\em The Computer Journal}, 40\penalty0 (7):\penalty0
416--434, ????   1997. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/7/416.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_40/Issue_07/
Vol40_07.body.html#AbstractHarris|}\fi.

\bibitem[HRS93]{Heintz:1993:TPC} Joos Heintz, Marie-Fran{
\c{c}}oise Roy, and Pablo Solern{\'o}. \newblock On the
theoretical and practical complexity of the existential
theory   of reals. \newblock {\em The Computer Journal}
, 36\penalty0 (5):\penalty0 427--431,   October 1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/5/427.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_05/Vol36_05.body.html#AbstractHeintz|}
\fi.
\bibitem[HRW95]{Hermanns:1995:FCI} Holger Hermanns, Michael
Rettelbach, and Thorsten Weiss. \newblock Formal characterisation
of immediate actions in {SPA} with   nondeterministic
branching. \newblock {\em The Computer Journal}, 38\penalty0
(7):\penalty0 530--541, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/7/530.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_07/Vol38_07.body.html#AbstractHermanns|}
\fi.
\bibitem[HS93]{Horn:1993:PSM} F.~Horn and J.~B. Stefani.
\newblock On programming and supporting multimedia object
synchronisation. \newblock {\em The Computer Journal}
, 36\penalty0 (1):\penalty0 4--18, February   1993. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/1/4.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/Vol36_01.body.html#AbstractHorn|}
\fi.
\bibitem[HSW92]{Hwang:1992:NAC} J.-J. Hwang, B.-M. Shao,
and P.~C. Wang. \newblock A new access control method
using prime factorisation. \newblock {\em The Computer
Journal}, 35\penalty0 (1):\penalty0 16--20,   February
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/1/16.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_01/Vol35_01.body.
html#AbstractHwang|}\fi.

\bibitem[HT92]{Hou:1992:ESD} Y-C. Hou and J-C. Tsay. \
newblock On equivalent systolic designs of {$LU$} decomposition
and its   algebraic representation. \newblock {\em The
Computer Journal}, 35\penalty0 (6):\penalty0 662--666,
 December 1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/6/662.full.pdf+html|}\fi.
\bibitem[Hut93a]{Hutchinson:1993:SIE} D.~Hutchinson. \
newblock Special issue editorial: Distributed multimedia
systems. \newblock {\em The Computer Journal}, 36\penalty0
(1):\penalty0 2--3, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/1/2.
full.pdf+html|}\fi.
\bibitem[Hut93b]{Hutchison:1993:SIE} David Hutchison.
\newblock Special issue editorial: Distributed multimedia
systems. \newblock {\em The Computer Journal}, 36\penalty0
(1):\penalty0 2--??, February   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://www3.oup.co.uk/computer_journal/Volume_36/
Issue_01/Vol36_01.body.html#AbstractHuthchison|}\fi.
\bibitem[HV92]{Hankins:1992:BRB} Christopher Hankins and
John Vince. \newblock Book reviews: {{\booktitle{The SGML
Handbook}}, by Charles Goldfarb,   1990}; {{\booktitle{
Fractal Creations}}, by Timothy Wegner and Mark Peterson,
 1991}. \newblock {\em The Computer Journal}, 35\penalty0
(1):\penalty0 40, February   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/1/40.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_35/Issue_01/tiff/40.tif|}\fi.
\bibitem[HW90]{Hunter:1990:NOL} A.~Hunter and P.~J. Willis.
\newblock A note on the optimal labelling of quadtree
nodes. \newblock {\em The Computer Journal}, 33\penalty0
(5):\penalty0 398--401,   October 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/398.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/398.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/399.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/400.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/401.tif|}\fi.
\bibitem[HW92]{Harper:1992:EEC} David~J. Harper and Andrew
D.~M. Walker. \newblock {ECLAIR}: an extensible class
library for information retrieval. \newblock {\em The
Computer Journal}, 35\penalty0 (3):\penalty0 256--267,
June   1992. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/3/256.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_03/Vol35_03.body.
html#AbstractHarper|}\fi.
\bibitem[HZ95]{He:1995:CSO} Weidong He and Chaochen Zhou.
\newblock A case study of optimization. \newblock {\em
The Computer Journal}, 38\penalty0 (9):\penalty0 734-
-746, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/9/734.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_09/
Vol38_09.body.html#AbstractHe|}\fi.
\bibitem[IB96]{Ingram:1996:ALT} R.~Ingram and S.~Benford.
\newblock The application of legibility techniques to
enhance information   visualizations. \newblock {\em The
Computer Journal}, 39\penalty0 (10):\penalty0 819--836,
????   1996. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/39/10/819.full.pdf+html;   http://www.oup.
co.uk/computer_journal/Volume_39/Issue_10/Vol39_10.body.
html#AbstractIngram;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_10/Vol39_10.body.html#AbstractIngram|}
\fi.
\bibitem[IHH95]{Ibbett:1995:HFT} R.~N. Ibbett, P.~E. Heywood,
and F.~W. Howell. \newblock {HASE}: a flexible toolset
for computer architects. \newblock {\em The Computer Journal}
, 38\penalty0 (10):\penalty0 755--764, ????   1995. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL

{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/10/755.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_38/Issue_10/Vol38_10.body.html#AbstractIbbett|}
\fi.
\bibitem[ILPP98]{Ierotheou:1998:SDM} C.~Ierotheou, C.
~H. Lai, C.~J. Palansuriya, and K.~A. Pericleous. \newblock
Simulation of {$2$-D} metal cutting by means of a distributed
 algorithm. \newblock {\em The Computer Journal}, 41\
penalty0 (1):\penalty0 57--63, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/1/
57.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_01/Vol41_01.body.html#AbstractIerotheou|}
\fi.
\bibitem[Imb93]{Imbert:1993:VED} Jean-Louis Imbert. \
newblock Variable elimination for disequations in generalized
linear   constraint systems. \newblock {\em The Computer
Journal}, 36\penalty0 (5):\penalty0 473--484,   October
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/5/473.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_05/Vol36_05.body.
html#AbstractImbert|}\fi.
\bibitem[Inc92a]{Ince:1992:CT} D.~Ince. \newblock Case
technology. \newblock {\em The Computer Journal}, 35\
penalty0 (5):\penalty0 430, October   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/5/
430.full.pdf+html|}\fi.
\bibitem[Inc92b]{Ince:1992:FSU} D.~Ince. \newblock Formal
specification using {Z}. \newblock {\em The Computer Journal}
, 35\penalty0 (2):\penalty0 147a--147,   April 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/2/147-a.full.pdf+html|}\fi.
\bibitem[Inc93]{Ince:1993:BR} D.~Ince. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 596, ???? 1993. \newblock \showCODEN{CMPJA6}

. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/596.1.full.pdf+html|}
\fi.
\bibitem[Ire93]{Ireland:1993:ESM} Andrew Ireland. \newblock
On exploiting the structure of {Martin-Lof}'s {Theory
of Types}. \newblock {\em The Computer Journal}, 36\penalty0
(4):\penalty0 387--398,   August 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/4/387.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_04/Vol36_04.body.html#AbstractIreland|}
\fi.
\bibitem[Isa95]{Isaac:1995:UID} Colin Isaac. \newblock
User interface design {\&} user interface evaluation.
\newblock {\em The Computer Journal}, 38\penalty0 (3)
:\penalty0 265b--265, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/265-
b.full.pdf+html|}\fi.
\bibitem[Jag91]{Jagiello:1991:BR} J.~Jagiello. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (3):\penalty0 194, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/194.
1.full.pdf+html|}\fi.
\bibitem[Jal95a]{Jalote:1995:BR} P.~Jalote. \newblock
Book reviews. \newblock {\em The Computer Journal}, 38\
penalty0 (3):\penalty0 265, ???? 1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/265.
full.pdf+html|}\fi.
\bibitem[Jal95b]{Jalote:1995:IDA} P.~Jalote. \newblock
Introduction to distributed algorithms. \newblock {\em
The Computer Journal}, 38\penalty0 (3):\penalty0 265a-
-265, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.

oxfordjournals.org/content/38/3/265-a.full.pdf+html|}
\fi.
\bibitem[Jam95a]{Jamieson:1995:BRa} M.~J. Jamieson. \
newblock Book review. \newblock {\em The Computer Journal}
, 38\penalty0 (10):\penalty0 834--835, ????   1995. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/10/834.full.pdf+html|}\fi.
\bibitem[Jam95b]{Jamieson:1995:BRb} M.~J. Jamieson. \
newblock Book review. \newblock {\em The Computer Journal}
, 38\penalty0 (10):\penalty0 835--836, ????   1995. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/10/835.full.pdf+html|}\fi.
\bibitem[JB94]{Jackson:1994:DUP} W.~Ken Jackson and F.
~Warren Burton. \newblock A definite and unfoldable, partially
deterministic language. \newblock {\em The Computer Journal}
, 37\penalty0 (8):\penalty0 711--714, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
711.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_08/Vol37_08.body.html#AbstractJackson|}
\fi.
\bibitem[JC97]{Ju:1997:USO} Jialin Ju and Vipin Chaudhary.
\newblock Unique sets oriented parallelization of loops
with non-uniform   dependences. \newblock {\em The Computer
Journal}, 40\penalty0 (6):\penalty0 322--339, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/6/322.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_06/Vol40_06.body.html#AbstractJu;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
06/Vol40_06.body.html#AbstractJu|}\fi.
\bibitem[JCCJ99]{Jia:1999:MPB} Weijia Jia, Jiannong Cao,
To-Yat Cheung, and Xiaohua Jia. \newblock A multicast
protocol based on a single logical ring using a virtual
 token and logical clocks. \newblock {\em The Computer
Journal}, 42\penalty0 (3):\penalty0 202--220, ????   1999.

\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/3/202.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_03/420202.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_03/pdf/420202.pdf|}\fi.
\bibitem[Jen99]{Jennings:1999:DAF} Esther Jennings. \
newblock Distributed algorithms for finding central paths
in tree networks. \newblock {\em The Computer Journal}
, 42\penalty0 (7):\penalty0 609--612, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/7/
609.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_07/420609.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_07/
pdf/420609.pdf|}\fi.
\bibitem[JES98]{Jones:1998:SUG} B.~F. Jones, D.~E. Eyres,
and H.~H. Sthamer. \newblock A strategy for using genetic
algorithms to automate branch and   fault-based testing.
\newblock {\em The Computer Journal}, 41\penalty0 (2)
:\penalty0 98--107, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/2/98.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_02/Vol41_02.body.html#AbstractJones2|}
\fi.
\bibitem[JI93]{Jesshope:1993:MNC} C.~R. Jesshope and C.
~Izu. \newblock The {MP1} network chip and its application
to parallel computers. \newblock {\em The Computer Journal}
, 36\penalty0 (8):\penalty0 763--777, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/8/
763.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractJesshope|}
\fi.
\bibitem[Jir94a]{Jirotka:1994:BRa} M.~Jirotka. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (1):\penalty0 77--78, ????   1994. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/1/
77.2.full.pdf+html|}\fi.
\bibitem[Jir94b]{Jirotka:1994:BRb} M.~Jirotka. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (1):\penalty0 77--78, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/1/
77.3.full.pdf+html|}\fi.
\bibitem[JKW94]{Jones:1994:ITI} Christopher~B. Jones,
David~B. Kidner, and J.~Mark Ware. \newblock The implicit
triangulated irregular network and multiscale spatial
 databases. \newblock {\em The Computer Journal}, 37\
penalty0 (1):\penalty0 43--57, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/1/
43.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractJones|}
\fi.
\bibitem[JLV99]{Jiang:1999:NAI} Tao Jiang, Ming Li, and
Paul Vit{\'a}nyi. \newblock New applications of the incompressibility
method. \newblock {\em The Computer Journal}, 42\penalty0
(4):\penalty0 287--293, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/4/287.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420287.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420287.pdf|}\fi.
\bibitem[JM96]{Jackson:1996:PPF} David~Jeff Jackson and
Wagdy Mahmoud. \newblock Parallel pipelined fractal image
compression using quadtree   recomposition. \newblock
{\em The Computer Journal}, 39\penalty0 (1):\penalty0
1--13, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/1/1.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
01/390001.sgm.abs.html;   http://www3.oup.co.uk/computer_

journal/Volume_39/Issue_01/Vol39_01.body.html#AbstractJackson|}
\fi.
\bibitem[JO92]{Jamison:1992:SNI} B.~Jamison and S.~Olariu.
\newblock Short notes: Indexing for multi-attribute retrieval.
\newblock {\em The Computer Journal}, 35\penalty0 (5)
:\penalty0 535--537,   October 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/5/535.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/350535.sgm.abs.html; http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/tiff/535.tif;
 http://www3.oup.co.uk/computer_journal/Volume_35/Issue_
05/tiff/536.tif;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/tiff/537.tif|}\fi.
\bibitem[Joh94]{Johnson:1994:BR} R.~C. Johnson. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 731, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/731.
2.full.pdf+html|}\fi.
\bibitem[Joh96a]{Johnson:1996:BRa} C.~Johnson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (1):\penalty0 93, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/1/93.
2.full.pdf+html|}\fi.
\bibitem[Joh96b]{Johnson:1996:BRb} C.~Johnson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (4):\penalty0 357, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/4/357.
1.full.pdf+html|}\fi.
\bibitem[Joh96c]{Johnson:1996:BR} C.~Johnson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (7):\penalty0 643, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/39/7/643.
2.full.pdf+html|}\fi.
\bibitem[Jok90]{Jokinen:1990:PTA} M.~O. Jokinen. \newblock
Parameter transmission abstractions. \newblock {\em The
Computer Journal}, 33\penalty0 (2):\penalty0 133--139,
April   1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/2/133.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/133.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/134.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/135.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/136.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/137.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/139.tif|}\fi.
\bibitem[Jon94a]{Jones:1994:BRb} J.~Jones. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(8):\penalty0 726--727, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/726.
full.pdf+html|}\fi.
\bibitem[Jon94b]{Jones:1994:BRa} S.~B. Jones. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (7):\penalty0 647--648, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
647.2.full.pdf+html|}\fi.
\bibitem[Jon96]{Jones:1996:BR} C.~B. Jones. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (6):\penalty0 555, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/6/555.
full.pdf+html|}\fi.
\bibitem[Jor90]{Jordan:1990:EC} A.~Jordan. \newblock Editorial
--- communications. \newblock {\em The Computer Journal}

, 33\penalty0 (4):\penalty0 289, ???? 1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/4/
289.full.pdf+html|}\fi.
\bibitem[Jos94]{Joseph:1994:VDQ} Mathai Joseph. \newblock
The visual display of quantitative information {\&} envisioning
 information. \newblock {\em The Computer Journal}, 37\
penalty0 (6):\penalty0 551--552, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
551.full.pdf+html|}\fi.
\bibitem[Jos96]{Joseph:1996:BR} M.~Joseph. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(3):\penalty0 252, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/3/252.full.pdf+html|}\fi.
\bibitem[Joy93a]{Joy:1993:BRa} M.~S. Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 600, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/600.1.full.pdf+html|}
\fi.
\bibitem[Joy93b]{Joy:1993:BRb} M.~S. Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 600--601, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/600.
2.full.pdf+html|}\fi.
\bibitem[Joy93c]{Joy:1993:BR} M.~S. Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 601, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/601.1.full.pdf+html|}
\fi.

\bibitem[Joy94a]{Joy:1994:BRa} M.~Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(5):\penalty0 471, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/5/471.1.full.pdf+html|}
\fi.
\bibitem[Joy94b]{Joy:1994:BRc} M.~Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(8):\penalty0 730--731, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/730.
2.full.pdf+html|}\fi.
\bibitem[Joy94c]{Joy:1994:BRd} M.~Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(9):\penalty0 825--826, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/9/825.
2.full.pdf+html|}\fi.
\bibitem[Joy94d]{Joy:1994:BRe} M.~Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(9):\penalty0 826, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/9/826.1.full.pdf+html|}
\fi.
\bibitem[Joy94e]{Joy:1994:BRB} Mike Joy. \newblock Book
review: {{\booktitle{Advanced Motif Programming Technique}
}}. \newblock {\em The Computer Journal}, 37\penalty0
(6):\penalty0 552, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/552.full.pdf+html|}\fi.
\bibitem[Joy94f]{Joy:1994:IXW} Mike Joy. \newblock Introduction
to the {X} window system. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 549, ???? 1994. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL

{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/6/549.full.pdf+html|}\fi.
\bibitem[Joy94g]{Joy:1994:JX} Mike Joy. \newblock The
joy of {X}. \newblock {\em The Computer Journal}, 37\
penalty0 (6):\penalty0 548a--549, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
548-a.full.pdf+html|}\fi.
\bibitem[Joy95a]{Joy:1995:LCG} Mike Joy. \newblock Lambda
calculi: a guide for computer scientists. \newblock {
\em The Computer Journal}, 38\penalty0 (1):\penalty0 78a-
-79, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/1/78-a.full.pdf+html|}\
fi.
\bibitem[Joy95b]{Joy:1995:PCI} Mike Joy. \newblock Programming
with {C}: an introduction. \newblock {\em The Computer
Journal}, 38\penalty0 (3):\penalty0 263a--264, ????
1995. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/3/263-a.full.pdf+html|}\fi.
\bibitem[Joy96a]{Joy:1996:BRa} M.~Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(9):\penalty0 815, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/9/815.1.full.pdf+html|}
\fi.
\bibitem[Joy96b]{Joy:1996:BRb} M.~Joy. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(9):\penalty0 817, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/9/817.full.pdf+html|}\fi.
\bibitem[Joy96c]{Joy:1996:BR} M.~Joy. \newblock Book review.
\newblock {\em The Computer Journal}, 39\penalty0 (10)
:\penalty0 891--892, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/10/891.
2.full.pdf+html|}\fi.
\bibitem[JSKCG94]{Jambor-Sadeghi:1994:SAC} Kamyar Jambor-
Sadeghi, Mohammad~A. Ketabchi, Junjie Chue, and M.~Ghiassi.
\newblock A systematic approach to corrective maintenance.
\newblock {\em The Computer Journal}, 37\penalty0 (9)
:\penalty0 764--778, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/9/764.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_09/Vol37_09.body.html#AbstractSadeghi|}
\fi.
\bibitem[JW98]{Jones:1998:PST} Christopher~B. Jones and
J.~Mark Ware. \newblock Proximity search with a triangulated
spatial model. \newblock {\em The Computer Journal}, 41\
penalty0 (2):\penalty0 71--83, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/2/
71.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_02/Vol41_02.body.html#AbstractJones|}
\fi.
\bibitem[KA98]{Kumar:1998:HRC} Akhil Kumar and Noha Adly.
\newblock {HPP}: a reliable causal broadcast protocol
for large-scale   replication in wide area networks. \
newblock {\em The Computer Journal}, 41\penalty0 (2):
\penalty0 108--124, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/2/108.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_02/Vol41_02.body.html#AbstractKumar|}
\fi.
\bibitem[Kac91]{Kacsuk:1991:PPA} P.~Kacsuk. \newblock
A parallel {Prolog} abstract machine and its multi-{Transputer}
 implementation. \newblock {\em The Computer Journal}
, 34\penalty0 (1):\penalty0 52--63,   February 1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/1/52.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/55.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/56.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/58.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/59.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/60.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/62.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/63.tif|}\fi.
\bibitem[Kal91]{Kalinchenko:1991:DPB} P.~A. Kalinchenko.
\newblock Does a point belong to a polygon? \newblock
{\em The Computer Journal}, 34\penalty0 (6):\penalty0
502, December   1991. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/6/502.full.pdf+html|}\fi.
\bibitem[Kau92]{Kaur:1992:HHP} D.~Kaur. \newblock Hierarchical
hypernet for pipelining. \newblock {\em The Computer Journal}
, 35\penalty0 (1):\penalty0 81--87,   February 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/1/81.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_01/Vol35_01.body.html#AbstractKaur|}
\fi.
\bibitem[KBTB97]{Kunz:1997:PTS} T.~Kunz, J.~P. Black,
D.~J. Taylor, and T.~Basten. \newblock {Poet}: Target-
system independent visualizations of complex   distributed-
application executions. \newblock {\em The Computer Journal}
, 40\penalty0 (8):\penalty0 499--512, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/8/
499.full.pdf+html;   http://www3.oup.co.uk/computer_journal/

Volume_40/Issue_08/Vol40_08.body.html#AbstractKunz|}\
fi.
\bibitem[KC97]{Kilgore:1997:TDP} Richard~B. Kilgore and
Craig~M. Chase. \newblock Testing distributed programs
containing racing messages. \newblock {\em The Computer
Journal}, 40\penalty0 (8):\penalty0 489--498, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/8/489.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_08/Vol40_08.body.html#AbstractKilgore|}
\fi.
\bibitem[KCP99]{Kocberber:1999:OSF} S.~Kocberber, F.~
Can, and J.~M. Patton. \newblock Optimization of signature
file parameters for databases with varying   record lengths.
\newblock {\em The Computer Journal}, 42\penalty0 (1)
:\penalty0 11--23, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/1/11.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_01/420011.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_01/
pdf/420011.pdf|}\fi.
\bibitem[Ked96a]{Kedem:1996:BR} K.~Kedem. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(4):\penalty0 357--358, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/4/357.
2.full.pdf+html|}\fi.
\bibitem[Ked96b]{Kedem:1996:E} K.~Kedem. \newblock Erratum.
\newblock {\em The Computer Journal}, 39\penalty0 (6)
:\penalty0 560, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/6/560.full.pdf+html|}\fi.
\bibitem[Kel98]{Kelly:1998:ASL} Declan~P. Kelly. \newblock
{ARC}: a specification language for communicating systems.
\newblock {\em The Computer Journal}, 41\penalty0 (6)
:\penalty0 386--401, ????   1998. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/6/386.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_06/Vol41_06.body.html#AbstractKelly|}
\fi.
\bibitem[Kem90]{Kemp:1990:C} D.~A. Kemp. \newblock Correspondence.
\newblock {\em The Computer Journal}, 33\penalty0 (1)
:\penalty0 90--91, ????   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/90.
full.pdf+html|}\fi.
\bibitem[Ken96]{Kent:1996:BR} S.~J.~J. Kent. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 724, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/8/724.
1.full.pdf+html|}\fi.
\bibitem[Kha91]{Khanna:1991:LPS} S.~Khanna. \newblock
Logic programming for software verification and testing.
\newblock {\em The Computer Journal}, 34\penalty0 (4)
:\penalty0 350--357,   August 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/350.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/350.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/351.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/352.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/353.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/355.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/356.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/357.tif|}\fi.
\bibitem[Kha94]{Khan:1994:BR} A.~H. Khan. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(5):\penalty0 472, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)

}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/5/472.1.full.pdf+html|}
\fi.
\bibitem[Kho93]{Khoong:1993:SPR} C.~M. Khoong. \newblock
Shortest-path reconstruction algorithms. \newblock {\
em The Computer Journal}, 36\penalty0 (6):\penalty0 588-
-592,   December 1993. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/588.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_06/
Vol36_06.body.html#AbstractKhoong|}\fi.
\bibitem[Kil94]{Kilgour:1994:C} Alistair Kilgour. \newblock
Correspondence. \newblock {\em The Computer Journal},
37\penalty0 (6):\penalty0 560, ???? 1994. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
560.full.pdf+html|}\fi.
\bibitem[Kil96]{Kilgour:1996:BR} A.~C. Kilgour. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (2):\penalty0 170--171, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/2/
170.2.full.pdf+html|}\fi.
\bibitem[Kin94]{Kingman:1994:BR} K.~Kingman. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 146, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/146.
2.full.pdf+html|}\fi.
\bibitem[Kir94a]{Kirkham:1994:BRa} C.~C. Kirkham. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 227--228, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/3/
227.2.full.pdf+html|}\fi.
\bibitem[Kir94b]{Kirkham:1994:BRb} C.~C. Kirkham. \newblock
Book review. \newblock {\em The Computer Journal}, 37\

penalty0 (4):\penalty0 350--351, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
350.2.full.pdf+html|}\fi.
\bibitem[KK92]{Korah:1992:SNP} A.~P. Korah and M.~R. Kaimal.
\newblock A short note on perfectly balanced binary search
trees. \newblock {\em The Computer Journal}, 35\penalty0
(6):\penalty0 660--662,   December 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/660.
full.pdf+html|}\fi.
\bibitem[KK94]{Kinnersley:1994:EPT} Nancy~G. Kinnersley
and William~M. Kinnersley. \newblock An efficient polynomial-
time algorithm for three-track gate matrix   layout. \
newblock {\em The Computer Journal}, 37\penalty0 (5):
\penalty0 449--462, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/5/449.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_05/Vol37_05.body.html#AbstractKinnersley|}
\fi.
\bibitem[KKR96]{Kranakis:1996:MLL} Evangelos Kranakis,
Danny Krizanc, and S.~S. Ravi. \newblock On multi-label
linear interval routing schemes. \newblock {\em The Computer
Journal}, 39\penalty0 (2):\penalty0 133--139, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/2/133.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_02/390133.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_02/
Vol39_02.body.html#AbstractKranakis|}\fi.
\bibitem[KLB91]{Kingdon:1991:HMH} H.~Kingdon, D.~R. Lester,
and G.~L. Burn. \newblock The {HDG}-machine: a highly
distributed graph-reducer for a   {Transputer} network.
\newblock {\em The Computer Journal}, 34\penalty0 (4)
:\penalty0 290--301,   August 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/290.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/290.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/291.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/292.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/293.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/294.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/295.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/296.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/297.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/298.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/299.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/300.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/301.
t.
\bibitem[Kle97]{Klein:1997:EOR} S.~T. Klein. \newblock
Efficient optimal recompression. \newblock {\em The Computer
Journal}, 40\penalty0 (2--3):\penalty0 117--126,   ??
?? 1997. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/2_and_3/117.full.pdf+html;   http://www.
oup.co.uk/computer_journal/Volume_40/Issue_02/Vol40_02.
body.html#AbstractKlein;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/Vol40_03.body.html#AbstractKlein|}
\fi.
\bibitem[KM94]{Knight:1994:TDM} Brian Knight and Jixin
Ma. \newblock A temporal database model supporting relative
and absolute time. \newblock {\em The Computer Journal}
, 37\penalty0 (7):\penalty0 588--597, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
588.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_07/Vol37_07.body.html#AbstractKnight|}
\fi.
\bibitem[KMK95]{Kirstein:1995:COP} Peter Kirstein and
Goli Montasser-Kohsari. \newblock The {C-ODA} project:
Experiences and tools. \newblock {\em The Computer Journal}
, 38\penalty0 (8):\penalty0 670--680, ????   1995. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/8/
670.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_08/Vol38_08.body.html#AbstractKirstein|}
\fi.
\bibitem[KMR94]{Kerber:1994:CTD} Gerrit Kerber, Helmut
Meitner, and Friedemann Reim. \newblock The {Comandos}
toolset for distributed systems management. \newblock
{\em The Computer Journal}, 37\penalty0 (6):\penalty0
509--518, ????   1994. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/509.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_06/
Vol37_06.body.html#AbstractKerber|}\fi.
\bibitem[Kni93]{Knight:1993:ILT} B.~Knight. \newblock
Information loss in temporal knowledge representations.
\newblock {\em The Computer Journal}, 36\penalty0 (2)
:\penalty0 127--136, April   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/2/127.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_02/Vol36_02.body.html#AbstractKnight|}
\fi.
\bibitem[Kor94]{Korver:1994:TS} Henri~P. Korver. \newblock
A theory for simulators. \newblock {\em The Computer Journal}
, 37\penalty0 (4):\penalty0 279--287, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
279.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_04/Vol37_04.body.html#AbstractKorver|}
\fi.
\bibitem[KP92]{Kim:1992:ODA} M.~Ho Kim and S.~Pramanik.
\newblock Optimizing database accesses for parallel processing
of multikey   range searches. \newblock {\em The Computer
Journal}, 35\penalty0 (1):\penalty0 45--51,   February
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/1/45.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/Volume_35/Issue_01/Vol35_01.body.
html#AbstractKim|}\fi.
\bibitem[KP97a]{Kumar:1997:ALP} M.~Kumar and L.~M. Patnaik.
\newblock Automatic loop parallelization. \newblock {
\em The Computer Journal}, 40\penalty0 (6):\penalty0 301,
???? 1997. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/6/301.full.pdf+html|}\fi.
\bibitem[KP97b]{Kumar:1997:SIE} Mohan Kumar and L.~M.
Patnaik. \newblock Special issue editorial: Automatic
loop parallelization. \newblock {\em The Computer Journal}
, 40\penalty0 (6):\penalty0 301, ???? 1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www.oup.co.uk/computer_journal/Volume_40/
Issue_06/AbstractEditorial;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_06/Vol40_06.body.html#AbstractEditorial|}
\fi.
\bibitem[KP98]{Kwai:1998:TBD} Ding-Ming Kwai and Behrooz
Parhami. \newblock Tight bounds on the diameter of {Gaussian}
cubes. \newblock {\em The Computer Journal}, 41\penalty0
(1):\penalty0 52--56, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/1/52.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_01/Vol41_01.body.html#AbstractKwai|}\
fi.
\bibitem[KPP{\etalchar{+}}95]{Kapelios:1995:CRI} V.~Kapelios,
G.~Panagopoulou, G.~Papamichail, S.~Sirmakessis, and
A.~Tsakalidis. \newblock The {`Cross'} rectangle intersection
problem. \newblock {\em The Computer Journal}, 38\penalty0
(3):\penalty0 227--235, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/227.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_03/Vol38_03.body.html#AbstractKapelios|}
\fi.
\bibitem[KPT96]{Koziris:1996:OTE} Nectarios Koziris, George
Papakonstantinou, and Panayotis Tsanakas. \newblock Optimal

time and efficient space free scheduling for nested loops.
\newblock {\em The Computer Journal}, 39\penalty0 (5)
:\penalty0 439--448, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/5/439.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_05/390439.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_05/Vol39_
05.body.html#AbstractKoziris|}\fi.
\bibitem[KS91]{King:1991:DDI} P.~King and C.~Small. \
newblock Default databases and incomplete information.
\newblock {\em The Computer Journal}, 34\penalty0 (3)
:\penalty0 239--244, June   1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/239.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/239.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/240.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/241.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/242.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/243.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/244.tif|}\fi.
\bibitem[KS92]{Kluzniak:1992:C} Feliks Kluzniak and Stanislaw
Szpakowicz. \newblock Correspondence. \newblock {\em The
Computer Journal}, 35\penalty0 (3):\penalty0 313, June
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/3/313.full.pdf+html|}\fi.
\bibitem[KS97]{Kulkarni:1997:LEL} Dattatraya Kulkarni
and Micheal Stumm. \newblock Linear and extended linear
transformations for shared-memory   multiprocessors. \
newblock {\em The Computer Journal}, 40\penalty0 (6):
\penalty0 373--387, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/6/373.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_06/Vol40_06.body.html#AbstractKulkarni;

 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
06/Vol40_06.body.html#AbstractKulkarni|}\fi.
\bibitem[KU99]{Kayan:1999:ERT} Ersan Kayan and {\"O}zg{
\"u}r Ulusoy. \newblock An evaluation of real-time transaction
management issues in mobile   database systems. \newblock
{\em The Computer Journal}, 42\penalty0 (6):\penalty0
501--510, ????   1999. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/6/501.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
06/420501.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_06/pdf/420501.pdf|}\fi.
\bibitem[Kuh92]{Kuhn:1992:TAE} D.~R. Kuhn. \newblock A
technique for analyzing the effects of changes in formal
 specifications. \newblock {\em The Computer Journal}
, 35\penalty0 (6):\penalty0 574--578,   December 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/574.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/Vol35_06.body.html#AbstractKuhn|}
\fi.
\bibitem[Kul95]{Kulkarni:1995:BR} V.~Kulkarni. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (9):\penalty0 747--748, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/9/
747.3.full.pdf+html|}\fi.
\bibitem[Kum99]{Kumar:1999:SIE} Mohan Kumar. \newblock
Special issue editorial: Mobile computing. \newblock {
\em The Computer Journal}, 42\penalty0 (6):\penalty0 441,
???? 1999. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/42/6/441.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_42/Issue_06/420441.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_06/pdf/420441.pdf|}\fi.
\bibitem[KV95]{Koster:1995:HST} C.~H.~A. Koster and Th.
~P. {Van Der Weide}. \newblock Hairy search trees. \newblock

{\em The Computer Journal}, 38\penalty0 (8):\penalty0
691--694, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/8/691.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_08/
Vol38_08.body.html#AbstractKoster|}\fi.
\bibitem[KVKP95]{Kumar:1995:HDB} M.~J. Kumar, S.~Venkatesh,
D.~Kieronska, and L.~M. Patnaik. \newblock Hierarchical
directory-based shared memory architecture. \newblock
{\em The Computer Journal}, 38\penalty0 (3):\penalty0
207--216, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/3/207.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_03/
Vol38_03.body.html#AbstractKumar|}\fi.
\bibitem[Lah91]{Lahodynskyj:1991:SC} Lu~Lahodynskyj. \
newblock Structure clashes. \newblock {\em The Computer
Journal}, 34\penalty0 (6):\penalty0 566--568,   December
1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/34/6/566.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/566.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/567.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/568.tif|}\fi.
\bibitem[Lai94]{Lai:1994:DDD} C.-H. Lai. \newblock Diakoptics,
domain decomposition and parallel computing. \newblock
{\em The Computer Journal}, 37\penalty0 (10):\penalty0
840--846, ????   1994. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/10/840.full.pdf+html;
http://www3.oup.co.uk/computer_journal/Volume_37/Issue_
10/Vol37_10.body.html#AbstractLai|}\fi.
\bibitem[Lal99]{Lalmas:1999:MRR} Mounia Lalmas. \newblock
A model for representing and retrieving heterogeneous
structured   documents based on evidential reasoning.
\newblock {\em The Computer Journal}, 42\penalty0 (7)
:\penalty0 547--568, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/7/547.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_07/420547.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_07/
pdf/420547.pdf|}\fi.
\bibitem[Lan90]{Landsdown:1990:EC} J.~Landsdown. \newblock
Editorial --- communications. \newblock {\em The Computer
Journal}, 33\penalty0 (5):\penalty0 385, ???? 1990. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/5/385.full.pdf+html|}\fi.
\bibitem[Lar94a]{Larcombe:1994:BR} M.~H.~E. Larcombe.
\newblock Book review. \newblock {\em The Computer Journal}
, 37\penalty0 (5):\penalty0 468--469, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/5/
468.full.pdf+html|}\fi.
\bibitem[Lar94b]{Larner:1994:BRa} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 228, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/228.
1.full.pdf+html|}\fi.
\bibitem[Lar94c]{Larner:1994:BRb} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 228--229, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/3/
228.2.full.pdf+html|}\fi.
\bibitem[Lar94d]{Larner:1994:BRc} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 351, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/351.
1.full.pdf+html|}\fi.

\bibitem[Lar94e]{Larner:1994:BRd} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 351--352, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
351.2.full.pdf+html|}\fi.
\bibitem[Lar94f]{Larner:1994:BRe} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (5):\penalty0 471--472, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/5/
471.2.full.pdf+html|}\fi.
\bibitem[Lar94g]{Larner:1994:BRf} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (7):\penalty0 646, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/7/646.
1.full.pdf+html|}\fi.
\bibitem[Lar94h]{Larner:1994:BRg} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (9):\penalty0 824--825, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/9/
824.full.pdf+html|}\fi.
\bibitem[Lar94i]{Larner:1994:BRh} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (9):\penalty0 825, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/9/825.
1.full.pdf+html|}\fi.
\bibitem[Lar94j]{Larner:1994:CCP} Adrian Larner. \newblock
{C} for corporate programmers. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 554, ???? 1994. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/6/554.full.pdf+html|}\fi.

\bibitem[Lar94k]{Larner:1994:DUE} Adrian Larner. \newblock
Designing usable electronic text. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 554a--555, ????
1994. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/6/554-a.full.pdf+html|}\fi.
\bibitem[Lar94l]{Larner:1994:ISU} Adrian Larner. \newblock
Introducing specification using. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 555, ???? 1994. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/6/555.full.pdf+html|}\fi.
\bibitem[Lar94m]{Larner:1994:SAB} Adrian Larner. \newblock
Software assistance for business re-engineering. \newblock
{\em The Computer Journal}, 37\penalty0 (6):\penalty0
547, ???? 1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/547.full.pdf+html|}\fi.
\bibitem[Lar95a]{Larner:1995:BRa} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (10):\penalty0 836, ????   1995. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/10/
836.1.full.pdf+html|}\fi.
\bibitem[Lar95b]{Larner:1995:BRb} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (10):\penalty0 836--837, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/10/
836.2.full.pdf+html|}\fi.
\bibitem[Lar95c]{Larner:1995:SQ} Adrian Larner. \newblock
Software quality. \newblock {\em The Computer Journal}
, 38\penalty0 (3):\penalty0 264, ???? 1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/38/3/
264.full.pdf+html|}\fi.
\bibitem[Lar95d]{Larsson:1995:EAF} Mats Larsson. \newblock
An engineering approach to formal digital system design.
\newblock {\em The Computer Journal}, 38\penalty0 (2)
:\penalty0 101--110, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/2/101.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_02/Vol38_02.body.html#AbstractLarsson|}
\fi.
\bibitem[Lar96a]{Larner:1996:BRa} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (3):\penalty0 253--254, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/3/
253.2.full.pdf+html|}\fi.
\bibitem[Lar96b]{Larner:1996:BRb} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (4):\penalty0 358--359, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/4/
358.2.full.pdf+html|}\fi.
\bibitem[Lar96c]{Larner:1996:BRc} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (5):\penalty0 449--452, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/5/
449.2.full.pdf+html|}\fi.
\bibitem[Lar96d]{Larner:1996:BRd} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (5):\penalty0 449--452, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/5/
449.3.full.pdf+html|}\fi.
\bibitem[Lar96e]{Larner:1996:BRe} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 39\

penalty0 (7):\penalty0 645--646, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/7/
645.full.pdf+html|}\fi.
\bibitem[Lar96f]{Larner:1996:BRf} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 723, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/8/723.
1.full.pdf+html|}\fi.
\bibitem[Lar96g]{Larner:1996:BRg} A.~Larner. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 726--727, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/8/
726.2.full.pdf+html|}\fi.
\bibitem[Lar97]{Larner:1997:BR} A.~Larner. \newblock Book
review. \newblock {\em The Computer Journal}, 40\penalty0
(4):\penalty0 233--234, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/4/233.
full.pdf+html|}\fi.
\bibitem[Lau91]{Lau:1991:TSS} K.~K. Lau. \newblock Top-
down synthesis of sorting algorithm. \newblock {\em The
Computer Journal}, 34\penalty0 (6):\penalty0 A001--A007,
 December 1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\bibitem[LB91]{Lee:1991:RDS} W.~W.~L. Lee and D.~S. Bowers.
\newblock A relational database simulator. \newblock {
\em The Computer Journal}, 34\penalty0 (6):\penalty0 A061-
-A072,   December 1991. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}.
\bibitem[LB92]{Ling:1992:MMT} D.~H.~O. Ling and D.~A.
Bell. \newblock Modelling and managing time in database
systems. \newblock {\em The Computer Journal}, 35\penalty0
(4):\penalty0 332--341,   August 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/4/332.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_04/Vol35_04.body.html#AbstractLing|}\
fi.
\bibitem[LBC{\etalchar{+}}94]{Lamming:1994:DHM} M.~Lamming,
P.~Brown, K.~Carter, M.~Eldridge, M.~Flynn, G.~Louie,
P.~Robinson,   and A.~Sellen. \newblock The design of
a human memory prosthesis. \newblock {\em The Computer
Journal}, 37\penalty0 (3):\penalty0 153--163, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/3/153.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_03/tiff/153.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_37/Issue_
03/tiff/154.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_37/Issue_03/tiff/155.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_37/Issue_03/tiff/156.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
37/Issue_03/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_03/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_37/Issue_
03/tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_37/Issue_03/tiff/160.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_37/Issue_03/tiff/161.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
37/Issue_03/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_03/tiff/163.tif|}\fi.
\bibitem[LC90]{Lin:1990:DIS} F.-.~C. Lin and K.~Chen.
\newblock On the design of an integrated systolic array
for solving   simultaneous linear equations. \newblock
{\em The Computer Journal}, 33\penalty0 (3):\penalty0
252--260, June   1990. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/3/252.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/253.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/255.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_03/tiff/256.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/257.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/258.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/259.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/260.tif|}\fi.
\bibitem[LC93]{Lin:1993:GRM} Min~Sheng Lin and Deng~Jyi
Chen. \newblock General reduction methods for the reliability
analysis of distributed   computing systems. \newblock
{\em The Computer Journal}, 36\penalty0 (7):\penalty0
631--644, ????   1993. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/7/631.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_07/
Vol36_07.body.html#AbstractLin|}\fi.
\bibitem[LC94]{Low:1994:SAP} Marie~Rose Low and Bruce
Christianson. \newblock Self authenticating proxies. \
newblock {\em The Computer Journal}, 37\penalty0 (5):
\penalty0 422--428, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/5/422.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_05/Vol37_05.body.html#AbstractLow|}\fi.
\bibitem[LC95]{Leung:1995:EFP} C.~H.~C. Leung and E.~
Currie. \newblock The effect of failures on the performance
of long-duration database   transactions. \newblock {
\em The Computer Journal}, 38\penalty0 (6):\penalty0 471-
-478, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/6/471.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_06/
Vol38_06.body.html#AbstractLeung|}\fi.
\bibitem[LC98]{Lin:1998:DMF} Shun-Shii Lin and Shan-Tai
Chen. \newblock On detecting multiple faults in baseline
interconnection networks. \newblock {\em The Computer
Journal}, 41\penalty0 (4):\penalty0 254--269, ????   1998.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL

{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/4/254.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_41/Issue_04/Vol41_04.body.html#AbstractLin|}
\fi.
\bibitem[LC99]{Leung:1999:NPP} H.~Leung and T.~Y. Chen.
\newblock A new perspective of the proportional sampling
strategy. \newblock {\em The Computer Journal}, 42\penalty0
(8):\penalty0 693--698, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/8/693.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_08/420693.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_08/
pdf/420693.pdf|}\fi.
\bibitem[LCH99]{Lin:1999:RAD} Min-Sheng Lin, Deng-Jyi
Chen, and Maw-Sheng Horng. \newblock The reliability analysis
of distributed computing systems with   imperfect nodes.
\newblock {\em The Computer Journal}, 42\penalty0 (2)
:\penalty0 129--141, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/2/129.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_02/420129.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_02/
pdf/420129.pdf|}\fi.
\bibitem[LCL92]{Lin:1992:ROC} C.~H. Lin, C.~C. Chang,
and R.~C.~T. Lee. \newblock A record-oriented cryptosystem
for database sharing. \newblock {\em The Computer Journal}
, 35\penalty0 (6):\penalty0 658--660,   December 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/658.full.pdf+html|}\fi.
\bibitem[Lee92]{Leeming:1992:IPS} A.~Leeming. \newblock
Informatization in the public sector. \newblock {\em The
Computer Journal}, 35\penalty0 (5):\penalty0 477, October
 1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/477.full.pdf+html|}\fi.

\bibitem[Lep95]{Leppanen:1995:PWO} Ville Lepp{\"a}nen.
\newblock Performance of work-optimal {PRAM} simulation
algorithms on coated   meshes. \newblock {\em The Computer
Journal}, 38\penalty0 (10):\penalty0 801--810, ????
1995. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/10/801.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_38/Issue_10/Vol38_10.body.
html#AbstractLeppanen|}\fi.
\bibitem[Lev96]{Levene:1996:MCI} Mark Levene. \newblock
Maintaining consistency of imprecise relations. \newblock
{\em The Computer Journal}, 39\penalty0 (2):\penalty0
114--123, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/2/114.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
02/390114.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_02/Vol39_02.body.html#AbstractLevene|}
\fi.
\bibitem[Lev99]{Levin:1999:RMI} Leonid.~A. Levin. \newblock
Robust measures of information. \newblock {\em The Computer
Journal}, 42\penalty0 (4):\penalty0 284--286, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/4/284.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_04/420284.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_04/pdf/420284.pdf|}\fi.
\bibitem[LF97]{Lee:1997:CMP} Hyuk-Jae Lee and Jos{\'e}
A.~B. Fortes. \newblock Communication-minimal partitioning
and data alignment for affine   nested loops. \newblock
{\em The Computer Journal}, 40\penalty0 (6):\penalty0
302--310, ????   1997. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/6/302.full.pdf+html;   http:
//www.oup.co.uk/computer_journal/Volume_40/Issue_06/Vol40_
06.body.html#AbstractLee;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_06/Vol40_06.body.html#AbstractLee|}
\fi.

\bibitem[LH95]{Lam:1995:CCT} Kam-Yiu Lam and Sheung-Lun
Hung. \newblock Concurrency control for time-constrained
transactions in distributed   databases systems. \newblock
{\em The Computer Journal}, 38\penalty0 (9):\penalty0
704--716, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/9/704.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_09/
Vol38_09.body.html#AbstractLam|}\fi.
\bibitem[Li98]{Li:1998:OAJ} Keqin Li. \newblock Optimizing
average job response time via decentralized probabilistic
 job dispatching in heterogeneous multiple computer systems.
\newblock {\em The Computer Journal}, 41\penalty0 (4)
:\penalty0 223--230, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/4/223.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_04/Vol41_04.body.html#AbstractLi|}\fi.
\bibitem[Lid90]{Liddell:1990:EPC} H.~Liddell. \newblock
Editorial --- parallel computing. \newblock {\em The Computer
Journal}, 33\penalty0 (3):\penalty0 193, ???? 1990. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/3/193.full.pdf+html|}\fi.
\bibitem[LJ98]{Li:1998:CTD} Y.~H. Li and A.~K. Jain. \
newblock Classification of text documents. \newblock {
\em The Computer Journal}, 41\penalty0 (8):\penalty0 537-
-546, ????   1998. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/8/537.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_08/
Vol41_08.body.html#AbstractLi|}\fi.
\bibitem[LL90]{Levene:1990:NRT} M.~Levene and G.~Loizou.
\newblock The nested relation type model: an application
of domain theory to   databases. \newblock {\em The Computer
Journal}, 33\penalty0 (1):\penalty0 19--30,   February
1990. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock

\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/33/1/19.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/19.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/20.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/21.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/22.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/24.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/25.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/26.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/27.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/29.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/30.tif|}\fi.
\bibitem[LL94]{Ladkin:1994:CCP} P.~B. Ladkin and S.~Leue.
\newblock Correspondence: Comments on a proposed semantics
for basic message   sequence charts. \newblock {\em The
Computer Journal}, 37\penalty0 (9):\penalty0 814--817,
????   1994. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/9/814.full.pdf+html|}\fi.
\bibitem[LL97]{Lau:1997:DHM} S.-W. Lau and J.~C.~S. Lui.
\newblock Designing a hierarchical multimedia storage
server. \newblock {\em The Computer Journal}, 40\penalty0
(9):\penalty0 529--540, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/9/529.
full.pdf+html|}\fi.
\bibitem[LO99]{Liu:1999:MAH} Kecheng Liu and Tina Ong.
\newblock A modelling approach for handling business rules
and exceptions. \newblock {\em The Computer Journal},
42\penalty0 (3):\penalty0 221--231, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/3/
221.full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_42/Issue_03/420221.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_03/
pdf/420221.pdf|}\fi.
\bibitem[Lod94a]{Lodaya:1994:BRa} K.~Lodaya. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 225--226, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/3/
225.full.pdf+html|}\fi.
\bibitem[Lod94b]{Lodaya:1994:BRb} K.~Lodaya. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 348--349, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
348.full.pdf+html|}\fi.
\bibitem[LOKM99]{Loucif:1999:PMB} S.~Loucif, M.~Ould-
Khaoua, and L.~M. Mackenzie. \newblock On the performance
merits of bypass channels in hypermeshes and   $k$-ary
$n$-cubes. \newblock {\em The Computer Journal}, 42\penalty0
(1):\penalty0 62--72, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/1/62.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_01/420067.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_01/
pdf/420062.pdf|}\fi.
\bibitem[Lop94]{Lopriore:1994:SCM} Lanfranco Lopriore.
\newblock Stack cache memory for block-structured programs.
\newblock {\em The Computer Journal}, 37\penalty0 (7)
:\penalty0 610--620, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/7/610.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_07/Vol37_07.body.html#AbstractLopriore|}
\fi.
\bibitem[Low95]{Lowe:1995:SOM} Gavin Lowe. \newblock Scheduling-
oriented models for real-time systems. \newblock {\em
The Computer Journal}, 38\penalty0 (6):\penalty0 443-

-456, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/6/443.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_06/
Vol38_06.body.html#AbstractLowe|}\fi.
\bibitem[LP92]{Larsen:1992:SNE} P.~G. Larsen and N.~Plat.
\newblock Standards for non-executable specification languages.
\newblock {\em The Computer Journal}, 35\penalty0 (6)
:\penalty0 567--573,   December 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/567.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_06/Vol35_06.body.html#AbstractLarsen|}
\fi.
\bibitem[LPS94]{Lorentzos:1994:IUO} Nikos~A. Lorentzos,
Alexandra Poulovassilis, and Carol Small. \newblock Implementation
of update operations for interval relations. \newblock
{\em The Computer Journal}, 37\penalty0 (3):\penalty0
164--176, ????   1994. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/3/164.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_03/
Vol37_03.body.html#AbstractLorentzos|}\fi.
\bibitem[LPSC99]{Lam:1999:REC} Kam-Yiu Lam, Chung-Leung
Pang, Sang~H. Son, and Jiannong Cao. \newblock Resolving
executing-committing conflicts in distributed real-time
 database systems. \newblock {\em The Computer Journal}
, 42\penalty0 (8):\penalty0 674--692, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/8/
674.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_08/420674.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_08/
pdf/420674.pdf|}\fi.
\bibitem[LQ96]{Loia:1996:OSD} Vincenzo Loia and Michel
Quaggetto. \newblock The {OPLA} system: designing complex
systems in an object-oriented   logic programming framework.
\newblock {\em The Computer Journal}, 39\penalty0 (1)
:\penalty0 20--35, ????   1996. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/1/20.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_01/390020.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_01/Vol39_
01.body.html#AbstractLoia|}\fi.
\bibitem[LS93a]{Liska:1993:AQE} Richard Liska and Stanly
Steinberg. \newblock Applying quantifier elimination to
stability analysis of difference   schemes. \newblock
{\em The Computer Journal}, 36\penalty0 (5):\penalty0
497--503,   October 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/5/497.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_05/
Vol36_05.body.html#AbstractLiska|}\fi.
\bibitem[LS93b]{Lougher:1993:DSS} Phillip Lougher and
Doug Shepherd. \newblock The design of a storage server
for continuous media. \newblock {\em The Computer Journal}
, 36\penalty0 (1):\penalty0 32--42,   February 1993. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/1/32.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/Vol36_01.body.html#AbstractLougher|}
\fi.
\bibitem[LS97]{Leong:1997:DCA} Hong~Va Leong and Antonio
Si. \newblock Database caching over the air-storage. \
newblock {\em The Computer Journal}, 40\penalty0 (7):
\penalty0 401--415, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/7/401.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_40/Issue_07/Vol40_07.body.html#AbstractLeong|}
\fi.
\bibitem[Luc94a]{Lucas:1994:BRa} K.~Lucas. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(2):\penalty0 150--151, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/37/2/150.
full.pdf+html|}\fi.
\bibitem[Luc94b]{Lucas:1994:BRb} K.~M. Lucas. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (9):\penalty0 826--827, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/9/
826.2.full.pdf+html|}\fi.
\bibitem[Luc94c]{Luck:1994:AI} Michael Luck. \newblock
Artificial intelligence. \newblock {\em The Computer Journal}
, 37\penalty0 (6):\penalty0 553, ???? 1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
553.full.pdf+html|}\fi.
\bibitem[Luc95a]{Lucas:1995:BR} K.~Lucas. \newblock Book
review. \newblock {\em The Computer Journal}, 38\penalty0
(9):\penalty0 748, ???? 1995. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/9/748.1.full.pdf+html|}
\fi.
\bibitem[Luc95b]{Lucas:1995:BAP} K.~M. Lucas. \newblock
Building applications with pexlib. \newblock {\em The
Computer Journal}, 38\penalty0 (3):\penalty0 268, ???
? 1995. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/3/268.full.pdf+html|}\fi.
\bibitem[Luc95c]{Luck:1995:BR} M.~Luck. \newblock Book
review. \newblock {\em The Computer Journal}, 38\penalty0
(9):\penalty0 748--749, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/9/748.
2.full.pdf+html|}\fi.
\bibitem[Luf94]{Luff:1994:BR} A.~P. Luff. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(7):\penalty0 644--645, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/37/7/644.
2.full.pdf+html|}\fi.
\bibitem[Luk99]{Luk:1999:NOH} R.~W.~P. Luk. \newblock
Near optimal $\beta$ heap. \newblock {\em The Computer
Journal}, 42\penalty0 (5):\penalty0 391--399, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/5/391.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_05/420391.sgm.abs.html|}\
fi.
\bibitem[Lun96]{Lunzer:1996:BR} A.~Lunzer. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(9):\penalty0 812--813, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/9/812.
full.pdf+html|}\fi.
\bibitem[LW93]{Loos:1993:ALQ} R{\"u}diger Loos and Volker
Weispfenning. \newblock Applying linear quantifier elimination.
\newblock {\em The Computer Journal}, 36\penalty0 (5)
:\penalty0 450--462,   October 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/5/450.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_05/Vol36_05.body.html#AbstractLoos|}\
fi.
\bibitem[Lyo93]{Lyons:1993:BR} D.~Lyons. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 599--600, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/599.
2.full.pdf+html|}\fi.
\bibitem[LZ95]{Lowe:1995:RCS} Gavin Lowe and Hussein Zedan.
\newblock Refinement of complex systems: a case study.
\newblock {\em The Computer Journal}, 38\penalty0 (10)
:\penalty0 785--800, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/10/785.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_10/Vol38_10.body.html#AbstractLowe|}\
fi.
\bibitem[Mac91]{Macleod:1991:QLR} I.~A. Macleod. \newblock
A query language for retrieving information from hierarchic
text   structures. \newblock {\em The Computer Journal}
, 34\penalty0 (3):\penalty0 254--264, June   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/3/
254.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/254.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/255.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/256.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/257.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/258.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/259.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/260.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/261.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/263.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/264.tif|}\fi.
\bibitem[Mac92]{MacCallum:1992:BR} M.~A.~H. MacCallum.
\newblock Book review. \newblock {\em The Computer Journal}
, 35\penalty0 (6):\penalty0 650a--650,   December 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/6/650-a.full.pdf+html|}\fi.
\bibitem[Mac93]{MacCallum:1993:BR} I.~R. MacCallum. \
newblock Book review. \newblock {\em The Computer Journal}
, 36\penalty0 (3):\penalty0 290, ???? 1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/3/
290.1.full.pdf+html|}\fi.
\bibitem[Mad97]{Madria:1997:SCC} Sanjay~Kumar Madria.
\newblock A study of the concurrency control and recovery
algorithms in nested   transaction environment. \newblock

{\em The Computer Journal}, 40\penalty0 (10):\penalty0
630--639, ????   1997. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/10/630.full.pdf+html;
http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
10/Vol40_10.body.html#AbstractMadria|}\fi.
\bibitem[Mad98]{Madria:1998:HMC} Sanjay~Kumar Madria.
\newblock Handling of mutual conflicts in distributed
databases using   timestamps. \newblock {\em The Computer
Journal}, 41\penalty0 (6):\penalty0 376--385, ????   1998.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/6/376.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_41/Issue_06/Vol41_06.body.html#AbstractMadria|}
\fi.
\bibitem[Maj94]{Majumdar:1994:NIA} A.~A.~K. Majumdar.
\newblock A note on the iterative algorithm for the {
Reve}'s puzzle. \newblock {\em The Computer Journal},
37\penalty0 (5):\penalty0 463--464, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/5/
463.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_05/Vol37_05.body.html#AbstractMajumdar|}
\fi.
\bibitem[Mak91a]{Makinen:1991:EGR} E.~Makinen. \newblock
Efficient generation of rotational-admissible codewords
for binary   trees. \newblock {\em The Computer Journal}
, 34\penalty0 (4):\penalty0 379, August   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/379.tif|}\fi.
\bibitem[M{\"a}k91b]{Makinen:1991:SN} E.~M{\"a}kinen.
\newblock Short notes. \newblock {\em The Computer Journal}
, 34\penalty0 (4):\penalty0 379, ???? 1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/4/
379.full.pdf+html|}\fi.

\bibitem[M{\"a}k91c]{Makinen:1991:SBT} E.~M{\"a}kinen.
\newblock A survey on binary tree codings. \newblock {
\em The Computer Journal}, 34\penalty0 (5):\penalty0 438-
-443,   October 1991. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/5/438.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/438.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/439.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/440.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/441.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/442.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/443.tif|}\fi.
\bibitem[Mak92]{Mak:1992:CRP} J.~Mak. \newblock {CD-ROMs}
in print: an international guide to {CD-ROM}, {CD-I},
 {CDTV} {\&} electronic book product. \newblock {\em The
Computer Journal}, 35\penalty0 (6):\penalty0 629, December
 1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/6/629.full.pdf+html|}\fi.
\bibitem[Mak95]{Makur:1995:BR} A.~Makur. \newblock Book
review. \newblock {\em The Computer Journal}, 38\penalty0
(10):\penalty0 831, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/10/831.
1.full.pdf+html|}\fi.
\bibitem[M{\"a}k98]{Makinen:1998:BTC} E.~M{\"a}kinen.
\newblock Binary tree code words as context-free languages.
\newblock {\em The Computer Journal}, 41\penalty0 (6)
:\penalty0 422--424, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/6/422.
full.pdf+html|}\fi.
\bibitem[M{\"a}k01]{Makinen:2001:CFM} Erkki M{\"a}kinen.
\newblock Comment on {`A Framework for Modelling Trojans
and Computer Virus   Infection'}. \newblock {\em The Computer

Journal}, 44\penalty0 (4):\penalty0 321--323, ????   2001.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://www3.oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/440321.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_44/Issue_04/
pdf/440321.pdf|}\fi. \newblock See \cite{Thimbleby:1998:
FMT,Thimbleby:2001:RCF,Cohen:2001:RCF}.
\bibitem[Mal96]{Malard:1996:BR} J.~Malard. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(2):\penalty0 169--170, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/2/169.
2.full.pdf+html|}\fi.
\bibitem[Man92]{Manolopoulos:1992:RCA} Y.~Manolopoulos.
\newblock Reverse chaining for answering temporal logical
queries. \newblock {\em The Computer Journal}, 35\penalty0
(6):\penalty0 666--668,   December 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/666.
full.pdf+html|}\fi.
\bibitem[Man93]{Manocha:1993:EAM} Dinesh Manocha. \newblock
Efficient algorithms for {MultiPolynomial} resultant.
\newblock {\em The Computer Journal}, 36\penalty0 (5)
:\penalty0 485--496,   October 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/5/485.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_05/Vol36_05.body.html#AbstractManocha|}
\fi.
\bibitem[Man95]{Mansharamani:1995:PCU} Rajesh~K. Mansharamani.
\newblock Parallel computing using the prefix problem.
\newblock {\em The Computer Journal}, 38\penalty0 (3)
:\penalty0 264a--265, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/264-
a.full.pdf+html|}\fi.

\bibitem[Man96]{Mannion:1996:BR} M.~Mannion. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (1):\penalty0 95, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/1/95.
1.full.pdf+html|}\fi.
\bibitem[Mar90]{Marcer:1990:C} P.~J. Marcer. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (1):\penalty0 91--92, ????   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/1/
91.full.pdf+html|}\fi.
\bibitem[Mar91a]{Marshall:1991:DNL} R.~Marshall. \newblock
Discourse and natural-language text analysis: functional
grammar   approach. \newblock {\em The Computer Journal}
, 34\penalty0 (6):\penalty0 A035--A060,   December 1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}.
\bibitem[Mar91b]{Marshall:1991:MFT} R.~Marshall. \newblock
Manipulating full-text scientific databases: a logic-
based   se\-man\-tico-prag\-matic approach. \newblock
{\em The Computer Journal}, 34\penalty0 (3):\penalty0
245--253, June   1991. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/3/245.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/246.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/247.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/248.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/249.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/250.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/251.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/252.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/253.tif|}\fi.
\bibitem[Mar92]{Mariani:1992:OOO} J.~A. Mariani. \newblock
{Oggetto}: An object oriented database layered on a triple

store. \newblock {\em The Computer Journal}, 35\penalty0
(2):\penalty0 108--118, April   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/2/108.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_02/Vol35_02.body.html#AbstractMariani|}
\fi.
\bibitem[Mar94a]{Marovac:1994:LAC} Nenad Marovac. \newblock
Link associated computation in {HyperNet}. \newblock {
\em The Computer Journal}, 37\penalty0 (2):\penalty0 88-
-95, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/2/88.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_02/
Vol37_02.body.html#AbstractMarovac|}\fi.
\bibitem[Mar94b]{Martin:1994:BRa} G.~R. Martin. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 232, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/232.
full.pdf+html|}\fi.
\bibitem[Mar94c]{Martin:1994:BRb} G.~R. Martin. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 355, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/355.
full.pdf+html|}\fi.
\bibitem[Mar96]{Martin:1996:BR} U.~Martin. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(6):\penalty0 558--559, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/6/558.
full.pdf+html|}\fi.
\bibitem[Mas92]{Massara:1992:BR} R.~E. Massara. \newblock
Book review. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 650, December   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)

, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
650.full.pdf+html|}\fi.
\bibitem[Mat94]{Matthews:1994:BR} S.~Matthews. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 720, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/720.
1.full.pdf+html|}\fi.
\bibitem[Mat98]{Matthews:1998:BR} B.~M. Matthews. \newblock
Book review. \newblock {\em The Computer Journal}, 41\
penalty0 (6):\penalty0 426--427, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/6/
426.full.pdf+html|}\fi.
\bibitem[May91]{Maybank:1991:BR} S.~Maybank. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 422, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/422.
1.full.pdf+html|}\fi.
\bibitem[MB90]{Mukkamala:1990:ESE} R.~Mukkamala and S.
~C. Bruell. \newblock Efficient schemes to evaluate transaction
performance in distributed   database systems. \newblock
{\em The Computer Journal}, 33\penalty0 (1):\penalty0
79--89,   February 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/1/79.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
01/tiff/79.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/80.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/81.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/83.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/84.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/86.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/87.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/88.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/89.tif|}\fi.
\bibitem[MB96]{Morrissey:1996:MDT} Joan~M. Morrissey and
W.~T. Bealor. \newblock Minimizing data transfers in distributed
query processing: a   comparative study and evaluation.
\newblock {\em The Computer Journal}, 39\penalty0 (8)
:\penalty0 675--687, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/8/675.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_08/390675.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_08/Vol39_
08.body.html#AbstractMorrissey|}\fi.
\bibitem[MC95]{Ma:1995:PAB} Ruey-Liang Ma and Chung-Ping
Chung. \newblock Periodic adaptive branch prediction and
its application in   superscalar processing in {Prolog}
. \newblock {\em The Computer Journal}, 38\penalty0 (
6):\penalty0 457--470, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/6/457.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_06/Vol38_06.body.html#AbstractMa|}\fi.
\bibitem[McC93]{McCallum:1993:SPS} Scott McCallum. \newblock
Solving polynomial strict inequalities using cylindrical
algebraic   decomposition. \newblock {\em The Computer
Journal}, 36\penalty0 (5):\penalty0 432--438,   October
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/5/432.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_05/Vol36_05.body.
html#AbstractMcCallum|}\fi.
\bibitem[MCC{\etalchar{+}}95]{Morrison:1995:EPL} Ron Morrison,
Richard C.~H. Connor, Quintin~I. Cutts, Vivienne~S. Dunstan,
and   Graham N.~C. Kirby. \newblock Exploiting persistent
linkage in software engineering environments. \newblock
{\em The Computer Journal}, 38\penalty0 (1):\penalty0

1--16, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/1/1.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_01/
Vol38_01.body.html#AbstractMorrison|}\fi.
\bibitem[McK94]{McKeag:1994:BR} R.~M. McKeag. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 149--150, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
149.full.pdf+html|}\fi.
\bibitem[MdlIDS99]{Mitchell:1999:BR} Richard Mitchell,
B.~de~la Iglesia, J.~C.~W. Debuse, and Mark Staples. \
newblock Book reviews. \newblock {\em The Computer Journal}
, 42\penalty0 (7):\penalty0 623--625, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/7/
623.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_07/420623.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_07/
pdf/420623.pdf|}\fi.
\bibitem[ME90]{Megson:1990:SAS} G.~M. Megson and D.~J.
Evans. \newblock A systolic array solution for the assignment
problem. \newblock {\em The Computer Journal}, 33\penalty0
(6):\penalty0 562--569,   December 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/562.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/562.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/563.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/564.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/565.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/566.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/567.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/568.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/569.tif|}\fi.

\bibitem[ME98]{Moon:1998:PIE} Soo-Mook Moon and Kemal
Ebcio{\u{g}}lu. \newblock The performance impact of exploiting
branch {ILP} with tree   representation of {ILP} code.
\newblock {\em The Computer Journal}, 41\penalty0 (1)
:\penalty0 26--44, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/1/26.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_01/Vol41_01.body.html#AbstractMoon|}\
fi.
\bibitem[Meg90]{Megson:1990:EDP} G.~M. Megson. \newblock
An efficient double pipe systolic array for matrix product.
\newblock {\em The Computer Journal}, 33\penalty0 (3)
:\penalty0 261--265, June   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/261.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/261.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/262.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/263.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/264.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/265.tif|}\fi.
\bibitem[Meg91]{Megson:1991:BRD} G.~M. Megson. \newblock
Book review: {D. Gassilloud and J. C. Grossetie (eds)
,   {\booktitle{Computing with Parallel Architectures
--- T. Node}}}. \newblock {\em The Computer Journal},
34\penalty0 (6):\penalty0 573b--573,   December 1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/6/573-b.full.pdf+html|}\fi.
\bibitem[Meg96]{Meghini:1996:LIM} Carlo Meghini. \newblock
Logical image modelling and retrieval. \newblock {\em
The Computer Journal}, 39\penalty0 (3):\penalty0 173-
-183, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/3/173.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_

03/390173.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_03/Vol39_03.body.html#AbstractMeghini|}
\fi.
\bibitem[Mel95]{Melham:1995:SIE} T.~F. Melham. \newblock
Special issue editorial: Higher order logic theorem proving
and its   applications. \newblock {\em The Computer Journal}
, 38\penalty0 (2):\penalty0 89--90, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/2/
89.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_02/Vol38_02.body.html#AbstractMelham|}
\fi.
\bibitem[Mer90a]{Merrett:1990:SOC} T.~H. Merrett. \newblock
Scoping is not orthogonal to completeness. \newblock {
\em The Computer Journal}, 33\penalty0 (4):\penalty0 381,
August   1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/381.
tif|}\fi.
\bibitem[Mer90b]{Merrett:1990:SNS} T.~H. Merrett. \newblock
Short note: Scoping is not orthogonal to completeness.
\newblock {\em The Computer Journal}, 33\penalty0 (4)
:\penalty0 381, ???? 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/4/381.full.pdf+html|}\fi.
\bibitem[Mer91]{Merry:1991:C} I.~W. Merry. \newblock Correspondence.
\newblock {\em The Computer Journal}, 34\penalty0 (3)
:\penalty0 264, ???? 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/3/264.full.pdf+html|}\fi.
\bibitem[Mer94]{Merali:1994:BR} Y.~Merali. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(5):\penalty0 472--473, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/37/5/472.
2.full.pdf+html|}\fi.
\bibitem[MES97]{Morris:1997:SBC} Derrick Morris, D.~Gareth
Evans, and Simon Schofield. \newblock Simulating the behaviour
of computer systems: co-simulation of   hardware\slash
software. \newblock {\em The Computer Journal}, 40\penalty0
(10):\penalty0 617--629, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/10/617.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_40/Issue_10/Vol40_10.body.html#AbstractMorris|}
\fi.
\bibitem[MG95]{Moore:1995:TUS} Simon~W. Moore and Brian~
T. Graham. \newblock Tagged up\slash down sorter --- {
A} hardware priority queue. \newblock {\em The Computer
Journal}, 38\penalty0 (9):\penalty0 695--703, ????   1995.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/9/695.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_38/Issue_09/Vol38_09.body.html#AbstractMoore|}
\fi.
\bibitem[MH94]{McFarlane:1994:ODR} I.~McFarlane and S.
~G. Hoggar. \newblock Optimal drivers for the `random'
iteration algorithm. \newblock {\em The Computer Journal}
, 37\penalty0 (7):\penalty0 629--640, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
629.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_07/Vol37_07.body.html#AbstractMcFarlane|}
\fi.
\bibitem[MH98]{Moreau:1998:ELH} Luc Moreau and Wendy Hall.
\newblock On the expressiveness of links in hypertext
systems. \newblock {\em The Computer Journal}, 41\penalty0
(7):\penalty0 459--473, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/7/459.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_07/Vol41_07.body.html#AbstractMoreau|}
\fi.

\bibitem[MHF93]{Millar:1993:MPU} R.~J. Millar, M.~Elizabeth~
C. Hull, and J.~H. Frazer. \newblock The {Millar} polyhedron
and its use in the construction of octrees. \newblock
{\em The Computer Journal}, 36\penalty0 (2):\penalty0
186--194, April   1993. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/2/186.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractMillar|}\fi.
\bibitem[MHLS93]{Mehrez:1993:ISP} Abraham Mehrez, Geoffry~
S. Howard, Yaakov Lugassi, and Peretz Shoval. \newblock
Information system planning and selection: a multiattribute
theoretic   approach. \newblock {\em The Computer Journal}
, 36\penalty0 (6):\penalty0 525--541,   December 1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/6/525.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_06/Vol36_06.body.html#AbstractMehrez|}
\fi.
\bibitem[Mic91]{Michie:1991:MML} D.~Michie. \newblock
Methodologies from machine learning in data analysis and
software. \newblock {\em The Computer Journal}, 34\penalty0
(6):\penalty0 559--565,   December 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/6/559.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/559.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/560.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/561.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/562.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/563.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/564.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/565.
tif|}\fi.
\bibitem[Mic97]{Mickle:1997:ECC} Marlin~H. Mickle. \newblock
An electrical circuit and computer system analogy. \newblock
{\em The Computer Journal}, 40\penalty0 (7):\penalty0
457--462, ????   1997. \newblock \showCODEN{CMPJA6}. \

newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/7/457.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_40/Issue_07/
Vol40_07.body.html#AbstractMickle|}\fi.
\bibitem[Mil95]{Mills:1995:LDW} Stella Mills. \newblock
To live or drown: when information systems become critical.
\newblock {\em The Computer Journal}, 38\penalty0 (6)
:\penalty0 413--417, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/6/413.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_06/Vol38_06.body.html#AbstractMills|}
\fi.
\bibitem[Mil96]{Mills:1996:BR} S.~Mills. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(10):\penalty0 891, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/10/891.
1.full.pdf+html|}\fi.
\bibitem[Mir98]{Mirkin:1998:LSS} Boris Mirkin. \newblock
Least-squares structuring, clustering and data processing
issues. \newblock {\em The Computer Journal}, 41\penalty0
(8):\penalty0 518--536, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/8/518.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_08/Vol41_08.body.html#AbstractMirkin|}
\fi.
\bibitem[Miy90]{Miyakoshi:1990:LCM} H.~Miyakoshi. \newblock
A literal code method for minimizing {Boolean} functions.
\newblock {\em The Computer Journal}, 33\penalty0 (4)
:\penalty0 375--380,   August 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/375.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/375.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/376.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

33/Issue_04/tiff/377.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/379.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/380.tif|}\fi.
\bibitem[MK94]{Ma:1994:GTT} Jixin Ma and Brian Knight.
\newblock A general temporal theory. \newblock {\em The
Computer Journal}, 37\penalty0 (2):\penalty0 114--123,
????   1994. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/2/114.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_37/Issue_02/Vol37_02.body.
html#AbstractMa|}\fi.
\bibitem[MK96]{Ma:1996:RTL} Jixin Ma and Brian Knight.
\newblock A reified temporal logic. \newblock {\em The
Computer Journal}, 39\penalty0 (9):\penalty0 800--807,
????   1996. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/39/9/800.full.pdf+html;   http://www.oup.
co.uk/computer_journal/Volume_39/Issue_09/Vol39_09.body.
html#AbstractMa;   http://www3.oup.co.uk/computer_journal/
Volume_39/Issue_09/Vol39_09.body.html#AbstractMa|}\fi.
\bibitem[MKP94]{Ma:1994:RTA} J.~Ma, B.~Knight, and M.
~Petridis. \newblock A revised theory of action and time
based on intervals and points. \newblock {\em The Computer
Journal}, 37\penalty0 (10):\penalty0 847--857, ????
1994. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/10/847.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_37/Issue_10/370858.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_37/Issue_10/tiff/847.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_37/Issue_10/tiff/848.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
37/Issue_10/tiff/849.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_10/tiff/850.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_37/Issue_
10/tiff/851.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_37/Issue_10/tiff/852.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_37/Issue_10/tiff/853.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
37/Issue_10/tiff/854.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_37/Issue_10/tiff/855.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_37/Issue_
10/tiff/856.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_37/Issue_10/ti.
\bibitem[MLR90]{Martin:1990:ESS} T.~P. Martin, K.~H. Lam,
and Judy~I. Russell. \newblock An evaluation of site selection
algorithms for distributed query   processing. \newblock
{\em The Computer Journal}, 33\penalty0 (1):\penalty0
61--70,   February 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/1/61.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
01/tiff/61.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/62.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/63.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/64.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/66.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/67.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/68.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/70.tif|}\fi.
\bibitem[MMG91]{Morgan:1991:WDM} K.~Morgan, R.~L. Morris,
and S.~Gibbs. \newblock When does a mouse become a rat?
or {\ldots{}comparing} performance   and preferences in
direct manipulation and command line environment. \newblock
{\em The Computer Journal}, 34\penalty0 (3):\penalty0
265--271, June   1991. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/3/265.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/265.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/266.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/267.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

34/Issue_03/tiff/268.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/270.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/271.tif|}\fi.
\bibitem[MMR{\etalchar{+}}98]{Mazzeo:1998:FSC} Antonino
Mazzeo, Nicola Mazzocca, Stefano Russo, Carlo Savy, and
Valeria   Vittorini. \newblock Formal specification of
concurrent systems: a structured approach. \newblock {
\em The Computer Journal}, 41\penalty0 (3):\penalty0 145-
-162, ????   1998. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/3/145.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_03/
Vol41_03.body.html#AbstractMazzeo|}\fi.
\bibitem[MN90]{Moltedo:1990:IBC} L.~Moltedo and A.~Noferini.
\newblock An implementation of $7$-bit character encoding
for standard computer   graphics environments. \newblock
{\em The Computer Journal}, 33\penalty0 (5):\penalty0
429--436,   October 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/5/429.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/429.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/430.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/431.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/432.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/433.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/434.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/435.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/436.
tif|}\fi.
\bibitem[Mof97a]{Moffat:1997:ESI} A.~Moffat. \newblock
Editorial for special issue on lossless compression. \
newblock {\em The Computer Journal}, 40\penalty0 (2--
3):\penalty0 65--66, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://www.oup.co.uk/computer_journal/Volume_40/

Issue_02/Vol40_02.body.html#AbstractEditorial;   http:
//www3.oup.co.uk/computer_journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractEditorial;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_02/Vol40_03.body.
html#AbstractEditorial|}\fi.
\bibitem[Mof97b]{Moffat:1997:LC} A.~Moffat. \newblock
Lossless compression. \newblock {\em The Computer Journal}
, 40\penalty0 (2--3):\penalty0 65--66, ????   1997. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/2_and_3/65.full.pdf+html|}\fi.
\bibitem[Moo92a]{Moody:1992:IFL} Ken Moody. \newblock
An introduction to formal languages and automata. \newblock
{\em The Computer Journal}, 35\penalty0 (2):\penalty0
176, April   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/2/176.full.pdf+html|}\fi.
\bibitem[Moo92b]{Moore:1992:PV} R.~Moore. \newblock Programming
with {VDM}. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 554, December   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
554.full.pdf+html|}\fi.
\bibitem[MP95]{Macias:1995:MCP} Benjamin Macias and Stephen~
G. Pulman. \newblock A method for controlling the production
of specifications in natural   language. \newblock {\
em The Computer Journal}, 38\penalty0 (4):\penalty0 310-
-318, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/4/310.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_04/
Vol38_04.body.html#AbstractMacias|}\fi.
\bibitem[MR94]{Mauw:1994:ASB} S.~Mauw and M.~A. Reniers.
\newblock An algebraic semantics of basic message sequence
charts. \newblock {\em The Computer Journal}, 37\penalty0
(4):\penalty0 269--277, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/4/269.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_04/Vol37_04.body.html#AbstractMauw|}\
fi.
\bibitem[MR96]{Mariani:1996:CIS} John~A. Mariani and Tom
Rodden. \newblock Cooperative information sharing: developing
a shared object service. \newblock {\em The Computer Journal}
, 39\penalty0 (6):\penalty0 455--470, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/6/
455.full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_06/390455.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_06/Vol39_
06.body.html#AbstractMariani|}\fi.
\bibitem[MRW90]{Mitchell:1990:SMA} C.~Mitchell, D.~Rush,
and M.~Walker. \newblock A secure messaging architecture
implementing the {X}.400-1988   security features. \newblock
{\em The Computer Journal}, 33\penalty0 (4):\penalty0
290--295,   August 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/4/290.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/290.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/291.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/292.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/293.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/294.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/295.tif|}\fi.
\bibitem[MS91]{Mancini:1991:FTR} L.~Mancini and S.~K.
Shrivastava. \newblock Fault-tolerant reference counting
for garbage collection in   distributed systems. \newblock
{\em The Computer Journal}, 34\penalty0 (6):\penalty0
503--513,   December 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/6/503.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/503.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_34/Issue_06/tiff/504.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/505.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/506.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/507.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/508.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/509.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/510.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/511.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/512.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/513.tif|}\fi.
\bibitem[MS92]{Manner:1992:FTD} R.~M{\"a}nner and O.~
Stucky. \newblock Fault-tolerant data transfer in a multiprocessor
system by forward   and backward hardware error recovery.
\newblock {\em The Computer Journal}, 35\penalty0 (4)
:\penalty0 361--368,   August 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/4/361.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_04/Vol35_04.body.html#AbstractManner|}
\fi.
\bibitem[MS97]{Muller:1997:TPL} Olaf M{\"u}ller and Konrad
Slind. \newblock Treating partiality in a logic of total
functions. \newblock {\em The Computer Journal}, 40\penalty0
(10):\penalty0 640--651, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/10/640.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_40/Issue_10/Vol40_10.body.html#AbstractMueller|}
\fi.
\bibitem[MSNW99]{Martin:1999:BTE} Keith~M. Martin, Rei
Safavi-Naini, and Huaxiong Wang. \newblock Bounds and
techniques for efficient redistribution of secret shares
 to new access structures. \newblock {\em The Computer
Journal}, 42\penalty0 (8):\penalty0 638--649, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/

42/8/638.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_08/420638.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_08/pdf/420638.pdf|}\fi.
\bibitem[MSS93]{Muchnick:1993:CIP} V.~B. Muchnick, A.
~V. Shafarenko, and C.~D. Sutton. \newblock {$F$}-code
and its implementation: a portable software platform for
 data parallelism. \newblock {\em The Computer Journal}
, 36\penalty0 (8):\penalty0 712--722, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/8/
712.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractMuchnick|}
\fi.
\bibitem[Mud94]{Mudur:1994:VTA} S.~P. Mudur. \newblock
Visual and technical aspects of type. \newblock {\em The
Computer Journal}, 37\penalty0 (6):\penalty0 549a--550,
????   1994. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/6/549-a.full.pdf+html|}\fi.
\bibitem[Muk94a]{Mukherjee:1994:BRa} P.~Mukherjee. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 226, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/226.
1.full.pdf+html|}\fi.
\bibitem[Muk94b]{Mukherjee:1994:BRb} P.~Mukherjee. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 226--227, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/3/
226.2.full.pdf+html|}\fi.
\bibitem[Muk94c]{Mukherjee:1994:BRc} P.~Mukherjee. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 349, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/37/4/349.
1.full.pdf+html|}\fi.
\bibitem[Muk94d]{Mukherjee:1994:BRd} P.~Mukherjee. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 349--350, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
349.2.full.pdf+html|}\fi.
\bibitem[Muk94e]{Mukherjee:1994:BRe} P.~Mukherjee. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (5):\penalty0 470, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/5/470.
1.full.pdf+html|}\fi.
\bibitem[Muk94f]{Mukherjee:1994:BRf} P.~Mukherjee. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 722, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/722.
1.full.pdf+html|}\fi.
\bibitem[Mul92]{Mullin:1992:HFH} J.~K. Mullin. \newblock
Hash functions for hash-based join methods. \newblock
{\em The Computer Journal}, 35\penalty0 (6):\penalty0
A499--A503,   December 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}.
\bibitem[Mul94]{Mullender:1994:BR} S.~J. Mullender. \
newblock Book review. \newblock {\em The Computer Journal}
, 37\penalty0 (8):\penalty0 727, ???? 1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
727.1.full.pdf+html|}\fi.
\bibitem[Mur93]{Murray:1993:BR} I.~R. Murray. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 294--295, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/36/3/
294.2.full.pdf+html|}\fi.
\bibitem[Mur94a]{Murakami:1994:BR} M.~Murakami. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 152, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/152.
2.full.pdf+html|}\fi.
\bibitem[Mur94b]{Murphy:1994:BRa} D.~Murphy. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 143--144, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
143.1.full.pdf+html|}\fi.
\bibitem[Mur94c]{Murphy:1994:BRb} D.~Murphy. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 143--144, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
143.2.full.pdf+html|}\fi.
\bibitem[Mur94d]{Murphy:1994:BR} D.~Murphy. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 143--144, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
143.3.full.pdf+html|}\fi.
\bibitem[Mur98a]{Murtagh:1998:CC} F.~Murtagh. \newblock
Clustering and classification. \newblock {\em The Computer
Journal}, 41\penalty0 (8):\penalty0 517, ???? 1998. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/8/517.full.pdf+html|}\fi.
\bibitem[Mur98b]{Murtagh:1998:FSI} F.~Murtagh. \newblock
Foreword to special issue on clustering and classification.
\newblock {\em The Computer Journal}, 41\penalty0 (8)
:\penalty0 517, ???? 1998. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)

}. \newblock \ifshowURL {\showURL   \path|http://www3.
oup.co.uk/computer_journal/Volume_41/Issue_08/Vol41_08.
body.html#AbstractEditorial|}\fi.
\bibitem[MVL92]{Misic:1992:FSD} V.~Mi{\v{s}}i{\'c}, D.
~Vela\v{s}evi{\'c}, and B.~Lazarevi{\'c}. \newblock Formal
specification of a data dictionary for an extended {ER}
data   model. \newblock {\em The Computer Journal}, 35\
penalty0 (6):\penalty0 611--622,   December 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/6/
611.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_06/Vol35_06.body.html#AbstractMisic|}
\fi. \newblock See \cite{Cooke:1992:EFMa}.
\bibitem[MW97]{Memon:1997:RDC} N.~Memon and X.~Wu. \newblock
Recent developments in context-based predictive techniques
for   lossless image compression. \newblock {\em The Computer
Journal}, 40\penalty0 (2--3):\penalty0 127--136,   ??
?? 1997. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/2_and_3/127.full.pdf+html;   http://www.
oup.co.uk/computer_journal/Volume_40/Issue_02/Vol40_02.
body.html#AbstractMemon;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/Vol40_03.body.html#AbstractMemon|}
\fi.
\bibitem[MWHC96]{Majewski:1996:FPH} Bohdan~S. Majewski,
Nicholas~C. Wormald, George Havas, and Zbigniew~J. Czech.
\newblock A family of perfect hashing methods. \newblock
{\em The Computer Journal}, 39\penalty0 (6):\penalty0
547--554, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/6/547.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
06/390547.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_06/Vol39_06.body.html#AbstractMajewski|}
\fi. \newblock This paper claims the discovery of order-
preserving perfect hashing   methods that run in linear
time.
\bibitem[MWR98]{Mitchell:1998:DAA} S.~E. Mitchell, A.
~J. Wellings, and A.~Romanovsky. \newblock Distributed
atomic actions in {Ada 95}. \newblock {\em The Computer

Journal}, 41\penalty0 (7):\penalty0 486--502, ????   1998.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/7/486.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_41/Issue_07/Vol41_07.body.html#AbstractMitchell|}
\fi.
\bibitem[Myc95]{Mycroft:1995:OOT} Alan Mycroft. \newblock
Object-oriented type systems. \newblock {\em The Computer
Journal}, 38\penalty0 (1):\penalty0 79a--80, ????   1995.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/1/79-a.full.pdf+html|}\fi.
\bibitem[MZ96]{McDonald:1996:AVP} A.~B. McDonald and T.
~Znati. \newblock {ARP} versus {ES-IS}: Performance evaluation
of neighbour-greeting   protocols. \newblock {\em The
Computer Journal}, 39\penalty0 (10):\penalty0 854--867,
????   1996. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/39/10/854.full.pdf+html;   http://www.oup.
co.uk/computer_journal/Volume_39/Issue_10/Vol39_10.body.
html#AbstractMcDonald;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_10/Vol39_10.body.html#AbstractMcDonald|}
\fi.
\bibitem[Nai97]{Naik:1997:DRT} Kshirasagar Naik. \newblock
Designing reliable test architectures for communication
protocols. \newblock {\em The Computer Journal}, 40\penalty0
(7):\penalty0 441--456, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/7/441.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_40/Issue_07/Vol40_07.body.html#AbstractNaik|}\
fi.
\bibitem[Nar96]{Narayanan:1996:CST} T.~S. Narayanan. \
newblock A class of semi-{$X$} tree-based dictionary machines.
\newblock {\em The Computer Journal}, 39\penalty0 (1)
:\penalty0 45--51, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/1/45.

full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_01/390045.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_01/Vol39_
01.body.html#AbstractNaryanan|}\fi.
\bibitem[Nau93]{Naur:1993:UTU} P.~Naur. \newblock Understanding
{Turing}'s universal machine --- personal style in   program
description. \newblock {\em The Computer Journal}, 36\
penalty0 (4):\penalty0 351--372,   August 1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/4/
351.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_04/Vol36_04.body.html#AbstractNaur|}\
fi.
\bibitem[Nau95]{Naur:1995:BR} P.~Naur. \newblock Book
review. \newblock {\em The Computer Journal}, 38\penalty0
(10):\penalty0 833--834, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/10/833.
full.pdf+html|}\fi.
\bibitem[NBL95]{Nie:1995:IRC} Jian-Yun Nie, Martin Brisebois,
and Fran{\c{c}}oise Lepage. \newblock Information retrieval
as counterfactual. \newblock {\em The Computer Journal}
, 38\penalty0 (8):\penalty0 643--657, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/8/
643.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_08/Vol38_08.body.html#AbstractNie|}\fi.
\bibitem[NCR97]{Nixon:1997:BPD} Paddy Nixon, Vinny Cahill,
and Fethi Rabhi. \newblock Building parallel and distributed
systems. \newblock {\em The Computer Journal}, 40\penalty0
(8):\penalty0 463--464, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/8/463.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_40/Issue_08/Vol40_08.body.html#AbstractEditorial|}
\fi.

\bibitem[NDD96]{Nash:1996:SSQ} Jonathan~M. Nash, Peter~
M. Dew, and Martin~E. Dyer. \newblock A scalable shared
queue on a distributed memory machine. \newblock {\em
The Computer Journal}, 39\penalty0 (6):\penalty0 483-
-495, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/6/483.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
06/390483.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_06/Vol39_06.body.html#AbstractNash|}
\fi.
\bibitem[New93]{Newton:1993:BR} M.~Newton. \newblock Book
review. \newblock {\em The Computer Journal}, 36\penalty0
(6):\penalty0 598, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/6/598.1.full.pdf+html|}
\fi.
\bibitem[NF94]{Nievergelt:1994:SIE} Jurg Nievergelt and
Mike Freeston. \newblock Special issue editorial: Spatial
data: applications, concepts,   techniques. \newblock
{\em The Computer Journal}, 37\penalty0 (1):\penalty0
1--2, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/1/1.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_01/
Vol37_01.body.html#AbstractNievergelt|}\fi.
\bibitem[Nip91]{Nipkow:1991:CR} T.~Nipkow. \newblock Constructive
rewriting. \newblock {\em The Computer Journal}, 34\penalty0
(1):\penalty0 34--41,   February 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/34.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/34.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/35.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/37.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/38.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_34/Issue_01/tiff/39.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/40.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/41.tif|}\fi.
\bibitem[NL90]{Ng:1990:CMP} K.~W. Ng and H.~F. Leung.
\newblock Competition: a model of {AND-Parallel} execution
of logic programs. \newblock {\em The Computer Journal}
, 33\penalty0 (3):\penalty0 215--218, June   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/3/
215.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/215.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/216.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/217.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/218.tif|}\fi.
\bibitem[NM90]{Neelamkavil:1990:SGA} F.~Neelamkavil and
O.~Mullarney. \newblock Separating graphics from application
in the design of user   interfaces. \newblock {\em The
Computer Journal}, 33\penalty0 (5):\penalty0 437--443,
 October 1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/5/437.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/437.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/438.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/439.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/440.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/441.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/442.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/443.tif|}\fi.
\bibitem[NMM91]{Naor:1991:PPQ} D.~Naor, C.~U. Martel,
and N.~S. Matloff. \newblock Performance of priority queue
structures in a virtual memory   environment. \newblock
{\em The Computer Journal}, 34\penalty0 (5):\penalty0
428--437,   October 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/5/428.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/428.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/429.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/430.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/431.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/432.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/433.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/434.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/435.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/436.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/437.tif|}\fi.
\bibitem[NMW97]{Nevill-Manning:1997:CEU} C.~G. Nevill-
Manning and I.~H. Witten. \newblock Compression and explanation
using hierarchical grammars. \newblock {\em The Computer
Journal}, 40\penalty0 (2--3):\penalty0 103--116,   ??
?? 1997. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/2_and_3/103.full.pdf+html;   http://www.
oup.co.uk/computer_journal/Volume_40/Issue_02/Vol40_02.
body.html#AbstractNevill;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/Vol40_02.body.html#AbstractNevill;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_03.body.html#AbstractNevill|}\fi.
\bibitem[NN97]{Novack:1997:RDL} Steven Novack and Alexandru
Nicolau. \newblock Resource-directed loop pipelining:
Exposing just enough parallelism. \newblock {\em The Computer
Journal}, 40\penalty0 (6):\penalty0 311--321, ????   1997.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
40/6/311.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_06/Vol40_06.body.html#AbstractNovack|}
\fi.
\bibitem[NO92a]{Neelamkavil:1992:IEP} F.~Neelamkavil and
S.~O'Shea. \newblock Interfacing {Eiffel} and {PCTE}.
\newblock {\em The Computer Journal}, 35\penalty0 (5)
:\penalty0 A439--??,   October 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://www3.oup.co.uk/computer_journal/Volume_35/
Issue_05/35000A439.sgm.abs.html;   http://www3.oup.co.
uk/computer_journal/Volume_35/Issue_05/tiff/A439.tif|}
\fi.
\bibitem[NO92b]{Ng:1992:SNS} D.~T.~H. Ng and B.~J. Oommen.
\newblock Short notes: a short note on doubly-linked list
re-organizing   heuristics. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 533--535,   October
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/533.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/350533.sgm.
abs.html; http://www3.oup.co.uk/computer_journal/Volume_
35/Issue_05/tiff/533.tif;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/tiff/534.tif;   http://www3.
oup.co.uk/computer_journal/Volume_35/Issue_05/tiff/535.
tif|}\fi.
\bibitem[Nob91]{Noble:1991:SWD} F.~Noble. \newblock Seven
ways to develop office systems: a managerial comparison
of   office system development methodologies. \newblock
{\em The Computer Journal}, 34\penalty0 (2):\penalty0
113--121, April   1991. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/2/113.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/113.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/114.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/115.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/117.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/118.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/119.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/120.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/121.tif|}\fi.
\bibitem[Nob92]{Noble:1992:OIS} Faith Noble. \newblock
Office information systems. \newblock {\em The Computer
Journal}, 35\penalty0 (1):\penalty0 70, February   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-

4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/1/70.full.pdf+html|}\fi.
\bibitem[Nor94]{Nori:1994:BR} K.~V. Nori. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(8):\penalty0 725--726, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/725.
2.full.pdf+html|}\fi.
\bibitem[Nor95a]{Nori:1995:BRa} K.~V. Nori. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (9):\penalty0 751--752, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/9/
751.full.pdf+html|}\fi.
\bibitem[Nor95b]{Nori:1995:BRb} K.~V. Nori. \newblock
Book review. \newblock {\em The Computer Journal}, 38\
penalty0 (10):\penalty0 832--833, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/10/
832.full.pdf+html|}\fi.
\bibitem[NP90]{Neal:1990:YMC} L.~R. Neal and M.~L.~V.
Pitteway. \newblock Yet more circle generators. \newblock
{\em The Computer Journal}, 33\penalty0 (5):\penalty0
408--411,   October 1990. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/5/408.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/408.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/409.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/410.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/411.tif|}\fi.
\bibitem[NR92]{Nevalainen:1992:IHS} O.~Nevalainen and
T.~Raita. \newblock An internal hybrid sort algorithm
revisited. \newblock {\em The Computer Journal}, 35\penalty0
(2):\penalty0 177--183, April   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/2/177.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_02/Vol35_02.body.html#AbstractNevalainen|}
\fi.
\bibitem[NW93]{Nievergelt:1993:GFS} J.~Nievergelt and
P.~Widmayer. \newblock Guard files: Stabbing and intersection
queries on fat spatial   objects. \newblock {\em The Computer
Journal}, 36\penalty0 (2):\penalty0 107--116, April
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/2/107.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_02/Vol36_02.body.
html#AbstractNievergelt|}\fi.
\bibitem[Obe92]{Obeid:1992:NRL} N.~Obeid. \newblock Nonmonotonic
reasoning: Logical foundation of commonsense. \newblock
{\em The Computer Journal}, 35\penalty0 (2):\penalty0
147, April   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/2/147.full.pdf+html|}\fi.
\bibitem[OD91]{ODocherty:1991:MIS} M.~H. O'Docherty and
C.~N. Daskalakis. \newblock Multimedia information systems
--- the management and semantic   retrieval of all electronic
data types. \newblock {\em The Computer Journal}, 34\
penalty0 (3):\penalty0 225--238, June   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/3/
225.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/225.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/226.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/227.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/228.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/229.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/230.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/231.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/232.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_03/tiff/233.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/234.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/235.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/236.
t.
\bibitem[OF93]{Oommen:1993:FLA} B.~John Oommen and Chris
Fothergill. \newblock Fast learning automaton-based image
examination and retrieval. \newblock {\em The Computer
Journal}, 36\penalty0 (6):\penalty0 542--553,   December
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/6/542.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_06/Vol36_06.body.
html#AbstractOommen|}\fi.
\bibitem[Oka98]{Oka:1998:SMC} Ryuichi Oka. \newblock Spotting
method for classification of real world data. \newblock
{\em The Computer Journal}, 41\penalty0 (8):\penalty0
559--565, ????   1998. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/8/559.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_08/
Vol41_08.body.html#AbstractOka|}\fi.
\bibitem[OKMS96]{Ould-Khaoua:1996:CEH} M.~Ould-Khaoua,
L.~M. MacKenzie, and R.~Sotudeh. \newblock Comparative
evaluation of hypermesh and multi-stage interconnection
 networks. \newblock {\em The Computer Journal}, 39\penalty0
(3):\penalty0 232--240, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/3/232.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_03/390232.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_03/Vol39_
03.body.html#AbstractOuld|}\fi.
\bibitem[Ole91]{Olerup:1991:DAC} A.~Olerup. \newblock
Design approaches: a comparative study of information
system design   and architectural design. \newblock {
\em The Computer Journal}, 34\penalty0 (3):\penalty0 215-
-224, June   1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.

\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/3/215.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/215.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/216.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/217.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/218.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/220.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/221.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/222.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/224.tif|}\fi.
\bibitem[OM91]{Oommen:1991:SAS} B.~J. Oommen and D.~C.
~Y. Ma. \newblock Stochastic automata solutions to the
object partitioning problem. \newblock {\em The Computer
Journal}, 34\penalty0 (6):\penalty0 A105--A120,   December
1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}.
\bibitem[ON90]{Oommen:1990:GRP} B.~J. Oommen and D.~T.
~H. Ng. \newblock On generating random permutations with
arbitrary distributions. \newblock {\em The Computer Journal}
, 33\penalty0 (4):\penalty0 368--374,   August 1990. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/4/368.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/368.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/369.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/370.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/371.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/372.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/373.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/374.tif|}\fi.
\bibitem[O'N92]{ONeill:1992:SNA} G.~O'Neill. \newblock
Short note: Automatic translation of {VDM} specifications
into   {Standard ML} programs. \newblock {\em The Computer

Journal}, 35\penalty0 (6):\penalty0 623--624,   December
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/6/623.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_35/Issue_06/tiff/623.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
35/Issue_06/tiff/624.tif|}\fi.
\bibitem[ON94]{Oommen:1994:NTE} B.~J. Oommen and D.~T.
~H. Ng. \newblock A new technique for enhancing linked-
list data retrieval: Reorganize   data using artificially
synthesized queries. \newblock {\em The Computer Journal}
, 37\penalty0 (7):\penalty0 598--609, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
598.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_07/Vol37_07.body.html#AbstractOommen|}
\fi.
\bibitem[Ong98]{Ong:1998:BR} C.-H.~L. Ong. \newblock Book
review. \newblock {\em The Computer Journal}, 41\penalty0
(1):\penalty0 65--66, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/1/65.
full.pdf+html|}\fi.
\bibitem[Oom93]{Oommen:1993:TIC} B.~J. Oommen. \newblock
Transforming ill-conditioned constrained problems using
projections. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 282--285, June   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/3/
282.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_03/Vol36_03.body.html#AbstractOommen|}
\fi.
\bibitem[OOW91]{Olariu:1991:MDE} S.~Olariu, C.~M. Overstreet,
and Z.~Wenn. \newblock A mergeable double-ended priority
queue. \newblock {\em The Computer Journal}, 34\penalty0
(5):\penalty0 423--427,   October 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/423.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/423.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/424.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/425.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/426.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/427.tif|}\fi.
\bibitem[Orm91]{Orman:1991:CMD} L.~Orman. \newblock Constraint
maintenance as a data model design criterion. \newblock
{\em The Computer Journal}, 34\penalty0 (1):\penalty0
73--79,   February 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/1/73.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
01/tiff/73.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/75.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/76.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/78.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/79.tif|}\fi.
\bibitem[O'S94]{OShea:1994:SVV} Greg O'Shea. \newblock
On the specification, validation and verification of security
in   access control systems. \newblock {\em The Computer
Journal}, 37\penalty0 (5):\penalty0 437--448, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/5/437.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/Vol37_05.body.html#AbstractO'
Shea|}\fi.
\bibitem[Osb91a]{Osborne:1991:BRE} Chris Osborne. \newblock
Book review: {Ernest S. Colantonio, {\booktitle{A Laboratory
Course   in DOS, Wordperfect 5.0, Lotus 1-2-3, dBASE IV}
}}. \newblock {\em The Computer Journal}, 34\penalty0
(6):\penalty0 573a--573,   December 1991. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/34/6/
573-a.full.pdf+html|}\fi.
\bibitem[Osb91b]{Osborne:1991:BRH} Chris Osborne. \newblock
Book review: {Helene G. Kershner, \booktitle{Introduction
to Computer   Literacy}}. \newblock {\em The Computer
Journal}, 34\penalty0 (6):\penalty0 573, December   1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/6/573.full.pdf+html|}\fi.
\bibitem[Osb93]{Osborne:1993:PIL} Chris Osborne. \newblock
Planning for {IT} literacy in an institution of higher
education: a   case study. \newblock {\em The Computer
Journal}, 36\penalty0 (4):\penalty0 343--350,   August
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/4/343.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_04/Vol36_04.body.
html#AbstractOsborne|}\fi.
\bibitem[OT91]{Ou:1991:HSU} Seng~Fat Ou and Alan~L. Tharp.
\newblock High storage utilisation for single-probe retrieval
linear hashing. \newblock {\em The Computer Journal},
34\penalty0 (5):\penalty0 455--468,   October 1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/5/455.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/455.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/456.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/457.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/458.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/459.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/460.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/461.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/462.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/463.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/464.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/465.tif;   http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/466.t.
\bibitem[OV92]{Oman:1992:UBS} P.~W. Oman and K.~{Van Houten}
. \newblock On the use of binary strings to represent
tree structures. \newblock {\em The Computer Journal}
, 35\penalty0 (5):\penalty0 A463--??,   October 1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/35000A463.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/Volume_35/Issue_05/tiff/A463.
tif|}\fi.
\bibitem[Oxb91]{Oxborrow:1991:BR} E.~Oxborrow. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 468, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/468.
1.full.pdf+html|}\fi.
\bibitem[PA90]{Philbrow:1990:EEH} P.~C. Philbrow and M.
~P. Atkinson. \newblock Events and exception handling
in {PS}-algol. \newblock {\em The Computer Journal}, 33\
penalty0 (2):\penalty0 108--125, April   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/2/
108.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/108.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/109.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/110.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/111.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/113.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/114.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/115.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/116.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/117.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/118.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/119.
t.
\bibitem[Paa91]{Paakki:1991:PPC} J.~Paakki. \newblock
{Prolog} in practical compiler writing. \newblock {\em
The Computer Journal}, 34\penalty0 (1):\penalty0 64--
72,   February 1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/1/64.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
01/tiff/64.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/65.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/66.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/67.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/68.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/69.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/71.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/72.tif|}\fi.
\bibitem[Pad91]{Padawitz:1991:RNH} P.~Padawitz. \newblock
Reduction and narrowing for {Horn} clause theories. \
newblock {\em The Computer Journal}, 34\penalty0 (1):
\penalty0 42--51,   February 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/42.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/43.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/45.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/46.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/48.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_01/tiff/49.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/50.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/

tiff/51.tif|}\fi.
\bibitem[Pai99]{Paige:1999:IPD} Richard~F. Paige. \newblock
Integrating a program design calculus and a subset of
{UML}. \newblock {\em The Computer Journal}, 42\penalty0
(2):\penalty0 82--99, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/2/82.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_02/420082.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_02/
pdf/420082.pdf|}\fi.
\bibitem[Pal90]{Palmer:1990:C} S.~W. Palmer. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (2):\penalty0 188, ???? 1990. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/2/
188.4.full.pdf+html|}\fi.
\bibitem[Par92]{Parhami:1992:SNR} B.~Parhami. \newblock
Systolic number radix converters. \newblock {\em The Computer
Journal}, 35\penalty0 (4):\penalty0 405--409,   August
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/4/405.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_04/Vol35_04.body.
html#AbstractParhami|}\fi.
\bibitem[Par93]{Parkin:1993:BR} G.~I. Parkin. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 292, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/3/292.
1.full.pdf+html|}\fi.
\bibitem[Par94]{Parberry:1994:LPC} Ian Parberry. \newblock
Lectures on parallel computation. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 556a--557, ????
1994. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/6/556-a.full.pdf+html|}\fi.

\bibitem[Par96a]{Parhami:1996:EVS} Behrooz Parhami. \
newblock Extreme-value search and general selection algorithms
for fully   parallel associative memories. \newblock {
\em The Computer Journal}, 39\penalty0 (3):\penalty0 241-
-250, ????   1996. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/3/241.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
03/390241.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_03/Vol39_03.body.html#AbstractParhami|}
\fi.
\bibitem[Par96b]{Parhami:1996:PTV} Behrooz Parhami. \
newblock Parallel threshold voting. \newblock {\em The
Computer Journal}, 39\penalty0 (8):\penalty0 692--700,
????   1996. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/39/8/692.full.pdf+html;   http://www.oup.
co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_08/390692.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
Volume_39/Issue_08/Vol39_08.body.html#AbstractParhami|}
\fi.
\bibitem[Pau92a]{Paulson:1992:LF} L.~Paulson. \newblock
Logical frameworks. \newblock {\em The Computer Journal}
, 35\penalty0 (2):\penalty0 156, April   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/2/
156.full.pdf+html|}\fi.
\bibitem[Pau92b]{Paulson:1992:FPH} L.~C. Paulson. \newblock
Functional programming with hope. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 491, October   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/5/491.full.pdf+html|}\fi.
\bibitem[Pau98]{Paulson:1998:BR} L.~C. Paulson. \newblock
Book review. \newblock {\em The Computer Journal}, 41\
penalty0 (6):\penalty0 425, ???? 1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/41/6/425.
full.pdf+html|}\fi.
\bibitem[PBF91]{Padget:1991:COO} J.~Padget, R.~Bradford,
and J.~Fitch. \newblock Concurrent object-oriented programming
in {Lisp}. \newblock {\em The Computer Journal}, 34\penalty0
(4):\penalty0 311--319,   August 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/311.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/311.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/312.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/315.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/316.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/317.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/318.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/319.tif|}\fi.
\bibitem[PBR{\etalchar{+}}90]{Plata:1990:ASP} O.~G. Plata,
J.~D. Bruguera, F.~F. Rivera, R.~Doallo, and E.~L. Zapata.
\newblock {ACLE}: a software package for {SIMD} computer
simulation. \newblock {\em The Computer Journal}, 33\
penalty0 (3):\penalty0 194--203, June   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/3/
194.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/194.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/195.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/196.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/197.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/198.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/199.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/200.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/201.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/202.tif;   http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/203.tif|}\fi.
\bibitem[PD96a]{Pan:1996:FBR} Si~Pan and R.~Geoff Dromey.
\newblock A formal basis for removing {\tt goto} statements.
\newblock {\em The Computer Journal}, 39\penalty0 (3)
:\penalty0 203--214, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/3/203.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_03/390203.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_03/Vol39_
03.body.html#AbstractPan2|}\fi.
\bibitem[PD96b]{Pan:1996:REL} Si~Pan and R.~Geoff Dromey.
\newblock Re-engineering loops. \newblock {\em The Computer
Journal}, 39\penalty0 (3):\penalty0 184--202, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/3/184.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_03/390184.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_03/
Vol39_03.body.html#AbstractPan1|}\fi.
\bibitem[Pel95]{Peled:1995:TLM} D.~Peled. \newblock Temporal
logic: Mathematical foundations and computational aspects,
 volume 1. \newblock {\em The Computer Journal}, 38\penalty0
(3):\penalty0 260--261, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/260.
full.pdf+html|}\fi.
\bibitem[Per94]{Perrott:1994:BR} R.~Perrott. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 147, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/147.
full.pdf+html|}\fi.
\bibitem[PG90]{Paton:1990:OED} N.~W. Paton and P.~M.~
D. Gray. \newblock Optimising and executing {DAPLEX} queries
using {Prolog}. \newblock {\em The Computer Journal},
33\penalty0 (6):\penalty0 547--555,   December 1990. \

newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/6/547.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/547.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/548.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/549.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/550.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/551.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/552.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/553.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/554.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/555.
tif|}\fi.
\bibitem[PGY97]{Patrick:1997:OCB} D.~G. Patrick, P.~R.
Green, and T.~A. York. \newblock Occam- and {C}-based
multiprocessor environments for {Unix} clusters. \newblock
{\em The Computer Journal}, 40\penalty0 (1):\penalty0
12--21, ????   1997. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/1/12.full.pdf+html|}\fi.
\bibitem[Pha90]{Pham:1990:ACC} B.~Pham. \newblock Algorithms
for calculating cubic {Beta-splines}. \newblock {\em The
Computer Journal}, 33\penalty0 (5):\penalty0 412--416,
 October 1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/5/412.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/412.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/413.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/414.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/415.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/416.tif|}\fi.
\bibitem[Phi94]{Philippou:1994:APP} A.~Philippou. \newblock
The art of parallel programming {\&} parallel programming,
an   introduction. \newblock {\em The Computer Journal}

, 37\penalty0 (6):\penalty0 546--547, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
546.full.pdf+html|}\fi.
\bibitem[PHS96]{Pan:1996:SDM} Yi~Pan, Mounir Hamdi, and
Gurdip Singh. \newblock Selection on $k$-dimensional meshes
with multiple broadcasting. \newblock {\em The Computer
Journal}, 39\penalty0 (2):\penalty0 140--144, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/2/140.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_02/390140.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_02/
Vol39_02.body.html#AbstractHamdi|}\fi.
\bibitem[Pie97]{Pierce:1997:BR} B.~C. Pierce. \newblock
Book review. \newblock {\em The Computer Journal}, 40\
penalty0 (5):\penalty0 297--298, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/5/
297.full.pdf+html|}\fi.
\bibitem[Pit91]{Pitteway:1991:BR} M.~L.~V. Pitteway. \
newblock Book review. \newblock {\em The Computer Journal}
, 34\penalty0 (4):\penalty0 357, ???? 1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/4/
357.full.pdf+html|}\fi.
\bibitem[PL94a]{Poix:1994:GRT} Christophe Poix and Robert
Laurini. \newblock A geometric reasoning tool based on
a successive improvement   approach. \newblock {\em The
Computer Journal}, 37\penalty0 (5):\penalty0 377--384,
????   1994. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/5/377.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_37/Issue_05/Vol37_05.body.
html#AbstractPoix|}\fi.
\bibitem[PL94b]{Poo:1994:OOS} Danny C.~C. Poo and Shwu-
Yi Lee. \newblock An object-oriented systems modelling

method based on the {Jackson}   approach. \newblock {
\em The Computer Journal}, 37\penalty0 (8):\penalty0 669-
-682, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/8/669.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_08/
Vol37_08.body.html#AbstractPoo|}\fi.
\bibitem[PM90]{Pan:1990:SCB} Y.~Pan and R.~Melhem. \newblock
Short circuits in buffered multi-stage interconnection
networks. \newblock {\em The Computer Journal}, 33\penalty0
(4):\penalty0 323--329,   August 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/4/323.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/323.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/324.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_04/tiff/325.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_04/tiff/326.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
04/tiff/327.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/328.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/329.
tif|}\fi.
\bibitem[PMC90]{Park:1990:C} J.~Won Park, S.~R. Maeng,
and J.~Wan Cho. \newblock Correspondence. \newblock {
\em The Computer Journal}, 33\penalty0 (4):\penalty0 382,
???? 1990. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/33/4/382.full.pdf+html|}\fi.
\bibitem[Pon91]{Pong:1991:PIG} M.-C. Pong. \newblock {
I}-pigs: an interactive graphical environment for concurrent
 programming. \newblock {\em The Computer Journal}, 34\
penalty0 (4):\penalty0 320--330,   August 1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/4/
320.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/320.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/321.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/322.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/323.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/324.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/325.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/326.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/327.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/328.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/329.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/330.tif|}\fi.
\bibitem[Poo91]{Poo:1991:RBP} C.-C.~D. Poo. \newblock
Representing business policies in the {Jackson} system
development   method. \newblock {\em The Computer Journal}
, 34\penalty0 (2):\penalty0 122--131, April   1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/2/122.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/122.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/123.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/124.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/125.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/126.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/127.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/128.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/129.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/130.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/131.tif|}\fi.
\bibitem[Pou92]{Poulovassilis:1992:IFF} A.~Poulovassilis.
\newblock The implementation of {FDL}, a functional database
language. \newblock {\em The Computer Journal}, 35\penalty0
(2):\penalty0 119--128, April   1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/2/119.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

Volume_35/Issue_02/Vol35_02.body.html#AbstractPoulovassilis|}
\fi.
\bibitem[Pou93]{Poulovassilis:1993:PMA} A.~Poulovassilis.
\newblock A pattern-matching algorithm for functional
databases. \newblock {\em The Computer Journal}, 36\penalty0
(2):\penalty0 195--199, April   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/2/195.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_02/Vol36_02.body.html#AbstractPoulovassilis|}
\fi.
\bibitem[PP98]{Panholzer:1998:TMP} Alois Panholzer and
Helmut Prodinger. \newblock Towards a more precise analysis
of an algorithm to generate binary   trees: a tutorial.
\newblock {\em The Computer Journal}, 41\penalty0 (3)
:\penalty0 201--204, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/3/201.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_03/Vol41_03.body.html#AbstractPanholzer|}
\fi.
\bibitem[PR98]{Prasse:1998:WCT} Michael Prasse and Peter
Rittgen. \newblock Why {Church}'s {Thesis} still holds.
some notes on {Peter Wegner}'s   tracts on interaction
and computability. \newblock {\em The Computer Journal}
, 41\penalty0 (6):\penalty0 357--362, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/6/
357.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_06/Vol41_06.body.html#AbstractPrasse|}
\fi.
\bibitem[Pri95]{Priami:1995:SC} Corrado Priami. \newblock
Stochastic $\pi$-calculus. \newblock {\em The Computer
Journal}, 38\penalty0 (7):\penalty0 578--589, ????   1995.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/7/578.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_38/Issue_07/Vol38_07.body.html#AbstractPriami|}
\fi.

\bibitem[PRLW93]{Pliskin:1993:PVA} N.~Pliskin, T.~Romm,
A.~S. Lee, and Y.~Weber. \newblock Presumed versus actual
organizational culture: managerial   implications for
implementation of information systems. \newblock {\em
The Computer Journal}, 36\penalty0 (2):\penalty0 143-
-152, April   1993. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/2/143.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_02/
Vol36_02.body.html#AbstractPliskin|}\fi.
\bibitem[PRS96]{Puliafito:1996:EPP} A.~Puliafito, S.~
Riccobene, and M.~Scarpa. \newblock Evaluation of performability
parameters in client-server   environments. \newblock
{\em The Computer Journal}, 39\penalty0 (8):\penalty0
647--662, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/8/647.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
08/390647.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_08/Vol39_08.body.html#AbstractPuliafito|}
\fi.
\bibitem[PS90]{Price:1990:SPC} Camille~C. Price and M.
~A. Salama. \newblock Scheduling of precedence-constrained
tasks on multiprocessors. \newblock {\em The Computer
Journal}, 33\penalty0 (3):\penalty0 219--229, June   1990.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/3/219.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/219.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/220.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/221.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/222.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/226.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/227.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/228.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/229.tif|}\fi.
\bibitem[PS92]{Paschos:1992:EEC} V.~T.~H. Paschos and
A.~N. Stafylopatis. \newblock Evaluation of the execution
cost of recursive definitions. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 A429--??,   October
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/35000A429.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_05/
tiff/A429.tif|}\fi.
\bibitem[PS93]{Pemberton:1993:EMI} Jonathan~D. Pemberton
and Joanne~M. Smailes. \newblock Educating the masses
--- is {IT} a tricky business? \newblock {\em The Computer
Journal}, 36\penalty0 (4):\penalty0 335--342,   August
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/4/335.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_04/Vol36_04.body.
html#AbstractPemberton|}\fi.
\bibitem[PS96]{Peng:1996:APD} Shietung Peng and Stanislav
Sedukhin. \newblock Array processors design for division-
free linear system solving. \newblock {\em The Computer
Journal}, 39\penalty0 (8):\penalty0 713--722, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/8/713.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_08/390713.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_08/
Vol39_08.body.html#AbstractPeng|}\fi.
\bibitem[PSK96]{Paakki:1996:AHS} Jukka Paakki, Airi Salminen,
and Jussi Koskinen. \newblock Automated hypertext support
for software maintenance. \newblock {\em The Computer
Journal}, 39\penalty0 (7):\penalty0 577--597, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/7/577.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_07/390577.sgm.abs.html;   http:

//www3.oup.co.uk/computer_journal/Volume_39/Issue_07/
Vol39_07.body.html#AbstractPaakki|}\fi.
\bibitem[PT90]{Puliti:1990:ITD} P.~Puliti and G.~Tascini.
\newblock Interpreting technical drawings. \newblock {
\em The Computer Journal}, 33\penalty0 (5):\penalty0 417-
-423,   October 1990. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/5/417.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/417.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/418.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/419.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/420.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/421.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/422.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/423.tif|}\fi.
\bibitem[PTV97]{Pan:1997:SEA} Yi~Pan, Jerry~L. Trathan,
and Ramachandran Vaidyanathan. \newblock A scalable and
efficient algorithm for computing the city block   distance
transform on reconfigurable meshes. \newblock {\em The
Computer Journal}, 40\penalty0 (7):\penalty0 435--440,
????   1997. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/7/435.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_07/Vol40_07.body.
html#AbstractPan|}\fi.
\bibitem[PW90]{Petre:1990:LMP} Marian Petre and R.~Winder.
\newblock On languages, models and programming styles.
\newblock {\em The Computer Journal}, 33\penalty0 (2)
:\penalty0 173--180, April   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/173.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/173.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/174.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/176.tif;   http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/177.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/178.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/179.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/180.tif|}\fi.
\bibitem[PW93]{Pavey:1993:DES} D.~J. Pavey and L.~A. Winsborrow.
\newblock Demonstrating equivalence of source code and
{PROM} contents. \newblock {\em The Computer Journal}
, 36\penalty0 (7):\penalty0 654--667, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/7/
654.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_07/Vol36_07.body.html#AbstractPavey|}
\fi.
\bibitem[PW94]{Patterson:1994:CAA} J.~W. Patterson and
P.~J. Willis. \newblock Computer assisted animation: {
$2$D} or not {$2$D}? \newblock {\em The Computer Journal}
, 37\penalty0 (10):\penalty0 829--839, ????   1994. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/10/829.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_10/Vol37_10.body.html#AbstractPatterson|}
\fi.
\bibitem[Qui92]{Quinlan:1992:NNP} P.~Quinlan. \newblock
Neural network parallel computing. \newblock {\em The
Computer Journal}, 35\penalty0 (5):\penalty0 450, October
 1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/450.full.pdf+html|}\fi.
\bibitem[QW96]{Qian:1996:EFC} Haifeng Qian and Jie Wu.
\newblock Enhanced {Fibonacci} cubes. \newblock {\em The
Computer Journal}, 39\penalty0 (4):\penalty0 331--345,
????   1996. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/39/4/331.full.pdf+html;   http://www.oup.
co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_04/390331.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/

Volume_39/Issue_04/Vol39_04.body.html#AbstractQian|}\
fi.
\bibitem[Rag94]{Raghunathan:1994:POS} V.~Raghunathan.
\newblock Prototyping-oriented software developmen. \
newblock {\em The Computer Journal}, 37\penalty0 (6):
\penalty0 548, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/548.full.pdf+html|}\fi.
\bibitem[Ram94a]{Ramani:1994:BR} S.~Ramani. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (5):\penalty0 467--468, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/5/
467.2.full.pdf+html|}\fi.
\bibitem[Ram94b]{Ramanujam:1994:BRa} R.~Ramanujam. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 223, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/223.
2.full.pdf+html|}\fi.
\bibitem[Ram94c]{Ramanujam:1994:BRb} R.~Ramanujam. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 346--347, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
346.2.full.pdf+html|}\fi.
\bibitem[Ram94d]{Ramesh:1994:BR} S.~Ramesh. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 151, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/151.
full.pdf+html|}\fi.
\bibitem[Ram95]{Ramesh:1995:PAF} S.~Ramesh. \newblock
A programming approach to formal methods. \newblock {
\em The Computer Journal}, 38\penalty0 (3):\penalty0 261-
-262, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock

\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/3/261.full.pdf+html|}\fi.
\bibitem[Ram99]{Ramsey:1999:ESE} Norman Ramsey. \newblock
Eliminating spurious error messages using exceptions,
polymorphism,   and higher-order functions. \newblock
{\em The Computer Journal}, 42\penalty0 (5):\penalty0
360--372, ????   1999. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/5/360.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
05/420360.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_05/Appendix.html|}\fi.
\bibitem[Ran94]{Randell:1994:ILP} D.~A. Randell. \newblock
Inductive logic programming. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 547a--548, ????
1994. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/6/547-a.full.pdf+html|}\fi.
\bibitem[Rao94a]{Rao:1994:BRaa} P.~V.~S. Rao. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (3):\penalty0 230--231, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/3/
230.full.pdf+html|}\fi.
\bibitem[Rao94b]{Rao:1994:BRbb} P.~V.~S. Rao. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (4):\penalty0 353--354, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/4/
353.full.pdf+html|}\fi.
\bibitem[Rao94c]{Rao:1994:BR} P.~V.~S. Rao. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 732--734, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/37/8/
732.2.full.pdf+html|}\fi.
\bibitem[RBCM97]{Rakotonirainy:1997:DOD} Andry Rakotonirainy,
Andrew Berry, Stephen Crawley, and Zoran Milosevic. \
newblock Describing open distributed systems: a foundation.
\newblock {\em The Computer Journal}, 40\penalty0 (8)
:\penalty0 479--488, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/8/479.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_40/Issue_08/Vol40_08.body.html#AbstractRakotonirainy|}
\fi.
\bibitem[RDPS95]{Ruiz-Delgado:1995:ROO} Antonio Ruiz-
Delgado, David Pitt, and Colin Smythe. \newblock A review
of object-oriented approaches in formal methods. \newblock
{\em The Computer Journal}, 38\penalty0 (10):\penalty0
777--784, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/10/777.full.pdf+html;
http://www3.oup.co.uk/computer_journal/Volume_38/Issue_
10/Vol38_10.body.html#AbstractDelgado|}\fi.
\bibitem[Ret95]{Rettelbach:1995:PBM} Michael Rettelbach.
\newblock Probabilistic branching in {Markovian} process
algebras. \newblock {\em The Computer Journal}, 38\penalty0
(7):\penalty0 590--599, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/7/590.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_07/Vol38_07.body.html#AbstractRettelbach|}
\fi.
\bibitem[Rey90]{Reynolds:1990:CAI} C.~F. Reynolds. \newblock
{CODIL}: The architecture of an information language.
\newblock {\em The Computer Journal}, 33\penalty0 (2)
:\penalty0 155--163, April   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/2/155.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/155.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/156.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/158.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/160.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/161.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/163.tif|}\fi.
\bibitem[RG95]{Reiter:1995:SCR} Michael Reiter and Li~
Gong. \newblock Securing causal relationships in distributed
systems. \newblock {\em The Computer Journal}, 38\penalty0
(8):\penalty0 633--642, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/8/633.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_08/Vol38_08.body.html#AbstractReiter|}
\fi.
\bibitem[RHHR98]{Robinson:1998:PSD} Hugh Robinson, Pat
Hall, Fiona Hovenden, and Janet Rachel. \newblock Postmodern
software development. \newblock {\em The Computer Journal}
, 41\penalty0 (6):\penalty0 363--375, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/6/
363.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_06/Vol41_06.body.html#AbstractRobinson|}
\fi.
\bibitem[Rho95]{Rhodes:1995:EFM} Paul~C. Rhodes. \newblock
Essentials of fuzzy modelling and control. \newblock {
\em The Computer Journal}, 38\penalty0 (1):\penalty0 86-
-87, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/1/86.full.pdf+html|}\fi.
\bibitem[Rib95]{Ribaudo:1995:ATS} Marina Ribaudo. \newblock
On the aggregation techniques in stochastic {Petri} nets
and   stochastic process algebras. \newblock {\em The
Computer Journal}, 38\penalty0 (7):\penalty0 600--611,
????   1995. \newblock \showCODEN{CMPJA6}. \newblock \

showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/7/600.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_38/Issue_07/Vol38_07.body.
html#AbstractRibaudo|}\fi.
\bibitem[Ris99a]{Rissanen:1999:HST} J.~Rissanen. \newblock
Hypothesis selection and testing by the {MDL} principle.
\newblock {\em The Computer Journal}, 42\penalty0 (4)
:\penalty0 260--269, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/4/260.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420260.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420260.pdf|}\fi.
\bibitem[Ris99b]{Rissanen:1999:DPM} Jorma Rissanen. \
newblock Discussion of paper {`Minimum Message Length
and Kolmogorov   Complexity' by C. S. Wallace and D. L.
Dowe}. \newblock {\em The Computer Journal}, 42\penalty0
(4):\penalty0 327--329, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/4/327.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420327.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420327.pdf|}\fi.
\bibitem[Ris99c]{Rissanen:1999:R} Jorma Rissanen. \newblock
Rejoinder. \newblock {\em The Computer Journal}, 42\penalty0
(4):\penalty0 343--344, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/4/343.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420343.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420343.pdf|}\fi.
\bibitem[RK92]{Robson:1992:SBS} A.~P. Robson and D.~J.
Kinniment. \newblock Simstrict: a behavioural simulator
for use with the strict hardware   description language.
\newblock {\em The Computer Journal}, 35\penalty0 (6)
:\penalty0 651--654,   December 1992. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/6/651.
full.pdf+html|}\fi.
\bibitem[RKA92]{Rosenberg:1992:AML} J.~Rosenberg, J.~
L. Keedy, and D.~Abramson. \newblock Addressing mechanisms
for large virtual memories. \newblock {\em The Computer
Journal}, 35\penalty0 (4):\penalty0 369--375,   August
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/4/369.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_04/Vol35_04.body.
html#AbstractRosenberg|}\fi. \newblock Australia.
\bibitem[RLD94]{Ranai:1994:SMF} Kisenchand Ranai, Koon~
Seng Lim, and R.~H. Deng. \newblock Simulating a multimedia
{FDDI} backbone network. \newblock {\em The Computer Journal}
, 37\penalty0 (3):\penalty0 188--198, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/3/
188.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_03/Vol37_03.body.html#AbstractRanai|}
\fi.
\bibitem[RLT99]{Rook:1999:BR} Jim Rook, Ian~J. Lloyd,
and Kenneth~J. Turner. \newblock Book reviews. \newblock
{\em The Computer Journal}, 42\penalty0 (2):\penalty0
150--152, ????   1999. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/2/150.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
02/420150.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_02/pdf/420150.pdf|}\fi.
\bibitem[Rob90]{Roberts:1990:ESC} H.~J. Roberts. \newblock
Expert systems clubs: Design methods. \newblock {\em The
Computer Journal}, 33\penalty0 (6):\penalty0 556--561,
 December 1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/6/556.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/556.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_33/Issue_06/tiff/557.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/558.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/559.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/560.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/561.tif|}\fi.
\bibitem[Rob93]{Robinson:1993:BR} H.~Robinson. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (6):\penalty0 597--598, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/6/
597.2.full.pdf+html|}\fi.
\bibitem[Roc97a]{Rock:1997:BRa} S.~Rock. \newblock Book
review. \newblock {\em The Computer Journal}, 40\penalty0
(4):\penalty0 231--232, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/4/231.
2.full.pdf+html|}\fi.
\bibitem[Roc97b]{Rock:1997:BRb} S.~Rock. \newblock Book
review. \newblock {\em The Computer Journal}, 40\penalty0
(5):\penalty0 299, ???? 1997. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/5/299.full.pdf+html|}\fi.
\bibitem[Rog94]{Roger:1994:LCM} R.~E. Roger. \newblock
Letters: Correspondence. on the {Millar} polyhedron and
its use in   the construction of octrees. \newblock {
\em The Computer Journal}, 37\penalty0 (5):\penalty0 465-
-466, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/5/465.full.pdf+html|}\fi.
\bibitem[Rog96a]{Rogerson:1996:BRa} S.~Rogerson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (1):\penalty0 95--96, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/39/1/
95.2.full.pdf+html|}\fi.
\bibitem[Rog96b]{Rogerson:1996:BRb} S.~Rogerson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (2):\penalty0 170, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/2/170.
1.full.pdf+html|}\fi.
\bibitem[Roh90]{Rohl:1990:RIL} J.~S. Rohl. \newblock Reversing
is not inherent in lexicographical permutation generation.
\newblock {\em The Computer Journal}, 33\penalty0 (3)
:\penalty0 274--278, June   1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/3/274.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/274.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/275.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/277.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/278.tif|}\fi.
\bibitem[Roh91]{Rohl:1991:OSP} J.~S. Rohl. \newblock {
Ord-Smith}'s pseudo-lexicographical permutation procedure
is the   {Tompkins-Paige} algorithm. \newblock {\em The
Computer Journal}, 34\penalty0 (6):\penalty0 569--570,
 December 1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/6/569.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/569.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/570.tif|}\fi.
\bibitem[Roj95]{Rojas:1995:CCS} Isabel Rojas. \newblock
Compositional construction of {SWN} models. \newblock
{\em The Computer Journal}, 38\penalty0 (7):\penalty0
612--621, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/7/612.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/Volume_38/Issue_07/
Vol38_07.body.html#AbstractRojas|}\fi.
\bibitem[Rom99]{Romanovsky:1999:SCC} Alexander Romanovsky.
\newblock On structuring cooperative and competitive concurrent
systems. \newblock {\em The Computer Journal}, 42\penalty0
(8):\penalty0 627--637, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/8/627.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_08/420627.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_08/
pdf/420627.pdf|}\fi.
\bibitem[Ros93]{Rossi:1993:PDE} G.~Rossi. \newblock Programs
as data in an extended {Prolog}. \newblock {\em The Computer
Journal}, 36\penalty0 (3):\penalty0 217--226, June   1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/3/217.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/Vol36_03.body.html#AbstractRossi|}
\fi.
\bibitem[Ros95]{Rosbottom:1995:MMU} J.~Rosbottom. \newblock
The muse method for usability engineering. \newblock {
\em The Computer Journal}, 38\penalty0 (3):\penalty0 266-
-267, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/3/266.full.pdf+html|}\fi.
\bibitem[Rou94]{Routen:1994:BR} T.~W. Routen. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (7):\penalty0 645--646, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/7/
645.full.pdf+html|}\fi.
\bibitem[RS93]{Rayward-Smith:1993:BR} V.~J. Rayward-Smith.
\newblock Book review. \newblock {\em The Computer Journal}
, 36\penalty0 (3):\penalty0 293--294, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/36/3/
293.full.pdf+html|}\fi.
\bibitem[Rum94]{Rumsby:1994:EC} S.~Rumsby. \newblock The
evolution of {C}++. \newblock {\em The Computer Journal}
, 37\penalty0 (6):\penalty0 557, ???? 1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
557.full.pdf+html|}\fi.
\bibitem[Ru{\v{z}}91]{Ruzicka:1991:SNN} P.~Ru{\v{z}}i{
\v{c}}ka. \newblock Short note: a note on the efficiency
of an interval routing   algorithm. \newblock {\em The
Computer Journal}, 34\penalty0 (5):\penalty0 475--476,
 October 1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/5/475.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/475.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/476.tif|}\fi.
\bibitem[RV93]{Ramanathan:1993:FTI} S.~Ramanathan and
P.~{Venkat Rangan}. \newblock Feedback techniques for
intra-media continuity and inter-media   synchronization
in distributed multimedia systems. \newblock {\em The
Computer Journal}, 36\penalty0 (1):\penalty0 19--31,
February 1993. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/1/19.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractRamanathan|}\fi.
\bibitem[Sad92]{Sadri:1992:DDD} F.~Sadri. \newblock Data
dependencies in deductive databases. \newblock {\em The
Computer Journal}, 35\penalty0 (5):\penalty0 A415--??
,   October 1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://www3.oup.
co.uk/computer_journal/hdb/Volume_35/Issue_05/tiff/A415.
tif;   http://www3.oup.co.uk/computer_journal/Volume_
35/Issue_05/35000A415.sgm.abs.html|}\fi.
\bibitem[San92]{Sanderson:1992:CIR} Mark Sanderson. \
newblock Conceptual information retrieval --- a case study

in adaptive partial   parsing. \newblock {\em The Computer
Journal}, 35\penalty0 (3):\penalty0 278, June 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/3/278.full.pdf+html|}\fi.
\bibitem[San94]{Sanders:1994:BR} J.~W. Sanders. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (2):\penalty0 148--149, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
148.2.full.pdf+html|}\fi.
\bibitem[San95a]{Sanders:1995:CFS} J.~W. Sanders. \newblock
The construction of formal specifications. \newblock {
\em The Computer Journal}, 38\penalty0 (1):\penalty0 80-
-81, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/1/80.full.pdf+html|}\fi.
\bibitem[San95b]{Sanyal:1995:CAS} S.~Sanyal. \newblock
Computer arithmetic systems. \newblock {\em The Computer
Journal}, 38\penalty0 (1):\penalty0 79, ???? 1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
79.full.pdf+html|}\fi.
\bibitem[San99]{Sanderson:1999:BR} Mark Sanderson. \newblock
Book reviews. \newblock {\em The Computer Journal}, 42\
penalty0 (8):\penalty0 713--715, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/8/
713.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_08/420713.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_08/
pdf/420713.pdf|}\fi.
\bibitem[Sar90]{Sarda:1990:AQL} N.~L. Sarda. \newblock
Algebra and query language for {A} historical data model.
\newblock {\em The Computer Journal}, 33\penalty0 (1)
:\penalty0 11--18,   February 1990. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/11.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/12.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/13.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/14.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/17.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/18.tif|}\fi.
\bibitem[Sar96]{Sarbo:1996:BSK} J.~J. Sarbo. \newblock
Building sub-knowledge bases using concept lattices. \
newblock {\em The Computer Journal}, 39\penalty0 (10)
:\penalty0 868--875, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/10/868.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_39/Issue_10/Vol39_10.body.html#AbstractSarbo;
http://www3.oup.co.uk/computer_journal/Volume_39/Issue_
10/Vol39_10.body.html#AbstractSarbo|}\fi.
\bibitem[Sas95]{Sasikumar:1995:EIP} M.~Sasikumar. \newblock
Exploring interior point linear programming: Algorithms
and software. \newblock {\em The Computer Journal}, 38\
penalty0 (1):\penalty0 83--84, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
83.full.pdf+html|}\fi.
\bibitem[SB99]{Schieder:1999:ACL} Birgit Schieder and
Manfred Broy. \newblock Adapting calculational logic to
the undefined. \newblock {\em The Computer Journal}, 42\
penalty0 (2):\penalty0 73--81, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/2/
73.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_02/420073.sgm.abs.html;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_02/
pdf/420073.pdf|}\fi.
\bibitem[SBRS94]{Sommerville:1994:CSD} Ian Sommerville,
Richard Bentley, Tom Rodden, and Peter Sawyer. \newblock
Cooperative systems design. \newblock {\em The Computer
Journal}, 37\penalty0 (5):\penalty0 357--366, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/5/357.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/Vol37_05.body.html#AbstractSommerville|}
\fi.
\bibitem[SC93]{Strain-Clark:1993:BR} P.~Strain-Clark.
\newblock Book review. \newblock {\em The Computer Journal}
, 36\penalty0 (6):\penalty0 595--596, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/6/
595.2.full.pdf+html|}\fi.
\bibitem[SC95]{Sherrell:1995:FIS} Linda~B. Sherrell and
Doris~L. Carver. \newblock {FunZ}: An intermediate specification
language. \newblock {\em The Computer Journal}, 38\penalty0
(3):\penalty0 193--206, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/193.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_03/Vol38_03.body.html#AbstractSherrell|}
\fi.
\bibitem[Sca90]{Scarrott:1990:C} G.~G. Scarrott. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (1):\penalty0 93, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/93.
1.full.pdf+html|}\fi.
\bibitem[Sca92]{Scarrott:1992:LI} G.~Scarrott. \newblock
Logic and information. \newblock {\em The Computer Journal}
, 35\penalty0 (2):\penalty0 118, April   1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/35/2/
118.full.pdf+html|}\fi.
\bibitem[SCC{\etalchar{+}}96]{Seed:1996:OOG} G.~M. Seed,
D.~E.~R. Clark, J.~R. Corney, R.~Tuttle, and G.~Little.
\newblock Object-oriented graph-based geometric feature
recognition. \newblock {\em The Computer Journal}, 39\
penalty0 (9):\penalty0 808--811, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
808.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_39/Issue_09/Vol39_09.body.html#AbstractSeed;
http://www3.oup.co.uk/computer_journal/Volume_39/Issue_
09/Vol39_09.body.html#AbstractSeed|}\fi.
\bibitem[Sch94]{Schorn:1994:DGC} Peter Schorn. \newblock
Degeneracy in geometric computation and the perturbation
approach. \newblock {\em The Computer Journal}, 37\penalty0
(1):\penalty0 35--42, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/1/35.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractSchorn|}
\fi.
\bibitem[Sch95]{Schneider:1995:BR} S.~A. Schneider. \
newblock Book review. \newblock {\em The Computer Journal}
, 38\penalty0 (10):\penalty0 837--838, ????   1995. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/10/837.full.pdf+html|}\fi.
\bibitem[SDR96]{Sekharan:1996:STT} Chandran~N. Sekharan,
John~G. {Del Greco}, and Sailesh Rathi. \newblock Space-
time trade-offs in the relative unranking of binary trees.
\newblock {\em The Computer Journal}, 39\penalty0 (1)
:\penalty0 36--44, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/1/36.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_01/390036.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_01/Vol39_
01.body.html#AbstractSekharan|}\fi.

\bibitem[Seb96]{Sebastiani:1996:BR} F.~Sebastiani. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 728, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/8/728.
1.full.pdf+html|}\fi.
\bibitem[Seb97]{Sebastiani:1997:BR} F.~Sebastiani. \newblock
Book review. \newblock {\em The Computer Journal}, 40\
penalty0 (5):\penalty0 298--299, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/5/
298.full.pdf+html|}\fi.
\bibitem[Seb98]{Sebastiani:1998:BR} F.~Sebastiani. \newblock
Book review. \newblock {\em The Computer Journal}, 41\
penalty0 (4):\penalty0 280--281, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/4/
280.full.pdf+html|}\fi.
\bibitem[Ser95]{Sereno:1995:TPF} Matteo Sereno. \newblock
Towards a product form solution for stochastic process
algebras. \newblock {\em The Computer Journal}, 38\penalty0
(7):\penalty0 622--632, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/7/622.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_07/Vol38_07.body.html#AbstractSereno|}
\fi.
\bibitem[SFD92]{Semmens:1992:ISA} L.~T. Semmens, R.~B.
France, and T.~W.~G. Docker. \newblock Integrated structured
analysis and formal specification techniques. \newblock
{\em The Computer Journal}, 35\penalty0 (6):\penalty0
600--610,   December 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/6/600.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_06/
Vol35_06.body.html#AbstractSemmens|}\fi. \newblock See
\cite{Cooke:1992:EFMa}.

\bibitem[SH97]{Storer:1997:LIC} J.~A. Storer and H.~Helfgott.
\newblock Lossless image compression by block matching.
\newblock {\em The Computer Journal}, 40\penalty0 (2-
-3):\penalty0 137--145,   ???? 1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/2_and_
3/137.full.pdf+html;   http://www.oup.co.uk/computer_
journal/Volume_40/Issue_02/Vol40_02.body.html#AbstractStorer;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_03.body.html#AbstractStorer|}\fi.
\bibitem[Sha92]{Shaw:1992:HFI} Mildred L.~G. Shaw. \newblock
Human factors for informatics usability. \newblock {\
em The Computer Journal}, 35\penalty0 (1):\penalty0 29,
February   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/1/29.full.pdf+html|}\fi.
\bibitem[Sha94]{Sharman:1994:SRO} R.~A. Sharman. \newblock
Speech recognition in the office: how the technology supports
 dictation. \newblock {\em The Computer Journal}, 37\
penalty0 (9):\penalty0 735--744, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/9/
735.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_09/Vol37_09.body.html#AbstractSharman|}
\fi.
\bibitem[Sha95]{Shafarenko:1995:SDP} A.~Shafarenko. \
newblock Symmetries in data parallelism. \newblock {\
em The Computer Journal}, 38\penalty0 (5):\penalty0 365-
-380, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/5/365.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_05/
Vol38_05.body.html#AbstractShafarenko|}\fi.
\bibitem[SHCP91]{Stewart:1991:ADF} A.~Stewart, C.~Holt,
M.~Clint, and R.~Perrott. \newblock Algorithm 128: a {
DAP FORTRAN} subroutine for thinning digitised   pictures.
\newblock {\em The Computer Journal}, 34\penalty0 (4)
:\penalty0 377--379,   August 1991. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/377.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/378.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/379.tif|}\fi.
\bibitem[She91]{Shepherd:1991:BR} D.~Shepherd. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 474, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/474.
full.pdf+html|}\fi.
\bibitem[She92a]{Shearer:1992:RUD} I.~Shearer. \newblock
Retrospective update: Data as it was believed to be. \
newblock {\em The Computer Journal}, 35\penalty0 (2):
\penalty0 184, April   1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/2/184.full.pdf+html|}\fi.
\bibitem[She92b]{Shen:1992:SAM} H.~Shen. \newblock Self-
adjusting mapping: a heuristic mapping algorithm for mapping
 parallel programs on to {Transputer} networks. \newblock
{\em The Computer Journal}, 35\penalty0 (1):\penalty0
71--80,   February 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/1/71.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_01/
Vol35_01.body.html#AbstractShen|}\fi.
\bibitem[She95a]{Shen:1995:EPS} Hong Shen. \newblock Efficient
parallel $k$-set chain range-join in hypercubes. \newblock
{\em The Computer Journal}, 38\penalty0 (3):\penalty0
217--226, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/3/217.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_03/
Vol38_03.body.html#AbstractShen|}\fi.

\bibitem[She95b]{Shepherdson:1995:LLP} J.~C. Shepherdson.
\newblock From logic to logic programming. \newblock {
\em The Computer Journal}, 38\penalty0 (1):\penalty0 78,
???? 1995. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/1/78.full.pdf+html|}\fi.
\bibitem[She97]{Shene:1997:ATP} C.-K. Shene. \newblock
An analysis of two in-place array rotation algorithms.
\newblock {\em The Computer Journal}, 40\penalty0 (9)
:\penalty0 541--546, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/9/541.
full.pdf+html|}\fi.
\bibitem[She99]{Shen:1999:DKC} Alexander Shen. \newblock
Discussion on {Kolmogorov} complexity and statistical
analysis. \newblock {\em The Computer Journal}, 42\penalty0
(4):\penalty0 340--342, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/4/340.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420340.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420340.pdf|}\fi.
\bibitem[Shi90]{Shizgal:1990:APS} I.~Shizgal. \newblock
The {Amoeba-Prolog} system. \newblock {\em The Computer
Journal}, 33\penalty0 (6):\penalty0 508--517,   December
1990. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/33/6/508.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/508.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/509.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/510.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/511.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/512.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/513.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/514.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_33/Issue_06/tiff/515.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/516.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/517.tif|}\fi.
\bibitem[Sim96]{Simpson:1996:BR} A.~Simpson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (6):\penalty0 556, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/6/556.
1.full.pdf+html|}\fi.
\bibitem[Sin97]{Sinharoy:1997:OTC} Balaram Sinharoy. \
newblock Optimized thread creation for processor multithreading.
\newblock {\em The Computer Journal}, 40\penalty0 (6)
:\penalty0 388--400, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/6/388.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_06/Vol40_06.body.html#AbstractSinharoy;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
06/Vol40_06.body.html#AbstractSinharoy|}\fi.
\bibitem[SK93a]{Schwederski:1993:E} T.~Schwederski and
R.~Kober. \newblock Editorial. \newblock {\em The Computer
Journal}, 36\penalty0 (8):\penalty0 683--684, ????   1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/8/683.full.pdf+html|}\fi.
\bibitem[SK93b]{Schwederski:1993:SIE} Thomas Schwederski
and Rudolph Kober. \newblock Special issue editorial:
Parallel processing. \newblock {\em The Computer Journal}
, 36\penalty0 (8):\penalty0 683, ???? 1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://www3.oup.co.uk/computer_journal/Volume_
36/Issue_08/Vol36_08.body.html#AbstractSchwederski|}\
fi.
\bibitem[SK95]{Sutton:1995:IIF} D.~R. Sutton and P.~J.
~H. King. \newblock Incomplete information and the functional
data model. \newblock {\em The Computer Journal}, 38\
penalty0 (1):\penalty0 31--42, ????   1995. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
31.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_01/Vol38_01.body.html#AbstractSutton|}
\fi.
\bibitem[Ski91]{Skillicorn:1991:PCP} D.~B. Skillicorn.
\newblock Practical concurrent programming for parallel
machines. \newblock {\em The Computer Journal}, 34\penalty0
(4):\penalty0 302--310,   August 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/4/302.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/302.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/303.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/304.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/305.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/306.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/307.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/308.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/309.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/310.tif|}\fi.
\bibitem[Ski96]{Skillicorn:1996:UDH} D.~B. Skillicorn.
\newblock Using distributed hypermedia for collaborative
learning in   universities. \newblock {\em The Computer
Journal}, 39\penalty0 (6):\penalty0 471--482, ????   1996.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
39/6/471.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_06/390471.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_06/
Vol39_06.body.html#AbstractSkillicorn|}\fi.
\bibitem[SM94]{Stephen:1994:WST} Graham~A. Stephen and
Paul Mather. \newblock What is {SP}? (theory of computing)
. \newblock {\em The Computer Journal}, 37\penalty0 (
9):\penalty0 745--752, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/37/9/745.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_09/Vol37_09.body.html#AbstractStephen|}
\fi.
\bibitem[SM97]{Singhai:1997:PLF} Sharad~K. Singhai and
Kathryn~S. McKinley. \newblock A parametrized loop fusion
algorithm for improving parallelism and   cache locality.
\newblock {\em The Computer Journal}, 40\penalty0 (6)
:\penalty0 340--355, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/6/340.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_06/Vol40_06.body.html#AbstractSinghai;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
06/Vol40_06.body.html#AbstractSinghai|}\fi.
\bibitem[Sme92]{Smeaton:1992:PAN} Alan~F. Smeaton. \newblock
Progress in the application of natural language processing
to   information retrieval tasks. \newblock {\em The Computer
Journal}, 35\penalty0 (3):\penalty0 268--278, June   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/3/268.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_03/Vol35_03.body.html#AbstractSmeaton|}
\fi.
\bibitem[SMFM97]{Sanna:1997:NAR} A.~Sanna, P.~Montuschi,
A.~Fisone, and B.~Montrucchio. \newblock A new algorithm
for the rendering of {CSG} scenes. \newblock {\em The
Computer Journal}, 40\penalty0 (9):\penalty0 555--564,
????   1997. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/40/9/555.full.pdf+html|}\fi.
\bibitem[SMR98]{Sanna:1998:FAM} Andrea Sanna, Paolo Montuschi,
and Massimo Rossi. \newblock A flexible algorithm for
multiprocessor ray tracing. \newblock {\em The Computer
Journal}, 41\penalty0 (7):\penalty0 503--516, ????   1998.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/7/503.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/Volume_41/Issue_07/Vol41_07.body.html#AbstractSanna|}
\fi.
\bibitem[Smu94]{Smurthwaite:1994:BR} M.~Smurthwaite. \
newblock Book review. \newblock {\em The Computer Journal}
, 37\penalty0 (2):\penalty0 152, ???? 1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
152.1.full.pdf+html|}\fi.
\bibitem[Smy91]{Smyth:1991:MBM} W.~F. Smyth. \newblock
{Mu-balancing M-way} search trees. \newblock {\em The
Computer Journal}, 34\penalty0 (5):\penalty0 406--414,
 October 1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/5/406.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/406.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/407.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/408.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/409.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/410.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/411.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/412.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/413.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/414.tif|}\fi.
\bibitem[SMY92]{Sun:1992:QOD} W.~Sun, W.~Meng, and C.
~Yu. \newblock Query optimisation in distributed object-
oriented database systems. \newblock {\em The Computer
Journal}, 35\penalty0 (2):\penalty0 98--107, April   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/2/98.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_02/Vol35_02.body.html#AbstractSun|}
\fi.
\bibitem[SMY95]{Stoyenko:1995:LCR} Alexander~D. Stoyenko,
Thomas~J. Marlowe, and Mohamed~F. Younis. \newblock A
language for complex real-time systems. \newblock {\em

The Computer Journal}, 38\penalty0 (4):\penalty0 319-
-338, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/4/319.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_04/
Vol38_04.body.html#AbstractStoyenko|}\fi.
\bibitem[Sob94a]{Sobczak:1994:BR} R.~Sobczak. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (9):\penalty0 823--824, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/9/
823.full.pdf+html|}\fi.
\bibitem[Sob94b]{Sobczak:1994:FHS} R.~Sobczak. \newblock
Fddi: a high speed network. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 542a--542, ????
1994. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/6/542-a.full.pdf+html|}\fi.
\bibitem[Sob94c]{Sobczak:1994:TRN} R.~Sobczak. \newblock
Token ring networks. \newblock {\em The Computer Journal}
, 37\penalty0 (6):\penalty0 542, ???? 1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/6/
542.full.pdf+html|}\fi.
\bibitem[Sol93a]{Solomon:1993:BRa} A.~Solomon. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 289--290, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/3/
289.2.full.pdf+html|}\fi.
\bibitem[Sol93b]{Solomon:1993:BRb} A.~Solomon. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (6):\penalty0 602, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/602.
2.full.pdf+html|}\fi.

\bibitem[Sol99]{Solomonoff:1999:TKP} Ray Solomonoff. \
newblock Two kinds of probabilistic induction. \newblock
{\em The Computer Journal}, 42\penalty0 (4):\penalty0
256--259, ????   1999. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/42/4/256.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_
04/420256.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_04/pdf/420256.pdf|}\fi.
\bibitem[SP91]{Son:1991:ADF} N.~T. Son and Y.~Paker. \
newblock Adaptive deadlock-free packet routing in transputer-
based   multiprocessor interconnection networks. \newblock
{\em The Computer Journal}, 34\penalty0 (6):\penalty0
493--502,   December 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/6/493.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/493.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/494.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/495.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/496.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/497.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/498.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/499.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/500.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/501.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/502.tif|}\fi.
\bibitem[Spe94a]{Speed:1994:BRa} C.~Speed. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(2):\penalty0 145--146, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/145.
2.full.pdf+html|}\fi.
\bibitem[Spe94b]{Speed:1994:BRb} C.~Speed. \newblock Book
review. \newblock {\em The Computer Journal}, 37\penalty0
(2):\penalty0 146--147, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/146.
3.full.pdf+html|}\fi.
\bibitem[Spe95]{Speed:1995:A} Claire Speed. \newblock
{It}'s alive! \newblock {\em The Computer Journal}, 38\
penalty0 (3):\penalty0 269, ???? 1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/3/269.
full.pdf+html|}\fi.
\bibitem[Spr92]{Sprugnoli:1992:PBT} R.~Sprugnoli. \newblock
Properties of binary trees related to position. \newblock
{\em The Computer Journal}, 35\penalty0 (4):\penalty0
395--404,   August 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/4/395.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_04/
Vol35_04.body.html#AbstractSprugnoli|}\fi.
\bibitem[Sri91]{Srinivasan:1991:AOT} B.~Srinivasan. \
newblock An adaptive overflow technique to defer splitting
in {B}-trees. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 397--405,   October 1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/5/
397.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/397.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/398.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/399.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
05/tiff/401.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/402.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_05/tiff/403.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_05/tiff/404.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_05/tiff/405.tif|}\fi.
\bibitem[Sri94]{Srivas:1994:BR} M.~K. Srivas. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (9):\penalty0 821--823, ????   1994. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/9/
821.full.pdf+html|}\fi.
\bibitem[SRS94]{Seshadri:1994:OAC} Sridhar Seshadri, Doron
Rotem, and Arie Segev. \newblock Optimal arrangements
of cartridges in carousel type mass storage   systems.
\newblock {\em The Computer Journal}, 37\penalty0 (10)
:\penalty0 873--887, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/10/873.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_10/Vol37_10.body.html#AbstractSeshadri|}
\fi.
\bibitem[SS92]{Sondergaard:1992:NDF} Harald S{\o}ndergaard
and Peter Sestoft. \newblock Non-determinism in functional
languages. \newblock {\em The Computer Journal}, 35\penalty0
(5):\penalty0 514--523,   October 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/5/514.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_05/Vol35_05.body.html#AbstractSondergaard|}
\fi. \newblock Accepted --- Melbourne Technical Report
88/18.
\bibitem[SS96a]{Saiedian:1996:EAC} Hossein Saiedian and
Thomas Spencer. \newblock An efficient algorithm to compute
the candidate keys of a relational   database schema.
\newblock {\em The Computer Journal}, 39\penalty0 (2)
:\penalty0 124--132, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/2/124.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_02/390124.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_02/Vol39_
02.body.html#AbstractSaiedian|}\fi.
\bibitem[SS96b]{Sekerinski:1996:TPC} Emil Sekerinski and
Kaisa Sere. \newblock A theory of prioritizing composition.
\newblock {\em The Computer Journal}, 39\penalty0 (8)
:\penalty0 701--712, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/8/701.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_08/390701.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_08/Vol39_
08.body.html#AbstractSekerinski|}\fi.
\bibitem[ST94]{Shawe-Taylor:1994:BR} J.~Shawe-Taylor.
\newblock Book review. \newblock {\em The Computer Journal}
, 37\penalty0 (8):\penalty0 723--725, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
723.full.pdf+html|}\fi.
\bibitem[Sta92]{Stamper:1992:TCS} Ronald Stamper. \newblock
A theory of computer semiotics. \newblock {\em The Computer
Journal}, 35\penalty0 (4):\penalty0 368, August   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/4/368.full.pdf+html|}\fi.
\bibitem[Sta94]{Stavridou:1994:FMV} Victoria Stavridou.
\newblock Formal methods and {VLSI} engineering practice.
\newblock {\em The Computer Journal}, 37\penalty0 (2)
:\penalty0 96--113, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/2/96.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_02/Vol37_02.body.html#AbstractStavridou|}
\fi.
\bibitem[Sta97]{Stavridou:1997:BR} V.~Stavridou. \newblock
Book review. \newblock {\em The Computer Journal}, 40\
penalty0 (1):\penalty0 58--59, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/1/
58.2.full.pdf+html|}\fi.
\bibitem[Sti96]{Stirling:1996:BR} C.~Stirling. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (2):\penalty0 171--172, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/39/2/
171.full.pdf+html|}\fi.
\bibitem[Sto90]{Stonier:1990:C} T.~Stonier. \newblock
Correspondence. \newblock {\em The Computer Journal},
33\penalty0 (1):\penalty0 92--93, ????   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/1/
92.3.full.pdf+html|}\fi.
\bibitem[Str94a]{Stroud:1994:BRa} R.~Stroud. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 727--728, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
727.2.full.pdf+html|}\fi.
\bibitem[Str94b]{Stroud:1994:BRb} R.~Stroud. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 728--730, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
728.full.pdf+html|}\fi.
\bibitem[Suf90]{Suffern:1990:QAC} K.~G. Suffern. \newblock
Quadtree algorithms for contouring functions of two variables.
\newblock {\em The Computer Journal}, 33\penalty0 (5)
:\penalty0 402--407,   October 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/402.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/402.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_05/tiff/403.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_05/tiff/404.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_05/tiff/405.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
05/tiff/406.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_05/tiff/407.tif|}\fi.
\bibitem[SW91a]{Slape:1991:AMS} J.~K. Slape and P.~J.
~L. Wallis. \newblock Algorithm 127: a modification of
{Sale}'s algorithm to accommodate   {FORTRAN 77}. \newblock

{\em The Computer Journal}, 34\penalty0 (4):\penalty0
373--376,   August 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://www3.
oup.co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/
373.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/374.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_34/Issue_04/tiff/375.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_34/
Issue_04/tiff/376.tif|}\fi.
\bibitem[SW91b]{Sutcliffe:1991:IHC} A.~G. Sutcliffe and
I.~Wang. \newblock Integrating human computer interaction
with {Jackson} system   development. \newblock {\em The
Computer Journal}, 34\penalty0 (2):\penalty0 132--142,
April   1991. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/2/132.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/132.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/133.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/134.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/135.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/136.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/137.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/138.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/139.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/140.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/141.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/142.tif|}\fi.
\bibitem[SWA97]{Sjoberg:1997:SCL} Dag I.~K. Sj{\"o}berg,
Ray Welland, and Malcolm~P. Atkinson. \newblock Software
constraints for large application systems. \newblock {
\em The Computer Journal}, 40\penalty0 (10):\penalty0
598--616, ????   1997. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/10/598.full.pdf+html;
http://www3.oup.co.uk/computer_journal/Volume_40/Issue_

10/Vol40_10.body.html#AbstractSjoberg|}\fi.
\bibitem[SWZ93]{Schneider:1993:SLS} M.~Schneider, U.~
Wever, and Q.~Zheng. \newblock Solving large and sparse
linear equations in analog circuit   simulation on a cluster
of workstations. \newblock {\em The Computer Journal}
, 36\penalty0 (8):\penalty0 685--689, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/8/
685.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractSchneider|}
\fi.
\bibitem[Sym95]{Symvonis:1995:OSM} Antonios Symvonis.
\newblock Optimal stable merging. \newblock {\em The Computer
Journal}, 38\penalty0 (8):\penalty0 681--690, ????   1995.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/8/681.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_38/Issue_08/Vol38_08.body.html#AbstractSymvonis|}
\fi.
\bibitem[SZNM94]{Sousa:1994:OPH} Pedro Sousa, Andr{\'
e} Z{\'u}quete, Nuno Neves, and Jos{\'e}~Alves Marques.
\newblock Orthogonal persistence in a heterogeneous distributed
object-oriented   environment. \newblock {\em The Computer
Journal}, 37\penalty0 (6):\penalty0 531--541, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/6/531.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_06/Vol37_06.body.html#AbstractSousa|}
\fi.
\bibitem[TAC98]{Thimbleby:1998:FMT} Harold Thimbleby,
Stuart Anderson, and Paul Cairns. \newblock A framework
for modelling {Trojans} and computer virus infection.
\newblock {\em The Computer Journal}, 41\penalty0 (7)
:\penalty0 444--458, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/7/444.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_07/Vol41_07.body.html#AbstractThimbleby|}

\fi. \newblock See comments   \cite{Makinen:2001:CFM,
Thimbleby:2001:RCF,Cohen:2001:RCF}.
\bibitem[TAC01]{Thimbleby:2001:RCF} H.~Thimbleby, S.~
O. Anderson, and P.~A. Cairns. \newblock Reply to {``
Comment on `A Framework for Modelling Trojans and   Computer
Virus Infection' by E. M{\"a}kinen''}. \newblock {\em
The Computer Journal}, 44\penalty0 (4):\penalty0 324-
-325, ????   2001. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://www3.oup.
co.uk/computer_journal/hdb/Volume_44/Issue_04/440324.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/440324.pdf|}\fi. \newblock
See \cite{Thimbleby:1998:FMT,Makinen:2001:CFM,Cohen:2001:
RCF}.
\bibitem[Tak96]{Takaoka:1996:LRP} Tadao Takaoka. \newblock
A left-to-right preprocessing computation for the {Boyer-
-Moore}   string matching algorithm. \newblock {\em The
Computer Journal}, 39\penalty0 (5):\penalty0 413--416,
????   1996. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/39/5/413.full.pdf+html;   http://www.oup.
co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_05/390413.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
Volume_39/Issue_05/Vol39_05.body.html#AbstractTakaoka|}
\fi.
\bibitem[Tak99]{Takaoka:1999:TAC} Tadao Takaoka. \newblock
{$O(1)$} time algorithms for combinatorial generation
by tree   traversal. \newblock {\em The Computer Journal}
, 42\penalty0 (5):\penalty0 400--408, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/5/
400.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_05/420400.sgm.abs.html|}\fi.
\bibitem[Tay90]{Taylor:1990:LOD} M.~C. Taylor. \newblock
Logical optimisation of distributed knowledge base queries.
\newblock {\em The Computer Journal}, 33\penalty0 (1)
:\penalty0 49--60,   February 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/1/49.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/50.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/51.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/52.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/55.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_01/tiff/56.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_01/tiff/57.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_01/
tiff/58.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_01/tiff/59.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_01/tiff/60.
tif|}\fi.
\bibitem[{Tay}96]{Taylor:1996:SFF} John~M. {Taylor, OBE,
F. Eng.} \newblock Special feature: Forward with foresight.
\newblock {\em The Computer Journal}, 39\penalty0 (7)
:\penalty0 561--565, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/7/561.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_07/390561.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_07/Vol39_
07.body.html#AbstractTaylor|}\fi.
\bibitem[TB94]{Tindell:1994:FPS} K.~Tindell and A.~Burns.
\newblock Fixed priority scheduling of hard real-time
multi-media disk traffic. \newblock {\em The Computer
Journal}, 37\penalty0 (8):\penalty0 691--697, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/8/691.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_08/Vol37_08.body.html#AbstractTindell|}
\fi.
\bibitem[TC92]{Turtle:1992:CTR} Howard~R. Turtle and W.
~Bruce Croft. \newblock A comparison of text retrieval
models. \newblock {\em The Computer Journal}, 35\penalty0
(3):\penalty0 279--290, June   1992. \newblock \showCODEN{

CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/3/279.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_03/Vol35_03.body.html#AbstractTurtle|}
\fi.
\bibitem[TCCH94]{Tangney:1994:RPP} Brendan Tangney, Andrew
Condon, Vinny Cahill, and Neville Harris. \newblock Requirements
for parallel programming in object-oriented distributed
 systems. \newblock {\em The Computer Journal}, 37\penalty0
(6):\penalty0 499--508, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/6/499.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_06/Vol37_06.body.html#AbstractTangney|}
\fi.
\bibitem[TCM94]{Taylor:1994:COO} Paul Taylor, Vinny Cahill,
and Michael Mock. \newblock Combining object-oriented
systems and open transaction processing. \newblock {\
em The Computer Journal}, 37\penalty0 (6):\penalty0 487-
-498, ????   1994. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/487.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_37/Issue_06/
Vol37_06.body.html#AbstractTaylor|}\fi.
\bibitem[TDW{\etalchar{+}}99]{Tomov:1999:AET} Neven Tomov,
Euan Dempster, M.~Howard Williams, Peter J.~B. King, and
Albert   Burger. \newblock Approximate estimation of transaction
response time. \newblock {\em The Computer Journal}, 42\
penalty0 (3):\penalty0 241--250, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/3/
241.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_03/420241.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_03/
pdf/420241.pdf|}\fi.
\bibitem[TF96]{Theel:1996:BRC} Oliver~E. Theel and Brett~
D. Fleisch. \newblock The boundary-restricted coherence
protocol for scalable and highly   available distributed

shared memory systems. \newblock {\em The Computer Journal}
, 39\penalty0 (6):\penalty0 496--510, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/6/
496.full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_06/390496.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_06/Vol39_
06.body.html#AbstractTheel|}\fi.
\bibitem[TGE96]{Tubella:1996:MAP} J.~Tubella, A.~Gonz{
\'a}lez, and Eduard Elias. \newblock The multipath architecture
for {Prolog} programs. \newblock {\em The Computer Journal}
, 39\penalty0 (9):\penalty0 780--792, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
780.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_39/Issue_09/Vol39_09.body.html#AbstractTubella;
 http://www3.oup.co.uk/computer_journal/Volume_39/Issue_
09/Vol39_09.body.html#AbstractTubella|}\fi.
\bibitem[TH98]{Tsaur:1998:ESE} Woei-Jiunn Tsaur and Shi-
Jinn Horng. \newblock Establishing secure {Ethernet LANs}
using intelligent switching hubs   in {Internet} environments.
\newblock {\em The Computer Journal}, 41\penalty0 (2)
:\penalty0 125--133, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/2/125.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_02/Vol41_02.body.html#AbstractTsaur|}
\fi.
\bibitem[Tha92]{Thainisch:1992:RTT} P.~Thainisch. \newblock
Real-time transaction processin. \newblock {\em The Computer
Journal}, 35\penalty0 (5):\penalty0 430a--430,   October
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/5/430-a.full.pdf+html|}\fi.
\bibitem[{The}91a]{ofHalsbury:1991:IFS} {The Rt. Hon.
The Earl of Halsbury}. \newblock `innovation for failure
--- some reflections on the work of the   {NRDC} relevant
to the early history of the computer industry in the {
UK}'. \newblock {\em The Computer Journal}, 34\penalty0

(3):\penalty0 272--274, June   1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/272.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/272.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/273.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/274.tif|}\fi.
\bibitem[{The}91b]{ofHalsbury:1991:TYC} {The Rt. Hon.
The Earl of Halsbury}. \newblock Ten years of computer
development. \newblock {\em The Computer Journal}, 34\
penalty0 (3):\penalty0 275--279, June   1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/3/
275.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/275.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/276.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/277.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/278.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/279.tif|}\fi.
\bibitem[The91c]{Theaker:1991:E} C.~J. Theaker. \newblock
Editorial. \newblock {\em The Computer Journal}, 34\penalty0
(2):\penalty0 97, ???? 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/2/97.full.pdf+html|}\fi.
\bibitem[THHS93]{Tawbi:1993:VCS} W.~Tawbi, F.~Horn, E.
~Horlait, and J.-B. St{\'e}fani. \newblock Video compression
standards and quality of service. \newblock {\em The Computer
Journal}, 36\penalty0 (1):\penalty0 43--54,   February
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/1/43.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_01/Vol36_01.body.
html#AbstractTawbi|}\fi.
\bibitem[Thi95]{Thimbleby:1995:NCW} Harold Thimbleby.
\newblock A new calculator and why it is necessary. \

newblock {\em The Computer Journal}, 38\penalty0 (6):
\penalty0 418--433, ????   1995. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/38/6/418.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_06/Vol38_06.body.html#AbstractThimbleby|}
\fi.
\bibitem[Tho91]{Thomas:1991:ETR} M.~Thomas. \newblock
Editorial --- term rewriting. \newblock {\em The Computer
Journal}, 34\penalty0 (1):\penalty0 1, ???? 1991. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/34/1/
1.full.pdf+html|}\fi.
\bibitem[Tho93a]{Thomas:1993:BRa} M.~Thomas. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 289, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/3/289.
1.full.pdf+html|}\fi.
\bibitem[Tho93b]{Thomas:1993:BRb} M.~Thomas. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (7):\penalty0 679--680, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/7/
679.2.full.pdf+html|}\fi.
\bibitem[Tho93c]{Thompson:1993:CIP} Peter Thompson. \
newblock Concurrent interconnect for parallel systems.
\newblock {\em The Computer Journal}, 36\penalty0 (8)
:\penalty0 778--784, ????   1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/8/778.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_08/Vol36_08.body.html#AbstractThompson|}
\fi.
\bibitem[Tho98]{Thomas:1998:MY} Martyn Thomas. \newblock
Myths of the {Year 2000}. \newblock {\em The Computer
Journal}, 41\penalty0 (2):\penalty0 67--70, ????   1998.

\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
41/2/67.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_41/Issue_02/Vol41_02.body.html#AbstractThomas|}
\fi.
\bibitem[tHO99]{terHofstede:1999:CSV} A.~H.~M. ter Hofstede
and M.~E. Orlowska. \newblock On the complexity of some
verification problems in process control   specifications.
\newblock {\em The Computer Journal}, 42\penalty0 (5)
:\penalty0 349--359, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/5/349.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_05/420349.sgm.abs.html|}\fi.
\bibitem[Tin94]{Tindell:1994:PRT} K.~Tindell. \newblock
Programming real-time multicomputers for signal processing.
\newblock {\em The Computer Journal}, 37\penalty0 (6)
:\penalty0 556, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/556.full.pdf+html|}\fi.
\bibitem[TK95]{Tahar:1995:FSV} Sofi{\`e}ne Tahar and Ramayya
Kumar. \newblock Formal specification and verification
techniques for {RISC} pipeline   conflicts. \newblock
{\em The Computer Journal}, 38\penalty0 (2):\penalty0
111--120, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/2/111.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractTahar|}\fi.
\bibitem[TK99]{Tsuchiya:1999:AEQ} Tatsuhiro Tsuchiya and
Tohru Kikuno. \newblock Availability evaluation of quorum-
based mutual exclusion schemes in   general topology networks.
\newblock {\em The Computer Journal}, 42\penalty0 (7)
:\penalty0 613--622, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/7/613.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_42/Issue_07/420613.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_07/
pdf/420613.pdf|}\fi.
\bibitem[TL92]{Thompson:1992:CMC} S.~J. Thompson and R.
~D. Lins. \newblock The {Categorical Multi-Combinator
Machine}: {CMCM}. \newblock {\em The Computer Journal}
, 35\penalty0 (2):\penalty0 170--176, April   1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/2/170.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_02/Vol35_02.body.html#AbstractThompson|}
\fi.
\bibitem[TL97]{Tsai:1997:MPR} W.-J. Tsai and S.-Y. Lee.
\newblock Multi-partition {RAID}: a new method for improving
performance of   disk arrays under failure. \newblock
{\em The Computer Journal}, 40\penalty0 (1):\penalty0
30--42, ????   1997. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/1/30.full.pdf+html;   http:
//www.oup.co.uk/computer_journal/Volume_40/Issue_01/Vol40_
01.body.html#AbstractTsai;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_01/Vol40_01.body.html#AbstractTsai|}
\fi.
\bibitem[TL98]{Tse:1998:MEI} Savio S.~H. Tse and Francis
C.~M. Lau. \newblock More on the efficiency of interval
routing. \newblock {\em The Computer Journal}, 41\penalty0
(4):\penalty0 238--242, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/4/238.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_04/Vol41_04.body.html#AbstractTse|}\fi.
\bibitem[TLF96]{TerHofstede:1996:CDM} A.~H.~M. {Ter Hofstede}
, E.~Lippe, and P.~J.~M. Frederiks. \newblock Conceptual
data modelling from a categorical perspective. \newblock
{\em The Computer Journal}, 39\penalty0 (3):\penalty0
215--231, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/3/215.full.pdf+html;   http:

//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
03/390215.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_03/Vol39_03.body.html#AbstractHofstede|}
\fi.
\bibitem[TM94]{Thomson:1994:SPP} Helen~E. Thomson and
Pam Mayhew. \newblock The software process: a perspective
on improvement. \newblock {\em The Computer Journal},
37\penalty0 (8):\penalty0 683--690, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/8/
683.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_08/Vol37_08.body.html#AbstractThomson|}
\fi.
\bibitem[TM95]{Turpin:1995:PLL} Andrew Turpin and Alistair
Moffat. \newblock Practical length-limited coding for
large alphabets. \newblock {\em The Computer Journal}
, 38\penalty0 (5):\penalty0 339--347, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/5/
339.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_05/Vol38_05.body.html#AbstractTurpin|}
\fi.
\bibitem[Tow98]{Townsend:1998:L} H.~R.~A. Townsend. \
newblock Letters. \newblock {\em The Computer Journal}
, 41\penalty0 (4):\penalty0 278, ???? 1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/4/
278.full.pdf+html|}\fi.
\bibitem[TP91]{Tse:1991:ERS} T.~H. Tse and L.~Pong. \
newblock An examination of requirements specification
languages. \newblock {\em The Computer Journal}, 34\penalty0
(2):\penalty0 143--152, April   1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/2/143.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/143.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/144.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/145.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_34/Issue_02/tiff/146.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/147.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/148.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/149.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/150.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/152.tif|}\fi.
\bibitem[TPvdW96]{TerHofstede:1996:QFI} A.~H.~M. {Ter
Hofstede}, H.~A. Proper, and Th.~P. van~der Weide. \newblock
Query formulation as an information retrieval problem.
\newblock {\em The Computer Journal}, 39\penalty0 (4)
:\penalty0 255--274, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/4/255.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_04/390255.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_04/Vol39_
04.body.html#AbstractHofstede|}\fi.
\bibitem[TR93]{Teuhola:1993:AFS} Jukka Teuhola and Timo
Raita. \newblock Application of a finite-state model to
text compression. \newblock {\em The Computer Journal}
, 36\penalty0 (7):\penalty0 607--614, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/7/
607.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_07/Vol36_07.body.html#AbstractTeuhola|}
\fi.
\bibitem[Tre91]{Trenouth:1991:SES} J.~Trenouth. \newblock
A survey of exploratory software development. \newblock
{\em The Computer Journal}, 34\penalty0 (2):\penalty0
153--163, April   1991. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/2/153.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/153.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/154.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/155.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/156.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/157.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/158.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_02/tiff/159.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_02/tiff/160.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_02/tiff/161.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_02/tiff/162.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
02/tiff/163.tif|}\fi.
\bibitem[Tre92]{Tredoux:1992:MES} G.~Tredoux. \newblock
Mechanizing execution sequence semantics in {HOL}. \newblock
{\em The Computer Journal}, 7\penalty0 (??):\penalty0
81--86, ????   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock Proceedings of the 7th Southern African Computer
Research Symposium,   Johannesburg, South Africa.
\bibitem[Tri97]{Trinder:1997:BR} P.~W. Trinder. \newblock
Book review. \newblock {\em The Computer Journal}, 40\
penalty0 (1):\penalty0 58, ???? 1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/1/58.
1.full.pdf+html|}\fi.
\bibitem[TU96]{Turkan:1996:EMM} Yasemin Turkan and {\
"O}zg{\"u}r Ulusoy. \newblock An efficient memory management
algorithm for client-server database   management systems.
\newblock {\em The Computer Journal}, 39\penalty0 (5)
:\penalty0 374--385, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/5/374.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_05/390374.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_05/Vol39_
05.body.html#AbstractTurkan|}\fi.
\bibitem[Tur92a]{Turega:1992:CAS} M.~Turega. \newblock
A computer architecture to support neural net simulation.
\newblock {\em The Computer Journal}, 35\penalty0 (4)
:\penalty0 353--360,   August 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-

2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/4/353.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_04/Vol35_04.body.html#AbstractTurega|}
\fi.
\bibitem[Tur92b]{Turega:1992:NN} M.~A. Turega. \newblock
Neural networks. \newblock {\em The Computer Journal}
, 35\penalty0 (3):\penalty0 290, June 1992. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/35/3/
290.full.pdf+html|}\fi.
\bibitem[Tur93a]{Turega:1993:BRa} M.~Turega. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (3):\penalty0 290--291, ????   1993. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/3/
290.2.full.pdf+html|}\fi.
\bibitem[Tur93b]{Turega:1993:BRb} M.~Turega. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (6):\penalty0 602, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/602.
1.full.pdf+html|}\fi.
\bibitem[Tur97]{Turner:1997:BR} R.~Turner. \newblock Book
review. \newblock {\em The Computer Journal}, 40\penalty0
(4):\penalty0 232--233, ????   1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/4/232.
full.pdf+html|}\fi.
\bibitem[TUW98]{Tayeb:1998:QBD} Jamel Tayeb, {\"O}zg{
\"u}r Ulusoy, and Ouri Wolfson. \newblock A quadtree-
based dynamic attribute indexing method. \newblock {\
em The Computer Journal}, 41\penalty0 (3):\penalty0 185-
-200, ????   1998. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/3/185.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/Volume_41/Issue_03/
Vol41_03.body.html#AbstractTayeb|}\fi.
\bibitem[TWMG93]{Tischer:1993:CBL} P.~E. Tischer, R.~
T. Worley, A.~J. Maeder, and M.~Goodwin. \newblock Context-
based lossless image compression. \newblock {\em The Computer
Journal}, 36\penalty0 (1):\penalty0 68--77,   February
1993. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/1/68.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_01/Vol36_01.body.
html#AbstractTischer|}\fi.
\bibitem[UC90]{Ueda:1990:DKL} Kazunori Ueda and Takashi~
C. Chikayama. \newblock Design of the kernel language
for the parallel inference machine. \newblock {\em The
Computer Journal}, 33\penalty0 (6):\penalty0 494--500,
 December 1990. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/6/494.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/494.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/495.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/496.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/497.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/498.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/499.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/500.tif|}\fi.
\bibitem[VA92]{VanRijsbergen:1992:ECI} Keith {Van Rijsbergen}
and Maristella Agosti. \newblock Editorial --- the context
of information retrieval. \newblock {\em The Computer
Journal}, 35\penalty0 (3):\penalty0 193, June 1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/3/193.full.pdf+html|}\fi.
\bibitem[Van91]{Vanthienen:1991:ALP} J.~Vanthienen. \
newblock Algorithm 124: a longest path algorithm to display
decision tables. \newblock {\em The Computer Journal}
, 34\penalty0 (4):\penalty0 358--362,   August 1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-

4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/358.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/359.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/360.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/362.tif|}\fi.
\bibitem[{van}93]{vanRijsbergen:1993:W} Keith {van Rijsbergen,
Editor-in-Chief}. \newblock Welcome. \newblock {\em The
Computer Journal}, 36\penalty0 (1):\penalty0 1, ???? 1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_01/Vol36_01.body.html#AbstractRijsbergen|}
\fi.
\bibitem[Vau95]{Vaughan:1995:HPD} John~G. Vaughan. \newblock
A hierarchical protocol for decentralising information
dissemination   in distributed systems. \newblock {\em
The Computer Journal}, 38\penalty0 (1):\penalty0 57--
70, ????   1995. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/1/57.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_01/
Vol38_01.body.html#AbstractVaughan|}\fi.
\bibitem[vBFvdW96]{vanBommel:1996:OOM} P.~van Bommel,
P.~J.~M. Frederiks, and Th.~P. van~der Weide. \newblock
Object-oriented modelling based on logbooks. \newblock
{\em The Computer Journal}, 39\penalty0 (9):\penalty0
793--799, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/9/793.full.pdf+html;   http:
//www.oup.co.uk/computer_journal/Volume_39/Issue_09/Vol39_
09.body.html#AbstractBommel|}\fi.
\bibitem[VCB93]{Vin:1993:CMD} H.~M. Vin, M.-S. Chen, and
T.~Barzilai. \newblock Collaboration management in {DiCE}
. \newblock {\em The Computer Journal}, 36\penalty0 (
1):\penalty0 87--96,   February 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/36/1/87.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_36/Issue_01/Vol36_01.body.html#AbstractVin|}\fi.
\bibitem[vdG96]{vanderGaag:1996:BBN} Linda~C. van~der
Gaag. \newblock {Bayesian} belief networks: Odds and ends.
\newblock {\em The Computer Journal}, 39\penalty0 (2)
:\penalty0 97--113, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/2/97.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_02/390097.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_02/Vol39_
02.body.html#AbstractGaag|}\fi.
\bibitem[vdL92]{vandeLiefvoort:1992:IAR} A.~van~de Liefvoort.
\newblock An iterative algorithm for the {Reve}'s puzzle.
\newblock {\em The Computer Journal}, 35\penalty0 (1)
:\penalty0 91--92,   February 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/1/91.
full.pdf+html|}\fi.
\bibitem[vdWHvB98]{vanderWeide:1998:ISS} Th.~P. van~der
Weide, T.~W.~C. Huibers, and P.~van Bommel. \newblock
The incremental searcher satisfaction model for information
 retrieval. \newblock {\em The Computer Journal}, 41\
penalty0 (5):\penalty0 311--318, ????   1998. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/41/5/
311.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_05/Vol41_05.body.html#AbstractVanDerWeide|}
\fi.
\bibitem[Vel93]{Veloso:1993:BR} P.~A.~S. Veloso. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (6):\penalty0 603, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/603.
1.full.pdf+html|}\fi.

\bibitem[Ver91]{Verrall:1991:UDI} M.~S. Verrall. \newblock
Unity doesn't imply unification or overcoming heterogeneity
problems   in distributed software engineering environments.
\newblock {\em The Computer Journal}, 34\penalty0 (6)
:\penalty0 522--533,   December 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/6/522.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/522.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/523.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/524.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/525.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/526.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/527.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/528.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/529.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_06/tiff/530.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
06/tiff/531.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/532.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/533.
t.
\bibitem[VG99]{Vovk:1999:CAP} Vladimir Vovk and Alexander
Gammerman. \newblock Complexity approximation principle.
\newblock {\em The Computer Journal}, 42\penalty0 (4)
:\penalty0 318--322, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/4/318.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420318.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420318.pdf|}\fi.
\bibitem[VH99]{Vee:1999:ACP} Voon-Yee Vee and Wen-Jing
Hsu. \newblock Applying {Cilk} in provably efficient task
scheduling. \newblock {\em The Computer Journal}, 42\
penalty0 (8):\penalty0 699--712, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL

 \path|http://comjnl.oxfordjournals.org/content/42/8/
699.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_08/420699.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_08/
pdf/420699.pdf|}\fi.
\bibitem[Vic94]{Vickers:1994:BR} S.~Vickers. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (9):\penalty0 818, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/9/818.
1.full.pdf+html|}\fi.
\bibitem[Vij95a]{Vijayaraman:1995:BRa} T.~M. Vijayaraman.
\newblock Book review. \newblock {\em The Computer Journal}
, 38\penalty0 (9):\penalty0 753, ???? 1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/9/
753.1.full.pdf+html|}\fi.
\bibitem[Vij95b]{Vijayaraman:1995:BRb} T.~M. Vijayaraman.
\newblock Book review. \newblock {\em The Computer Journal}
, 38\penalty0 (9):\penalty0 753--754, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/9/
753.2.full.pdf+html|}\fi.
\bibitem[Vin91]{Vince:1991:BR} J.~A. Vince. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (1):\penalty0 15, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/15.
full.pdf+html|}\fi.
\bibitem[Vin92]{Vince:1992:FC} John Vince. \newblock Fractal
creations. \newblock {\em The Computer Journal}, 35\penalty0
(1):\penalty0 40a--40,   February 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/1/40-
a.full.pdf+html|}\fi.
\bibitem[vLW93]{vanLeeuwen:1993:IH} J.~van Leeuwen and
D.~Wood. \newblock Interval heaps. \newblock {\em The

Computer Journal}, 36\penalty0 (3):\penalty0 209--216,
June   1993. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/36/3/209.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_03/Vol36_03.body.
html#AbstractLeeuwen|}\fi.
\bibitem[VM93]{Vazirgiannis:1993:OOM} M.~Vazirgiannis
and C.~Mourlas. \newblock An object-oriented model for
interactive multimedia presentations. \newblock {\em The
Computer Journal}, 36\penalty0 (1):\penalty0 78--86,
February 1993. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/1/78.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractVazirgiannis|}\fi.
\bibitem[VM94]{Vadera:1994:EFS} Sunil Vadera and Farid
Meziane. \newblock From {English} to formal specifications.
\newblock {\em The Computer Journal}, 37\penalty0 (9)
:\penalty0 753--763, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/9/753.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_09/Vol37_09.body.html#AbstractVadera|}
\fi.
\bibitem[vR92a]{vanRijsbergen:1992:PRR} C.~J. van Rijsbergen.
\newblock Probabilistic retrieval revisited. \newblock
{\em The Computer Journal}, 35\penalty0 (3):\penalty0
291--298, June   1992. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/3/291.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_03/
Vol35_03.body.html#AbstractRijsbergen|}\fi.
\bibitem[VR92b]{Vassallo:1992:ABS} M.~Vassallo and A.
~Ralston. \newblock Algorithms for {De Bruijn} sequences
--- {A} case study in the   empirical analysis of algorithms.
\newblock {\em The Computer Journal}, 35\penalty0 (1)
:\penalty0 88--90,   February 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/35/1/88.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_01/Vol35_01.body.html#AbstractVassallo|}
\fi.
\bibitem[vR93]{vanRijsbergen:1993:E} K.~van Rijsbergen.
\newblock Editorial. \newblock {\em The Computer Journal}
, 36\penalty0 (1):\penalty0 1, ???? 1993. \newblock \
showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/36/1/
1.full.pdf+html|}\fi.
\bibitem[vSFV92]{vanSinderen:1992:PDI} M.~van Sinderen,
L.~{Ferreira Pires}, and C.~A. Vissers. \newblock Protocol
design and implementation using formal methods. \newblock
{\em The Computer Journal}, 35\penalty0 (5):\penalty0
478--491,   October 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/5/478.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractSinderen|}\fi.
\bibitem[vSHBT98]{vanSteen:1998:ADG} Maarten van Steen,
Franz~J. Hauck, Gerco Ballintijn, and Andrew~S. Tanenbaum.
\newblock Algorithmic design of the globe wide-area location
service. \newblock {\em The Computer Journal}, 41\penalty0
(5):\penalty0 297--310, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/5/297.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_05/Vol41_05.body.html#AbstractVanSteen|}
\fi.
\bibitem[VTV92]{VanDerWeide:1992:UDS} Th.~P. {Van Der
Weide}, A.~H.~M. {Ter Hofstede}, and P.~{Van Bommel}.
\newblock {Uniquest}: Determining the semantics of complex
uniqueness   constraints. \newblock {\em The Computer
Journal}, 35\penalty0 (2):\penalty0 148--156, April
1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/2/148.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_02/Vol35_02.body.
html#AbstractWeide|}\fi.

\bibitem[VV94]{Varley:1994:RET} D.~A. Varley and M.~Visvalingam.
\newblock Road extraction and topographic data validation
using area topology. \newblock {\em The Computer Journal}
, 37\penalty0 (1):\penalty0 3--15, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/1/
3.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractVarley|}
\fi.
\bibitem[VVW96]{Visvalingam:1996:CAR} M.~Visvalingam,
D.~Varley, and C.~Wright. \newblock A cognitive approach
to road recognition with novel feature   indicators. \
newblock {\em The Computer Journal}, 39\penalty0 (9):
\penalty0 751--756, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/9/751.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_39/Issue_09/Vol39_09.body.html#AbstractVisvalingham;
 http://www3.oup.co.uk/computer_journal/Volume_39/Issue_
09/Vol39_09.body.html#AbstractVisvalingham|}\fi.
\bibitem[VW95a]{Verdon:1995:CBU} Frank~P. Verdon and Mike
Wells. \newblock Computing in {British} universities:
the {Computer Board} 1966--1991. \newblock {\em The Computer
Journal}, 38\penalty0 (10):\penalty0 822--830, ????
1995. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/10/822.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_38/Issue_10/Vol38_10.body.
html#AbstractVerdon|}\fi.
\bibitem[vW95b]{vonWright:1995:RHO} J.~von Wright. \newblock
Representing higher-order logic proofs in {HOL}. \newblock
{\em The Computer Journal}, 38\penalty0 (2):\penalty0
171--179, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/2/171.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_38/Issue_02/
Vol38_02.body.html#AbstractWright|}\fi.
\bibitem[VW96]{Verco:1996:PMC} K.~L. Verco and M.~J. Wise.
\newblock Plagiarism {\`a} la mode: a comparison of automated

systems for   detecting suspected plagiarism. \newblock
{\em The Computer Journal}, 39\penalty0 (9):\penalty0
741--750, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/9/741.full.pdf+html;   http:
//www.oup.co.uk/computer_journal/Volume_39/Issue_09/Vol39_
09.body.html#AbstractVerco;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_09/Vol39_09.body.html#AbstractVerco|}
\fi.
\bibitem[V'y99]{Vyugin:1999:ACS} V.~V. V'yugin. \newblock
Algorithmic complexity and stochastic properties of finite
binary   sequences. \newblock {\em The Computer Journal}
, 42\penalty0 (4):\penalty0 294--317, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/4/
294.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420294.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420294.pdf|}\fi.
\bibitem[Vyt94]{Vytopil:1994:BR} J.~Vytopil. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (8):\penalty0 732, ???? 1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/8/732.
1.full.pdf+html|}\fi.
\bibitem[WA91]{Wah:1991:SMM} W.~Wah and S.~G. Akl. \newblock
Simulating multiple memory accesses in logarithmic time
and linear   space. \newblock {\em The Computer Journal}
, 34\penalty0 (6):\penalty0 A085--A088,   December 1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}.
\bibitem[Wai93]{Waite:1993:C} W.~M. Waite. \newblock Correspondence.
\newblock {\em The Computer Journal}, 36\penalty0 (7)
:\penalty0 681, ???? 1993. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/7/681.2.full.pdf+html|}
\fi.

\bibitem[Wai97]{Waite:1997:BR} C.~Waite. \newblock Book
review. \newblock {\em The Computer Journal}, 40\penalty0
(5):\penalty0 296, ???? 1997. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/5/296.1.full.pdf+html|}
\fi.
\bibitem[Wak93]{Wakeman:1993:PVO} Ian Wakeman. \newblock
Packetized video: options for interaction between the
user, the   network and the codec. \newblock {\em The
Computer Journal}, 36\penalty0 (1):\penalty0 55--67,
February 1993. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/1/55.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_01/
Vol36_01.body.html#AbstractWakeman|}\fi.
\bibitem[Wal92a]{Wallis:1992:MSS} Peter Wallis. \newblock
Mathematical structures for software engineering. \newblock
{\em The Computer Journal}, 35\penalty0 (1):\penalty0
80, February   1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/1/80.full.pdf+html|}\fi.
\bibitem[Wal92b]{Wallis:1992:SRG} Peter Wallis. \newblock
Software reuse guidelines and method. \newblock {\em The
Computer Journal}, 35\penalty0 (6):\penalty0 668, December
 1992. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/35/6/668.full.pdf+html|}\fi.
\bibitem[Wal98]{Wallace:1998:ICS} C.~S. Wallace. \newblock
Intrinsic classification of spatially correlated data.
\newblock {\em The Computer Journal}, 41\penalty0 (8)
:\penalty0 602--611, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/8/602.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_08/Vol41_08.body.html#AbstractWallace|}
\fi.

\bibitem[War94]{Ward:1994:RET} M.~P. Ward. \newblock Reverse
engineering through formal transformation: {Knuth's
``Polynomial Addition''} algorithm. \newblock {\em The
Computer Journal}, 37\penalty0 (9):\penalty0 795--813,
????   1994. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/37/9/795.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_37/Issue_09/Vol37_09.body.
html#AbstractWard|}\fi.
\bibitem[War96]{Ward:1996:PAF} M.~P. Ward. \newblock Program
analysis by formal transformation. \newblock {\em The
Computer Journal}, 39\penalty0 (7):\penalty0 598--618,
????   1996. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/39/7/598.full.pdf+html;   http://www.oup.
co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_07/390598.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
Volume_39/Issue_07/Vol39_07.body.html#AbstractWard|}\
fi.
\bibitem[Wat91a]{Watson:1991:EDS} David Watson. \newblock
Editorial: Data structures. \newblock {\em The Computer
Journal}, 34\penalty0 (5):\penalty0 385, October   1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_05/tiff/385.tif|}\fi.
\bibitem[Wat91b]{Watson:1991:ECP} Ian Watson. \newblock
Editorial --- concurrent programming. \newblock {\em The
Computer Journal}, 34\penalty0 (4):\penalty0 289, August
 1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/34/4/289.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/289.
tif|}\fi.
\bibitem[Wat91c]{Watt:1991:MDP} D.~A. Watt. \newblock
Modular description of programming languages. \newblock
{\em The Computer Journal}, 34\penalty0 (6):\penalty0
A009--A028,   December 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}.

\bibitem[Wat95a]{Watson:1995:RE} Tim Watson. \newblock
Rules of encounter. \newblock {\em The Computer Journal}
, 38\penalty0 (3):\penalty0 268a--269, ????   1995. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/3/268-a.full.pdf+html|}\fi.
\bibitem[Wat95b]{Watt:1995:STP} David~A. Watt. \newblock
The structure of typed programming languages. \newblock
{\em The Computer Journal}, 38\penalty0 (3):\penalty0
262--263, ????   1995. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/38/3/262.full.pdf+html|}\fi.
\bibitem[Wat96a]{Watson:1996:BRa} T.~Watson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (1):\penalty0 93--94, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/1/
93.3.full.pdf+html|}\fi.
\bibitem[Wat96b]{Watson:1996:BRb} T.~Watson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (7):\penalty0 644, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/7/644.
1.full.pdf+html|}\fi.
\bibitem[Wat96c]{Watson:1996:BR} T.~Watson. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (8):\penalty0 724--725, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/8/
724.2.full.pdf+html|}\fi.
\bibitem[Wat96d]{Watt:1996:BR} D.~A. Watt. \newblock Book
review. \newblock {\em The Computer Journal}, 39\penalty0
(8):\penalty0 725, ???? 1996. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.

oxfordjournals.org/content/39/8/725.full.pdf+html|}\fi.
\bibitem[WB96]{Wellings:1996:PRS} A.~J. Wellings and A.
~Burns. \newblock Programming replicated systems in {
Ada 95}. \newblock {\em The Computer Journal}, 39\penalty0
(5):\penalty0 361--373, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/5/361.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_05/390361.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_05/Vol39_
05.body.html#AbstractWellings|}\fi.
\bibitem[WB98]{Witter:1998:DLS} Dian~I. Witter and Michael~
W. Berry. \newblock Downdating the latent semantic indexing
model for conceptual   information retrieval. \newblock
{\em The Computer Journal}, 41\penalty0 (8):\penalty0
589--601, ????   1998. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/41/8/589.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_41/Issue_08/
Vol41_08.body.html#AbstractWitter|}\fi.
\bibitem[WB99]{Ward:1999:RRI} Martin~P. Ward and Keith~
H. Bennett. \newblock Recursion removal\slash introduction
by formal transformation: An aid   to program development
and program comprehension. \newblock {\em The Computer
Journal}, 42\penalty0 (8):\penalty0 650--673, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/8/650.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_08/420650.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_08/pdf/420650.pdf|}\fi.
\bibitem[WBM{\etalchar{+}}94]{Witten:1994:SGM} Ian~H.
Witten, Timothy~C. Bell, Alistair Moffat, Craig~G. Nevill-
Manning,   Tony~C. Smith, and Harold Thimbleby. \newblock
Semantic and generative models for lossy text compression.
\newblock {\em The Computer Journal}, 37\penalty0 (2)
:\penalty0 83--87, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \

path|http://comjnl.oxfordjournals.org/content/37/2/83.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_02/Vol37_02.body.html#AbstractWitten|}
\fi.
\bibitem[WC91]{Wu:1991:SNN} J.-L. Wu and S.-T. Cheng.
\newblock Short note: a neural network simulating system.
\newblock {\em The Computer Journal}, 34\penalty0 (1)
:\penalty0 92--93,   February 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/1/92.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/93.
tif|}\fi.
\bibitem[WC94]{Wade:1994:DCG} Greg~A. Wade and Jiang-
Hsing Chu. \newblock Drawability of complete graphs using
a minimal slope set. \newblock {\em The Computer Journal}
, 37\penalty0 (2):\penalty0 139--142, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/2/
139.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_02/Vol37_02.body.html#AbstractWade|}\
fi.
\bibitem[WC98]{Wu:1998:DBG} I-Chen Wu and Loon-Been Chen.
\newblock On detection of bounded global predicates. \
newblock {\em The Computer Journal}, 41\penalty0 (4):
\penalty0 231--237, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/4/231.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_04/Vol41_04.body.html#AbstractWu2|}\fi.
\bibitem[WD96]{Wright:1996:BR} P.~Wright and A.~Dearden.
\newblock Book review. \newblock {\em The Computer Journal}
, 39\penalty0 (9):\penalty0 815--816, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/9/
815.2.full.pdf+html|}\fi.

\bibitem[WD99a]{Wallace:1999:MML} C.~S. Wallace and D.
~L. Dowe. \newblock Minimum message length and {Kolmogorov}
complexity. \newblock {\em The Computer Journal}, 42\
penalty0 (4):\penalty0 270--283, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/4/
270.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420270.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420270.pdf|}\fi.
\bibitem[WD99b]{Wallace:1999:RMM} C.~S. Wallace and D.
~L. Dowe. \newblock Refinements of {MDL} and {MML} coding.
\newblock {\em The Computer Journal}, 42\penalty0 (4)
:\penalty0 330--337, ????   1999. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/42/4/330.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_04/420330.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_04/
pdf/420330.pdf|}\fi.
\bibitem[WD99c]{Wallace:1999:R} C.~S. Wallace and D.~
L. Dowe. \newblock Rejoinder. \newblock {\em The Computer
Journal}, 42\penalty0 (4):\penalty0 345--347, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/4/345.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_04/420345.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_04/pdf/420345.pdf|}\fi.
\bibitem[Weg93]{Wegener:1993:SMX} I.~Wegener. \newblock
A simple modification of {Xunrang} and {Yuzhang}'s {HEAPSORT}
variant   improving its complexity significantly. \newblock
{\em The Computer Journal}, 36\penalty0 (3):\penalty0
286--288, June   1993. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/3/286.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractWegener|}\fi. \newblock See
\cite{Xunrang:1990:SNN}.

\bibitem[Wei91]{Weiss:1991:SNE} M.~A. Weiss. \newblock
Short note: Empirical study of the expected running time
of   {Shellsort}. \newblock {\em The Computer Journal}
, 34\penalty0 (1):\penalty0 88--91,   February 1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/1/88.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_01/tiff/88.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/89.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/90.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_01/tiff/91.
tif|}\fi.
\bibitem[Wei93]{Weiss:1993:FHB} M.~A. Weiss. \newblock
On finding the height of a binary search tree. \newblock
{\em The Computer Journal}, 36\penalty0 (3):\penalty0
280--281, June   1993. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/36/3/280.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/Volume_36/Issue_03/
Vol36_03.body.html#AbstractWeiss|}\fi.
\bibitem[WFD98]{Wu:1998:OFS} Jie Wu, Eduardo~B. Fernandex,
and Donglai Dai. \newblock Optimal fault-secure scheduling.
\newblock {\em The Computer Journal}, 41\penalty0 (4)
:\penalty0 207--222, ????   1998. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/41/4/207.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_04/Vol41_04.body.html#AbstractWu1|}\fi.
\bibitem[WG92]{Wichmann:1992:E} Brian Wichmann and Philip
Giles. \newblock Editorial. \newblock {\em The Computer
Journal}, 35\penalty0 (1):\penalty0 1--2, February   1992.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/1/1.full.pdf+html|}\fi.
\bibitem[Wha90]{Whale:1990:IPS} G.~Whale. \newblock Identification
of program similarity in large populations. \newblock
{\em The Computer Journal}, 33\penalty0 (2):\penalty0

140--146, April   1990. \newblock \showCODEN{CMPJA6}.
\newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/33/2/140.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/140.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/141.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/142.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/143.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/144.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/145.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/146.tif|}\fi.
\bibitem[Wic91]{Wichmann:1991:BR} B.~Wichmann. \newblock
Book review. \newblock {\em The Computer Journal}, 34\
penalty0 (5):\penalty0 422, ???? 1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/5/422.
2.full.pdf+html|}\fi.
\bibitem[Wic92]{Wichmann:1992:NUF} B.~A. Wichmann. \newblock
A note on the use of floating point in critical systems.
\newblock {\em The Computer Journal}, 35\penalty0 (1)
:\penalty0 41--44,   February 1992. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/35/1/41.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_35/Issue_01/Vol35_01.body.html#AbstractWichmann|}
\fi.
\bibitem[Wic94]{Wichmann:1994:BR} B.~Wichmann. \newblock
Book review. \newblock {\em The Computer Journal}, 37\
penalty0 (5):\penalty0 470--471, ????   1994. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/37/5/
470.2.full.pdf+html|}\fi.
\bibitem[Wie95]{Wieringa:1995:CSD} R.~J. Wieringa. \newblock
Combining static and dynamic modelling methods: a comparison
of four   methods. \newblock {\em The Computer Journal}
, 38\penalty0 (1):\penalty0 17--30, ????   1995. \newblock

\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/1/
17.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_38/Issue_01/Vol38_01.body.html#AbstractWieringa|}
\fi.
\bibitem[Wil90]{Wills:1990:FPS} H.~Wills. \newblock Fundamentals
of pricing and scheduling computer services and   investment
in computer equipment. \newblock {\em The Computer Journal}
, 33\penalty0 (3):\penalty0 266--273, June   1990. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/33/3/
266.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/266.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/267.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/268.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/269.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/270.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/271.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/272.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_03/tiff/273.tif|}\fi.
\bibitem[Wil91a]{Wilkes:1991:EMV} M.~V. Wilkes. \newblock
Editorial --- the maturing of {VLSI} --- and its consequences
for the   computer industry. \newblock {\em The Computer
Journal}, 34\penalty0 (3):\penalty0 193--194, June   1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/3/193.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/194.tif|}\fi.
\bibitem[Wil91b]{Wilkinson:1991:CPO} B.~Wilkinson. \newblock
Comparative performance of overlapping connectivity multiprocessor
 interconnection networks. \newblock {\em The Computer
Journal}, 34\penalty0 (3):\penalty0 207--214, June   1991.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/

34/3/207.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/207.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/208.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/209.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/210.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/211.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/212.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
03/tiff/213.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/214.tif|}\fi.
\bibitem[Wil95]{Wilson:1995:SVH} R.~Wilson. \newblock
Spatial vision in humans and robots. \newblock {\em The
Computer Journal}, 38\penalty0 (1):\penalty0 84, ????
1995. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}. \newblock
\ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/1/84.full.pdf+html|}\fi.
\bibitem[Wil96a]{Willett:1996:BR} P.~Willett. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (3):\penalty0 253, ???? 1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/3/253.
1.full.pdf+html|}\fi.
\bibitem[Wil96b]{Williams:1996:BR} M.~R. Williams. \newblock
Book review. \newblock {\em The Computer Journal}, 39\
penalty0 (10):\penalty0 890--891, ????   1996. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/39/10/
890.full.pdf+html|}\fi.
\bibitem[Wis90]{Wiseman:1990:BR} N.~E. Wiseman. \newblock
Book review. \newblock {\em The Computer Journal}, 33\
penalty0 (5):\penalty0 436, ???? 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/5/436.
full.pdf+html|}\fi.
\bibitem[WJW95]{Wang:1995:PAC} Hu~Jun Wang, Yuan-Da Jiang,
and Hui-Zhong Wu. \newblock Performance analysis of cluster-

based {PPMB} multiprocessor systems. \newblock {\em The
Computer Journal}, 38\penalty0 (5):\penalty0 355--364,
????   1995. \newblock \showCODEN{CMPJA6}. \newblock \
showISSN{0010-4620 (print), 1460-2067 (electronic)}. \
newblock \ifshowURL {\showURL   \path|http://comjnl.oxfordjournals.
org/content/38/5/355.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_38/Issue_05/Vol38_05.body.
html#AbstractWang|}\fi.
\bibitem[WL96]{Wang:1996:BAC} Hu-Jun Wang and Jian Li.
\newblock Bandwidth analysis for {A} class of bus-based
systems. \newblock {\em The Computer Journal}, 39\penalty0
(4):\penalty0 346--352, ????   1996. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/39/4/346.
full.pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_04/390346.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_04/Vol39_
04.body.html#AbstractWang|}\fi.
\bibitem[WM95]{Wilson:1995:IAC} Stephen~P. Wilson and
John~A. McDermid. \newblock Integrated analysis of complex
safety critical systems. \newblock {\em The Computer Journal}
, 38\penalty0 (10):\penalty0 765--776, ????   1995. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/10/765.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_38/Issue_10/Vol38_10.body.html#AbstractWilson|}
\fi.
\bibitem[Wol90]{Wolff:1990:SPS} J.~G. Wolff. \newblock
Simplicity and power --- some unifying ideas in computing.
\newblock {\em The Computer Journal}, 33\penalty0 (6)
:\penalty0 518--534,   December 1990. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/33/6/518.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/518.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/519.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/520.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/521.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_

06/tiff/522.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/523.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/524.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/525.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/526.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/527.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/528.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/529.
t.
\bibitem[Wol94]{Wolff:1994:DAO} Gerry Wolff. \newblock
Data abstraction, the object-oriented approach using {
C}++. \newblock {\em The Computer Journal}, 37\penalty0
(6):\penalty0 544, ???? 1994. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/37/6/544.full.pdf+html|}\fi.
\bibitem[Wol95]{Wolff:1995:WSR} J.~Gerard Wolff. \newblock
`what is sp?': a reply. \newblock {\em The Computer Journal}
, 38\penalty0 (3):\penalty0 253--256, ????   1995. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/38/3/
253.full.pdf+html|}\fi.
\bibitem[Won99]{Wong:1999:SLS} W.~F. Wong. \newblock Source
level static branch prediction. \newblock {\em The Computer
Journal}, 42\penalty0 (2):\penalty0 142--149, ????   1999.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
42/2/142.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_42/Issue_02/420142.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_42/
Issue_02/pdf/420142.pdf|}\fi.
\bibitem[Woo92]{Woodcock:1992:RAR} J.~C.~P. Woodcock.
\newblock The rudiments of algorithm refinement. \newblock
{\em The Computer Journal}, 35\penalty0 (5):\penalty0
441--450,   October 1992. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/35/5/441.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/Volume_35/Issue_05/
Vol35_05.body.html#AbstractWoodcock|}\fi. \newblock See
\cite{Cooke:1992:EFMa}.
\bibitem[Woo93]{Woodman:1993:BR} M.~Woodman. \newblock
Book review. \newblock {\em The Computer Journal}, 36\
penalty0 (6):\penalty0 599, ???? 1993. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/36/6/599.
1.full.pdf+html|}\fi.
\bibitem[Wor94]{Worboys:1994:UMS} Michael~F. Worboys.
\newblock A unified model for spatial and temporal information.
\newblock {\em The Computer Journal}, 37\penalty0 (1)
:\penalty0 26--34, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/1/26.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractWorboys|}
\fi.
\bibitem[WTW91]{Wicovsky:1991:SNI} K.~Wicovsky, P.~F.
Thanisch, and M.~H. Williams. \newblock Short notes: Improved
recursion handling through integrity   constraints. \
newblock {\em The Computer Journal}, 34\penalty0 (3):
\penalty0 282--285, June   1991. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/34/3/282.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_03/tiff/282.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_03/tiff/283.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/284.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_03/tiff/285.tif|}\fi.
\bibitem[Wu96]{Wu:1996:BDR} Xindong Wu. \newblock A {
Bayesian} discretizer for real-valued attributes. \newblock
{\em The Computer Journal}, 39\penalty0 (8):\penalty0
688--691, ????   1996. \newblock \showCODEN{CMPJA6}. \
newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/39/8/688.full.pdf+html;   http:
//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
08/390688.sgm.abs.html;   http://www3.oup.co.uk/computer_

journal/Volume_39/Issue_08/Vol39_08.body.html#AbstractWu|}
\fi.
\bibitem[Wu97]{Wu:1997:HIN} X.~Wu. \newblock Hybrid interpretation
of `no match' and `multiple match' in   induction. \newblock
{\em The Computer Journal}, 40\penalty0 (1):\penalty0
50--57, ????   1997. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.
\newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/40/1/50.full.pdf+html;   http:
//www.oup.co.uk/computer_journal/Volume_40/Issue_01/Vol40_
01.body.html#AbstractWu;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_01/Vol40_01.body.html#AbstractWu|}
\fi.
\bibitem[WWK{\etalchar{+}}90]{Waugh:1990:DSE} K.~G. Waugh,
M.~H. Williams, Q.~Kong, S.~Salvini, and G.~Chen. \newblock
Designing {SQUIRREL}: an extended {SQL} for a deductive
database   system. \newblock {\em The Computer Journal}
, 33\penalty0 (6):\penalty0 535--546,   December 1990.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/6/535.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/536.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/537.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/538.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/539.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/540.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/541.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/542.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_06/tiff/543.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_06/tiff/544.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_06/tiff/545.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
06/tiff/546.t.
\bibitem[WWN91]{Wang:1991:SGM} C.~J. Wang, C.~H. Wu, and
V.~P. Nelson. \newblock A study of the generalized multiple
bus-connected parallel computer. \newblock {\em The Computer
Journal}, 34\penalty0 (6):\penalty0 A089--A094,   December

1991. \newblock \showCODEN{CMPJA6}. \newblock \showISSN{
0010-4620 (print), 1460-2067 (electronic)}.
\bibitem[WZ99]{Williams:1999:CIF} Hugh~E. Williams and
Justin Zobel. \newblock Compressing integers for fast
file access. \newblock {\em The Computer Journal}, 42\
penalty0 (3):\penalty0 193--201, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/3/
193.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_03/420193.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_03/
pdf/420193.pdf|}\fi.
\bibitem[XTU97]{Xiang:1997:GOE} L.~Xiang, C.~Tang, and
K.~Ushijima. \newblock Grammar-oriented enumeration of
binary trees. \newblock {\em The Computer Journal}, 40\
penalty0 (5):\penalty0 278--291, ????   1997. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/40/5/
278.full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_05/Vol40_05.body.html#AbstractXiang;
http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
05/Vol40_05.body.html#AbstractXiang|}\fi.
\bibitem[XY90]{Xunrang:1990:SNN} G.~Xunrang and Z.~Yuzhang.
\newblock Short note: a new {HEAPSORT} algorithm and the
analysis of its   complexity. \newblock {\em The Computer
Journal}, 33\penalty0 (3):\penalty0 281--282, June   1990.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
33/3/281.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_03/tiff/281.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
03/tiff/282.tif|}\fi. \newblock See modification \cite{
Wegener:1993:SMX}.
\bibitem[YAB92]{Yannakoudakis:1992:CNL} E.~J. Yannakoudakis
and H.~A. Attar-Bashi. \newblock Constructing natural
language sentences from database tuples. \newblock {\
em The Computer Journal}, 35\penalty0 (5):\penalty0 A445-
-??,   October 1992. \newblock \showCODEN{CMPJA6}. \newblock
\showISSN{0010-4620 (print), 1460-2067 (electronic)}.

\newblock \ifshowURL {\showURL   \path|http://www3.oup.
co.uk/computer_journal/Volume_35/Issue_05/35000A445.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/Volume_
35/Issue_05/tiff/A445.tif|}\fi.
\bibitem[YCC92]{Yang:1992:MSF} J.~S.~H. Yang, Y.~H. Chin,
and C.~G. Chung. \newblock Many-sorted first-order logic
database language. \newblock {\em The Computer Journal}
, 35\penalty0 (2):\penalty0 129--137, April   1992. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
35/2/129.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_02/Vol35_02.body.html#AbstractYang|}
\fi.
\bibitem[YG93]{Yau:1993:PLD} S.~S. Yau and J.~V. Gore.
\newblock Paradigm lost: discovering the intersection
of assertion-guided   program construction and constraint-
driven re-evaluation. \newblock {\em The Computer Journal}
, 36\penalty0 (2):\penalty0 137--142, April   1993. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/2/137.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_02/Vol36_02.body.html#AbstractYau|}
\fi.
\bibitem[YJ91]{Yue:1991:ESF} K-B. Yue and R.~T. Jacob.
\newblock An efficient starvation-free semaphore solution
for the graphical   mutual exclusion problem. \newblock
{\em The Computer Journal}, 34\penalty0 (4):\penalty0
345--349,   August 1991. \newblock \showCODEN{CMPJA6}
. \newblock \showISSN{0010-4620 (print), 1460-2067 (electronic)
}. \newblock \ifshowURL {\showURL   \path|http://comjnl.
oxfordjournals.org/content/34/4/345.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/345.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/346.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/347.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/348.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/349.tif|}\fi.
\bibitem[YJ94]{Yue:1994:OAR} Kwok-Bun Yue and R.~T. Jacob.
\newblock An optimal algorithm for reducing edge-solvable
mutual exclusion   graphs. \newblock {\em The Computer

Journal}, 37\penalty0 (2):\penalty0 129--138, ????   1994.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
37/2/129.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_02/Vol37_02.body.html#AbstractYue|}
\fi.
\bibitem[Yok97]{Yokoo:1997:DCU} Hidetoshi Yokoo. \newblock
Data compression using a sort-based context similarity
measure. \newblock {\em The Computer Journal}, 40\penalty0
(2--3):\penalty0 94--102,   ???? 1997. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/40/2_and_
3/94.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_40/Issue_02/Vol40_02.body.html#AbstractYokoo;
 http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_03.body.html#AbstractYokoo|}\fi.
\bibitem[YS91]{Yannakoudakis:1991:CEM} E.~J. Yannakoudakis
and A.~A. Shahid. \newblock A controlled environment for
meta-message processing. \newblock {\em The Computer Journal}
, 34\penalty0 (4):\penalty0 331--344,   August 1991. \
newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
34/4/331.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/331.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/332.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/333.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/334.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/336.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/338.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/339.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_04/tiff/340.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/341.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/342.t.

\bibitem[ZG99]{Zhou:1999:MRR} Wanlei Zhou and Andrzej
Goscinski. \newblock Managing replicated remote procedure
call transactions. \newblock {\em The Computer Journal}
, 42\penalty0 (7):\penalty0 592--608, ????   1999. \newblock
\showCODEN{CMPJA6}. \newblock \showISSN{0010-4620 (print)
, 1460-2067 (electronic)}. \newblock \ifshowURL {\showURL
 \path|http://comjnl.oxfordjournals.org/content/42/7/
592.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_42/Issue_07/420592.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_42/Issue_07/
pdf/420592.pdf|}\fi.
\bibitem[ZH93]{Zviran:1993:CPT} M.~Zviran and W.~J. Haga.
\newblock A comparison of password techniques for multilevel
authentication   mechanisms. \newblock {\em The Computer
Journal}, 36\penalty0 (3):\penalty0 227--237, June   1993.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
36/3/227.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_03/Vol36_03.body.html#AbstractZviran|}
\fi.
\bibitem[ZHS94]{Zheng:1994:RSS} Yuliang Zheng, T.~Hardjono,
and J.~Seberry. \newblock Reusing shares in secret sharing
schemes. \newblock {\em The Computer Journal}, 37\penalty0
(3):\penalty0 199--205, ????   1994. \newblock \showCODEN{
CMPJA6}. \newblock \showISSN{0010-4620 (print), 1460-
2067 (electronic)}. \newblock \ifshowURL {\showURL   \
path|http://comjnl.oxfordjournals.org/content/37/3/199.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_03/Vol37_03.body.html#AbstractZheng|}
\fi.
\bibitem[ZM95]{Zhang:1995:VGV} Kang Zhang and Gaurav Marwaha.
\newblock {Visputer} --- a graphical visualization tool
for parallel   programming. \newblock {\em The Computer
Journal}, 38\penalty0 (8):\penalty0 658--669, ????   1995.
\newblock \showCODEN{CMPJA6}. \newblock \showISSN{0010-
4620 (print), 1460-2067 (electronic)}. \newblock \ifshowURL
{\showURL   \path|http://comjnl.oxfordjournals.org/content/
38/8/658.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_38/Issue_08/Vol38_08.body.html#AbstractZhang|}
\fi.
\bibitem[ZS96]{Zhang:1996:ANR} N.~Zhang and Q.~Shi. \
newblock Achieving non-repudiation of receipt. \newblock